### v1

# Deformation quantization and perverse sheaves

Abstract.
Kashiwara, Polesello, Schapira and D’Agnolo defined canonical deformation quantizations of a holomorphic symplectic manifold and a holomorphic Lagrangian submanifold equipped with an orientation data. The goal of this paper is to describe the RHom complex between two such deformation quantization modules associated to a pair of Lagrangian submanifolds in terms of the derived geometry of the Lagrangian intersection. Namely, we identify the RHom complex with the DT sheaf associated to the d-critical structure on the Lagrangian intersection.
Introduction
Deformation quantization modules
Let $Y$ be a holomorphic contact manifold. Kashiwara [Kas96] has defined a filtered $\mathbf{C}$-algebroid $\mathsf{E}_{Y}$ on $Y$. It provides a deformation quantization of $Y$ in the sense that $\operatorname{gr}\mathsf{E}_{Y}$ is isomorphic to $\oplus_{m\in\mathbf{Z}}\mathcal{O}_{Y}(m)$ as a sheaf of Poisson algebras. Locally $Y$ is isomorphic to an open subset of the projectivized cotangent bundle $\mathrm{P}^{*}X$ of a complex manifold $X$ in which case $\mathsf{E}_{Y}$ reduces to the algebra of microdifferential operators introduced by Sato, Kawai and Kashiwara [SKK73]. Given a holomorphic Legendrian submanifold $\Lambda\subset Y$ there is, moreover, a classification of regular holonomic $\mathsf{E}_{Y}$-modules supported on $\Lambda$ (see [Kas96, Proposition 4] and its corrected version [DS07, Corollary 6.4]) in terms of twisted local systems on the $\mathbf{C}^{\times}$-torsor $\tilde{\Lambda}\rightarrow\Lambda$ constructed from the symplectization $\tilde{Y}\rightarrow Y$.
These results were generalized from the contact to the symplectic setting in [PS04, DS07]. Namely, given a holomorphic symplectic manifold $S$ Polesello and Schapira [PS04] have defined a $\mathbf{C}(\!(\hbar)\!)$-algebroid $\mathsf{W}_{S}$ together with a $\mathbf{C}[\![\hbar]\!]$-lattice $\mathsf{W}_{S}(0)\subset\mathsf{W}_{S}$. It again provides a deformation quantization of $S$ in the sense that $\mathsf{W}_{S}(0)/\hbar$ is isomorphic to $\mathcal{O}_{S}$ as a sheaf of Poisson algebras. The construction of $\mathsf{W}_{S}$ is based on the following property: given a contactification $\rho\colon Y\rightarrow S$ of the symplectic manifold $S$, there is an embedding $\mathsf{W}_{S}\subset\rho_{*}\mathsf{E}_{Y}$ which identifies $\mathsf{W}_{S}$ with microdifferential operators $\mathsf{E}_{Y}$ on $Y$ commuting with $\partial_{t}$ (the translation vector field along the fibers of $\rho$). Under this correspondence $\hbar=\partial_{t}^{-1}$. Given a holomorphic Lagrangian submanifold $L\subset S$ equipped with a twisted local system and a number $\lambda\in\mathbf{C}/\mathbf{Z}$ there is a regular holonomic $\mathsf{W}_{S}$-module supported on $L$. In a contactification $\rho\colon Y\rightarrow S$ with a lift of $L\subset S$ to a holomorphic Legendrian $\Lambda\subset Y$ it corresponds to a regular holonomic $\mathsf{E}_{Y}$-module supported on $\Lambda$ associated to a twisted local system on $\tilde{\Lambda}\cong L\times\mathbf{C}^{\times}$ with the monodromy $\exp(2\pi i\lambda)$ along the $\mathbf{C}^{\times}$ direction. In particular, if $L$ is equipped with a square root $K_{L}^{1/2}$ of the canonical bundle, there is a simple $\mathsf{W}_{S}$-module $\mathsf{M}^{\lambda}_{L}$ supported on $L$.
Kashiwara and Schapira [KS12] have developed a general theory of deformation quantization modules. In particular, they have shown (see [KS12, Theorem 7.2.3]) that for any regular holonomic $\mathsf{W}_{S}$-modules $\mathcal{M}_{1},\mathcal{M}_{2}$ the complexes of sheaves $\mathbb{R}\mathcal{H}om_{\mathsf{W}_{S}}(\mathcal{M}_{1},\mathcal{M}_{2})[\dim S/2]$ and $\mathcal{M}_{1}\otimes^{\mathbb{L}}_{\mathsf{W}_{S}}\mathcal{M}_{2}$ are perverse sheaves of $\mathbf{C}(\!(\hbar)\!)$-vector spaces. The main goal of this paper is to describe these perverse sheaves in terms of the derived geometry of the Lagrangian intersection $L\cap M$ of the supports of $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ when both $L$ and $M$ are smooth holomorphic Lagrangian submanifolds of $S$.
Microlocal sheaves and Fukaya categories
In this section we outline connections between the theory of deformation quantization modules and Fukaya categories, on the one hand, and microlocal sheaves, on the other hand, to place our results in a broader context.
As the first goal let us describe the relevant category of deformation quantization modules associated to a holomorphic symplectic manifold $S$ (we review this construction in Section 4). Fix a (possibly singular) complex Lagrangian subvariety $L\subset S$. There is a canonical contactification $\rho\colon Y\rightarrow S$ of a neighborhood of $L$ with a lift of $L\subset S$ to a Legendrian subvariety $\Lambda\subset Y$ (see 1.12). As in [DK11] we define the abelian stack $\mathrm{RH}_{S,L}$ (i.e. a sheaf of $\mathbf{C}$-linear abelian categories) of regular holonomic DQ modules on $S$ supported on $L$ to be the stack of regular holonomic $\mathsf{E}_{Y}$-modules supported on $\Lambda$, i.e.
$\mathrm{RH}_{S,L}=\rho_{*}\mathrm{Mod}_{\Lambda,\mathrm{rh}}(\mathsf{E}_{Y}).$
Lagrangian subvarieties of $S$ form a filtered system and we define $\mathrm{RH}_{S}$ as the filtered colimit
$\mathrm{RH}_{S}=\operatorname*{colim}_{L}\mathrm{RH}_{S,L}.$
In particular, if $L\subset S$ is a smooth Lagrangian submanifold equipped with a square root line bundle $K_{L}^{1/2}$ (orientation data) and $\lambda\in\mathbf{C}/\mathbf{Z}$ we have the objects $\mathsf{M}^{\lambda}_{L}\in\mathrm{RH}_{S,L}$ mentioned before.
Let us now describe a topological avatar of the stack $\mathrm{RH}_{S}$. Let $W$ be an exact holomorphic symplectic manifold. The authors of [NS20, Côt+22] have introduced a sheaf of stable $\infty$-categories $\mu\mathrm{sh}_{W,\mathbf{C}\text{-}c}$ on $W$ by globalizing the microlocal theory of sheaves of Kashiwara and Schapira [KS94]. For a holomorphic contact manifold $Y$ one then defines the stack $\mathbb{P}\mu\mathrm{sh}_{Y,\mathbf{C}\text{-}c}=\gamma_{*}\mu\mathrm{sh}_{\tilde{Y},\mathbf{C}\text{-}c}$, where $\gamma\colon\tilde{Y}\rightarrow Y$ is the symplectization. One has
$\mathbb{P}\mu\mathrm{sh}_{Y,\mathbf{C}\text{-}c}=\operatorname*{colim}_{\Lambda}\mathbb{P}\mu\mathrm{sh}_{\Lambda},$
where $\mathbb{P}\mu\mathrm{sh}_{\Lambda}\subset\mathbb{P}\mu\mathrm{sh}_{Y,\mathbf{C}\text{-}c}$ is the subcategory of objects supported on a Legendrian subvariety $\Lambda\subset Y$. Moreover, the authors of [Côt+22] construct a $t$-structure on $\mathbb{P}\mu\mathrm{sh}_{Y,\mathbf{C}\text{-}c}$ (respectively, $\mathbb{P}\mu\mathrm{sh}_{\Lambda}$) whose heart is the stack $\mu\mathrm{Perv}_{Y}$ (respectively, $\mu\mathrm{Perv}_{Y,\Lambda}$) of microlocal perverse sheaves as in [And96, Was04].
For $Y=\mathrm{P}^{*}X$ a projectivized cotangent bundle and $\Lambda\subset Y$ a fixed Legendrian subvariety Waschkies [Was05] has defined a microlocal Riemann–Hilbert correspondence which identifies
$\mu\mathrm{Perv}_{Y}\cong\mathrm{Mod}_{\mathrm{rh}}(\mathsf{E}_{Y}),\qquad\mu\mathrm{Perv}_{Y,\Lambda}\cong\mathrm{Mod}_{\Lambda,\mathrm{rh}}(\mathsf{E}_{Y}),$
which explains a connection between the stack $\mathrm{RH}_{S}$ and the stacks of microlocal sheaves.
If the exact holomorphic symplectic manifold $\tilde{Y}$ is Weinstein, then Ganatra, Pardon and Shende [GPS18] identify $\mu\mathrm{sh}_{\tilde{Y}}$ with the ind-completion of the wrapped Fukaya category of $\tilde{Y}$. Thus, we have three independent ways to approach the stack $\mathrm{RH}_{S}$: in terms of deformation quantization modules, in terms of microlocal sheaf theory and in terms of Fukaya categories.
D-critical loci and DQ modules
The relevant derived structure on a Lagrangian intersection that we will need is the structure of a d-critical locus introduced in [Joy15]. We recall the definition in Section 6.2, but let us simply mention that a d-critical locus is locally isomorphic to the critical locus $\mathrm{Crit}(f)$ of a holomorphic function $f\colon U\rightarrow\mathbf{C}$ on a complex manifold $U$ as a complex-analytic space. Any d-critical locus has an orientation $\mathbf{Z}/2$-gerbe whose trivialization is the choice of an orientation data. The key result of [Bra+15] is a construction of a canonical perverse sheaf $\phi_{X}$ (of $\mathbf{C}$-vector spaces), which we call the DT sheaf, on an oriented d-critical locus globalizing the sheaf of vanishing cycles $\phi_{f}$. (More canonically, we can think of $\phi_{X}$ as a sheaf twisted by the orientation gerbe similar to the fundamental class of a closed manifold being a homology class twisted by the orientation local system.)
As explained in [Bus14], if $S$ is a holomorphic symplectic manifold and $L,M\subset S$ are holomorphic Lagrangian submanifolds, the Lagrangian intersection $L\cap M$ has a canonical structure of a d-critical locus. Moreover, if $L$ and $M$ are equipped with square roots $K_{L}^{1/2},K_{M}^{1/2}$ of the canonical bundles, then the d-critical locus $L\cap M$ is oriented; in particular, there is the DT sheaf $\phi_{L\cap M}$ on the Lagrangian intersection. We may then state the main result (7.1) as follows.
Theorem A.
For any $\lambda,\mu\in\mathbf{C}$ there is an isomorphism of $\mathbf{C}(\!(\hbar)\!)$-linear perverse sheaves
$\mathsf{M}^{\lambda}_{L}\otimes^{\mathbb{L}}_{\mathsf{W}_{S}}\mathsf{M}^{\mu}_{M}\cong\phi_{L\cap M}\otimes_{\mathbf{C}}\mathbf{C}(\!(\hbar)\!).$
When the Lagrangians $L,M$ intersect cleanly, the right-hand side reduces to an explicit $\mathbf{Z}/2$-torsor described in Section 7.4. This fact should be compared with a similar computation of Lagrangian Floer homology of clean intersections [Poź99, Sch16].
In fact, we prove a strengthening of the above statement which explains the origin of the $\hbar$ parameter. For this, recall that the DT sheaf $\phi_{L\cap M}$ carries a canonical monodromy automorphism. The Riemann–Hilbert correspondence identifies perverse sheaves $P$ of $\mathbf{C}$-vector spaces with an automorphism and perverse sheaves $\mathrm{RH}^{-1}(P)$ of $\mathbf{C}(\!(\hbar)\!)$-vector spaces equipped with a connection in the $\hbar$-direction with regular singularities (see 2.11). We call perverse sheaves of $\mathbf{C}(\!(\hbar)\!)$-vector spaces equipped with connections in the $\hbar$-direction differential perverse sheaves. We also equip the derived tensor product $\mathsf{M}^{\lambda}_{L}\otimes^{\mathbb{L}}_{\mathsf{W}_{S}}\mathsf{M}^{\mu}_{M}$ with a natural connection in the $\hbar$-direction coming from the fact that the $\mathsf{W}_{S}$-modules $\mathsf{M}^{\lambda}_{L},\mathsf{M}^{\mu}_{M}$ locally come from $\mathsf{E}_{Y}$-modules on a contactification. The stronger statement of the main result (7.1) is then as follows.
Theorem B.
For any $\lambda,\mu\in\mathbf{C}$ there is an isomorphism of differential perverse sheaves
$\mathsf{M}^{\lambda}_{L}\otimes^{\mathbb{L}}_{\mathsf{W}_{S}}\mathsf{M}^{\mu}_{M}\cong\mathrm{RH}^{-1}(\phi_{L\cap M}\otimes\mathbf{C}_{\mu+\lambda}),$
where $\mathbf{C}_{\mu+\lambda}$ is the trivial one-dimensional vector space equipped with the automorphism $\exp(2\pi i(\mu+\lambda))$.
One motivation for this statement is a proposal [Bra+15, Remark 6.15] for a Fukaya-like category associated to a holomorphic symplectic manifold $S$:
  • Holomorphic Lagrangian submanifolds $L\subset S$ equipped with square roots $K_{L}^{1/2}$ should define objects.
  • The graded Hom space between objects associated to oriented Lagrangians $L,M$ is given by the hypercohomology $\mathrm{H}^{\bullet-\dim S/2}(L\cap M,\phi_{L\cap M})$.
  • Composition of Homs is given by a conjectural functoriality of the DT sheaf, see [JS19, Conjecture 1.1].
We refer to [SV19] and [DR22, Section B.2] for a connection between this category and Lagrangian Floer homology. 7.1 shows that the stack $\mathrm{RH}_{S}$ provides such a category. Indeed, there is an isomorphism $\mathcal{M}_{1}\otimes^{\mathbb{L}}_{\mathsf{W}_{S}}\mathcal{M}_{2}\cong\mathbb{R}\mathcal{H}om_{\mathsf{W}_{S}}(\mathbb{D}(\mathcal{M}_{1}),\mathcal{M}_{2})[\dim S/2]$ valid for any holonomic $\mathsf{W}_{S}$-modules $\mathcal{M}_{1},\mathcal{M}_{2}$, where the duality functor is $\mathbb{D}(\mathcal{M}_{1})=\mathbb{R}\mathcal{H}om(\mathcal{M}_{1},\mathsf{W}_{S})[\dim S/2]$. Moreover, the modules $\mathsf{M}^{\lambda}_{L}$ are self-dual.
The format of the statement of 7.1 is inspired by a related result [SS14] of Sabbah and Saito (see also [Sab10]) relating the twisted de Rham complex
$\mathrm{DR}_{f}=(\Omega^{\bullet}_{U}(\!(\hbar)\!),d+\hbar^{-1}df)|_{f^{-1}(0)}$
and the sheaf of vanishing cycles $\phi_{f}$. (Note that we restrict to the zero set of $f$ to simplify the presentation.) Namely, there is a natural connection on $\mathrm{DR}_{f}$ in the $\hbar$ direction given by $\partial_{\hbar}-\hbar^{-2}f$, which equips $\mathrm{DR}_{f}$ with the structure of a differential perverse sheaf. Sabbah and Saito [SS14] identify the cohomology sheaves $\mathcal{H}^{k}\mathrm{DR}_{f}$ and $\mathcal{H}^{k}\mathrm{RH}^{-1}(\phi_{f})$. We show that their isomorphism lifts to an isomorphism of differential perverse sheaves (see 5.11).
Theorem C.
Let $f\colon U\rightarrow\mathbf{C}$ be a holomorphic function on a complex manifold $U$. Then there is an isomorphism of differential perverse sheaves
$(\mathrm{DR}_{f},\hbar\partial_{\hbar}-\hbar^{-1}f)\cong\mathrm{RH}^{-1}(\phi_{f}).$
A closely related result has recently been shown by Schefers in [Sch23, Proposition 9.20] using different methods.
Skein modules and perverse sheaves
Another motivation for this work is a conjectural relationship between skein modules and the DT sheaf formulated in [GJS23]. Let us recall the statement. Let $(M,x)$ be a connected closed oriented 3-manifold and $G$ a reductive algebraic group equipped with some extra data described in [GJS23, Section 1.3]. Then one can define the skein module $\mathrm{Sk}^{\mathrm{gen}}_{G}(M)$ of $M$ for generic quantum parameters, which is a $\mathbf{C}(q^{1/d})$-vector space for some integer $d$. For instance, for $G=\mathrm{SL}_{2}$ we obtain the Kauffman bracket skein module, which is a $\mathbf{C}(A)$-vector space for $A=q^{1/2}$. The representation variety
$R_{G}(M)=\mathrm{Hom}(\pi_{1}(M,x),G)$
carries a natural structure of a d-critical locus. In the case $G=\mathrm{SL}_{2}$ Abouzaid and Manolescu [AM20] denote the hypercohomology of the DT sheaf $\phi_{R_{G}(M)}$ on $R_{G}(M)$ by $\mathrm{HP}^{\bullet}_{\sharp}(M)$, the framed complexified instanton Floer homology of $M$. The conjectural relationship between the two objects is as follows.
Conjecture D.
There is an isomorphism of $\mathbf{C}(q^{1/d})$-vector spaces
$\mathrm{Sk}^{\mathrm{gen}}_{G}(M)\cong\mathrm{H}^{0}(R_{G}(M),\phi_{R_{G}(M)})\otimes_{\mathbf{C}}\mathbf{C}(q^{1/d}).$
We refer to [DKS23] for some checks of this conjecture for the Kauffman bracket skein module.
The above conjecture is related to A as follows. Choose a Heegaard splitting $M=N_{1}\cup_{\Sigma}N_{2}$ of $M$ as a union of two handlebodies $N_{1},N_{2}$ along the common boundary $\Sigma$. In this case we have the following:
  • The representation variety $R_{G}(M)$ becomes an intersection $L_{1}\cap L_{2}$ of two complex algebraic Lagrangian submanifolds $L_{1},L_{2}$ (representation varieties of the handlebodies $N_{1},N_{2}$) in a complex symplectic variety $S$ (an open subset of the representation variety of the open surface $\Sigma^{*}=\Sigma\setminus D$). We refer to [GJS23, Section 5.3] for details.
  • The generic skein module $\mathrm{Sk}^{\mathrm{gen}}_{G}(M)\otimes_{\mathbf{C}(q^{1/d})}\mathbf{C}(\!(\hbar)\!)$ defined with respect to the formal variable $\hbar$ with $q=\exp(\hbar)$ is a relative tensor product
    $\mathrm{Sk}^{\mathrm{gen}}_{G}(M)\otimes_{\mathbf{C}(q^{1/d})}\mathbf{C}(\!(\hbar)\!)\cong\mathrm{Sk}^{\mathrm{int}}_{G}(N_{1})\otimes_{\mathrm{SkAlg}^{\mathrm{int}}_{G}(\Sigma^{*})}\mathrm{Sk}^{\mathrm{int}}_{G}(N_{2})$
    of algebraic deformation quantization modules $\mathrm{Sk}^{\mathrm{int}}_{G}(N_{i})$ quantizing $L_{i}\subset S$ over the deformation quantization algebra $\mathrm{SkAlg}^{\mathrm{int}}_{G}(\Sigma^{*})$ providing an algebraic deformation quantization of the symplectic variety $S$.
Thus, in order to prove D using A, one must in particular relate the tensor product of the algebraic deformation quantization provided by skein theory to the corresponding tensor product of the canonical analytic deformation quantization modules. We plan to address this problem in future work.
Conventions
  • For a complex manifold $X$ and a ring $R$ we denote by $\mathbf{D}(X;R)$ the derived category of sheaves of complexes of $R$-modules.
  • $\mathbf{D}_{c}(X;R)\subset\mathbf{D}(X;R)$ denotes the derived category of constructible sheaves of $R$-modules on $X$.
  • $\mathrm{Perv}(X;R)\subset\mathbf{D}_{c}(X;R)$ denotes the subcategory of perverse sheaves.
  • Given a closed immersion $i\colon Y\hookrightarrow X$, perverse sheaves on $Y$ are identified with perverse sheaves on $X$ supported on $Y$ via the functor $i_{*}$. This identification will often be implicit.
  • We consider the subset $G=\{-1<\operatorname{Re}\lambda\leq 0\}\subset\mathbf{C}$ which defines a splitting of the projection $\mathbf{C}\rightarrow\mathbf{C}/\mathbf{Z}$.
Acknowledgements
We would like to thank Claude Sabbah, Pierre Schapira and Nick Sheridan for useful conversations.
1. Background on symplectic geometry
In this section we recall a relationship between holomorphic symplectic and holomorphic contact manifolds, their local models as well as Maslov indices of Lagrangians.
1.1. Symplectic and contact manifolds
In this paper we will deal with holomorphic symplectic manifolds $(S,\omega)$, i.e. complex manifolds $S$ equipped with holomorphic two-forms $\omega$ such that
$\mathrm{vol}_{S}=\frac{\omega^{\dim S/2}}{(\dim S/2)!}$
is a volume form. As an example, if $X$ is a complex manifold, its cotangent bundle $\mathrm{T}^{*}X$ is an exact symplectic manifold with $\omega=d\lambda_{X}$ for the Liouville one-form $\lambda_{X}$. We will also be interested in the open subset
$\mathring{\mathrm{T}}^{*}X=\mathrm{T}^{*}X\setminus\mathrm{T}^{*}_{X}X,$
the cotangent bundle minus the zero section. Given a symplectic manifold $(S,\omega)$, a Lagrangian subvariety is a locally closed subvariety $L\subset S$ of dimension $\dim L=\dim S/2$ such that $\omega|_{L_{\mathrm{reg}}}=0$, where $L_{\mathrm{reg}}$ is the smooth locus of $L$. We say $L\subset S$ is a Lagrangian submanifold if $L$ is a smooth Lagrangian subvariety.
Example 1.1.
If $Z\subset X$ is a submanifold, its conormal bundle $\mathrm{T}^{*}_{Z}X\subset\mathrm{T}^{*}X$ defines a (homogeneous) Lagrangian submanifold. If $\alpha$ is a one-form, its graph $\Gamma_{\alpha}\subset\mathrm{T}^{*}X$ defines a Lagrangian submanifold if, and only if, $\alpha$ is closed.
Given a symplectomorphism $\phi\colon S_{1}\rightarrow S_{2}$, its graph defines a Lagrangian submanifold $\Gamma_{\phi}\subset S_{1}\times\overline{S}_{2}$, where $\overline{S}_{2}$ refers to $S_{2}$ with the opposite symplectic structure. For instance, the graph of the identity defines the diagonal Lagrangian $\Delta_{S}\subset S\times\overline{S}$.
We will use natural symplectomorphisms
$\mathrm{T}^{*}X\times\mathrm{T}^{*}Y\cong\mathrm{T}^{*}(X\times Y),\qquad\mathrm{T}^{*}X\times\overline{\mathrm{T}^{*}Y}\cong\mathrm{T}^{*}(X\times Y)$
given by $(x,p,x^{\prime},p^{\prime})\mapsto(x,x^{\prime},p,p^{\prime})$ and $(x,p,x^{\prime},p^{\prime})\mapsto(x,x^{\prime},p,-p^{\prime})$.
Besides holomorphic symplectic manifolds we will also encounter holomorphic contact manifolds. Recall that a holomorphic contact manifold $Y$ is a complex manifold equipped with a line bundle $\mathcal{O}_{Y}(1)$ and a twisted one-form $\alpha\in\Gamma(Y,\Omega^{1}_{Y}\otimes\mathcal{O}_{Y}(1))$ satisfying a nondegeneracy condition. Alternatively, a contact structure can be specified by a hyperplane distribution in $\mathrm{T}_{Y}$ given by the kernel of $\alpha$.
Let $\gamma\colon\tilde{Y}\rightarrow Y$ be the $\mathbf{C}^{\times}$-torsor associated to the inverse of $\mathcal{O}_{Y}(1)$. Then the twisted 1-form $\alpha$ on $Y$ defines a 1-form $\alpha_{\tilde{Y}}$ on $\tilde{Y}$ linear along the fibers. The nondegeneracy condition on $\alpha$ is equivalent to the condition that $(\tilde{Y},\omega_{\tilde{Y}}=-d\alpha_{\tilde{Y}})$ is an exact symplectic manifold. We call $\tilde{Y}$ the symplectization of $Y$. It carries a natural Euler vector field given by the derivative of the natural $\mathbf{C}^{\times}$-action with respect to which $\omega_{\tilde{Y}}$ has weight $1$.
A contactomorphism $\psi\colon Y_{1}\rightarrow Y_{2}$ is the same as a homogeneous symplectomorphism $\tilde{\psi}\colon\tilde{Y}_{1}\rightarrow\tilde{Y}_{2}$. A Legendrian subvariety of a contact manifold $Y$ is a locally closed subvariety $\Lambda\subset Y$ of dimension $\dim\Lambda=(\dim(Y)-1)/2$ such that $\alpha|_{\Lambda_{\mathrm{reg}}}=0$. A Legendrian subvariety lifts to a $\mathbf{C}^{\times}$-invariant (alias, conic) Lagrangian subvariety $\tilde{\Lambda}\subset\tilde{Y}$ in the symplectization.
Example 1.2.
Consider the projectivized cotangent bundle of a complex manifold $X$ which is the quotient
$\mathrm{P}^{*}X=\mathring{\mathrm{T}}^{*}X/\mathbf{C}^{\times}.$
There is a natural contact structure on $\mathrm{P}^{*}X$ with respect to which $\mathrm{T}^{*}X\setminus X\rightarrow\mathrm{P}^{*}X$ is the symplectization. The corresponding one-form $\alpha_{\tilde{Y}}=-\lambda_{X}$ minus the Liouville one-form. A submanifold $Z\subset X$ of positive codimension defines a homogeneous Lagrangian submanifold $\mathring{\mathrm{T}}^{*}_{Z}X\subset\mathring{\mathrm{T}}^{*}X$ and a Legendrian submanifold $\mathrm{P}^{*}_{Z}X\subset\mathrm{P}^{*}X$.
Example 1.3.
Consider the first jet bundle $\mathrm{J}^{1}X=\mathrm{T}^{*}X\times\mathbf{C}$ of a complex manifold $X$. There is a natural contact structure on $\mathrm{J}^{1}X$ given by $\alpha=dt-\lambda_{X}$, where $t$ is the coordinate along $\mathbf{C}$ and $\lambda_{X}$ the Liouville one-form. Given a holomorphic function $f\colon X\rightarrow\mathbf{C}$ its one-jet $j^{1}f\colon X\rightarrow\mathrm{J}^{1}X$ defines a Legendrian submanifold. There is a homogeneous symplectic embedding $\tilde{\mathrm{J}^{1}X}\hookrightarrow\mathring{\mathrm{T}}^{*}(X\times\mathbf{C})$ given by $(x,p,t,\tau)\mapsto(x,t,p\tau,-\tau)$, where $(x,p)$ are Darboux coordinates on $\mathrm{T}^{*}X$. Under this embedding the conic Lagrangian submanifold $\tilde{j^{1}f}\subset\tilde{\mathrm{J}^{1}X}$ is sent to $\mathring{\mathrm{T}}^{*}_{\Gamma_{f}}(X\times\mathbf{C})\subset\mathring{\mathrm{T}}^{*}(X\times\mathbf{C})$, where $\Gamma_{f}\subset X\times\mathbf{C}$ is the submanifold defined by $t=f(x)$.
1.2. Contactification
Let us now study the process converse to symplectization.
Definition 1.4.
Let $(S,\omega)$ be a symplectic manifold. A contactification of $S$ is a principal $\mathbf{C}$-bundle $\rho\colon Y\rightarrow S$ together with a connection whose curvature is equal to $-\omega$. Let $L\subset S$ be a Lagrangian subvariety. A contactification of $L\subset S$ is a choice of a contactification $\rho\colon Y\rightarrow S$ of $S$ together with a subvariety $\Lambda\subset Y$ (the Legendrian lift), such that $\rho\colon\Lambda\rightarrow L$ is a homeomorphism and an analytic isomorphism over the smooth locus $\Lambda_{\mathrm{reg}}$, on which the connection one-form restricts to zero.
Remark 1.5.
It is not in general possible to find a contactification with $\rho\colon\Lambda\rightarrow L$ an analytic isomorphism everywhere, see [DK11, Example 2.5.4].
Given a symplectomorphism $\phi\colon S_{1}\rightarrow S_{2}$ and contactifications $Y_{1}\rightarrow S_{1}$ and $Y_{2}\rightarrow S_{2}$, an isomorphism of contactifications is a diffeomorphism $\psi\colon Y_{1}\rightarrow Y_{2}$ which is $\mathbf{C}$-equivariant and which intertwines the connection one-forms $\alpha_{i}$.
A contactification $\rho\colon Y\rightarrow S$ has a natural contact structure with $\alpha$ the connection one-form on $Y$. A contactification of a Lagrangian subvariety of $S$ is then a Legendrian subvariety of $Y$. It is easy to see that contactifications of a given symplectic manifold $(S,\omega)$ form a $\mathbf{C}$-gerbe $\mathsf{C}_{S}$ whose characteristic class is $[\omega]\in\mathrm{H}^{2}(S;\mathbf{C})$.
Given a contactification $\rho\colon Y\rightarrow S$ of a symplectic manifold $S$, $\rho\colon\overline{Y}\rightarrow\overline{S}$ is a contactification of the opposite symplectic manifold, where $\overline{Y}$ is $Y$ equipped with the opposite contact form $\alpha$ and opposite $\mathbf{C}$-action.
Example 1.6.
Suppose $(S,\omega=d\lambda)$ is an exact symplectic manifold. Then $Y=S\times\mathbf{C}$ with $\rho$ the projection to the first factor, $\mathcal{O}_{Y}(1)$ trivial line bundle and $\alpha=dt-\lambda$, where $t$ is the coordinate along $\mathbf{C}$, is a contactification of $S$.
We will be particularly interested in the following case of the previous example.
Example 1.7.
The projection $\rho\colon\mathrm{J}^{1}X\rightarrow\mathrm{T}^{*}X$ is a contactification. Given a closed one-form $\alpha$ on $X$, Legendrian lifts of the Lagrangian submanifold $\Gamma_{\alpha}\subset\mathrm{T}^{*}X$ are $j^{1}f\subset\mathrm{J}^{1}X$ for some holomorphic function $f\colon X\rightarrow\mathbf{C}$ with $\alpha=df$.
Consider a contactification $\rho\colon Y\rightarrow S$ and its symplectization $\gamma\colon\tilde{Y}=Y\times\mathbf{C}^{\times}\rightarrow Y$. The $\mathbf{C}$-action on $Y$ lifts to a Hamiltonian $\mathbf{C}$-action on $\tilde{Y}$ with Hamiltonian given by the natural coordinate $\tau$ along $\mathbf{C}^{\times}$. In this case we may identify $S$ with the Hamiltonian reduction $\tilde{Y}/\!/\mathbf{C}$ at $\tau=1$. An isomorphism of contactifications $\psi\colon Y_{1}\rightarrow Y_{2}$ covering a symplectomorphism $\phi\colon S_{1}\rightarrow S_{2}$ gives rise to a conic symplectomorphism $\tilde{\psi}\colon\tilde{Y}_{1}\rightarrow\tilde{Y}_{2}$ compatible the Hamiltonian $\mathbf{C}$-actions.
Given contactifications $Y_{1}\rightarrow S_{1}$ and $Y_{2}\rightarrow S_{2}$ we denote
$Y_{1}\times^{\ast}Y_{2}=(Y_{1}\times Y_{2})/\mathbf{C}\rightarrow S_{1}\times S_{2},$
where the quotient is taken with respect to the antidiagonal action of $\mathbf{C}$ on $Y_{1}\times Y_{2}$. Given Legendrians $\Lambda_{1}\subset Y_{1}$ and $\Lambda_{2}\subset Y_{2}$ we denote by
$\Lambda_{1}\times^{\ast}\Lambda_{2}\subset Y_{1}\times^{\ast}Y_{2}$
the image of $\Lambda_{1}\times\Lambda_{2}$ under the projection $Y_{1}\times Y_{2}\rightarrow Y_{1}\times^{\ast}Y_{2}$.
Proposition 1.8.
Let $S_{1},S_{2}$ be complex symplectic manifolds and $\rho\colon Y_{i}\rightarrow S_{i}$ their contactifications.
  • $Y_{1}\times^{\ast}Y_{2}\rightarrow S_{1}\times S_{2}$ is a contactification.
  • Given a pair of Lagrangians $L_{1}\subset S_{1},L_{2}\subset S_{2}$ together with their Legendrian lifts $\Lambda_{1}\subset Y_{1},\Lambda_{2}\subset Y_{2}$ the submanifold $\Lambda_{1}\times^{\ast}\Lambda_{2}\subset Y_{1}\times^{\ast}Y_{2}$ provides a contactification of the Lagrangian $L_{1}\times L_{2}\subset S_{1}\times S_{2}$.
  • Given a commutative diagram
[…]
    where $\phi\colon S_{1}\rightarrow S_{2}$ is a symplectomorphism and $\psi\colon Y_{1}\rightarrow Y_{2}$ is an isomorphism of contactifications, the image $\Gamma_{\psi}\subset Y_{1}\times^{\ast}\overline{Y}_{2}$ of the graph of the contactomorphism $\psi$ under $Y_{1}\times\overline{Y}_{2}\rightarrow Y_{1}\times^{\ast}\overline{Y}_{2}$ defines a Legendrian lift of the graph $\Gamma_{\phi}\subset S_{1}\times\overline{S}_{2}$ of the symplectomorphism $\phi$.
Proof.
The one-forms $\alpha_{i}$ on $Y_{i}$ satisfy $g^{*}\alpha_{i}=\alpha_{i}$ for any $g\in\mathbf{C}$ as well as $\iota_{R}\alpha_{i}=1$, where $R$ is the vector field generating the $\mathbf{C}$-action. Thus, $\alpha_{1}+\alpha_{2}$ is basic, so it descends along the projection $Y_{1}\times Y_{2}\rightarrow(Y_{1}\times Y_{2})/\mathbf{C}$. The diagonal action of $\mathbf{C}$ on $Y_{1}\times Y_{2}$ descends to an action on $(Y_{1}\times Y_{2})/\mathbf{C}$ which defines the structure of a principal $\mathbf{C}$-bundle on $(Y_{1}\times Y_{2})/\mathbf{C}\rightarrow S_{1}\times S_{2}$. The one-form $\alpha_{1}+\alpha_{2}$ on $(Y_{1}\times Y_{2})/\mathbf{C}$ defines a connection one-form whose curvature is $\omega_{1}+\omega_{2}$, so $(Y_{1}\times Y_{2})/\mathbf{C}\rightarrow S_{1}\times S_{2}$ is a contactification.
The one-form $\alpha_{1}+\alpha_{2}$ vanishes on $\Lambda_{1}\times\Lambda_{2}$ and so, by counting dimensions, $\Lambda_{1}\times\Lambda_{2}\subset(Y_{1}\times Y_{2})/\mathbf{C}$ is Legendrian.
It is easy to see that the composite $\Gamma_{\psi}\rightarrow(Y_{1}\times Y_{2})/\mathbf{C}\rightarrow Y_{1}/\mathbf{C}=S_{1}$ is a diffeomorphism. The fact that $\psi^{*}\alpha_{2}=\alpha_{1}$ implies that $\alpha_{1}-\alpha_{2}$ vanishes on $\Gamma_{\psi}\subset(Y_{1}\times Y_{2})/\mathbf{C}$. ∎
Remark 1.9.
We may identify the symplectization of $Y_{1}\times^{\ast}Y_{2}$ with the Hamiltonian reduction $(\tilde{Y}_{1}\times\tilde{Y}_{2})/\!/\mathbf{C}$ at $\tau_{1}=\tau_{2}$. The graph $\Gamma_{\tilde{\psi}}\subset\tilde{Y}_{1}\times\overline{\tilde{Y}_{2}}$ of the conic symplectomorphism $\tilde{\psi}\colon\tilde{Y}_{1}\rightarrow\tilde{Y}_{2}$ is $\mathbf{C}$-equivariant and we may identify $\tilde{\Gamma_{\psi}}\subset(\tilde{Y}_{1}\times\tilde{Y}_{2})/\!/\mathbf{C}$ with $\Gamma_{\tilde{\psi}}/\mathbf{C}$.
Example 1.10.
Given a contactification $\rho\colon Y\rightarrow S$, the contactification of the diagonal Lagrangian $\Delta_{S}\subset S\times\overline{S}$ is given by the Legendrian $\Delta_{Y}\subset Y\times^{\ast}\overline{Y}$ given by the image of the diagonal of $Y$.
Example 1.11.
Let $X_{1},X_{2}$ be complex manifolds. Then
$\mathrm{J}^{1}X_{1}\times^{\ast}\mathrm{J}^{1}X_{2}\longrightarrow\mathrm{J}^{1}(X_{1}\times X_{2})$
given by $(x,p,t),(x^{\prime},p^{\prime},t^{\prime})\mapsto(x,x^{\prime},p,p^{\prime},t+t^{\prime})$ is a contactomorphism. Given holomorphic functions $f_{1}\colon X_{1}\rightarrow\mathbf{C}$ and $f_{2}\colon X_{2}\rightarrow\mathbf{C}$, under this isomorphism the Legendrian
$j^{1}f_{1}\times^{\ast}j^{1}f_{2}\subset\mathrm{J}^{1}X_{1}\times^{\ast}\mathrm{J}^{1}X_{2}$
is identified with the Legendrian $j^{1}(f_{1}\boxplus f_{2})\subset\mathrm{J}^{1}(X_{1}\times X_{2})$, where $f_{1}\boxplus f_{2}\colon X_{1}\times X_{2}\rightarrow\mathbf{C}$ is the function $(x,x^{\prime})\mapsto f_{1}(x)+f_{2}(x^{\prime})$. Similarly,
$\mathrm{J}^{1}X_{1}\times^{\ast}\overline{\mathrm{J}^{1}X_{2}}\longrightarrow\mathrm{J}^{1}(X_{1}\times X_{2})$
given by $(x,p,t),(x^{\prime},p^{\prime},t^{\prime})\mapsto(x,x^{\prime},p,-p^{\prime},t-t^{\prime})$ is a contactomorphism.
We will now show that near a Lagrangian subvariety $L\subset S$ there is a canonical contactification.
Proposition 1.12.
Let $L\subset S$ be a Lagrangian subvariety. Then there is a canonical section of $\mathsf{C}_{\omega}|_{L}$. Explicitly, there is a unique contactification $\rho\colon Y\rightarrow S$ in a neighborhood of $L$ together with a lift of $L$ to a Legendrian $\Lambda\subset Y$ such that $\rho\colon\Lambda\rightarrow L$ is a homeomorphism and, moreover, is an analytic isomorphism over the smooth locus of $L$.
Proof.
The existence of such a contactification is shown in [DK11, Proposition 2.5.3]. Let us now show that any two such contactifications are uniquely isomorphic. By [Gie64] we may choose a triangulation of $S$ with $L\subset S$ as a simplicial subcomplex. Therefore, there is a neighborhood $U$ of $L$ which deformation retracts to $L$.
Choose a cover $\{U_{i}\}$ of $U$. Suppose $\{\theta^{1}_{i},f^{1}_{i},\varphi^{1}_{ij}\}$ and $\{\theta^{2}_{i},f^{2}_{i},\varphi^{2}_{ij}\}$ are two descent data on $\{U_{i}\}$ defining two contactifications $Y^{1},Y^{2}$ of $S$. Here $\theta^{n}_{i}\in\Omega^{1}(U_{i})$, $f^{n}_{i}$ is a continuous function on $L\cap U_{i}$ holomorphic on the smooth locus and $\varphi^{n}_{ij}$ is a holomorphic function on $U_{ij}$ which together satisfy
$d\theta^{n}_{i}=\omega,\qquad df^{n}_{i}|_{L_{\mathrm{reg}}}=\theta^{n}_{i}|_{L_{\mathrm{reg}}},\qquad d\varphi^{n}_{ij}=\theta^{n}_{i}-\theta^{n}_{j},\qquad\varphi^{n}_{ij}|_{L}=f^{n}_{i}-f^{n}_{j},\qquad\varphi^{n}_{ij}+\varphi^{n}_{jk}+\varphi^{n}_{ki}=0.$
Consider their difference:
$\theta_{i}=\theta^{1}_{i}-\theta^{2}_{i},\qquad f_{i}=f^{1}_{i}-f^{2}_{i},\qquad\varphi_{ij}=\varphi^{1}_{ij}-\varphi^{2}_{ij}.$
Then
$d\theta_{i}=0,\qquad df_{i}|_{L_{\mathrm{reg}}}=\theta_{i}|_{L_{\mathrm{reg}}},\qquad d\varphi_{ij}=\theta_{i}-\theta_{j},\qquad\varphi_{ij}|_{L}=f_{i}-f_{j},\qquad\varphi_{ij}+\varphi_{jk}+\varphi_{ki}=0.$
Possibly refining the cover, we choose holomorphic functions $g_{i}$ on $U_{i}$ such that $dg_{i}=\theta_{i}$. Then $g_{i}-g_{j}-\varphi_{ij}$ define locally constant functions on $U_{ij}$ which determine a cohomology class in $\mathrm{H}^{1}(U;\mathbf{C})$. The restriction of these functions to $L$ is $(g_{i}-f_{i})-(g_{j}-f_{j})$. Note that $d(g_{i}-f_{i})|_{L_{\mathrm{reg}}}=0$, i.e. $g_{i}-f_{i}$ is locally constant on $L_{\mathrm{reg}}\cap U_{i}$ and hence, by density, on $L\cap U_{i}$ (see [DK11, Sublemma 2.5.2]). Thus, under the restriction map $\mathrm{H}^{1}(U;\mathbf{C})\rightarrow\mathrm{H}^{1}(L;\mathbf{C})$ the class $[g_{i}-g_{j}-\varphi_{ij}]$ restricts to zero. Since $U$ deformation retracts to $L$, this implies that there are locally constant functions $h_{i}$ on $U_{i}$ such that
$\varphi_{ij}=(g_{i}-h_{i})-(g_{j}-h_{j}),\qquad(g_{i}-h_{i})|_{L}=f_{i}.$
Thus, the functions $g_{i}-h_{i}$ define an isomorphism of the contactifications of $S$ which preserves the Legendrian lift of $L$. ∎
1.3. Orientation torsors
Given a triple of Lagrangian subspaces of a symplectic vector space, Wall and Kashiwara have associated a certain quadratic form whose index (in the real case) is the Maslov index of the triple of Lagrangians. In this section we will present a variant of this construction.
Throughout this section we fix a field $k$. Let $W$ be a $k$-vector space equipped with a quadratic form $q$. There is an induced quadratic form on the determinant line $\det(W)$. We denote by $P_{q}$ the $\mathbf{Z}/2$-torsor of elements $\omega\in\det(W)$ of norm $1$. In other words, $P_{q}$ is the torsor of orientations of the quadratic vector space $(W,q)$. If $W$ is a quadratic vector bundle over a manifold, the definition of $P_{q}$ extends to define a $\mathbf{Z}/2$-torsor over the base manifold.
Given a pair $(W_{1},q_{1})$, $(W_{2},q_{2})$ of quadratic vector spaces, we have a natural isomorphism
(1) $P_{q_{1}}\otimes P_{q_{2}}\cong P_{q_{1}\oplus q_{2}}$
given by $\omega_{1}\otimes\omega_{2}\mapsto\omega_{1}\wedge\omega_{2}$, where we denote by $q_{1}\oplus q_{2}\colon W_{1}\oplus W_{2}\rightarrow k$ is the orthogonal direct sum of quadratic forms.
Now let $V$ be a symplectic $k$-vector space and $L_{1},L_{2}\subset V$ are two Lagrangian subspaces which are transverse. The transversality condition may be equivalently stated as the fact that the map
$f_{L_{1}L_{2}}\colon L_{1}\longrightarrow L_{2}^{*}$
given by $f_{L_{1}L_{2}}(v)=(w\mapsto\omega(v,w))$ is an isomorphism.
Definition 1.13.
Let $L_{1},L_{2},\dots,L_{n}\subset V$ be a sequence of Lagrangian subspaces, where any two subsequent subspaces are transverse.
  • If $n$ is even, we define the isomorphism $C(L_{1},\dots,L_{n})\colon L_{1}\rightarrow L_{n}^{*}$ to be given by the composite
    $C(L_{1},\dots,L_{n})\colon L_{1}\xrightarrow{f_{L_{1}L_{2}}}L_{2}^{*}\xrightarrow{f_{L_{3}L_{2}}^{-1}}L_{3}\rightarrow\dots\xrightarrow{f_{L_{n-1}L_{n}}}L_{n}^{*}$
  • If $n$ is odd, we define the isomorphism $C(L_{1},\dots,L_{n})\colon L_{1}\rightarrow L_{n}$ to be given by the composite
    $C(L_{1},\dots,L_{n})\colon L_{1}\xrightarrow{f_{L_{1}L_{2}}}L_{2}^{*}\xrightarrow{f_{L_{3}L_{2}}^{-1}}L_{3}\rightarrow\dots\xrightarrow{f_{L_{n-1}L_{n}}^{-1}}L_{n}$
Example 1.14.
Consider pairwise transverse Lagrangian subspaces $L_{1},L_{2},L_{3}\subset V$. The isomorphism
$C(L_{1},L_{2},L_{3},L_{1})\colon L_{1}\longrightarrow L_{1}^{*}$
is self-dual and comes from a quadratic form
$q(L_{1},L_{2},L_{3})\colon L_{1}\rightarrow k$
which we call the Maslov quadratic form. For $k=\mathbf{R}$ its signature is the Maslov index $\tau(L_{1},L_{2},L_{3})$ of the triple of Lagrangians as defined in [KS94, Chapter A.3].
Example 1.15.
Consider a finite-dimensional vector space $L$ and the corresponding symplectic vector space $V=L\oplus L^{*}$ with the symplectic form
$\omega(x+\eta,y+\xi)=\eta(y)-\xi(x).$
It carries two transverse Lagrangian subspaces $L,L^{*}\subset V$. Consider a linear map $A\colon L\rightarrow L^{*}$ and the corresponding half-dimensional subspace $M\subset V$ given by
$M=\{(l,A(l))\mid l\in L\}.$
The subspace $M\subset V$ is Lagrangian precisely if $A$ is self-dual, i.e. $A$ comes from a quadratic form $q\colon L\rightarrow k$.
The subspace $M\subset V$ is transverse to $L\subset V$ (equivalently, to $L^{*}\subset V$) precisely if $A$ is an isomorphism, i.e. $q$ is nondegenerate.
In this case $C(L,M,L^{*},L)$ is given by the composite
$L\xrightarrow{f_{LM}}M^{*}\xleftarrow{f_{L^{*}M}}L^{*}\xrightarrow{f_{L^{*}L}}L^{*}.$
It is easy to work out it to be $-A$, so that the Maslov quadratic form $q(L,M,L^{*})$ is equal to $-q$.
We have the following obvious properties of the isomorphisms $C(L_{1},\dots,L_{n})$.
Proposition 1.16.
  • Suppose $n$ is odd. Then
    $C(L_{n},L_{n+1},\dots,L_{n+m})\circ C(L_{1},\dots,L_{n})=C(L_{1},\dots,L_{n},\dots,L_{n+m}).$
  • Consider a sequence $L_{1},\dots,L_{n-1},L_{n},L_{n-1},\dots,L_{m}$ of Lagrangian subspaces, where the subsequent ones are transverse to each other. Then
    $C(L_{1},\dots,L_{n-1},L_{n},L_{n-1},\dots,L_{m})=C(L_{1},\dots,L_{n-1},\widehat{L_{n}},\widehat{L_{n-1}},\dots,L_{m}).$
Definition 1.17.
Let $V$ be a symplectic space and $L\subset V$ a Lagrangian subspace. An orientation data on $L$ is the choice of the square root line $\det(L)^{1/2}$. Let $L\subset S$ be a Lagrangian submanifold of a symplectic manifold. An orientation data on $L$ is the choice of the square root line bundle $K_{L}^{1/2}$ of the canonical bundle of $L$.
Definition 1.18.
Let $L_{1},\dots,L_{n}\subset V$ be a sequence of Lagrangian subspaces, where any two subsequent subspaces are transverse. Moreover, suppose $L_{1}$ and $L_{n}$ are equipped with an orientation data. An orientation of the sequence $L_{1},\dots,L_{n}$, where $n$ is odd, is the choice of an isomorphism
$\det(L_{1})^{1/2}\longrightarrow\det(L_{n})^{1/2}$
whose square is
$\det(C(L_{1},\dots,L_{n}))\colon\det(L_{1})\longrightarrow\det(L_{n}).$
In the case of $n$ even the definition is analogous. The choices of an orientation of the sequence define a $\mathbf{Z}/2$-torsor $P_{L_{1},\dots,L_{n}}$.
Remark 1.19.
The notion of an orientation of the sequence $L_{1},L_{2},L_{3},L_{1}$ is independent of the choice of the orientation data on $L_{1}$. In this case we have
$P_{L_{1},L_{2},L_{3},L_{1}}\cong P_{q(L_{1},L_{2},L_{3})}$
for the Maslov quadratic form $q(L_{1},L_{2},L_{3})\colon L_{1}\rightarrow k$.
These definitions extend in an obvious way to symplectic bundles and sequences of Lagrangian subbundles.
1.4. Charts on symplectic manifolds
In this section we describe convenient local models of symplectic manifolds. Let us first recall the notion of a polarization.
Definition 1.20.
A polarization of $S$ is a holomorphic Lagrangian fibration $\pi\colon S\rightarrow E$.
Remark 1.21.
The Darboux theorem ensures that polarizations of symplectic manifolds exist locally.
Definition 1.22.
A polarization $\pi\colon S\rightarrow E$ is transverse to a Lagrangian $L\subset S$ if the composite $L\subset S\rightarrow E$ is a local diffeomorphism.
When a polarization $\pi$ is transverse to a Lagrangian $L$ we obtain a projection map $\pi^{S\rightarrow L}\colon S\rightarrow L$ locally on $L$.
Example 1.23.
Consider a symplectic manifold $S$ equipped with a polarization $\pi$ and two Lagrangian submanifolds $L,M\subset S$ transverse to $\pi$. At any point $x\in L\cap M$ we have a sequence $T_{L,x},T_{\pi,x},T_{M,x}$ of Lagrangian subspaces of $T_{S,x}$, where the subsequent subspaces are transverse. Thus, from 1.13 we obtain an isomorphism
$C(\mathrm{T}_{L,x},\mathrm{T}_{\pi,x},\mathrm{T}_{M,x})\colon\mathrm{T}_{L,x}\longrightarrow\mathrm{T}_{M,x}.$
Unpacking this isomorphism, it is given by the composite of the isomorphisms
[…]
Example 1.24.
If $S=\mathrm{T}^{*}E$ is a cotangent bundle, there is a canonical vertical polarization $\mathrm{T}^{*}E\rightarrow E$ given by the projection. It is transverse to the zero section $E\subset\mathrm{T}^{*}E$.
The following three statements provide a local description of symplectic manifolds equipped with polarizations.
Proposition 1.25.
Let $S$ be a symplectic manifold, $\rho\colon Y\rightarrow S$ a contactification $\pi\colon S\rightarrow E$ a polarization, $L\subset S$ a Lagrangian submanifold transverse to $\pi$ and $\Lambda_{L}\subset Y$ is its Legendrian lift. For every point $y\in\Lambda_{L}$ there is an open neighborhood $U\subset S$ of $\rho(y)$, a commutative diagram
[…]
with $i_{\pi}$ a contact embedding and $\overline{i}_{\pi}$ a symplectic embedding which identifies $L\subset S$ with the zero section of $\mathrm{T}^{*}L$ and $\Lambda_{L}\subset Y$ with $j^{1}0\subset\mathrm{J}^{1}L$. Moreover, the maps $i_{\pi}$ and $\overline{i}_{\pi}$ are uniquely determined by these properties.
Proof.
For the statement on the level of symplectic manifolds, see [Wei71, Theorem 7.1]. By Example 1.7 $j^{1}0\subset\mathrm{J}^{1}L$ is a contactification of the zero section $L\subset\mathrm{T}^{*}L$ and by 1.12 contactifications of Lagrangians are unique. This implies the statement about the local form of the contactification. ∎
Remark 1.26.
Given a symplectic manifold $S$ with a polarization $\pi$ transverse to a Lagrangian submanifold $L\subset S$ we may apply 1.25 to both $S$ and $\overline{S}$ to get symplectic embeddings $U\hookrightarrow\mathrm{T}^{*}L$ and $\overline{U}\hookrightarrow\mathrm{T}^{*}L$. The two differ by negation along the fibers of the cotangent bundle.
In the case of two Lagrangians we have the following statement.
Proposition 1.27.
Let $S$ be a symplectic manifold, $\rho\colon Y\rightarrow S$ a contactification, $\pi\colon S\rightarrow E$ a polarization, $L,M\subset S$ Lagrangian submanifolds transverse to $\pi$ and $\Lambda_{L},\Lambda_{M}\subset Y$ their Legendrian lifts. For every point $y\in\Lambda_{L}\cap\Lambda_{M}$ there is an open neighborhood $U\subset S$ of $\rho(y)$, a commutative diagram
[…]
where $i_{\pi}$ is a contact embedding and $\overline{i}_{\pi}$ is a symplectic embedding and a locally defined holomorphic function $f\colon L\rightarrow\mathbf{C}$ vanishing on $(L\cap M)^{\mathrm{red}}$, such that:
  • Under the embedding $\overline{i}_{\pi}\colon U\hookrightarrow\mathrm{T}^{*}L$ the polarization $\pi$ identifies with the vertical polarization $\mathrm{T}^{*}L\rightarrow L$.
  • The Legendrian $\Lambda_{L}\subset Y$ covering $L\subset S$ identifies with $j^{1}0\subset\mathrm{J}^{1}L$ covering the zero section $L\subset\mathrm{T}^{*}L$.
  • The Legendrian $\Lambda_{M}\subset Y$ covering $M\subset S$ identifies with $j^{1}f\subset\mathrm{J}^{1}L$ covering the graph $\Gamma_{df}\subset\mathrm{T}^{*}L$ of $df$.
Moreover, the maps $i_{\pi}$, $\overline{i}_{\pi}$ as well as the function $f$ are uniquely determined by these conditions.
Proof.
Using 1.25 we find a neighborhood $U$ of $\rho(y)$ and a contact embedding $\rho^{-1}(U)\hookrightarrow\mathrm{J}^{1}L$ covering a symplectic embedding $U\subset\mathrm{T}^{*}L$ under which $\Lambda_{L}\subset Y$ covering $L\subset S$ identifies with $j^{1}0\subset\mathrm{J}^{1}L$ covering the zero section $L\subset\mathrm{T}^{*}L$.
Since $M$ is transverse to $\pi$, the composite $M\subset S\rightarrow E$ is a local diffeomorphism. Therefore, under $U\subset\mathrm{T}^{*}L$ the submanifold $M\subset S$ identifies with the graph of a locally defined one-form $\alpha$ on $L$ which is closed since $M$ is Lagrangian. Lifts of $\Gamma_{\alpha}\subset\mathrm{T}^{*}L$ to the contactification $\mathrm{J}^{1}L\rightarrow\mathrm{T}^{*}L$ are given by choosing a function $f\colon M\rightarrow\mathbf{C}$ with $\alpha=df$, so that $\Lambda_{M}\subset Y$ is identified with $j^{1}f\subset\mathrm{J}^{1}L$. The function $f$ vanishes at $y\in\Lambda_{L}\cap\Lambda_{M}$ and it is locally constant on $(L\cap M)^{\mathrm{red}}$. Therefore, shrinking the neighborhood $U$, we may assume $f$ vanishes on all of $(L\cap M)^{\mathrm{red}}$. ∎
Remark 1.28.
It is not in general possible to arrange $f$ to be zero on $L\cap M$ when this intersection is non-reduced, see [Joy15, Example 2.13].
Finally, we will need to describe the change of local models under a change of polarizations. Recall the symplectomorphism $\mathrm{T}^{*}L\times\overline{\mathrm{T}^{*}M}\cong\mathrm{T}^{*}(L\times M)$ and the contactomorphism $\mathrm{J}^{1}L\times^{\ast}\overline{\mathrm{J}^{1}M}\cong\mathrm{J}^{1}(L\times M)$ from Example 1.11.
Proposition 1.29.
Let $S$ be a symplectic manifold, $\rho\colon Y\rightarrow S$ a contactification, $\pi_{i}\colon S\rightarrow E_{i}$ for $i=1,2$ transverse polarizations, $L,M\subset S$ Lagrangian submanifolds transverse to both $\pi_{1}$ and $\pi_{2}$ and $\Lambda_{L},\Lambda_{M}\subset Y$ their Legendrian lifts. For every point $y\in\Lambda_{L}\cap\Lambda_{M}$ there is an open neighborhood $U\subset S$ of $\rho(y)$, commutative diagrams
[…]
where $i_{\pi_{1}},i_{\pi_{2}}$ are contact embeddings and $\overline{i}_{\pi_{1}},\overline{i}_{\pi_{2}}$ are symplectic embeddings and locally defined functions $h\colon L\times M\rightarrow\mathbf{C}$, $f\colon L\rightarrow\mathbf{C}$, $g\colon M\rightarrow\mathbf{C}$ vanishing on $(L\cap M)^{\mathrm{red}}$, such that:
  • Under the embedding $U\hookrightarrow\mathrm{T}^{*}L$ the polarization $\pi_{1}$ identifies with the vertical polarization $\mathrm{T}^{*}L\rightarrow L$.
  • Under the contact embedding $\rho^{-1}(U)\hookrightarrow\mathrm{J}^{1}L$ the Legendrian $\Lambda_{L}\subset Y$ covering $L\subset S$ identifies with $j^{1}0\subset\mathrm{J}^{1}L$ covering the zero section $L\subset\mathrm{T}^{*}L$. Under the same embedding the Legendrian $\Lambda_{M}\subset Y$ covering $M\subset S$ identifies with $j^{1}f\subset\mathrm{J}^{1}L$ covering $\Gamma_{df}\subset\mathrm{T}^{*}L$.
  • Under the embedding $U\hookrightarrow\mathrm{T}^{*}M$ the polarization $\pi_{2}$ identifies with the vertical polarization $\mathrm{T}^{*}M\rightarrow M$.
  • Under the contact embedding $\rho^{-1}(U)\hookrightarrow\mathrm{J}^{1}M$ the Legendrian $\Lambda_{M}\subset Y$ covering $M\subset S$ identifies with $j^{1}0\subset\mathrm{J}^{1}M$ covering the zero section $M\subset\mathrm{T}^{*}M$. Under the same embedding the Legendrian $\Lambda_{L}\subset Y$ covering $L\subset S$ identifies with $j^{1}g\subset\mathrm{J}^{1}M$ covering $\Gamma_{dg}\subset\mathrm{T}^{*}M$.
  • Under the embedding $U\times\overline{U}\hookrightarrow\mathrm{T}^{*}(L\times M)$ the polarization $\pi_{1}\times\pi_{2}$ identifies with the vertical polarization $\mathrm{T}^{*}(L\times M)\rightarrow L\times M$.
  • Under the contact embedding $\rho^{-1}(U)\times^{\ast}\overline{\rho^{-1}(U)}\hookrightarrow\mathrm{J}^{1}(L\times M)$ the Legendrian $\Lambda_{L}\times^{\ast}\Lambda_{M}\subset Y\times^{\ast}\overline{Y}$ covering $L\times M\subset S\times\overline{S}$ identifies with $j^{1}0\subset\mathrm{J}^{1}(L\times M)$ covering the zero section $L\times M\subset\mathrm{T}^{*}(L\times M)$. Under the same embedding the Legendrian $\Delta_{Y}\subset Y\times^{\ast}\overline{Y}$ covering the diagonal $\Delta_{S}\subset S\times\overline{S}$ identifies with $j^{1}h\subset\mathrm{J}^{1}(L\times M)$ covering $\Gamma_{dh}\subset\mathrm{T}^{*}(L\times M)$.
  • Consider the local diffeomorphism $\pi^{S\rightarrow L}_{1}\colon M\rightarrow L$. Under the locally defined embedding
    $\Xi=(\mathrm{id}\times(\pi^{S\rightarrow L}_{1})^{-1})\colon L\hookrightarrow L\times M$
    we have $\Xi^{*}h=f$.
  • Consider the local diffeomorphism $\pi^{S\rightarrow M}_{2}\colon L\rightarrow M$. Under the locally defined embedding
    $\Upsilon=((\pi^{S\rightarrow M}_{2})^{-1}\times\mathrm{id})\colon M\hookrightarrow L\times M$
    we have $\Upsilon^{*}h=-g$.
Proof.
Applying 1.27 to $(Y\rightarrow S,\pi_{1},L,M)$ we get a commutative diagram
[…]
together with a locally defined function $f\colon L\rightarrow\mathbf{C}$, such that under the embeddings the polarization $\pi_{1}$ goes to the vertical polarization, $L$ goes to the zero section, $\Lambda_{L}$ goes to $j^{1}0$, $M$ goes to $\Gamma_{df}\subset\mathrm{T}^{*}L$ and $\Lambda_{M}$ goes to $j^{1}f\subset\mathrm{J}^{1}L$.
Applying 1.27 to $(Y\rightarrow S,\pi_{2},M,L)$ (and possibly shrinking $U$) we get a commutative diagram
[…]
together with a locally defined function $g\colon M\rightarrow\mathbf{C}$, such that under the embeddings the polarization $\pi_{2}$ goes to the vertical polarization, $M$ goes to the zero section, $\Lambda_{M}$ goes to $j^{1}0$, $L$ goes to $\Gamma_{dg}\subset\mathrm{T}^{*}M$ and $\Lambda_{L}$ goes to $j^{1}g\subset\mathrm{J}^{1}M$.
Consider the product $S\times\overline{S}$. By 1.8 it has a contactification $Y\times^{\ast}\overline{Y}$. We have a Lagrangian submanifold $L\times M\subset S\times\overline{S}$ with a Legendrian lift $\Lambda_{L}\times\Lambda_{M}$. The diagonal $\Delta_{S}\subset S\times\overline{S}$ also defines a Lagrangian submanifold with a Legendrian lift $\Delta_{Y}\subset Y\times^{\ast}\overline{Y}$. We have a polarization $\pi_{1}\times\pi_{2}$ on $S\times\overline{S}$ which is transverse to both $L\times M$ and $\Delta$. Therefore, we may apply 1.27 to $(Y\times^{\ast}\overline{Y}\rightarrow S\times\overline{S},\pi_{1}\times\pi_{2},L\times M,\Delta_{S})$ (and again by shrinking $U$) we get a commutative diagram
[…]
together with a locally defined function $h\colon L\times M\rightarrow\mathbf{C}$, such that under the embeddings the polarization $\pi_{1}\times\pi_{2}$ goes to the vertical polarization, $L\times M$ goes to the zero section, $\Lambda_{L}\times\Lambda_{M}$ goes to $j^{1}0$, $\Delta_{S}$ goes to $\Gamma_{dh}\subset\mathrm{T}^{*}(L\times M)$ and $\Delta_{Y}$ goes to $j^{1}h\subset\mathrm{J}^{1}(L\times M)$. By uniqueness of the contact embeddings, we see that the top map is $i_{1}\times i_{2}\colon\rho^{-1}(U)\times^{\ast}\overline{\rho^{-1}(U)}\rightarrow\mathrm{J}^{1}L\times^{\ast}\overline{\mathrm{J}^{1}M}\cong\mathrm{J}^{1}(L\times M)$ and the bottom map is $\overline{i}_{1}\times\overline{i}_{2}\colon U\times\overline{U}\rightarrow\mathrm{T}^{*}L\times\overline{\mathrm{T}^{*}M}\cong\mathrm{T}^{*}(L\times M)$.
By construction we have a commutative diagram
[…]
Its commutativity implies that $(\pi_{1}^{S\rightarrow L}\times\mathrm{id})^{*}h=(\pi_{1}^{S\rightarrow L})^{*}f$ and hence $\Xi^{*}h=f$. The claim about $\Upsilon$ is proven similarly. ∎
In the setup of the previous proposition we get a local symplectomorphism $\mathrm{T}^{*}L\cong\mathrm{T}^{*}M$ and a local contactomorphism $\mathrm{J}^{1}L\cong\mathrm{J}^{1}M$ covering it. Their graphs may be described as follows:
  • The correspondence
[…]
    where $p_{1}\colon\mathrm{T}^{*}(L\times M)\rightarrow\mathrm{T}^{*}L$ is the projection on the first factor and $p_{2}^{a}\colon\mathrm{T}^{*}(L\times M)\rightarrow\mathrm{T}^{*}M$ is the projection on the second factor post-composed with negation along the fibers is the graph of the symplectomorphism between $U\hookrightarrow\mathrm{T}^{*}L$ and $U\hookrightarrow\mathrm{T}^{*}M$.
  • Under the identification $\mathrm{J}^{1}L\times^{\ast}\overline{\mathrm{J}^{1}M}\cong\mathrm{J}^{1}(L\times M)$ from Example 1.7 the graph $\Gamma_{\psi}$ of the local contactomorphism $\mathrm{J}^{1}L\cong\mathrm{J}^{1}M$ corresponds to the Legendrian $j^{1}h\subset\mathrm{J}^{1}(L\times M)$.
  • The correspondence
[…]
    is the graph of the local homogeneous symplectomorphism between $\mathrm{T}^{*}(L\times\mathbf{C})$ and $\mathrm{T}^{*}(M\times\mathbf{C})$, where $Z=\{t_{L}-t_{M}=h\}\subset L\times\mathbf{C}\times M\times\mathbf{C}$.
Consider the map $\Xi\colon L\rightarrow L\times M$ from 1.29. By construction the critical loci of $f$ on $L$ and $h$ on $L\times M$ are both $L\cap M$. So, by [Joy15, Proposition 2.25] there is a canonical quadratic form $q_{\Xi}$ on the normal bundle $N_{\Xi}|_{L\cap M}$ of the embedding $\Xi\colon L\rightarrow L\times M$. It has the following alternative description. For any $x\in L\cap M$ we have a sequence $\mathrm{T}_{M,x},\mathrm{T}_{\pi_{1},x},\mathrm{T}_{\pi_{2},x},\mathrm{T}_{M,x}$ of Lagrangian subspaces of $\mathrm{T}_{S,x}$, where the subsequent subspaces are transverse. Therefore, we have the Maslov quadratic form $q(\mathrm{T}_{M},\mathrm{T}_{\pi_{1}},\mathrm{T}_{\pi_{2}})$ on $\mathrm{T}_{M}|_{L\cap M}$.
Proposition 1.30.
Under the natural isomorphism $N_{\Xi}\cong\mathrm{T}_{M}$ provided by the projection $L\times M\rightarrow M$ the quadratic form $q_{\Xi}$ on $N_{\Xi}|_{L\cap M}$ is equal to the Maslov quadratic form $q(\mathrm{T}_{M},\mathrm{T}_{\pi_{1}},\mathrm{T}_{\pi_{2}})$ on $\mathrm{T}_{M}|_{L\cap M}$. In particular, the quadratic form $q_{\Xi}$ on $\mathrm{T}_{M}|_{L\cap M}$ induces a quadratic form on $(\det\mathrm{T}^{*}_{M})^{\otimes 2}|_{L\cap M}$ given by
$\omega^{\otimes 2}\mapsto\frac{(\pi_{12}^{S\times S\rightarrow M\times M})^{*}(\omega\wedge\omega)}{\mathrm{vol}_{S}}|_{L\cap M},$
where $\pi_{12}^{S\times S\rightarrow M\times M}$ is the composite $S\hookrightarrow S\times S\xrightarrow{\pi_{1}\times\pi_{2}}M\times M$.
Proof.
Consider $x\in L\cap M$. Choose local coordinates $\{l_{a}\}$ on $L$ and $\{m_{i}\}$ on $M$. We have a local diffeomorphism $L\times M\rightarrow\mathrm{T}^{*}L$ near $(x,x)$ given by $(l,m)\mapsto(l,\frac{\partial h}{\partial l_{a}})$. Similarly, we have a local diffeomorphism $L\times M\rightarrow\mathrm{T}^{*}M$ near $(x,x)$ given by $(l,m)\mapsto(m,-\frac{\partial h}{\partial m_{i}})$. The induced maps on tangent spaces are given by the matrices
$\begin{pmatrix}\mathrm{id}&0\\ \frac{\partial^{2}h}{\partial l_{a}\partial l_{b}}&\frac{\partial^{2}h}{\partial l_{a}\partial m_{i}}\end{pmatrix}\colon\mathrm{T}_{L\times M,(x,x)}\longrightarrow\mathrm{T}_{\mathrm{T}^{*}L,(x,0)}$
and
$\begin{pmatrix}0&\mathrm{id}\\ -\frac{\partial^{2}h}{\partial l_{a}\partial m_{i}}&-\frac{\partial^{2}h}{\partial m_{i}\partial m_{j}}\end{pmatrix}\colon\mathrm{T}_{L\times M,(x,x)}\longrightarrow\mathrm{T}_{\mathrm{T}^{*}M,(x,0)}.$
By assumption the matrix $\frac{\partial^{2}h}{\partial l_{a}\partial m_{i}}(x)$ is invertible. Let us denote by $A_{ia}$ its inverse. Thus, the composite $\mathrm{T}_{\mathrm{T}^{*}L,(x,0)}\rightarrow\mathrm{T}_{L\times M,(x,x)}\rightarrow\mathrm{T}_{\mathrm{T}^{*}M,(x,0)}$ is given by
$\begin{pmatrix}0&\mathrm{id}\\ -\frac{\partial^{2}h}{\partial l_{a}\partial m_{i}}&-\frac{\partial^{2}h}{\partial m_{i}\partial m_{j}}\end{pmatrix}\begin{pmatrix}\mathrm{id}&0\\ -\sum_{b}\frac{\partial^{2}h}{\partial l_{a}\partial l_{b}}A_{bi}&A_{ai}\end{pmatrix}=\begin{pmatrix}-\sum_{b}\frac{\partial^{2}h}{\partial l_{a}\partial l_{b}}A_{bi}&A_{ai}\\ -\frac{\partial^{2}h}{\partial l_{a}\partial m_{i}}+\sum_{b,j}\frac{\partial^{2}h}{\partial m_{i}\partial m_{j}}\frac{\partial^{2}h}{\partial l_{a}\partial l_{b}}A_{bj}&-\sum_{j}\frac{\partial^{2}h}{\partial m_{i}\partial m_{j}}A_{aj}\end{pmatrix}$
Therefore, the subspace $\mathrm{T}_{\pi_{1},x}\subset\mathrm{T}_{\mathrm{T}^{*}M,(x,0)}\cong\mathrm{T}_{M,x}\oplus\mathrm{T}^{*}_{M,x}$ is given by
$\left\{\left(\sum_{a}A_{ai}v^{a},-\sum_{a,j}\frac{\partial^{2}h}{\partial m_{i}\partial m_{j}}A_{aj}v^{a}\right)\mid v^{a}\in\mathrm{T}_{M,x}\right\}.$
Comparing with Example 1.15 we deduce that the Maslov quadratic form $q(\mathrm{T}_{M,x},\mathrm{T}_{\pi_{1},x},\mathrm{T}_{\pi_{2},x})$ is given by $\frac{\partial^{2}h}{\partial m_{i}\partial m_{j}}$.
By construction the embedding $\Xi\colon L\hookrightarrow L\times M$ is given by the equation $\frac{\partial h}{\partial m_{i}}=0$. Therefore, the induced quadratic form $q_{\Xi}$ on the normal bundle is given by $\frac{\partial^{2}h}{\partial m_{i}\partial m_{j}}$ which coincides with the Maslov quadratic form $q(\mathrm{T}_{M,x},\mathrm{T}_{\pi_{1},x},\mathrm{T}_{\pi_{2},x})$. ∎
2. Monodromic sheaves
2.1. Monodromic $D$-modules
We begin with the following well-known result.
Proposition 2.1.
Let $\mathcal{C}$ be an idempotent complete $\mathbf{C}$-linear category and $M\in\mathcal{C}$ an object with $\mathrm{End}_{\mathcal{C}}(M)$ finite-dimensional. Given an endomorphism $\theta\colon M\rightarrow M$ there is a decomposition
$M\cong\bigoplus_{\lambda\in\mathbf{C}}M^{\lambda},$
where only finitely many $M^{\lambda}$ are nonzero, and where $\theta$ acts on $M^{\lambda}$ with generalized eigenvalue $\lambda$.
In this paper we apply the above result to the category of regular holonomic $D$-modules and perverse sheaves on a complex manifold.
For a complex manifold $Y$ denote by $\mathrm{Mod}_{\mathrm{rh}}(\mathcal{D}_{Y})$ the stack of regular holonomic $D$-modules on $Y$. Now suppose $Y$ is equipped with a $\mathbf{C}^{\times}$-action and denote by $\theta\in\mathcal{D}_{Y}$ the generating vector field. Suppose $\gamma\colon Y\rightarrow X$ is $\mathbf{C}^{\times}$-equivariant with respect to the trivial action on $X$.
Definition 2.2.
A (regular holonomic) $D$-module $\mathcal{M}$ on $Y$ is monodromic if it is weakly $\mathbf{C}^{\times}$-equivariant. Denote by $\mathrm{Mod}_{\mathrm{rh},\mathrm{mon}}(\mathcal{D}_{Y})\subset\gamma_{*}\mathrm{Mod}_{\mathrm{rh}}(\mathcal{D}_{Y})$ the stack of monodromic regular holonomic $D$-modules.
Assume that $Y=X\times\mathbf{C}^{\times}$ for a complex manifold $X$ with the obvious $\mathbf{C}^{\times}$-action on the second factor. Let $\gamma\colon Y\rightarrow X$ be the projection on the first factor. For a $\mathcal{D}_{Y}$-module $\mathcal{M}$ denote
$\mathcal{M}^{\lambda}=\bigcup_{l\geq 0}\ker((\theta-\lambda\cdot\mathrm{id})^{l}\colon\gamma_{*}\mathcal{M}\rightarrow\gamma_{*}\mathcal{M})$
its generalized eigenspace with eigenvalue $\lambda\in\mathbf{C}$. Let
$[\mathcal{M}]=\bigoplus_{\lambda\in G}\mathcal{M}^{\lambda}.$
For instance,
$[\mathcal{D}_{Y}]\cong\mathcal{D}_{X}[\theta].$
If $\mathcal{M}\in\mathrm{Mod}_{\mathrm{rh}}(\mathcal{D}_{Y})$, then $[\mathcal{M}]$ is a regular holonomic $\mathcal{D}_{X}$-module with an endomorphism $\theta$. For brevity we call such a module a regular holonomic $\mathcal{D}_{X}[\theta]$-module. The following statement is standard.
Proposition 2.3.
A $D$-module $\mathcal{M}\in\mathrm{Mod}_{\mathrm{rh}}(\mathcal{D}_{Y})$ is monodromic if, and only if, either of the following conditions holds:
  • $\mathcal{M}$ is lisse along the orbits of the $\mathbf{C}^{\times}$-action.
  • The natural map
    $\mathcal{D}_{Y}\otimes_{\gamma^{-1}[\mathcal{D}_{Y}]}\gamma^{-1}[\mathcal{M}]\longrightarrow\mathcal{M}$
    is an isomorphism.
The functor
$\mathrm{Mod}_{\mathrm{rh}}(\mathcal{D}_{X}[\theta])\longrightarrow\mathrm{Mod}_{\mathrm{rh},\mathrm{mon}}(\mathcal{D}_{Y})$
given by $\mathcal{M}\mapsto\mathcal{D}_{Y}\otimes_{\gamma^{-1}\mathcal{D}_{X}[\theta]}\gamma^{-1}\mathcal{M}$ identifies $\mathrm{Mod}_{\mathrm{rh},\mathrm{mon}}(\mathcal{D}_{Y})$ with the full subcategory of regular holonomic $\mathcal{D}_{X}[\theta]$-modules, where the eigenvalues of $\theta$ are contained in $G$.
It will be convenient to introduce another perspective on the category of monodromic $D$-modules. Consider the category $\mathrm{Mod}_{\mathrm{rh},\mathrm{aut}}(\mathcal{D}_{X})$ of regular holonomic $\mathcal{D}_{X}$-modules $\mathcal{M}$ equipped with an automorphism $T\colon\mathcal{M}\rightarrow\mathcal{M}$. There is a functor
$\mathrm{Mod}_{\mathrm{rh}}(\mathcal{D}_{X}[\theta])\longrightarrow\mathrm{Mod}_{\mathrm{rh},\mathrm{aut}}(\mathcal{D}_{X})$
given by setting $T=\exp(-2\pi i\theta)$. It also identifies $\mathrm{Mod}_{\mathrm{rh},\mathrm{aut}}(\mathcal{D}_{X})$ with the full subcategory of regular holonomic $\mathcal{D}_{X}[\theta]$-modules, where the eigenvalues of $\theta$ are contained in $G$.
The resulting equivalence
$\mathrm{Mod}_{\mathrm{rh},\mathrm{mon}}(\mathcal{D}_{Y})\cong\mathrm{Mod}_{\mathrm{rh},\mathrm{aut}}(\mathcal{D}_{X})$
may be understood as follows. Consider the inclusion $i\colon X\times\{1\}\hookrightarrow Y$. Then the composite
$\mathrm{Mod}_{\mathrm{rh},\mathrm{mon}}(\mathcal{D}_{Y})\cong\mathrm{Mod}_{\mathrm{rh},\mathrm{aut}}(\mathcal{D}_{X})\longrightarrow\mathrm{Mod}_{\mathrm{rh}}(\mathcal{D}_{X})$
is given by the pullback $i^{*}$. The automorphism $T$ is given by the monodromy along the $\mathbf{C}^{\times}$ direction.
2.2. Monodromic perverse sheaves
Let us now describe analogous objects on the other side of the Riemann–Hilbert correspondence. Denote by $\mathrm{Perv}_{\mathrm{aut}}(X)$ the stack of $\mathbf{C}$-linear perverse sheaves $\mathcal{F}$ on a complex manifold $X$ equipped with an automorphism $T\colon\mathcal{F}\rightarrow\mathcal{F}$. As in the case of $D$-modules, this stack has the following two equivalent presentations:
  • The stack of perverse sheaves $\mathcal{F}$ on $X$ equipped with an endomorphism $\theta\colon\mathcal{F}\rightarrow\mathcal{F}$ whose eigenvalues lie in $G$.
  • The stack of perverse sheaves on $X\times\mathbf{C}^{\times}$ lisse along $\mathbf{C}^{\times}$.
Given a regular holonomic left $\mathcal{D}_{X}$-module $\mathcal{L}$ and a regular holonomic right $\mathcal{D}_{X}$-module $\mathcal{M}$, the complex of sheaves $\mathcal{M}\otimes^{\mathbb{L}}_{\mathcal{D}_{X}}\mathcal{L}$ is a perverse sheaf on $X$. This statement generalizes to monodromic $D$-modules as follows. Let $Y=X\times\mathbf{C}^{\times}$ with $\gamma\colon Y\rightarrow X$ the projection.
Proposition 2.4.
Consider a regular holonomic left $\mathcal{D}_{X}[\theta]$-module $\mathcal{L}$, a regular holonomic right $\mathcal{D}_{X}[\theta]$-module $\mathcal{M}$ and the monodromic $\mathcal{D}_{Y}$-modules
$\mathcal{L}^{\prime}=\mathcal{D}_{Y}\otimes_{\gamma^{-1}\mathcal{D}_{X}[\theta]}\mathcal{L},\qquad\mathcal{M}^{\prime}=\mathcal{M}\otimes_{\gamma^{-1}\mathcal{D}_{X}[\theta]}\mathcal{D}_{Y}.$
Then:
  • The perverse sheaf
    $\mathcal{M}^{\prime}\otimes^{\mathbb{L}}_{\mathcal{D}_{Y}}\mathcal{L}^{\prime}$
    on $Y$ is lisse along $\mathbf{C}^{\times}$.
  • There is an isomorphism of perverse sheaves
    $\mathcal{M}\otimes_{\mathcal{D}_{X}}^{\mathbb{L}}\mathcal{L}\cong(\mathcal{M}^{\prime}\otimes^{\mathbb{L}}_{\mathcal{D}_{Y}}\mathcal{L}^{\prime})|_{X\times\{1\}}[-1].$
  • Consider the endomorphism $\theta$ on the left-hand side induced by the $\theta$-action on $\mathcal{L}$ and $\mathcal{M}$ and let $T=\exp(2\pi i\theta)$. Then it is equal to the monodromy of the sheaf on the right along the $\mathbf{C}^{\times}$ direction.
We will also consider perverse sheaves of $\mathbf{C}[\![\hbar]\!]$-modules as in [DGS11, Section 5]. We say a perverse sheaf $\mathcal{L}\in\mathrm{Perv}(X;\mathbf{C}[\![\hbar]\!])$ is torsion-free if the map $\hbar\colon\mathcal{L}\rightarrow\mathcal{L}$ is a monomorphism in the abelian category $\mathrm{Perv}(X;\mathbf{C}[\![\hbar]\!])$.
We have a natural functor
$\mathbf{D}_{c}(X;\mathbf{C}[\![\hbar]\!])\longrightarrow\mathbf{D}_{c}(X;\mathbf{C}[\hbar]/\hbar^{n})$
given by $\mathcal{L}\mapsto\mathcal{L}\otimes^{\mathbb{L}}_{\mathbf{C}[\![\hbar]\!]}\mathbf{C}[\hbar]/\hbar^{n}$ which we for brevity denote by $\mathcal{L}\mapsto\mathcal{L}/\hbar^{n}$. If $\mathcal{L}$ is a torsion-free perverse sheaf, then $\mathcal{L}/\hbar^{n}$ can be understood as the quotient in the abelian category of perverse sheaves.
Definition 2.5.
A differential perverse sheaf is a pair $(\mathcal{M},D)$, where $\mathcal{M}$ is a $\mathbf{C}(\!(\hbar)\!)$-linear perverse sheaf and $D$ a $\mathbf{C}$-linear endomorphism of the underlying complex of sheaves of $\mathbf{C}$-vector spaces satisfying
$D(f\eta)=\hbar\frac{df}{d\hbar}\eta+fD(\eta)$
for any local section $\eta$ of $\mathcal{M}$, and $f\in\mathbf{C}(\!(\hbar)\!)$. We denote by $\mathrm{Perv}_{\mathrm{diff}}(X)$ the stack of differential perverse sheaves.
Definition 2.6.
Let $\mathcal{M}\in\mathrm{Perv}_{\mathrm{diff}}(X)$ be a differential perverse sheaf. A lattice $\mathcal{L}\subset\mathcal{M}$ is a torsion-free $\mathbf{C}[\![\hbar]\!]$-perverse sheaf $\mathcal{L}$ together with an isomorphism $\mathcal{L}[\hbar^{-1}]\cong\mathcal{M}$ which is preserved by the action of $D$ with the following property: all eigenvalues of the action of $D$ on the $\mathbf{C}$-linear perverse sheaf $\mathcal{L}/\hbar$ lie in $G$. If $\mathcal{M}\in\mathrm{Perv}_{\mathrm{diff}}(X)$ has a lattice, we say that it has regular singularities.
2.3. The Riemann-Hilbert functor
Our next goal is to define an inverse Riemann–Hilbert functor
$\widehat{\mathrm{RH}}^{-1}\colon\mathrm{Perv}_{\mathrm{aut}}(X)\to\mathrm{Perv}_{\mathrm{diff}}(X)$
relating differential perverse sheaves and perverse sheaves equipped with an automorphism.
Definition 2.7.
Given a sheaf $\mathcal{F}\in\mathbf{D}(X;\mathbf{C})$, we denote by $\mathcal{F}[\![\hbar]\!]$ the inverse limit
$\mathcal{F}[\![\hbar]\!]=\lim_{m}\left(\mathcal{F}\otimes_{\mathbf{C}}\mathbf{C}[\hbar]/\hbar^{m}\right)\in\mathbf{D}(X;\mathbf{C}[\![\hbar]\!]).$
We write
$\mathcal{F}(\!(\hbar)\!)=\mathcal{F}[\![\hbar]\!]\otimes_{\mathbf{C}[\![\hbar]\!]}\mathbf{C}(\!(\hbar)\!)\in\mathbf{D}(X;\mathbf{C}(\!(\hbar)\!)).$
Remark 2.8.
As tensor products preserve colimits, there is a natural isomorphism
$\mathcal{F}(\!(\hbar)\!)=\operatorname*{colim}_{n}\left(\hbar^{-n}\mathcal{F}[\![\hbar]\!]\right)\cong\operatorname*{colim}_{n}\lim_{m}\left(\mathcal{F}\otimes_{\mathbf{C}}\hbar^{-n}\mathbf{C}[\hbar]/\hbar^{m}\right).$
We have the following two technical lemmas.
Lemma 2.9.
Suppose $\mathcal{F}\in\mathbf{D}_{c}(X;\mathbf{C})$ is a constructible complex. Then the natural maps
$\mathcal{F}\otimes_{\mathbf{C}}\mathbf{C}[\![\hbar]\!]\to\mathcal{F}[\![\hbar]\!]$
and
$\mathcal{F}\otimes_{\mathbf{C}}\mathbf{C}(\!(\hbar)\!)\to\mathcal{F}(\!(\hbar)\!)$
are isomorphisms.
Proof.
As $\mathcal{F}$ is a constructible complex, for any point $x\in X$, there is a basis of open neighborhoods $U_{x}$ of $x$ such that the map $\mathcal{F}_{x}\to\mathcal{F}(U_{x})$ is a quasi-isomorphism. In particular, taking the stalk commutes with the inverse limit. Thus, we are reduced to checking that the natural maps
$\mathcal{F}_{x}\otimes_{\mathbf{C}}\mathbf{C}[\![\hbar]\!]\to\lim_{m}\mathcal{F}_{x}\otimes_{\mathbf{C}}\mathbf{C}[\hbar]/\hbar^{m}$
are quasi-isomorphisms of complexes. The projective diagram of complexes is naturally split, so the inverse limit commutes with cohomology. The result then follows from the statement the corresponding statement for finite-dimensional vector spaces, which is well-known. The second statement follows immediately from the first. ∎
Lemma 2.10.
Suppose $\mathcal{F}\in\mathbf{D}_{c}(X;\mathbf{C}[\![\hbar]\!])$ is a constructible complex. Then the natural morphism
$\mathcal{F}\longrightarrow\lim_{n}\mathcal{F}/\hbar^{n}$
is an isomorphism.
Proof.
By [KS12, Proposition 7.1.7] a constructible complex $\mathcal{F}$ is cohomologically complete, i.e.
$\mathbb{R}\mathcal{H}om_{\mathbf{C}_{X}[\![\hbar]\!]}(\mathbf{C}_{X}(\!(\hbar)\!),\mathcal{F})=0.$
The fact that for a cohomologically complete $\mathbf{C}_{X}[\![\hbar]\!]$-module the map
$\mathcal{F}\longrightarrow\lim_{n}\mathcal{F}/\hbar^{n}$
is an isomorphism is the content of [PSY14, Theorem 7.8] in the case $X=\mathrm{pt}$ and the same proof works for sheaves. ∎
We are now ready to define the functor $\widehat{\mathrm{RH}}^{-1}$. Given a perverse sheaf $\mathcal{F}$ equipped with an automorphism $T$ we can write $T=\exp(-2\pi iM)$ for a unique endomorphism $M\colon\mathcal{F}\rightarrow\mathcal{F}$ whose eigenvalues lie in $G\subset\mathbf{C}$. Then we define $\widehat{\mathrm{RH}}^{-1}(\mathcal{F})$ to be the pair $(\mathcal{F}(\!(\hbar)\!),D)$, where
$D(f\otimes\eta)=f\otimes(M\eta)+\hbar\frac{df}{d\hbar}\otimes\eta$
for any local section $\eta$ of $\mathcal{F}$ and $f\in\mathbf{C}(\!(\hbar)\!)$.
Conversely, if $(\mathcal{M},D)\in\mathrm{Perv}_{\mathrm{diff}}(X)$ is a differential perverse sheaf with a lattice $\mathcal{L}\subset\mathcal{M}$, consider $\mathcal{F}=\mathcal{L}/\hbar\in\mathrm{Perv}(X;\mathbf{C})$. By assumption it is preserved by $D$ and we define the monodromy automorphism $T\colon\mathcal{F}\rightarrow\mathcal{F}$ to be $T=\exp(-2\pi iD)$. In this way we get an object $(\mathcal{F},T)\in\mathrm{Perv}_{\mathrm{aut}}(X)$.
Proposition 2.11.
The functor $\widehat{\mathrm{RH}}^{-1}\colon\mathrm{Perv}_{\mathrm{aut}}(X)\rightarrow\mathrm{Perv}_{\mathrm{diff}}(X)$ is fully faithful. Moreover, if $\mathcal{M}\in\mathrm{Perv}_{\mathrm{diff}}(U)$ is a differential perverse sheaf with a lattice $\mathcal{L}\subset\mathcal{M}$, then RH^1(/)superscriptnormal-^normal-$\mathcal{M}\cong\widehat{\mathrm{RH}}^{-1}(\mathcal{L}/\hbar)$. In particular, the essential image of $\widehat{\mathrm{RH}}^{-1}$ consists of differential perverse sheaves with regular singularities.
Proof.
Given a morphism $\mathcal{F}_{1}\rightarrow\mathcal{F}_{2}$ in $\mathrm{Perv}_{\mathrm{aut}}(X)$, a morphism $\widehat{\mathrm{RH}}^{-1}(\mathcal{F}_{1})\rightarrow\widehat{\mathrm{RH}}^{-1}(\mathcal{F}_{2})$ of $\mathbf{C}(\!(\hbar)\!)$-linear perverse sheaves is uniquely specified by its restriction $\mathcal{F}_{1}\rightarrow\mathcal{F}_{2}(\!(\hbar)\!)$ to $\mathcal{F}_{1}\hbar^{0}\subset\mathcal{F}_{1}(\!(\hbar)\!)$. Let $\mathcal{F}_{1}^{\lambda}\subset\mathcal{F}_{1}$ be a generalized eigenspace of $D$ with eigenvalue $m\in G$, where $\exp(-2\pi im)=\lambda$. The map $\mathcal{F}_{1}\rightarrow\mathcal{F}_{2}(\!(\hbar)\!)$ will therefore send $\mathcal{F}_{1}^{\lambda}$ to the $m$-generalized eigenspace of 2(())subscript2Planck-constant-over-$\mathcal{F}_{2}(\!(\hbar)\!)$ which is precisely $\mathcal{F}_{2}^{\lambda}\hbar^{0}\subset\mathcal{F}_{2}(\!(\hbar)\!)$. The resulting map $\mathcal{F}_{1}^{\lambda}\rightarrow\mathcal{F}_{2}^{\lambda}$ is compatible with the nilpotent part of $D$ precisely when it is compatible with the unipotent part of $T$. This proves fully faithfulness.
Now suppose $\mathcal{E}\in\mathrm{Perv}_{\mathrm{diff}}(X)$ has a lattice $\mathcal{L}\subset\mathcal{M}$ and consider the perverse sheaf $\mathcal{F}=\mathcal{L}/\hbar$ equipped with an automorphism. Consider the extension
$0\longrightarrow\mathcal{L}/\hbar\xrightarrow{\hbar}\mathcal{L}/\hbar^{2}\longrightarrow\mathcal{L}/\hbar\longrightarrow 0$
The perverse sheaves $\mathcal{F}=\mathcal{L}/\hbar$ and /2superscriptPlanck-constant-over-$\mathcal{L}/\hbar^{2}$ split into generalized eigenspaces for $D$. But since no two eigenvalues on $\mathcal{F}$ differ by $1$, the above exact sequence is canonically split, i.e. we can identify $\mathcal{L}/\hbar^{2}\cong\mathcal{F}\otimes_{\mathbf{C}}\mathbf{C}[\hbar]/\hbar^{2}$. Proceeding by induction we identify
$\mathcal{L}/\hbar^{n}\cong\mathcal{F}\otimes_{\mathbf{C}}\mathbf{C}[\hbar]/\hbar^{n}$
compatibly with the maps obtained by modding out by a power of $\hbar$. Therefore, by Lemma 2.9
$\lim_{n}\mathcal{L}/\hbar^{n}\cong\mathcal{F}[\![\hbar]\!].$
By Lemma 2.10 the natural morphism
$\mathcal{L}\longrightarrow\lim_{n}\mathcal{L}/\hbar^{n}$
is an isomorphism, so this identifies $\mathcal{L}\cong\mathcal{F}[\![\hbar]\!]$. Therefore, $\mathcal{E}\cong\mathcal{F}(\!(\hbar)\!)=\widehat{\mathrm{RH}}^{-1}(\mathcal{F})$. Under this isomorphism the action of $D$ on $\mathcal{E}$ coincides with the action of $D$ coming from the monodromy operator $T$ on $\mathcal{F}$. ∎
3. Microdifferential systems
Throughout this section $X$ denotes a complex manifold.
3.1. Microdifferential operators
In this section we recall some algebras of microdifferential operators that we will use. We refer to [Kas03, Kas86] for more details. Recall the punctured cotangent bundle $\mathring{\mathrm{T}}^{*}X=\mathrm{T}^{*}X\setminus\mathrm{T}^{*}_{X}X$ and the projectivized cotangent bundle $\mathrm{P}^{*}X=\mathring{\mathrm{T}}^{*}X/\mathbf{C}^{\times}$ with $\gamma\colon\mathring{\mathrm{T}}^{*}X\rightarrow\mathrm{P}^{*}X$ the natural projection. We denote the projections $\mathrm{T}^{*}X\rightarrow X$ and $\mathrm{P}^{*}X\rightarrow X$ by the same letter $\pi$.
Consider the sheaf ^$\widehat{\mathcal{E}}_{X}$ on $\mathrm{P}^{*}X$ of formal microdifferential operators introduced in [SKK73]. It has a filtration
^$\widehat{\mathcal{E}}_{X}=\bigcup_{m\in\mathbf{Z}}\widehat{\mathcal{E}}_{X}(m)$
by order of the microdifferential operator. Denote by
$\sigma_{m}\colon\widehat{\mathcal{E}}_{X}(m)\longrightarrow\mathcal{O}_{\mathrm{T}^{*}X}(m)$
the principal symbol, where $\mathcal{O}_{\mathrm{T}^{*}X}(m)\subset\gamma_{*}\mathcal{O}_{\mathring{\mathrm{T}}^{*}X}$ consists of functions on $\mathring{\mathrm{T}}^{*}X$ homogeneous of order $m$. The principal symbol establishes an isomorphism
$\operatorname{gr}\widehat{\mathcal{E}}_{X}\cong\bigoplus_{m\in\mathbf{Z}}\mathcal{O}_{\mathrm{T}^{*}X}(m).$
Remark 3.1.
Sometimes one considers the sheaf $\mathcal{E}_{X}$ of microdifferential operators; the sheaf ^$\widehat{\mathcal{E}}_{X}$ is obtained from $\mathcal{E}_{X}$ by completion with respect to the order. The theory of regular holonomic $\mathcal{E}_{X}$- and ^$\widehat{\mathcal{E}}_{X}$-modules coincide (see e.g. [KK83, Theorem 6.3]), so the difference is inessential for our purposes.
By construction there is a natural embedding
$\pi^{-1}\mathcal{D}_{X}\rightarrow\widehat{\mathcal{E}}_{X}$
of differential into microdifferential operators.
Definition 3.2.
A coherent ^$\widehat{\mathcal{E}}_{X}$-module is holonomic if its support is a Legendrian subvariety of $\mathrm{P}^{*}X$.
Let $\Lambda\subset\mathrm{P}^{*}X$ a Legendrian subvariety and $\tilde{\Lambda}\subset\mathring{\mathrm{T}}^{*}X$ the corresponding conic Lagrangian subvariety. Let ^$\widehat{\mathcal{E}}_{\Lambda/X}\subset\widehat{\mathcal{E}}_{X}|_{\Lambda}$ be the subalgebra generated by
$I_{\Lambda}=\{P\in\widehat{\mathcal{E}}_{X}(1)|_{\Lambda}\mid\sigma_{1}(P)|_{\Lambda}=0\}.$
Definition 3.3.
A holonomic ^$\widehat{\mathcal{E}}_{X}$-module $\mathcal{M}$ is regular holonomic if locally there is a coherent ^$\widehat{\mathcal{E}}_{X}(0)$-submodule $\mathcal{M}_{0}\subset\mathcal{M}$ and a Legendrian subvariety $\Lambda\subset\mathrm{P}^{*}X$ containing $\operatorname{supp}\mathcal{M}$ such that $I_{\Lambda}\mathcal{M}_{0}\subset\mathcal{M}_{0}$ and ^$\widehat{\mathcal{E}}_{X}\mathcal{M}_{0}=\mathcal{M}$.
Remark 3.4.
A $\mathcal{D}_{X}$-module $\mathcal{M}$ is (regular) holonomic if, and only if, the ^$\widehat{\mathcal{E}}_{X}$-module ^$\widehat{\mathcal{E}}_{X}\otimes_{\pi^{-1}\mathcal{D}_{X}}\pi^{-1}\mathcal{M}$ is (regular) holonomic.
3.2. Specialization and microlocalization
Let us now recall the theory of $V$-filtration for microdifferential operators. Let $Z\subset X$ be a submanifold with $J_{Z}\subset\mathcal{O}_{X}|_{Z}$ the vanishing ideal. Define the Kashiwara–Malgrange filtration of $\mathcal{D}_{X}|_{Z}$ by
$V^{k}_{Z}\mathcal{D}_{X}=\{P\in\mathcal{D}_{X}|_{Z}\mid PJ_{Z}^{j}\subset J_{Z}^{j+k},\ \forall j,j+k\geq 0\}.$
As shown in [Kas83], there is a natural embedding
$\operatorname{gr}_{Z}\mathcal{D}_{X}\subset\tau_{*}\mathcal{D}_{\mathrm{T}_{Z}X},$
where $\tau\colon\mathrm{T}_{Z}X\rightarrow Z$ is the projection, whose image $\mathcal{D}_{[\mathrm{T}_{Z}X]}$ consists of differential operators polynomial along the fibers of $\tau$. Let $\theta\in V^{0}_{Z}\mathcal{D}_{X}$ be any operator lifting the Euler vector field in $\operatorname{gr}_{Z}\mathcal{D}_{X}=\mathcal{D}_{[\mathrm{T}_{Z}X]}$ which acts on $J_{Z}/J_{Z}^{2}$ by the identity.
If $\mathcal{M}\in\mathrm{Mod}_{\mathrm{rh}}(\mathcal{D}_{X})$ is a regular holonomic $\mathcal{D}_{X}$-module and $u\in\mathcal{M}$ is a local nonvanishing section, then there is a polynomial $b_{u}(s)\in\mathbf{C}[s]$ such that $b_{u}(\theta)u\in V^{1}_{Z}(\mathcal{D}_{X})u$. The order $\mathrm{ord}_{Z}(u)$ of $u\in\mathcal{M}$ is the set of zeros of its $b$-function. The canonical $V$-filtration of $\mathcal{M}$ is given by the subspaces
$V^{k}_{Z}\mathcal{M}=\{u\in\mathcal{M}|_{Z}\mid\operatorname{Re}\mathrm{ord}_{Z}(u)\geq k-1\},\qquad k\in\mathbf{Z}.$
Denote by $\operatorname{gr}_{Z}\mathcal{M}$ the associated graded module with respect to the canonical $V$-filtration $V^{k}_{Z}\mathcal{M}$. We refer to [Kas83, Mon96] for details on the following functors.
Definition 3.5.
Let $Z\subset X$ be a submanifold.
  • The specialization functor
    $\nu^{dR}_{Z}\colon\mathrm{Mod}_{\mathrm{rh}}(\mathcal{D}_{X})\longrightarrow\tau_{*}\mathrm{Mod}_{\mathrm{rh},\mathrm{mon}}(\mathcal{D}_{\mathrm{T}_{Z}X})$
    is
    $\nu^{dR}_{Z}(\mathcal{M})=\mathcal{D}_{\mathrm{T}_{Z}X}\otimes_{\tau^{-1}\operatorname{gr}_{Z}\mathcal{D}_{X}}\tau^{-1}\operatorname{gr}_{Z}\mathcal{M},$
    where $\mathrm{Mod}_{\mathrm{rh},\mathrm{mon}}(\mathcal{D}_{\mathrm{T}_{Z}X})\subset\mathrm{Mod}_{\mathrm{rh}}(\mathcal{D}_{\mathrm{T}_{Z}X})$ is the subcategory of monodromic regular holonomic $D$-modules.
  • The microlocalization functor
    $\mu^{dR}_{Z}\colon\mathrm{Mod}_{\mathrm{rh}}(\mathcal{D}_{X})\longrightarrow\pi_{*}\mathrm{Mod}_{\mathrm{rh},\mathrm{mon}}(\mathcal{D}_{\mathrm{T}^{*}_{Z}X})$
    is the composite of the specialization functor $\nu^{dR}_{Z}$ with the formal Fourier transform, where $\pi\colon\mathrm{T}^{*}X\rightarrow X$ is the projection.
There is a microlocal analog of the above construction as follows. Let $\Lambda\subset\mathrm{P}^{*}X$ be a Legendrian submanifold and $\tilde{\Lambda}\subset\mathring{\mathrm{T}}^{*}X$ the corresponding conic Lagrangian submanifold. Consider the filtration on ^$\widehat{\mathcal{E}}_{X}|_{\Lambda}$ defined by
^$\widehat{\mathcal{E}}_{\Lambda/X}(m)=\widehat{\mathcal{E}}_{\Lambda/X}\widehat{\mathcal{E}}_{X}(m)|_{\Lambda}.$
Consider the line bundle
$K_{\tilde{\Lambda}|X}=K_{\tilde{\Lambda}}\otimes\pi^{*}K_{X}^{-1}$
on $\tilde{\Lambda}$. By [KK81, Lemma 1.5.1] there is a natural embedding
$L\colon\operatorname{gr}_{\Lambda}\widehat{\mathcal{E}}_{X}\hookrightarrow\gamma_{*}\mathcal{D}_{\tilde{\Lambda}}^{K^{1/2}_{\tilde{\Lambda}|X}},$
where $\gamma\colon\tilde{\Lambda}\rightarrow\Lambda$, whose image $\mathcal{D}_{[\tilde{\Lambda}]}^{K^{1/2}_{\tilde{\Lambda}|X}}$ consists of differential operators polynomial along the fibers of $\gamma$.
Consider the Euler vector field on $\mathrm{T}^{*}X$. Since $\tilde{\Lambda}\subset\mathrm{T}^{*}X$ is homogeneous, the Lie derivative along the Euler vector field defines an element $\theta^{\prime}\in\mathcal{D}_{\tilde{\Lambda}}^{K^{1/2}_{\tilde{\Lambda}|X}}$. We denote by $\theta^{\prime}\in\widehat{\mathcal{E}}_{X}(1)|_{\Lambda}$ any preimage (such a lift is characterized by [KK81, Equations (1.5.5) and (1.5.6)]). If $\mathcal{M}$ is a regular holonomic ^$\widehat{\mathcal{E}}_{X}$-module and $u\in\mathcal{M}$ is a local nonvanishing section, then by [KK80, Theorem 4.1.1] there is a polynomial $b_{u}(s)\in\mathbf{C}[s]$ such that $b_{u}(\theta^{\prime})u\in\widehat{\mathcal{E}}_{\Lambda/X}(-1)u$. The order $\mathrm{ord}_{\Lambda}(u)$ of $u\in\mathcal{M}$ is the set of zeros of its $b$-function. For a fixed $\alpha\in\mathbf{R}$ the canonical $V$-filtration of $\mathcal{M}$ is given by the subspaces
$V^{k}_{\Lambda}\mathcal{M}=\{u\in\mathcal{M}|_{\Lambda}\mid\operatorname{Re}\mathrm{ord}_{\Lambda}(u)\leq\alpha-k\},\qquad k\in\mathbf{Z}.$
Denote by $\operatorname{gr}_{\Lambda}\mathcal{M}$ the associated graded module with respect to the canonical $V$-filtration $V^{k}_{\Lambda}\mathcal{M}$.
Definition 3.6.
Let $\Lambda\subset\mathrm{P}^{*}X$ be a Legendrian submanifold and $\tilde{\Lambda}\subset\mathring{\mathrm{T}}^{*}X$ the corresponding conic Lagrangian. The microlocalization functor
$\mu^{dR}_{\Lambda}\colon\mathrm{Mod}_{\mathrm{rh}}(\widehat{\mathcal{E}}_{X})\longrightarrow\mathrm{Mod}_{\mathrm{rh},\mathrm{mon}}\left(\mathcal{D}^{K_{\tilde{\Lambda}|X}^{1/2}}_{\tilde{\Lambda}}\right)$
is
$\mu^{dR}_{\Lambda}(\mathcal{M})=\mathcal{D}^{K_{\tilde{\Lambda}|X}^{1/2}}_{\tilde{\Lambda}}\otimes_{\gamma^{-1}\operatorname{gr}_{\Lambda}\widehat{\mathcal{E}}_{X}}\gamma^{-1}\operatorname{gr}_{\Lambda}\mathcal{M}.$
Suppose $\mathcal{M}$ is a regular holonomic $\mathcal{D}_{X}$-module, $Z\subset X$ a submanifold and $\Lambda=\mathrm{P}^{*}_{Z}X\subset\mathrm{P}^{*}X$ its projectivized conormal bundle. If we denote by $K_{Z|X}=K_{Z}\otimes K_{X}^{-1}$ the relative canonical bundle, then there is a natural isomorphism
$K_{\tilde{\Lambda}|X}\cong\pi^{*}K_{Z|X}^{\otimes 2}.$
Consider the microlocalization
~=^$\tilde{\mathcal{M}}=\widehat{\mathcal{E}}_{X}\otimes_{\pi^{-1}\mathcal{D}_{X}}\pi^{-1}\mathcal{M}.$
Fix $\alpha=\operatorname{codim}Z/2+1$. Then we may consider the $V$-filtration on $\mathcal{M}$ along $Z$ or its microlocal version on $\tilde{\mathcal{M}}$ along $\Lambda$. They are related as follows (see [Mon96, Remark 1.2.4]): the natural map $\pi^{-1}\mathcal{M}\rightarrow\tilde{\mathcal{M}}$ induces an isomorphism
(2) $\pi^{-1}\operatorname{gr}^{k}_{Z}\mathcal{M}\cong K_{Z|X}^{-1}\otimes_{\pi^{-1}\mathcal{O}_{X}}\operatorname{gr}^{k}_{\Lambda}\tilde{\mathcal{M}}$
for any $k\leq 0$.
3.3. Half twists
Let $X$ be a complex manifold. The adjoint of a differential operator defines a map
$\ast\colon\mathcal{D}_{X}\longrightarrow K_{X}\otimes\mathcal{D}_{X}\otimes K_{X}^{-1}.$
One may define the sheaf of half-twisted differential operators $\mathcal{D}^{\sqrt{v}}_{X}$, so that if $K_{X}^{1/2}$ is a square root of the canonical bundle, one has
$\mathcal{D}^{\sqrt{v}}_{X}=K^{1/2}_{X}\otimes_{\mathcal{O}_{X}}\mathcal{D}_{X}\otimes_{\mathcal{O}_{X}}K^{-1/2}_{X}.$
The sheaf $\mathcal{D}^{\sqrt{v}}_{X}$ carries an antiinvolution $\ast$ given by passing to the adjoint differential operator. A vector field $\xi\in T_{X}$ defines a section of $\mathcal{D}^{\sqrt{v}}_{X}$ via Lie derivative. By construction $\xi^{*}=-\xi$.
Remark 3.7.
The sheaf $\mathcal{D}^{\sqrt{v}}_{X}$ is independent of the choice of the square root of the canonical bundle.
Similarly, let
^$\widehat{\mathcal{E}}^{\sqrt{v}}_{X}=\pi^{-1}K^{1/2}_{X}\otimes_{\pi^{-1}\mathcal{O}_{X}}\widehat{\mathcal{E}}_{X}\otimes_{\pi^{-1}\mathcal{O}_{X}}\pi^{-1}K^{-1/2}_{X}$
be the sheaf of half-twisted microdifferential operators. The antiinvolution $\ast$ on $\mathcal{D}^{\sqrt{v}}_{X}$ extends to one on ^$\widehat{\mathcal{E}}^{\sqrt{v}}_{X}$.
The previous constructions extend to the half-twisted setting in a straightforward way. For instance, for a Legendrian submanifold $\Lambda\subset\mathrm{P}^{*}X$ there is a subalgebra ^$\widehat{\mathcal{E}}^{\sqrt{v}}_{\Lambda/X}\subset\widehat{\mathcal{E}}^{\sqrt{v}}_{X}|_{\Lambda}$, an embedding
$L\colon\operatorname{gr}_{\Lambda}\widehat{\mathcal{E}}^{\sqrt{v}}_{X}\hookrightarrow\gamma_{*}\mathcal{D}_{\tilde{\Lambda}}^{\sqrt{v}}$
whose image $\mathcal{D}_{[\tilde{\Lambda}]}^{\sqrt{v}}$ consists of twisted differential operators polynomial along the fibers of $\gamma$ and a microlocalization functor
$\mu^{dR}_{\Lambda}\colon\mathrm{Mod}_{\mathrm{rh}}(\widehat{\mathcal{E}}^{\sqrt{v}}_{X})\longrightarrow\gamma_{*}\mathrm{Mod}_{\mathrm{rh},\mathrm{mon}}\left(\mathcal{D}^{\sqrt{v}}_{\tilde{\Lambda}}\right).$
Any left ^$\widehat{\mathcal{E}}_{X}^{\sqrt{v}}$-module $\mathcal{M}$ may be considered as a right ^$\widehat{\mathcal{E}}_{X}^{\sqrt{v}}$-module via the formula $uP:=P^{\ast}u$ for $P\in\widehat{\mathcal{E}}_{X}^{\sqrt{v}}$, $u\in\mathcal{M}$. We will freely pass between left and right modules using the antiinvolution $\ast$.
3.4. Microdifferential modules
Let $\Lambda\subset\mathrm{P}^{*}X$ be a Legendrian submanifold and $\tilde{\Lambda}\subset\mathring{\mathrm{T}}^{*}X$ the corresponding conic Lagrangian submanifold.
Definition 3.8.
Let $\mathcal{M}$ be a coherent ^$\widehat{\mathcal{E}}_{X}$-module supported on $\Lambda$. It is regular along $\Lambda$ if there is locally a coherent ^$\widehat{\mathcal{E}}_{\Lambda/X}$-submodule $\mathcal{M}_{0}\subset\mathcal{M}|_{\Lambda}$ generating $\mathcal{M}|_{\Lambda}$ over ^$\widehat{\mathcal{E}}_{X}|_{\Lambda}$.
If $\mathcal{M}$ is regular along a Legendrian $\Lambda\subset\mathrm{P}^{*}X$, then it is a regular holonomic ^$\widehat{\mathcal{E}}_{X}$-module and $\mu_{\Lambda}^{dR}(\mathcal{M})$ is a vector bundle of finite rank on $\tilde{\Lambda}$.
Definition 3.9.
An ^$\widehat{\mathcal{E}}_{X}$-module $\mathcal{M}$ is simple along $\Lambda$ if it is regular along $\Lambda$ and $\mu_{\Lambda}^{dR}(\mathcal{M})$ is a line bundle.
The notion of regular and simple ^$\widehat{\mathcal{E}}^{\sqrt{v}}_{X}$-modules is defined similarly. If $\mathcal{M}$ is an ^$\widehat{\mathcal{E}}^{\sqrt{v}}_{X}$-module simple along $\Lambda$, then for any choice of the square root $K^{1/2}_{\tilde{\Lambda}}$ we have a rank 1 local system
$\mathcal{H}om_{\mathcal{D}^{\sqrt{v}}_{\tilde{\Lambda}}}\left(\mu_{\Lambda}^{dR}(\mathcal{M}),K^{1/2}_{\tilde{\Lambda}}\right)$
which is a subsheaf of $K^{1/2}_{\tilde{\Lambda}}$. For any simple generator $u\in\mathcal{M}$ it allows us to define the symbol
$\sigma_{\Lambda}(u)\in K^{1/2}_{\tilde{\Lambda}}/\mathbf{C}^{\times}_{\tilde{\Lambda}}.$
It satisfies the following properties:
  • If $u$ has order $\lambda\in\mathbf{C}$, then the Euler vector field $\theta^{\prime}$ acts on $\sigma_{\Lambda}(u)$ by $\lambda$.
  • If $P\in\widehat{\mathcal{E}}^{\sqrt{v}}_{X}(m)$ is a section such that $\sigma_{m}(P)|_{\Lambda}$ never vanishes, then $Pu$ is also a simple generator and
    $\sigma_{\Lambda}(Pu)=\sigma_{m}(P)\sigma_{\Lambda}(u)$
    and $\mathrm{ord}(Pu)=\mathrm{ord}(u)+m$.
  • The order of simple ^$\widehat{\mathcal{E}}^{\sqrt{v}}_{X}$-module is $\mathrm{ord}(\mathcal{M})\in\mathbf{C}/\mathbf{Z}$ defined to be the equivalence class of $\mathrm{ord}(u)$ for a simple generator $u$.
We have the following important classification result of microdifferential modules regular along a Legendrian (see [Kas86, Theorem 10.3.1]). Let
$\mathrm{Mod}_{\mathrm{lisse}}\left(\mathcal{D}^{\sqrt{v}}_{\tilde{\Lambda}}\right)\subset\mathrm{Mod}_{\mathrm{rh}}\left(\mathcal{D}^{\sqrt{v}}_{\tilde{\Lambda}}\right)$
be the full substack of twisted $D$-modules on $\tilde{\Lambda}$ which are lisse, i.e. whose underlying $\mathcal{O}$-module is a finite rank vector bundle.
Theorem 3.10.
Let $\Lambda\subset\mathrm{P}^{*}X$ be a Legendrian submanifold. Microlocalization defines an equivalence of stacks
$\mu_{\Lambda}^{dR}\colon\mathrm{Mod}_{\Lambda,\mathrm{rh}}(\widehat{\mathcal{E}}^{\sqrt{v}}_{X})\xrightarrow{\sim}\gamma_{*}\mathrm{Mod}_{\mathrm{lisse}}\left(\mathcal{D}^{\sqrt{v}}_{\tilde{\Lambda}}\right)$
between the stack of ^$\widehat{\mathcal{E}}^{\sqrt{v}}_{X}$-modules regular along $\Lambda$ and lisse twisted $D$-modules on $\tilde{\Lambda}$.
Remark 3.11.
Locally on $\Lambda$ twisted $D$-modules on $\tilde{\Lambda}$ whose underlying $\mathcal{O}$-module is a line bundle are classified by an element $\mathbf{C}/\mathbf{Z}$ given by the eigenvalues of the Euler operator. So, simple ^$\widehat{\mathcal{E}}^{\sqrt{v}}_{X}$-modules along $\Lambda$ are locally classified by the order.
3.5. Quantized contact transformations
To define the stack of microdifferential modules for a general contact manifold $Y$, Kashiwara in [Kas96] uses the theory of quantized contact transformations developed in [SKK73] that we now recall.
Definition 3.12.
Suppose $V_{X}\subset\mathrm{P}^{*}X$ and $V_{Y}\subset\mathrm{P}^{*}Y$ are open subsets together with a contactomorphism $\psi\colon V_{X}\xrightarrow{\sim}V_{Y}$. A quantized contact transformation over $\psi$ is an isomorphism
$\Psi\colon\widehat{\mathcal{E}}_{Y}|_{V_{Y}}\xrightarrow{\sim}\psi_{*}\widehat{\mathcal{E}}_{X}|_{V_{X}}$
of sheaves of filtered $\mathbf{C}$-algebras which coincides with $\psi$ after passing to the associated graded algebras.
By [Kas03, Theorem 8.11] quantized contact transformations exist locally. We will also need a half-twisted version as follows.
Definition 3.13.
Suppose $V_{X}\subset\mathrm{P}^{*}X$ and $V_{Y}\subset\mathrm{P}^{*}Y$ are open subsets together with a contactomorphism $\psi\colon V_{X}\xrightarrow{\sim}V_{Y}$. A normal-$\ast$-preserving quantized contact transformation over $\psi$ is an isomorphism
$\Psi\colon\widehat{\mathcal{E}}^{\sqrt{v}}_{Y}|_{V_{Y}}\xrightarrow{\sim}\psi_{*}\widehat{\mathcal{E}}^{\sqrt{v}}_{X}|_{V_{X}}$
of sheaves of filtered $\mathbf{C}$-algebras which intertwines the antiinvolutions $\ast$ and which coincides with $\psi$ after passing to the associated graded algebras.
The construction of the microlocalization functor $\mu^{dR}_{\Lambda}$ involves the filtration of microdifferential operators by order and the involution $\ast$. Therefore, microlocalization (and, a posteriori, the notion of symbols and orders of simple ^$\widehat{\mathcal{E}}^{\sqrt{v}}_{X}$-modules) commutes with $\ast$-preserving quantized contact transformations.
Proposition 3.14.
Consider a normal-$\ast$-preserving quantized contact transformation $\Psi$ over $\psi\colon V_{X}\xrightarrow{\sim}V_{Y}$ as in the above definition. Suppose $\Lambda_{X}\subset V_{X}$ and $\Lambda_{Y}\subset V_{Y}$ are two Legendrians such that $\Lambda_{Y}=\psi(\Lambda_{X})$. Then a normal-$\ast$-preserving quantized contact transformation $\Psi\colon\widehat{\mathcal{E}}^{\sqrt{v}}_{Y}|_{V_{Y}}\xrightarrow{\sim}\psi_{*}\widehat{\mathcal{E}}^{\sqrt{v}}_{X}|_{V_{X}}$ restricts to an isomorphism
$\Psi\colon\widehat{\mathcal{E}}^{\sqrt{v}}_{\Lambda_{Y}/Y}|_{V_{Y}}\xrightarrow{\sim}\psi_{*}\widehat{\mathcal{E}}^{\sqrt{v}}_{\Lambda_{X}/X}|_{V_{X}}.$
Moreover, the diagram
${\operatorname{gr}_{\Lambda}\widehat{\mathcal{E}}^{\sqrt{v}}_{\Lambda_{Y}/Y}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\Psi}$${\gamma_{*}\mathcal{D}_{\tilde{\Lambda_{Y}}}^{\sqrt{v}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\psi}$${\operatorname{gr}_{\Lambda}\widehat{\mathcal{E}}^{\sqrt{v}}_{\Lambda_{X}/X}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\gamma_{*}\mathcal{D}_{\tilde{\Lambda_{X}}}^{\sqrt{v}}}$
is commutative. In particular, there is a commutative diagram
${\mathrm{Mod}_{\mathrm{rh}}(\widehat{\mathcal{E}}^{\sqrt{v}}_{Y})|_{V_{Y}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mu_{\Lambda_{Y}}^{dR}}$$\scriptstyle{\Psi}$${\gamma_{*}\mathrm{Mod}_{\mathrm{rh},\mathrm{mon}}\left(\mathcal{D}^{\sqrt{v}}_{\tilde{\Lambda_{Y}}}\right)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\psi}$${\psi_{*}\mathrm{Mod}_{\mathrm{rh}}(\widehat{\mathcal{E}}^{\sqrt{v}}_{X})|_{V_{X}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mu_{\Lambda_{X}}^{dR}}$${\gamma_{*}\mathrm{Mod}_{\mathrm{rh},\mathrm{mon}}\left(\mathcal{D}^{\sqrt{v}}_{\tilde{\Lambda_{X}}}\right)}$
It will be convenient to rephrase quantized contact transformations as follows. Let $U_{X}\subset\mathrm{T}^{*}X$ and $U_{Y}\subset\mathrm{T}^{*}Y$ be the preimages of $V_{X},V_{Y}$ and $\tilde{\psi}\colon U_{X}\xrightarrow{\sim}U_{Y}$ the homogeneous symplectic transformation lifting $\psi$. Let
$p_{1}\colon\mathrm{T}^{*}(X\times Y)\rightarrow\mathrm{T}^{*}X$
be the projection on the first factor and
$p_{2}^{a}\colon\mathrm{T}^{*}(X\times Y)\rightarrow\mathrm{T}^{*}Y$
be the projection on the second factor post-composed with the multiplication by $-1$ along the cotangent fibers.
We have the symplectic volume form $\mathrm{vol}_{\mathrm{T}^{*}Y}$ on $\mathrm{T}^{*}Y$. So, there is a canonical choice of the square root line bundle $K^{1/2}_{\mathrm{T}^{*}Y}$ with section $\sqrt{\mathrm{vol}_{\mathrm{T}^{*}Y}}$.
Proposition 3.15.
The data of a quantized contact transformation $\Psi\colon\widehat{\mathcal{E}}^{\sqrt{v}}_{Y}|_{V_{Y}}\xrightarrow{\sim}\psi_{*}\widehat{\mathcal{E}}^{\sqrt{v}}_{X}|_{V_{X}}$ over $\psi$ is equivalent to the data of a ^$\widehat{\mathcal{E}}^{\sqrt{v}}_{X\times Y}$-module $\mathcal{M}$ defined on an open subset $U\subset\mathrm{T}^{*}(X\times Y)$ which is simple along a Legendrian $\Lambda$ with a simple generator $K$ satisfying the following properties:
  • $p_{1}(U)\subset U_{X}$ and $p_{1}\colon\tilde{\Lambda}\rightarrow U_{X}$ is an isomorphism.
  • $p_{2}^{a}(U)\subset U_{Y}$ and $p^{a}_{2}\colon\tilde{\Lambda}\rightarrow U_{Y}$ is an isomorphism.
The relationship between the pair $(\mathcal{M},K)$ and the QCT $\Psi$ is given by
$\Psi(Q)K=Q^{*}K.$
If $\Psi$ is a normal-$\ast$-preserving quantized contact transformation, then
$\sigma_{\Lambda}(K)=\sqrt{\mathrm{vol}_{\mathrm{T}^{*}Y}},\qquad\mathrm{ord}(K)=(\dim Y)/2.$
Proof.
The correspondence between quantized contact transformations and pairs $(\mathcal{M},K)$ is established in [Kas03, Chapter 8.2].
Now assume $\Psi$ is a $\ast$-preserving quantized contact transformation. Consider $Q\in\widehat{\mathcal{E}}^{\sqrt{v}}_{Y}(m+1)$. By definition we have
$(\Psi(Q)-Q^{*})K=0.$
To show that $\sigma_{\Lambda}(K)=\sqrt{\mathrm{vol}_{\mathrm{T}^{*}Y}}$, using the notation of [Kas03, Chapter 8.6] we have to prove that
$(H_{\sigma_{m+1}(\Psi(Q))}+\sigma_{m}(\Psi(Q))-H_{\sigma_{m+1}(Q^{*})}-\sigma_{m}(Q^{*}))\sqrt{\mathrm{vol}_{\mathrm{T}^{*}Y}}=0.$
We have $\sigma_{m}(Q^{*})=(-1)^{m}\sigma_{m}(Q)=p_{2}^{a}\sigma_{m}(Q)$ and $\sigma_{m}(\Psi(Q))=\psi(\sigma_{m}(Q))$. Moreover, the function
$\psi(\sigma_{m}(Q))-(-1)^{m}\sigma_{m}(Q)$
vanishes along the Lagrangian $\tilde{\Lambda}$.
The vector field $H_{\sigma_{m+1}(\Psi(Q))}-H_{\sigma_{m+1}(Q^{*})}=H_{\psi(\sigma_{m+1}(Q)}+(-1)^{m}H_{\sigma_{m+1}(Q)}$ is tangent to the Lagrangian $\tilde{\Lambda}$. Identifying $\Lambda$ with $U_{X}$ using $p_{1}$ this vector field is identified with $H_{\psi(\sigma_{m+1}(Q))}$. Under the same identification the section $\sqrt{\mathrm{vol}_{\mathrm{T}^{*}Y}}$ of $K^{1/2}_{\tilde{\Lambda}}$ goes to the section $\sqrt{\mathrm{vol}_{\mathrm{T}^{*}X}}$ of $K^{1/2}_{U_{X}}$. By Liouville’s theorem $H_{\sigma_{m+1}(\Psi(Q))}\sqrt{\mathrm{vol}_{\mathrm{T}^{*}X}}=0$ which implies the result. ∎
3.6. Globalization
Recall that algebras form a 2-category as follows:
  • Its objects are algebras.
  • 1-morphisms from $A$ to $B$ are given by homomorphisms $f\colon A\rightarrow B$.
  • 2-morphisms from $f\colon A\rightarrow B$ to $g\colon A\rightarrow B$ are given by invertible elements $b\in B$ such that $g(a)=bf(a)b^{-1}$.
By an algebroid over a complex manifold we will mean a sheaf (i.e. a stack) of algebras viewed as objects of the above 2-category. For an algebroid it makes sense to consider the sheaf (i.e. a stack) of modules over it.
Let $Y$ be a contact manifold with a symplectization $\gamma\colon\tilde{Y}\rightarrow Y$. The papers [Kas96, PS04] have defined the algebroid $\mathsf{E}_{Y}$ over $Y$ uniquely determined by the following properties:
  • If $U\subset Y$ is an open subset together with a contact embedding $U\hookrightarrow\mathrm{P}^{*}X$ into a projectivized cotangent bundle, then
    $\mathsf{E}_{Y}|_{U}\cong\widehat{\mathcal{E}}^{\sqrt{v}}_{X}|_{U}.$
  • Given an intersection of two such charts, the isomorphism of the algebras of half-twisted microdifferential operators is given by a $\ast$-preserving quantized contact transformation.
Consider a Legendrian submanifold $\Lambda\subset Y$. As the notion of regularity of ^$\widehat{\mathcal{E}}^{\sqrt{v}}_{X}$-modules is local and invariant under quantized contact transformations, it extends to $\mathsf{E}_{Y}$-modules. Moreover, using 3.14 we obtain a global version of the microlocalization functor
$\mu^{dR}_{\Lambda}\colon\mathrm{Mod}_{\mathrm{rh}}(\mathsf{E}_{Y})\longrightarrow\gamma_{*}\mathrm{Mod}_{\mathrm{rh}}\left(\mathcal{D}^{\sqrt{v}}_{\tilde{\Lambda}}\right).$
Using this functor one obtains the following classification result (see [DS07, Corollary 6.4]).
Theorem 3.16.
Let $\Lambda\subset Y$ be a Legendrian submanifold. Microlocalization defines an equivalence of stacks
$\mu_{\Lambda}^{dR}\colon\mathrm{Mod}_{\Lambda,\mathrm{rh}}(\mathsf{E}_{Y})\xrightarrow{\sim}\gamma_{*}\mathrm{Mod}_{\mathrm{lisse}}\left(\mathcal{D}^{\sqrt{v}}_{\tilde{\Lambda}}\right)$
between the stack of $\mathsf{E}_{Y}$-modules regular along $\Lambda$ and the stack of lisse twisted $D$-modules on $\tilde{\Lambda}$.
The antiinvolution $\ast$ on ^$\widehat{\mathcal{E}}^{\sqrt{v}}_{X}$ globalizes to an isomorphism of algebroids $\mathsf{E}_{Y}\cong\mathsf{E}_{Y}^{\mathrm{op}}$ for any contact manifold $Y$. Thus, any left $\mathsf{E}_{Y}$-module $\mathsf{M}$ can be considered as a right $\mathsf{E}_{Y}$-module and vice versa.
4. DQ modules
In this section we define one of the main objects of the paper: the stack of deformation quantization modules on a holomorphic symplectic manifold.
4.1. WKB-differential operators
Let $X$ be a complex manifold. Consider the sheaf of algebras ^$\widehat{\mathcal{E}}_{X\times\mathbf{C}}$ on $\mathrm{J}^{1}X\subset\mathrm{P}^{*}(X\times\mathbf{C})$. The natural $\mathbf{C}$-action on $\mathrm{J}^{1}X$ lifts to a Hamiltonian $\mathbf{C}$-action on ^$\widehat{\mathcal{E}}_{X\times\mathbf{C}}$ with the quantum moment map given by $\partial_{t}$. We consider the sheaf
$\widehat{\mathcal{W}}_{X}\subset\rho_{*}(\widehat{\mathcal{E}}_{X\times\mathbf{C}})$
of $\mathbf{C}(\!(\hbar)\!)$-algebras of formal WKB-differential operators introduced in [PS04], where $\rho\colon\mathrm{J}^{1}X\rightarrow\mathrm{T}^{*}X$ and $\hbar=\partial_{t}^{-1}$.
Remark 4.1.
We may identify $\widehat{\mathcal{W}}_{X}$ with the quantum Hamiltonian reduction of ^$\widehat{\mathcal{E}}_{X\times\mathbf{C}}$ at the moment map value $\partial_{t}=\hbar^{-1}$ for $\hbar\in\mathbf{C}(\!(\hbar)\!)$ a formal variable.
Example 4.2.
There is an isomorphism of sheaves of $\mathbf{C}(\!(\hbar)\!)$-algebras
$\widehat{\mathcal{W}}_{X}|_{X}\cong\mathcal{D}_{X}(\!(\hbar)\!)$
on $X$.
Remark 4.3.
The filtration of ^$\widehat{\mathcal{E}}_{X\times\mathbf{C}}$ by order induces a filtration on $\widehat{\mathcal{W}}_{X}$. We have an isomorphism
$\widehat{\mathcal{W}}_{X}(0)/\hbar\cong\mathcal{O}_{\mathrm{T}^{*}X}$
of sheaves of algebras, so $\widehat{\mathcal{W}}_{X}(0)$ defines a deformation quantization of $\mathrm{T}^{*}X$.
Lemma 4.4.
The $\widehat{\mathcal{W}}_{X}$-module $\rho_{*}(\widehat{\mathcal{E}}_{X\times\mathbf{C}})$ is flat.
Proof.
To prove the claim it is enough to show that the $\widehat{\mathcal{W}}_{X}(0)$-module $\rho_{*}(\widehat{\mathcal{E}}_{X\times\mathbf{C}}(0))$ is flat as $\rho_{*}(\widehat{\mathcal{E}}_{X\times\mathbf{C}})=\rho_{*}(\widehat{\mathcal{E}}_{X\times\mathbf{C}}(0))[\hbar^{-1}]$.
We have that $\rho_{*}(\widehat{\mathcal{E}}_{X\times\mathbf{C}}(0))$ is $\hbar$-complete and has no $\hbar$-torsion. Therefore, by [KS12, Corollary 1.5.7] it is cohomologically $\hbar$-complete. By [KS12, Theorem 1.6.6] in this case it is enough to prove flatness modulo $\hbar$, i.e. it is enough to prove that $\rho_{*}\mathcal{O}_{\mathrm{J}^{1}X}$ is flat as a $\mathcal{O}_{\mathrm{T}^{*}X}$-module. Since the map $\mathrm{J}^{1}X\rightarrow\mathrm{T}^{*}X$ is open, it is flat by [BS76, Theorem V.2.13]. ∎
The half-twisted algebra
$\widehat{\mathcal{W}}^{\sqrt{v}}_{X}\subset\rho_{*}(\widehat{\mathcal{E}}^{\sqrt{v}}_{X\times\mathbf{C}})$
is defined similarly, where we note that $\hbar^{*}=-\hbar$. In this case the antiinvolution $\ast$ on ^$\widehat{\mathcal{E}}^{\sqrt{v}}_{X\times\mathbf{C}}$ restricts to one on $\widehat{\mathcal{W}}^{\sqrt{v}}_{X}$.
Let us now explain how to glue these sheaves of algebras. Suppose $U_{X}\subset\mathrm{T}^{*}X$ and $U_{Y}\subset\mathrm{T}^{*}Y$ are open subsets together with a symplectomorphism $\phi\colon U_{X}\xrightarrow{\sim}U_{Y}$. Suppose the symplectomorphism $\phi$ lifts to an isomorphism of contactifications $\psi\colon V_{X}=\rho^{-1}(U_{X})\xrightarrow{\sim}V_{Y}=\rho^{-1}(U_{Y})$, where $\rho\colon\mathrm{J}^{1}X\rightarrow\mathrm{T}^{*}X$ and $\rho\colon\mathrm{J}^{1}Y\rightarrow\mathrm{T}^{*}Y$ are the standard contactifications. Recall that in this case $\psi$ is compatible with the Hamiltonian $\mathbf{C}$-actions. The following is [PS04, Lemma 8.5].
Proposition 4.5.
Consider an isomorphism of contactifications $\psi\colon V_{X}=\rho^{-1}(U_{X})\xrightarrow{\sim}V_{Y}=\rho^{-1}(U_{Y})$ as above. Then there is a normal-$\ast$-preserving quantized contact transformation
$\Psi\colon\widehat{\mathcal{E}}^{\sqrt{v}}_{Y\times\mathbf{C}}|_{V_{Y}}\xrightarrow{\sim}\psi_{*}\widehat{\mathcal{E}}^{\sqrt{v}}_{X\times\mathbf{C}}|_{V_{X}}$
compatible with the Hamiltonian $\mathbf{C}$-actions, i.e. $\Psi(\partial_{t})=\partial_{t}$ and $\Psi$ is $\mathbf{C}$-equivariant.
Consider the setting of 4.5. Then we may restrict the $\ast$-preserving quantized contact transformation $\Psi$ to a normal-$\ast$-preserving quantized symplectic transformation
$\Phi\colon\widehat{\mathcal{W}}^{\sqrt{v}}_{Y}|_{U_{Y}}\xrightarrow{\sim}\phi_{*}\widehat{\mathcal{W}}^{\sqrt{v}}_{X}|_{U_{X}}.$
Let $\mathcal{M}$ be the ^$\widehat{\mathcal{E}}^{\sqrt{v}}_{X\times\mathbf{C}\times Y\times\mathbf{C}}$-module simple along a Legendrian $\Lambda$ with a generator $K$ corresponding to the quantized contact transformation $\Psi$ via 3.15. We have
$\sigma_{\Lambda}(K)=\sqrt{\tau^{\dim Y}\omega_{\mathrm{T}^{*}Y}\wedge d\tau\wedge dt}.$
Consider the Hamiltonian $\mathbf{C}$-action on ^$\widehat{\mathcal{E}}^{\sqrt{v}}_{X\times\mathbf{C}\times Y\times\mathbf{C}}$ given by $(t_{1},t_{2})\mapsto(t_{1}+\delta t,t_{2}+\delta t)$ with Hamiltonian $\partial_{t_{1}}+\partial_{t_{2}}$. We may identify the Hamiltonian reduction as
(^$(\widehat{\mathcal{E}}^{\sqrt{v}}_{X\times\mathbf{C}\times Y\times\mathbf{C}})/\!/\mathbf{C}\cong\widehat{\mathcal{E}}^{\sqrt{v}}_{X\times Y\times\mathbf{C}},$
where $t$ corresponds to $(t_{1}-t_{2})$ and $\partial_{t}$ corresponds to $\partial_{t_{1}}=-\partial_{t_{2}}$. By 4.5 the module $\mathcal{M}$ is strongly $\mathbf{C}$-equivariant, so after Hamiltonian reduction we obtain a ^$\widehat{\mathcal{E}}^{\sqrt{v}}_{X\times Y\times\mathbf{C}}$-module
=$\mathcal{M}^{\partial}=\{m\in\mathcal{M}\mid(\partial_{t_{1}}+\partial_{t_{2}})m=0\}.$
Proposition 4.6.
The ^$\widehat{\mathcal{E}}^{\sqrt{v}}_{X\times Y\times\mathbf{C}}$-module $\mathcal{M}^{\partial}$ is simple along the Legendrian $\Lambda^{\prime}=\Gamma_{\psi}\subset(V_{X}\times\overline{V_{Y}})/\mathbf{C}\hookrightarrow\mathrm{J}^{1}(X\times Y)$. Moreover,
$\sigma_{\Lambda^{\prime}}(K)=\sqrt{\tau^{\dim Y}\omega_{\mathrm{T}^{*}Y}\wedge d\tau},\qquad\mathrm{ord}(K)=(\dim Y+1)/2.$
Proof.
We may identify
$(\mathcal{M},K)\cong(\widehat{\mathcal{E}}^{\sqrt{v}}_{X\times\mathbf{C}}|_{V_{X}},1)$
using the projection $p_{1}\colon\mathrm{T}^{*}(X\times\mathbf{C}\times Y\times\mathbf{C})\rightarrow\mathrm{T}^{*}(X\times\mathbf{C})$. In a similar way, we may identify
$(\mathcal{M}^{\partial},K)\cong(\widehat{\mathcal{W}}^{\sqrt{v}}_{X}|_{U_{X}},1).$
To simplify the notation, we will omit the restrictions to open subsets $V_{X}$, $U_{X}$ and so on. Consider the ^$\widehat{\mathcal{E}}^{\sqrt{v}}_{X\times Y\times\mathbf{C}}(0)$-submodule $\mathcal{M}^{\partial}(n)\subset\mathcal{M}^{\partial}$ given by
$\mathcal{M}^{\partial}(n)=\widehat{\mathcal{E}}^{\sqrt{v}}_{X\times Y\times\mathbf{C}}(n)K.$
Note that we have $\widehat{\mathcal{W}}^{\sqrt{v}}_{X\times Y}(n)K=\widehat{\mathcal{W}}^{\sqrt{v}}_{X}(n)$. But $t\in\widehat{\mathcal{E}}^{\sqrt{v}}_{X\times Y\times\mathbf{C}}(0)$ acts on $\mathcal{M}^{\partial}\cong\widehat{\mathcal{W}}^{\sqrt{v}}_{X}$ by 2superscriptPlanck-constant-over-2-pi2subscriptPlanck-constant-over-$\hbar^{2}\partial_{\hbar}$, so $\mathcal{M}^{\partial}(n)=\widehat{\mathcal{W}}^{\sqrt{v}}_{X\times Y}(n)K=\widehat{\mathcal{W}}^{\sqrt{v}}_{X}(n)$.
Similarly, one defines the ^$\widehat{\mathcal{E}}^{\sqrt{v}}_{X\times\mathbf{C}\times Y\times\mathbf{C}}(0)$-submodules $\mathcal{M}(n)\subset\mathcal{M}$ which can be identified with ^$\widehat{\mathcal{E}}^{\sqrt{v}}_{X\times\mathbf{C}}(n)$. Note that such an identification implies an isomorphism $\mathcal{M}(n)^{\mathbf{C}}\cong\mathcal{M}^{\partial}(n)$. To show that $\mathcal{M}^{\partial}$ is simple along $\Lambda^{\prime}$, we have to show that the ^$\widehat{\mathcal{E}}^{\sqrt{v}}_{X\times Y\times\mathbf{C}}(0)/\widehat{\mathcal{E}}^{\sqrt{v}}_{X\times Y\times\mathbf{C}}(-1)$-module structure on $\mathcal{M}^{\partial}(0)/\mathcal{M}^{\partial}(-1)$ extends to a $\operatorname{gr}^{0}_{\Lambda^{\prime}}\widehat{\mathcal{E}}^{\sqrt{v}}_{X\times Y\times\mathbf{C}}$-module structure and the corresponding $\mathcal{O}_{\Lambda^{\prime}}$-module is locally free of rank 1.
Since the homogeneous symplectic transformation $\tilde{\psi}$ preserves $\tau$, we have $(\tau_{1}+\tau_{2})|_{\tilde{\Lambda}^{\prime}}=0$. In particular, $\partial_{t_{1}}+\partial_{t_{2}}\in\widehat{\mathcal{E}}^{\sqrt{v}}_{\Lambda/X\times\mathbf{C}\times Y\times\mathbf{C}}$. So, the Hamiltonian $\mathbf{C}$-action on ^$\widehat{\mathcal{E}}^{\sqrt{v}}_{X\times\mathbf{C}\times Y\times\mathbf{C}}$ restricts to one on ^$\widehat{\mathcal{E}}^{\sqrt{v}}_{\Lambda/X\times\mathbf{C}\times Y\times\mathbf{C}}$. Passing to the associated graded, we get the natural Hamiltonian $\mathbf{C}$-action on $\mathcal{D}^{\sqrt{v}}_{[\tilde{\Lambda}]}$ induced by the $\mathbf{C}$-action on $\Lambda$. As $\mathcal{D}^{\sqrt{v}}_{[\tilde{\Lambda}]}/\!/\mathbf{C}\cong\mathcal{D}^{\sqrt{v}}_{[\tilde{\Lambda^{\prime}}]}$, we get an isomorphism
^$\widehat{\mathcal{E}}^{\sqrt{v}}_{\Lambda/X\times\mathbf{C}\times Y\times\mathbf{C}}/\!/\mathbf{C}\cong\widehat{\mathcal{E}}^{\sqrt{v}}_{\Lambda^{\prime}/X\times Y\times\mathbf{C}},$
so that the Hamiltonian reduction of the homomorphism
$L\colon\widehat{\mathcal{E}}^{\sqrt{v}}_{\Lambda/X\times\mathbf{C}\times Y\times\mathbf{C}}\longrightarrow\mathcal{D}^{\sqrt{v}}_{[\tilde{\Lambda}]}$
is the homomorphism
$L\colon\widehat{\mathcal{E}}^{\sqrt{v}}_{\Lambda^{\prime}/X\times Y\times\mathbf{C}}\longrightarrow\mathcal{D}^{\sqrt{v}}_{[\tilde{\Lambda^{\prime}}]}.$
We have a pair of strongly $\mathbf{C}$-equivariant $\mathcal{D}^{\sqrt{v}}_{\tilde{\Lambda}}(0)$-modules $\mathcal{M}(0)/\mathcal{M}(-1)$ and $K^{1/2}_{\tilde{\Lambda}}$ whose Hamiltonian reduction gives $\mathcal{D}^{\sqrt{v}}_{\tilde{\Lambda^{\prime}}}(0)$-modules $(\mathcal{M}(0)/\mathcal{M}(-1))^{\mathbf{C}}$ and $(K^{1/2}_{\tilde{\Lambda}})^{\mathbf{C}}\cong K^{1/2}_{\tilde{\Lambda^{\prime}}}$.
We may identify
$\mathcal{M}(0)/\mathcal{M}(-1)\cong\widehat{\mathcal{E}}^{\sqrt{v}}_{X\times\mathbf{C}}(0)/\widehat{\mathcal{E}}^{\sqrt{v}}_{X\times\mathbf{C}}(-1)\cong\mathcal{O}_{\mathrm{J}^{1}X}$
and
$\mathcal{M}^{\partial}(0)/\mathcal{M}^{\partial}(-1)\cong\widehat{\mathcal{W}}^{\sqrt{v}}_{X}(0)/\widehat{\mathcal{W}}^{\sqrt{v}}_{X}(-1)\cong\mathcal{O}_{\mathrm{T}^{*}X}.$
In particular, the natural map $\mathcal{M}^{\partial}(0)/\mathcal{M}^{\partial}(-1)\rightarrow(\mathcal{M}(0)/\mathcal{M}(-1))^{\mathbf{C}}$ is an isomorphism. This implies that $\mathcal{M}$ is simple along the $\Lambda^{\prime}$ and
$\sigma_{\Lambda^{\prime}}(K)=\sigma_{\Lambda}(K)/\sqrt{dt}$
which, in conjunction with the computation of the symbol $\sigma_{\Lambda}(K)$ in 3.15, implies the result. ∎
4.2. Globalization
Let $S$ be a symplectic manifold. The authors of [PS04] have constructed a $\mathbf{C}(\!(\hbar)\!)$-algebroid $\mathsf{W}_{S}$ on $S$ with the following properties:
  • If $U\subset S$ is an open subset together with a symplectic embedding $U\hookrightarrow\mathrm{T}^{*}X$ into a cotangent bundle, then
    $\mathsf{W}_{S}|_{U}\cong\widehat{\mathcal{W}}^{\sqrt{v}}_{X}|_{U}.$
  • Given an intersection of two such charts, the isomorphism of the algebras of half-twisted WKB-differential operators is given by a $\ast$-preserving quantized symplectic transformation.
  • If $Y\rightarrow S$ is a contactification, then there is a natural embedding
    $\mathsf{W}_{S}\subset\rho_{*}\mathsf{E}_{Y}.$
The antiinvolution $\ast$ on $\widehat{\mathcal{W}}^{\sqrt{v}}_{X}$ globalizes to a $\mathbf{C}$-linear isomorphism of algebroids $\mathsf{W}_{S}\cong\mathsf{W}^{\mathrm{op}}_{S}$ which sends $\hbar\mapsto-\hbar$. As for $\mathsf{E}_{Y}$-modules, we freely pass between left and right $\mathsf{W}_{S}$-modules using this antiinvolution.
Definition 4.7.
A $\mathsf{W}_{S}$-module is holonomic if its support is a Lagrangian subvariety of $S$.
For a Lagrangian subvariety $L\subset S$ we denote by $\mathsf{W}_{L/S}\subset\mathsf{W}_{S}|_{L}$ generated by
$I_{L}=\{P\in\mathsf{W}_{S}(1)|_{L}\mid\sigma_{1}(P)|_{L}=0\}.$
Definition 4.8.
A holonomic $\mathsf{W}_{S}$-module $\mathcal{M}$ is regular holonomic if locally there is a coherent $\mathsf{W}_{S}(0)$-module $\mathcal{M}_{0}\subset\mathcal{M}$ and a Lagrangian subvariety $L\subset S$ containing $\operatorname{supp}\mathcal{M}$ such that $I_{L}\mathcal{M}_{0}\subset\mathcal{M}_{0}$ and $\mathsf{W}_{S}\mathcal{M}_{0}=\mathcal{M}$.
Example 4.9.
Let $L\subset S$ be a Lagrangian subvariety and $\Lambda\subset Y$ be its contactification with $\rho\colon Y\rightarrow S$ the projection map. If $\mathcal{M}$ is a $\mathsf{E}_{Y}$-module supported on $\Lambda$, then $\mathcal{M}$, regarded as a $\mathsf{W}_{S}$-module via the embedding $\mathsf{W}_{S}\subset\rho_{*}\mathsf{E}_{Y}$, is supported on $L$ and is, therefore, holonomic. This gives a forgetful functor
$\rho_{*}\mathrm{Mod}_{hol}(\mathsf{E}_{Y})\longrightarrow\mathrm{Mod}_{hol}(\mathsf{W}_{S})$
from holonomic $\mathsf{E}_{Y}$-modules to holonomic $\mathsf{W}_{S}$-modules. Since $\mathsf{W}_{L/S}\subset\rho_{*}\mathsf{E}_{\Lambda/Y}$, this forgetful functor restricts to a functor
$\rho_{*}\mathrm{Mod}_{\mathrm{rh}}(\mathsf{E}_{Y})\longrightarrow\mathrm{Mod}_{\mathrm{rh}}(\mathsf{W}_{S}).$
The authors of [DK11] have introduced the stack $\mathrm{RH}_{S}$ of regular holonomic modules on $S$ as follows:
  • Let $L\subset S$ be a Lagrangian subvariety and $\Lambda\subset Y$ its contactification with projection $\rho\colon Y\rightarrow S$. We define
    $\mathrm{RH}_{S,L}=\rho_{*}\mathrm{Mod}_{\Lambda,\mathrm{rh}}(\mathsf{E}_{Y})=\mathrm{Mod}_{\Lambda,\mathrm{rh}}(\rho_{*}\mathsf{E}_{Y}),$
    the stack of regular holonomic $\mathsf{E}_{Y}$-modules supported on $\Lambda$.
  • As the contactification of $L\subset S$ exists locally and is unique by 1.12, the stack $\mathrm{RH}_{S,L}$ only depends on $L\subset S$.
  • Lagrangian subvarieties of $S$ form a filtered system and we define
    $\mathrm{RH}_{S}=\operatorname*{colim}_{L}\mathrm{RH}_{S,L}.$
Example 4.10.
Let $L\subset S$ be a Lagrangian submanifold. The stack $\mathrm{RH}_{S,L}$ consists of $\mathsf{E}_{Y}$-modules regular along a contactification $\Lambda\subset Y$ of $L\subset S$. By 3.16 microlocalization establishes an equivalence
$\mathrm{RH}_{S,L}\cong\gamma_{*}\mathrm{Mod}_{\mathrm{lisse}}(\mathcal{D}^{\sqrt{v}}_{L\times\mathbf{C}^{\times}}),$
where $\gamma\colon L\times\mathbf{C}^{\times}\rightarrow L$ is the projection on the first factor. Half-twisted $D$-modules along $\mathbf{C}^{\times}$ whose underlying $\mathcal{O}$-module is a line bundle are classified by their order $\lambda\in\mathbf{C}/\mathbf{Z}$. In particular, if $L$ is equipped with an orientation data and given $\lambda\in\mathbf{C}/\mathbf{Z}$, there is a canonical module
$\mathsf{M}^{\lambda}_{L}\in\mathrm{RH}_{S,L}\subset\mathrm{RH}_{S}$
simple along $L$ with order $\lambda$.
4.3. Tensor product of DQ modules
Consider two objects $\mathcal{L},\mathcal{M}\in\mathrm{RH}_{S}$ which are regular holonomic $\rho_{*}\mathsf{E}_{Y}$-modules. In this section we will study the complex of $\mathbf{C}(\!(\hbar)\!)$-linear sheaves $\mathcal{L}\otimes^{\mathbb{L}}_{\mathsf{W}_{S}}\mathcal{M}$ on $X$. Our first goal is to endow it with extra structure.
Choose a contactification $\rho\colon Y\rightarrow S$ and assume that $\mathsf{E}_{Y}$ and $\mathsf{W}_{S}$ are represented by sheaves of algebras, there is an element $\hbar\in\mathsf{E}_{Y}(-1)$ with principal symbol $\tau^{-1}$ satisfying $\hbar^{*}=-\hbar$ and $\mathsf{W}_{S}$ is the subalgebra of $\rho_{*}\mathsf{E}_{Y}$ of elements commuting with $\hbar$.
Consider
$\mathsf{B}_{Y}=(\rho_{*}\mathsf{E}_{Y})\otimes_{\mathsf{W}_{S}}(\rho_{*}\mathsf{E}_{Y}),$
where the right action of $\mathsf{W}_{S}$ on the first copy of $\rho_{*}\mathsf{E}_{Y}$ is via $P\in\mathsf{W}_{S},Q\in\rho_{*}\mathsf{E}_{Y}\mapsto P^{*}Q$. It carries a natural bimodule structure as follows: $\rho_{*}\mathsf{E}_{Y}\otimes\rho_{*}\mathsf{E}_{Y}$ acts in an obvious way on the right and $\hbar$ acts on $Q_{1}\otimes Q_{2}\in\mathsf{B}_{Y}$ as $Q_{1}\otimes\hbar Q_{2}=-\hbar Q_{1}\otimes Q_{2}$.
Proposition 4.11.
$\mathsf{B}_{Y}$ has a canonical structure of a $(\mathbf{C}(\!(\hbar)\!)\langle D\rangle,\rho_{*}\mathsf{E}_{Y}\otimes\rho_{*}\mathsf{E}_{Y})$-bimodule, where $D\hbar=\hbar(D+1)$.
Proof.
Let $\partial_{t}=\hbar^{-1}\in\mathsf{E}_{Y}(1)$. Locally we may choose an element $t\in\mathsf{E}_{Y}(0)$ satisfying
$[\partial_{t},t]=1.$
If we let $D=\partial_{t}t$, we have
$[D,\hbar]=\hbar,\qquad[D^{*},\hbar]=-\hbar.$
This allows us to enhance $\mathsf{B}_{Y}$ to a $(\mathbf{C}(\!(\hbar)\!)\langle D\rangle,\rho_{*}\mathsf{E}_{Y}\otimes\rho_{*}\mathsf{E}_{Y})$-bimodule, where $D=\partial_{t}t$ acts on $Q_{1}\otimes Q_{2}\in\mathsf{B}_{Y}$ as $-D^{*}Q_{1}\otimes Q_{2}+Q_{1}\otimes DQ_{2}$.
Let us show that this enhancement is independent of the choice of $t$. If $t^{\prime}$ is another element of $\mathsf{E}_{Y}(0)$ satisfying $[\partial_{t},t^{\prime}]=1$, then $[\partial_{t},t-t^{\prime}]=0$, i.e. $[\hbar,t-t^{\prime}]=0$. Thus, there is an element $R\in\mathsf{W}_{S}(0)$ such that
$t^{\prime}=t+R.$
Then
$\partial_{t}t^{\prime}=\partial_{t}t+\partial_{t}R.$
We then have
$-(\partial_{t}t^{\prime})^{*}Q_{1}\otimes Q_{2}+Q_{1}\otimes\partial_{t}t^{\prime}Q_{2}=-(\partial_{t}t)^{*}Q_{1}\otimes Q_{2}+Q_{1}\otimes\partial_{t}tQ_{2}-(\partial_{t}R)^{*}Q_{1}\otimes Q_{2}+Q_{1}\otimes\partial_{t}RQ_{2}.$
But $\partial_{t}R\in\mathsf{W}_{S}$ and hence the last two terms cancel. ∎
Using the previous proposition we obtain the following result.
Theorem 4.12.
For any $\mathcal{L},\mathcal{M}\in\mathrm{RH}_{S}$ the sheaf $\mathcal{L}\otimes^{\mathbb{L}}_{\mathsf{W}_{S}}\mathcal{M}$ has a canonical structure of a differential perverse sheaf.
Proof.
Consider the duality functor $\mathbb{D}(\mathcal{L})=\mathbb{R}\mathcal{H}om_{\mathsf{W}_{S}}(\mathcal{L},\mathsf{W}_{S})[\dim S/2]$. By [KS12, Lemma 7.2.2] $\mathbb{D}(\mathcal{L})$ is a holonomic $\mathsf{W}_{S}$-module concentrated in degree 0 and we have
$\mathbb{R}\mathcal{H}om_{\mathsf{W}_{S}}(\mathbb{D}(\mathcal{L}),\mathcal{M})[\dim S/2]\cong\mathcal{L}\otimes^{\mathbb{L}}_{\mathsf{W}_{S}}\mathcal{M}.$
It is then shown in [KS12, Theorem 7.2.3] that $\mathbb{R}\mathcal{H}om_{\mathsf{W}_{S}}(\mathbb{D}(\mathcal{L}),\mathcal{M})[\dim S/2]$ is a perverse sheaf of $\mathbf{C}(\!(\hbar)\!)$-vector spaces.
$\rho_{*}\mathsf{E}_{Y}$ is flat as a $\mathsf{W}_{S}$-module by Lemma 4.4. Therefore,
$\mathcal{L}\otimes^{\mathbb{L}}_{\mathsf{W}_{S}}\mathcal{M}\cong\mathsf{B}_{Y}\otimes^{\mathbb{L}}_{\rho_{*}\mathsf{E}_{Y}\otimes\rho_{*}\mathsf{E}_{Y}}(\mathcal{L}\otimes\mathcal{M})$
which has a differential structure by 4.11. ∎
Our next goal is to show that the differential structure has regular singularities.
Theorem 4.13.
Consider two regular holonomic modules $\mathcal{L},\mathcal{M}\in\mathrm{RH}_{S}$ and assume $\mathcal{L}$ is simple along a Lagrangian submanifold $L\subset S$. Then
$\mathcal{L}\otimes^{\mathbb{L}}_{\mathsf{W}_{S}}\mathcal{M}$
is a differential perverse sheaf with regular singularities.
Proof.
Let $\rho\colon Y\rightarrow S$ be a contactification defined in a neighborhood of the union of supports of $\mathcal{L},\mathcal{M}$ and $\Lambda\subset Y$ a Legendrian lift of $L\subset S$.
To define the differential structure we may choose $t\in\mathsf{E}_{Y}(0)$ whose principal symbol $\sigma_{0}(t)\in\mathcal{O}_{Y}$ vanishes along $\Lambda$. With this choice $D=\partial_{t}t=\hbar\partial_{\hbar}\in\mathsf{E}_{\Lambda/Y}$. By assumption $\mathcal{L}$ is an $\mathsf{E}_{Y}$-module simple along $\Lambda$. Thus, there is an $\mathsf{E}_{\Lambda/Y}$-lattice $\mathcal{L}_{0}\subset\mathcal{L}$ generating $\mathcal{L}$ over $\mathsf{E}_{Y}$ and $D$ acts on
¯$\overline{\mathcal{L}}=\mathcal{L}_{0}/\mathsf{E}_{\Lambda/Y}(-1)\mathcal{L}_{0}$
via multiplication by $\lambda\in\mathbf{C}$.
Let $\mathcal{M}_{0}\subset\mathcal{M}|_{\Lambda}$ be a $\mathsf{E}_{\Lambda/Y}$-lattice such that no two eigenvalues of $D$ on
¯$\overline{\mathcal{M}}=\mathcal{M}_{0}/\mathsf{E}_{\Lambda/Y}(-1)\mathcal{M}_{0}$
differ by an integer.
By [KS12, Lemma 7.1.3]
$\mathcal{L}_{0}\otimes^{\mathbb{L}}_{\mathsf{W}_{L/S}}\mathcal{M}_{0}$
is a $\mathbf{C}[\![\hbar]\!]$-lattice in $\mathcal{L}\otimes^{\mathbb{L}}_{\mathsf{W}_{S}}\mathcal{M}$. Since $D\in\mathsf{E}_{\Lambda/Y}$, this lattice is preserved by $D$. Moreover, no two eigenvalues of $D$ on the $\mathbf{C}$-linear perverse sheaf
$(\mathcal{L}_{0}\otimes^{\mathbb{L}}_{\mathsf{W}_{L/S}}\mathcal{M}_{0})/\hbar\cong\overline{\mathcal{L}}_{0}\otimes_{\mathsf{W}_{L/S}/\hbar}\overline{\mathcal{M}}_{0}$
differ by an integer. ∎
We will now give a description of the above differential perverse sheaf using microlocalization. Consider the setting of 4.13 and suppose $L\subset S$ is lifted to a Legendrian $\Lambda\subset Y$ in a contactification $\rho\colon Y\rightarrow S$. In this case we may identify $\tilde{\Lambda}\cong L\times\mathbf{C}^{\times}$. Then $\mu^{dR}_{\Lambda}\mathcal{L},\mu^{dR}_{\Lambda}\mathcal{M}$ are twisted $D$-modules on $\tilde{\Lambda}$, monodromic with respect to the $\mathbf{C}^{\times}$-action. Therefore,
$\mu^{dR}_{\Lambda}\mathcal{L}\otimes^{\mathbb{L}}_{\mathcal{D}^{\sqrt{v}}_{\tilde{\Lambda}}}\mu^{dR}_{\Lambda}\mathcal{M}$
is a perverse sheaf on $\tilde{\Lambda}=L\times\mathbf{C}^{\times}$, lisse in the $\mathbf{C}^{\times}$-direction. Its restriction to $L\times\{1\}$ defines a perverse sheaf
$(\mu^{dR}_{\Lambda}\mathcal{L}\otimes^{\mathbb{L}}_{\mathcal{D}^{\sqrt{v}}_{\tilde{\Lambda}}}\mu^{dR}_{\Lambda}\mathcal{M})|_{L}[-1]$
equipped with a monodromy operator $T$.
Theorem 4.14.
Consider the setting of 4.13. There is an isomorphism of differential perverse sheaves
$\mathcal{L}\otimes^{\mathbb{L}}_{\mathsf{W}_{S}}\mathcal{M}\cong\mathrm{RH}^{-1}\left((\mu^{dR}_{\Lambda}\mathcal{L}\otimes^{\mathbb{L}}_{\mathcal{D}^{\sqrt{v}}_{\tilde{\Lambda}}}\mu^{dR}_{\Lambda}\mathcal{M})|_{L}[-1]\right)$
on $L$.
Proof.
As in the proof of 4.13 we choose $\mathsf{E}_{\Lambda/Y}$-submodules $\mathcal{L}_{0}\subset\mathcal{L}$ and $\mathcal{M}_{0}\subset\mathcal{M}|_{\Lambda}$ generating these modules over $\mathsf{E}_{Y}|_{\Lambda}$ as well as an element $D=\partial_{t}t\in\mathsf{E}_{\Lambda/Y}$.
We may identify
$\operatorname{gr}^{0}_{\Lambda}\mathsf{E}_{\Lambda/Y}\cong\mathcal{D}^{\sqrt{v}}_{L}[\theta],$
where $\theta$ is the Euler operator along $\Lambda$ which is the image of $D$ under the projection
$\mathsf{E}_{\Lambda/Y}\longrightarrow\operatorname{gr}^{0}_{\Lambda}\mathsf{E}_{\Lambda/Y}.$
Consider
¯$\overline{\mathcal{L}}_{0}=\mathcal{L}_{0}/\hbar,\qquad\overline{\mathcal{M}}_{0}=\mathcal{M}_{0}/\hbar.$
As $\hbar\mathsf{E}_{\Lambda/Y}(0)=\mathsf{E}_{\Lambda/Y}(-1)$, these are exactly the associated graded sheaves with respect to the microlocal $V$-filtration
¯$\overline{\mathcal{L}}_{0}\cong\operatorname{gr}^{0}_{\Lambda}\mathcal{L},\qquad\overline{\mathcal{M}}_{0}\cong\operatorname{gr}^{0}_{\Lambda}\mathcal{M}.$
The complex of sheaves
$\mathcal{L}_{0}\otimes^{\mathbb{L}}_{\mathsf{W}_{L/S}}\mathcal{M}_{0}$
defines a $\mathbf{C}[\![\hbar]\!]$-lattice in $\mathcal{L}\otimes_{\mathsf{W}_{S}}\mathcal{M}$ preserved by $D$. Therefore, by 2.11 the claim will follow once we establish an isomorphism
$\operatorname{gr}^{0}_{\Lambda}\mathcal{L}\otimes^{\mathbb{L}}_{\mathcal{D}^{\sqrt{v}}_{L}}\operatorname{gr}^{0}_{\Lambda}\mathcal{M}\cong(\mu^{dR}_{\Lambda}\mathcal{L}\otimes^{\mathbb{L}}_{\mathcal{D}^{\sqrt{v}}_{L\times\mathbf{C}^{\times}}}\mu^{dR}_{\Lambda}\mathcal{M})|_{L}[-1]$
of perverse sheaves on $L$ equipped with monodromy automorphisms.
We have
$\mu^{dR}_{\Lambda}\mathcal{L}=\mathcal{D}^{\sqrt{v}}_{L\times\mathbf{C}^{\times}}\otimes_{\gamma^{-1}\mathcal{D}^{\sqrt{v}}_{L}[\theta]}\gamma^{-1}\operatorname{gr}^{0}_{\Lambda}\mathcal{L}$
and similarly for $\mathcal{M}$ and so the claim follows from 2.4. ∎
The local model of a contactification $\rho\colon Y\rightarrow S$ is given by $Y=\mathrm{J}^{1}X\rightarrow S=\mathrm{T}^{*}X$ in which case we have
$\mathsf{B}_{Y}=(\rho_{*}\widehat{\mathcal{E}}^{\sqrt{v}}_{X\times\mathbf{C}})\otimes_{\widehat{\mathcal{W}}^{\sqrt{v}}_{X}}(\rho_{*}\widehat{\mathcal{E}}^{\sqrt{v}}_{X\times\mathbf{C}}).$
The involution $i\colon X\times\mathbf{C}\rightarrow X\times\mathbf{C}$ given by $(x,t)\mapsto(x,-t)$ induces an involution
(3) ():^$(-)^{\prime}\colon\widehat{\mathcal{E}}^{\sqrt{v}}_{X\times\mathbf{C}}\xrightarrow{\sim}i_{*}\widehat{\mathcal{E}}^{\sqrt{v}}_{X\times\mathbf{C}}$
which sends $\partial_{t}\mapsto-\partial_{t}$ and which intertwines the antiinvolution $\ast$. Therefore, we get an involution
$(-)^{\prime}\colon\mathsf{B}_{Y}\longrightarrow\mathsf{B}_{Y}$
given by $Q_{1}\otimes Q_{2}\mapsto Q_{2}^{\prime}\otimes Q_{1}^{\prime}$ of $(\mathbf{C}(\!(\hbar)\!)\langle D\rangle,\rho_{*}\mathsf{E}_{Y}\otimes\rho_{*}\mathsf{E}_{Y})$-bimodules.
For an object $\mathcal{M}\in\mathrm{RH}_{\mathrm{T}^{*}X}$ denote by $\mathcal{M}^{\prime}$ the same module with the ^$\widehat{\mathcal{E}}^{\sqrt{v}}_{X\times\mathbf{C}}$-action twisted by the involution $(-)^{\prime}$. Then the above involution of $\mathsf{B}_{Y}$ induces an isomorphism
(4) $\mathcal{M}_{1}\otimes^{\mathbb{L}}_{\widehat{\mathcal{W}}^{\sqrt{v}}_{X}}\mathcal{M}_{2}\cong\mathcal{M}_{2}^{\prime}\otimes^{\mathbb{L}}_{\widehat{\mathcal{W}}^{\sqrt{v}}_{X}}\mathcal{M}_{1}^{\prime}$
of differential perverse sheaves.
5. Vanishing cycles and DQ modules
In this section we investigate a relationship between the following four objects: vanishing cycles, $\mathcal{E}$-modules, twisted de Rham complex and $\mathcal{W}$-modules.
5.1. Vanishing cycles
Let $X$ be a complex manifold equipped with a holomorphic function $f\colon X\rightarrow\mathbf{C}$. Recall the perverse sheaf $\phi_{f}\in\mathrm{Perv}(f^{-1}(0);\mathbf{C})$ of vanishing cycles of the constant perverse sheaf $\mathbf{C}_{X}[\dim X]\in\mathrm{Perv}(X)$ [Dim04], which carries a monodromy automorphism $T$. In this section we give a microlocal description of this sheaf.
Consider $X\times\mathbf{C}$ with $t$ the coordinate on $\mathbf{C}$ and let
$\Gamma_{f}=\{(x,t)\in X\times\mathbf{C}\mid f(x)-t=0\}\subset X\times\mathbf{C}$
be the graph of $f$.
Consider the $\mathcal{D}_{X\times\mathbf{C}}$-module
$\mathcal{B}_{f}=i^{\mathcal{D}}_{*}\mathcal{O}_{X},$
where $i\colon X\hookrightarrow X\times\mathbf{C}$ is the embedding of the graph $\Gamma_{f}$ given by $x\mapsto(x,f(x))$ and $i^{\mathcal{D}}_{*}$ is the $D$-module direct image. Its microlocalization ^$\widehat{\mathcal{E}}_{X\times\mathbf{C}}\otimes_{\pi^{-1}\mathcal{D}_{X\times\mathbf{C}}}\pi^{-1}\mathcal{B}_{f}$ is supported on the open subset $\mathrm{J}^{1}X\subset\mathrm{P}^{*}(X\times\mathbf{C})$ and we denote by
$\mathcal{C}_{f}=\rho_{*}\left(\widehat{\mathcal{E}}_{X\times\mathbf{C}}\otimes_{\pi^{-1}\mathcal{D}_{X\times\mathbf{C}}}\pi^{-1}\mathcal{B}_{f}|_{\mathrm{J}^{1}X}\right)$
the corresponding $\rho_{*}\widehat{\mathcal{E}}_{X\times\mathbf{C}}$-module.
We may identify $\mathcal{B}_{f}=\mathcal{O}_{X}[\partial_{t}]$ as an $\mathcal{O}_{X}[\partial_{t}]$-module with generator $\delta(t-f)$. The action of vector fields $\xi$ on $X$ and $t$ are given by
$\xi(g\partial_{t}^{i}\delta(t-f))$ $=(\xi g)\partial_{t}^{i}\delta(t-f)-(\xi f)g\partial_{t}^{i+1}\delta(t-f)$
$t(g\partial_{t}^{i}\delta(t-f))$ $=fg\partial_{t}^{i}\delta(t-f)-ig\partial_{t}^{i-1}\delta(t-f).$
Example 5.1.
Consider the Legendrian $\Lambda=j^{1}f\subset\mathrm{J}^{1}X$. Then $\mathcal{C}_{f}$ is a simple $\rho_{*}\widehat{\mathcal{E}}_{X\times\mathbf{C}}$-module along $\Lambda$. As the Euler operator we may take
$\theta^{\prime}=-(t-f)\partial_{t}-1/2\in\widehat{\mathcal{E}}_{\Lambda/\tilde{X}}.$
Then $(\theta^{\prime}-1/2)\delta(t-f)=0$, so $\delta(t-f)\in\mathcal{C}_{f}$ is a simple generator of order $1/2$.
We may identify
$\mathcal{C}_{f}=\widehat{\mathcal{E}}_{X\times\mathbf{C}}/\widehat{\mathcal{E}}_{X\times\mathbf{C}}\left(t-f,\frac{\partial f}{\partial x_{i}}\partial_{t}+\partial_{x_{i}}\right).$
Let
$\mathcal{C}^{\lambda}_{f}=\widehat{\mathcal{E}}_{X\times\mathbf{C}}/\widehat{\mathcal{E}}_{X\times\mathbf{C}}\left(t-f+(\lambda-1/2)\partial^{-1}_{t},\frac{\partial f}{\partial x_{i}}\partial_{t}+\partial_{x_{i}}\right).$
We denote its generator by $\partial_{t}^{\lambda-1/2}\delta(t-f)$ and the computation similar to Example 5.1 shows that
$\mathrm{ord}(\partial_{t}^{\lambda-1/2}\delta(t-f))=\lambda.$
We have an obvious isomorphism $(\mathcal{C}^{\lambda}_{f})^{\prime}\cong\mathcal{C}^{\lambda}_{-f}$ which preserves the generators, where $(-)^{\prime}$ is the involution from (3).
Let us now introduce versions of these modules. We consider $\mathbf{C}$ with a canonical coordinate $t$. It allows us to trivialize $K_{\mathbf{C}}$ using the section $dt$ and hence we obtain a square root line bundle $K_{\mathbf{C}}^{1/2}$.
  • Let $X$ be a complex manifold. Let
    $\mathcal{C}^{\lambda,\mathrm{R}}_{f}=\pi^{-1}K_{X\times\mathbf{C}}\otimes_{\pi^{-1}\mathcal{O}_{X\times\mathbf{C}}}\mathcal{C}^{\lambda}_{f}$
    be the corresponding right $\rho_{*}\widehat{\mathcal{E}}_{X\times\mathbf{C}}$-module.
  • Let $X$ be a complex manifold together with a choice of the square root $K^{1/2}_{X}$. Let
    $\mathcal{C}^{\lambda,\sqrt{X}}_{f}=\pi^{-1}K^{1/2}_{X\times\mathbf{C}}\otimes_{\pi^{-1}\mathcal{O}_{X\times\mathbf{C}}}\mathcal{C}^{\lambda}_{f}$
    be the corresponding $\rho_{*}\widehat{\mathcal{E}}^{\sqrt{v}}_{X\times\mathbf{C}}$-module.
Remark 5.2.
We use the notation $(-)^{\sqrt{v}}$ for $\mathcal{E},\mathcal{W},\mathcal{D}$ and $(-)^{\sqrt{X}}$ for the modules to indicate that the half-twisted version of the algebras does not depend on the choice of a square root line bundle, while the half-twisted version of the modules does depend on this choice.
Example 5.3.
Choose a nonvanishing section $\sqrt{\omega}$ of $K_{X}^{1/2}$ and consider the simple generator
$\sqrt{\omega}\partial_{t}^{\lambda-1/2}\delta(t-f)\in\mathcal{C}^{\lambda,\sqrt{X}}_{f}.$
Then
$\sigma_{\Lambda}(\sqrt{\omega}\partial_{t}^{\lambda-1/2}\delta(t-f))=\tau^{\lambda-1/2}\sqrt{d\tau\wedge\omega},$
see [Kas03, Lemma 8.23]. Its order is
$\mathrm{ord}(\sqrt{\omega}\partial_{t}^{\lambda-1/2}\delta(t-f))=\lambda.$
The standard projection $\rho\colon\mathrm{J}^{1}X\rightarrow\mathrm{T}^{*}X$ provides a contactification, so we may regard $\mathcal{C}^{\lambda,\sqrt{X}}_{f}$ as an object of $\mathrm{RH}_{\mathrm{T}^{*}X}$. This is a simple module along the Lagrangian $\Gamma_{df}\subset\mathrm{T}^{*}X$ of order $\lambda$.
Consider the $V$-filtration along the hypersurface $\Gamma_{0}:t=0$ in $X\times\mathbf{C}$ and let
$\phi_{f}^{\mathcal{D}}=\operatorname{gr}^{0}_{V}\mathcal{B}_{f}\in\mathrm{Mod}_{\mathrm{rh}}(\mathcal{D}_{X}[\theta])$
be the associated graded $D$-module which carries an action of the Euler vector field $\theta=t\partial_{t}$.
Example 5.4.
$\phi_{0}^{\mathcal{D}}$ is isomorphic to $\mathcal{O}_{X}$ equipped with the endomorphism $\theta=0$.
The following statement is due to Kashiwara and Malgrange, see [Mal83, Théorème 6.2.6] and [Kas83, Theorem 2].
Proposition 5.5.
There is an isomorphism of perverse sheaves
$\mathbb{R}\mathcal{H}om_{\mathcal{D}_{X}}(\mathcal{O}_{X},\phi_{f}^{\mathcal{D}})[\dim X]\cong\phi_{f}.$
Under this isomorphism the monodromy operator on $\phi_{f}$ is given by $T=\exp(-2\pi i\partial_{t}t)$.
We will also work with the sheaves of vanishing cycles microlocally. For this consider the microlocal $V$-filtration along the Legendrian $\Lambda=\mathrm{P}^{*}_{\Gamma_{0}}(X\times\mathbf{C})\subset\mathrm{P}^{*}(X\times\mathbf{C})$, where we may identify $\Lambda\cong X$. Using the trivialization of $K_{\Gamma_{0}|X\times\mathbf{C}}$ by $dt$ the associated graded $\operatorname{gr}^{0}_{\Lambda}\mathcal{C}_{f}$ becomes a $\operatorname{gr}^{0}_{\Lambda}\widehat{\mathcal{E}}_{\tilde{X}}=\mathcal{D}_{X}[\theta]$-module and $\mu^{dR}_{\Lambda}\mathcal{C}_{f}$ a $\mathcal{D}_{\tilde{\Lambda}}$-module monodromic along $\mathbf{C}^{\times}$. Recall the correspondence between monodromic $D$-modules on $X\times\mathbf{C}^{\times}$ and $\mathcal{D}_{X}[\theta]$-modules from Section 2.1.
Proposition 5.6.
For any $\lambda\in\mathbf{C}$ there is an isomorphism of $\mathcal{D}_{X}[\theta]$-modules
$\operatorname{gr}^{0}_{\Lambda}\mathcal{C}^{\lambda}_{f}\cong(\phi_{f}^{\mathcal{D}},\theta+\lambda-1/2).$
Moreover,
$\mu^{dR}_{\Lambda}\mathcal{C}_{f}=\mathcal{D}_{\tilde{\Lambda}}\otimes_{\gamma^{-1}\mathcal{D}_{X}[\theta]}\gamma^{-1}\operatorname{gr}^{0}_{\Lambda}\mathcal{C}_{f}$
is the corresponding monodromic $D$-module on $X\times\mathbf{C}^{\times}$.
Proof.
The first claim for $\lambda=1/2$ follows from the comparison of the usual and microlocal $V$-filtration given by (2). The automorphism $t\mapsto t+(\lambda-1/2)\partial_{t}^{-1}$ of ^$\widehat{\mathcal{E}}_{X\times\mathbf{C}}|_{\mathrm{J}^{1}X}$ sends $\mathcal{C}_{f}$ to $\mathcal{C}^{\lambda}_{f}$ and $\theta\mapsto\theta+(\lambda-1/2)$ which proves the first claim for a general $\lambda$.
The second claim is merely the definition of the microlocalization functor. ∎
Using this microlocal perspective we can prove a relationship between vanishing cycles and DQ modules. Recall from 4.13 that
$\mathcal{C}^{\mathrm{R}}_{0}\otimes^{\mathbb{L}}_{\widehat{\mathcal{W}}_{X}}\mathcal{C}_{f}$
is a differential perverse sheaf.
Theorem 5.7.
For any $\lambda,\mu\in\mathbf{C}$ there is an isomorphism of differential perverse sheaves
$M_{f}\colon\mathcal{C}^{\lambda,\mathrm{R}}_{0}\otimes^{\mathbb{L}}_{\widehat{\mathcal{W}}_{X}}\mathcal{C}^{\mu}_{f}\cong\widehat{\mathrm{RH}}^{-1}(\phi_{f}\otimes\mathbf{C}_{\mu+\lambda}),$
where $\mathbf{C}_{\mu+\lambda}$ is the trivial one-dimensional vector space equipped with the automorphism $\exp(2\pi i(\mu+\lambda))$.
Proof.
Consider the zero section $L=X\subset\mathrm{T}^{*}X$ and its contactification $\Lambda=j^{1}0\subset\mathrm{J}^{1}X$. $\mathcal{C}_{0}$ as a $\widehat{\mathcal{W}}_{X}$-module is simple along $L$. Using 4.14 we get an isomorphism
$\mathcal{C}^{\lambda,\mathrm{R}}_{0}\otimes^{\mathbb{L}}_{\widehat{\mathcal{W}}_{X}}\mathcal{C}^{\mu}_{f}\cong\widehat{\mathrm{RH}}^{-1}\left((\mu^{dR}_{\Lambda}\mathcal{C}^{\lambda,\mathrm{R}}_{0}\otimes^{\mathbb{L}}_{\mathcal{D}_{\tilde{\Lambda}}}\mu^{dR}_{\Lambda}\mathcal{C}^{\mu}_{f})|_{L}[-1]\right)$
of differential perverse sheaves.
Here $\mu^{dR}_{L}\mathcal{C}_{f}$ is a twisted left $D$-module on $X\times\mathbf{C}^{\times}$ monodromic along $\mathbf{C}^{\times}$. By 5.6 and 2.4 we identify
$\mu^{dR}_{L}\mathcal{C}^{\lambda,\mathrm{R}}_{0}\otimes^{\mathbb{L}}_{\mathcal{D}_{X\times\mathbf{C}^{\times}}}\mu^{dR}_{L}\mathcal{C}^{\mu}_{f}\cong K_{X}\otimes^{\mathbb{L}}_{\mathcal{D}_{X}}\phi_{f}^{\mathcal{D}}$
with the monodromy automorphism $T=\exp(-2\pi i\partial_{t}t+2\pi i(\mu+\lambda))$ on the right. Finally, using 5.5 we identify
$K_{X}\otimes^{\mathbb{L}}_{\mathcal{D}_{X}}\phi_{f}^{\mathcal{D}}\cong\mathbb{R}\mathcal{H}om_{\mathcal{D}_{X}}(\mathcal{O}_{X},\phi^{\mathcal{D}}_{f})[\dim X]\cong\phi_{f}.$
Example 5.8.
Consider 5.7 for $X=\mathrm{pt}$. Namely, choose $\lambda,\mu\in\mathbf{C}$ and consider the differential perverse sheaf $\mathcal{C}^{\lambda,\mathrm{R}}_{0}\otimes^{\mathbb{L}}_{\mathbf{C}(\!(\hbar)\!)}\mathcal{C}^{\mu}_{0}$. From the definitions we obtain an isomorphism
$\mathcal{C}^{\lambda,\mathrm{R}}_{0}\otimes^{\mathbb{L}}_{\mathbf{C}(\!(\hbar)\!)}\mathcal{C}^{\mu}_{0}\cong\mathbf{C}(\!(\hbar)\!)$
where the differential structure on the right-hand side is given by $D(1)=\lambda+\mu+1$. Thus, there is an isomorphism
$\mathcal{C}^{\lambda,\mathrm{R}}_{0}\otimes^{\mathbb{L}}_{\mathbf{C}(\!(\hbar)\!)}\mathcal{C}^{\mu}_{0}\cong\mathrm{RH}^{-1}(\mathbf{C}_{\mu+\lambda}).$
5.2. Aside: twisted de Rham complex
Consider the twisted de Rham complex
$\mathrm{DR}_{f}=(\Omega^{\bullet}_{X}(\!(\hbar)\!),d+\hbar^{-1}df)|_{f^{-1}(0)}[\dim X].$
In this section we explain a relationship between DQ modules and the twisted de Rham complex.
Recall that
$\widehat{\mathcal{W}}_{X}|_{X}\cong\mathcal{D}_{X}(\!(\hbar)\!).$
Thus, $\mathcal{C}_{f}|_{X}$ is a certain $\mathcal{D}_{X}(\!(\hbar)\!)$-module. To describe it, let us first analyze $\mathcal{B}_{f}$. We may identify
$\mathcal{B}_{f}\cong\mathcal{O}_{X}[\partial_{t}]^{f},$
where the superscript indicates the $\mathcal{D}_{X\times\mathbf{C}}$-action as follows: the $\mathcal{O}_{X}$ and the $\mathbf{C}[\partial_{t}]$-action is the obvious one; a vector field $\xi\in\mathcal{D}_{X}$ acts by
$\xi(g\partial_{t}^{i})=(\xi g)\partial_{t}^{i}+(\xi f)g\partial_{t}^{i+1}$
and $t\in\mathcal{D}_{\mathbf{C}}$ acts by
$t(g\partial_{t}^{i})=-fg\partial_{t}^{i}-ig\partial_{t}^{i-1}.$
The restriction $\mathcal{B}_{f}|_{X}$ to the hypersurface $t=0$ is then
$\mathcal{B}_{f}|_{X}\cong\mathcal{O}_{X,f^{-1}(0)}[\partial_{t}]^{f}$
As before, let $\hbar=\partial_{t}^{-1}$ and consider the $\mathcal{D}_{X}(\!(\hbar)\!)$-module $\mathcal{O}_{X}(\!(\hbar)\!)^{f}$ with the $\mathcal{D}_{X}(\!(\hbar)\!)$-action as above. Rewriting the $t$-action in terms of the variable $\hbar$ we get
$t(g\hbar^{i})=-fg\hbar^{i}+ig\hbar^{i+1},$
so $t$ acts by $\hbar^{2}\partial_{\hbar}-f$.
Theorem 5.9.
For any $\lambda\in\mathbf{C}$ there is an isomorphism of differential perverse sheaves
$(\mathrm{DR}_{f},\hbar\partial_{\hbar}-\hbar^{-1}f+1/2-\lambda)\cong\mathcal{C}^{\mathrm{R}}_{0}\otimes^{\mathbb{L}}_{\widehat{\mathcal{W}}_{X}}\mathcal{C}^{\lambda}_{f}.$
Proof.
For simplicity we give a proof for $\lambda=1/2$ to simplify some shifts.
We may identify $\mathcal{C}^{\mathrm{R}}_{0}|_{X}\cong K_{X}(\!(\hbar)\!)$ with the obvious right $\mathcal{D}_{X}(\!(\hbar)\!)$-action. By definition
$\mathcal{C}_{f}=\widehat{\mathcal{E}}_{X\times\mathbf{C}}\otimes_{\mathcal{D}_{X\times\mathbf{C}}}\mathcal{B}_{f},$
where we omit obvious inverse image functors. Therefore, its restriction to $t=0$ is
$\mathcal{C}_{f}|_{X}\cong\widehat{\mathcal{E}}_{X\times\mathbf{C}}|_{X}\otimes_{\mathcal{D}_{X\times\mathbf{C}}|_{X}}\mathcal{B}_{f}|_{X}\cong\widehat{\mathcal{E}}_{\mathbf{C},0}\otimes_{\mathcal{D}_{\mathbf{C},0}}\mathcal{B}_{f}|_{X}.$
Consider the de Rham resolution
$\Omega^{\bullet}_{X}\otimes_{\mathcal{O}_{X}}\mathcal{D}_{X}(\!(\hbar)\!)[\dim X]$
of the right $\mathcal{D}_{X}(\!(\hbar)\!)$-module $K_{X}(\!(\hbar)\!)$. Using this resolution we obtain an isomorphism
$\mathcal{C}^{\mathrm{R}}_{0}\otimes^{\mathbb{L}}_{\widehat{\mathcal{W}}_{X}}\mathcal{C}_{f}$ $\cong K_{X}(\!(\hbar)\!)\otimes^{\mathbb{L}}_{\mathcal{D}_{X}(\!(\hbar)\!)}\mathcal{C}_{f}|_{X}$
$\cong\mathcal{C}_{f}|_{X}\otimes_{\mathcal{O}_{X}}\Omega^{\bullet}_{X}[\dim X],$
where the last complex of sheaves is equipped with the de Rham differential.
The $\mathcal{D}_{\mathbf{C},0}$-action on $\mathcal{O}_{X,f^{-1}(0)}(\!(\hbar)\!)^{f}$ extends to a ^$\widehat{\mathcal{E}}_{\mathbf{C},0}$-action, so the natural inclusion
$\mathcal{O}_{X,f^{-1}(0)}[\hbar^{-1}]^{f}\hookrightarrow\mathcal{O}_{X,f^{-1}(0)}(\!(\hbar)\!)^{f}$
extends to a morphism
$\mathcal{C}_{f}|_{X}=\widehat{\mathcal{E}}_{\mathbf{C},0}\otimes_{\mathcal{D}_{\mathbf{C},0}}\mathcal{O}_{X,f^{-1}(0)}[\hbar^{-1}]^{f}\longrightarrow\mathcal{O}_{X,f^{-1}(0)}(\!(\hbar)\!)^{f}.$
Under this morphism the action of $D=\partial_{t}t=\hbar^{-1}t$ on the left goes to the action of $D=\hbar^{-1}(\hbar^{2}\partial_{\hbar}-f)=\hbar\partial_{\hbar}-\hbar^{-1}f$ on the right.
By [SS14, Proposition 2] the induced morphism
$K_{X}(\!(\hbar)\!)\otimes^{\mathbb{L}}_{\mathcal{D}_{X}(\!(\hbar)\!)}\mathcal{C}_{f}|_{X}\longrightarrow\mathrm{DR}_{f}$
is a quasi-isomorphism. ∎
Example 5.10.
Let $W$ be a complex vector space equipped with a quadratic form $q\colon W\rightarrow\mathbf{C}$ and consider $\lambda\in\mathbf{C}$. The differential perverse sheaf $\mathcal{C}^{\mathrm{R}}_{0}\otimes^{\mathbb{L}}_{\widehat{\mathcal{W}}_{W}}\mathcal{C}^{\lambda}_{q}$ is supported on the critical locus of $q$, i.e. at the origin. Therefore, $(\mathcal{C}^{\mathrm{R}}_{0}\otimes^{\mathbb{L}}_{\widehat{\mathcal{W}}_{W}}\mathcal{C}^{\lambda}_{q})|_{0}$ is a $\mathbf{C}(\!(\hbar)\!)$-vector space with an action of the derivation $D$. Let us compute it.
We have an isomorphism
(5) $P_{q}\otimes_{\mathbf{Z}/2}\mathbf{C}(\!(\hbar)\!)\longrightarrow\mathrm{DR}_{f}|_{0}$
given by $\omega\otimes 1\mapsto\omega$. To compute the derivation $D$, choose an identification $W\cong\mathbf{C}^{n}$ under which $q\mapsto\sum_{i}x_{i}^{2}$. Take $\omega=dx_{1}\wedge\dots\wedge dx_{n}$. Then
$(\hbar\partial_{\hbar}-\hbar^{-1}q+1/2-\lambda)\omega=(1/2-\lambda)\omega-\hbar^{-1}q\omega=\left(d+2\hbar^{-1}\sum_{i}x_{i}dx_{i}\right)\alpha+\frac{n+1-2\lambda}{2}\omega,$
where
$\alpha=\sum_{i=1}^{n}(-1)^{i}x_{i}dx_{1}\wedge\dots\widehat{dx_{i}}\wedge\dots dx_{n}.$
So, (5) is an isomorphism of differential vector spaces if we equip $\mathbf{C}(\!(\hbar)\!)$ with the differential structure with
$D(1)=(\dim W+1)/2-\lambda.$
Under the isomorphism constructed in 5.9 the isomorphism (5) becomes the isomorphism
$P_{q}\otimes_{\mathbf{Z}/2}\mathbf{C}(\!(\hbar)\!)\longrightarrow(\mathcal{C}^{\mathrm{R}}_{0}\otimes^{\mathbb{L}}_{\widehat{\mathcal{W}}_{W}}\mathcal{C}^{\lambda}_{q})|_{0}$
given by
$\omega\otimes 1\mapsto\omega\delta(t)\otimes\partial_{t}^{\lambda-1/2}\delta(t-q)$
which is therefore also an isomorphism of differential vector spaces.
Corollary 5.11.
There is an isomorphism of differential perverse sheaves
$(\mathrm{DR}_{f},\hbar\partial_{\hbar}-\hbar^{-1}f)\cong\widehat{\mathrm{RH}}^{-1}(\phi_{f}).$
Proof.
The statement follows from a combination of 5.7 and 5.9. ∎
Remark 5.12.
Such an isomorphism was established in [Sab10, SS14] on the level of cohomology sheaves and was recently upgraded in [Sch23] to an isomorphism of differential perverse sheaves, by different methods than above.
5.3. Thom–Sebastiani theorem
Throughout this section we consider complex manifolds $X_{1},X_{2}$ and holomorphic functions $f_{i}\colon X_{i}\to\mathbf{C}$. Let $f_{1}\boxplus f_{2}\colon X_{1}\times X_{2}\rightarrow\mathbf{C}$ be the function $x_{1},x_{2}\mapsto f_{1}(x_{1})+f_{2}(x_{2})$. We obviously have $f_{1}^{-1}(0)\times f_{2}^{-1}(0)\subseteq(f_{1}\boxplus f_{2})^{-1}(0)$.
Consider the following external tensor product functors:
  • For a commutative ring $R$ and objects $\mathcal{M}_{i}\in\mathbf{D}(X_{i};R)$, $i=1,2$, we have their external tensor product $\mathcal{M}_{1}\boxtimes_{R}\mathcal{M}_{2}\in\mathbf{D}(X_{1}\times X_{2};R)$.
  • For $(\mathcal{F}_{i},T_{i})\in\mathrm{Perv}_{\mathrm{aut}}(X_{i})$ we define
    $(\mathcal{F}_{1},T_{1})\boxtimes(\mathcal{F}_{2},T_{2})=(\mathcal{F}_{1}\boxtimes_{\mathbf{C}}\mathcal{F}_{2},T_{1}\otimes T_{2}).$
  • For $(\mathcal{F}_{i},D_{i})\in\mathrm{Perv}_{\mathrm{diff}}(X_{i})$ we define $(\mathcal{F}_{1},D_{1})\boxtimes(\mathcal{F}_{2},D_{2})=(\mathcal{F}_{1}\boxtimes_{\mathbf{C}(\!(\hbar)\!)}\mathcal{F}_{2},D)$ with the differential structure $D(m_{1}\otimes m_{2})=Dm_{1}\otimes m_{2}+m_{1}\otimes Dm_{2}$.
By construction we have a commutative diagram
[…]
Let us introduce analogs of these tensor products for microdifferential modules. Consider the following external tensor products:
  • Given $\mathcal{M}_{i}\in\mathrm{Mod}(\mathcal{D}_{X_{i}})$, $i=1,2$, we define
    $\mathcal{M}_{1}\stackrel{{\scriptstyle\mathcal{D}}}{{\boxtimes}}\mathcal{M}_{2}=\mathcal{D}_{X_{1}\times X_{2}}\bigotimes_{\mathcal{D}_{X_{1}}\boxtimes_{\mathbf{C}}\mathcal{D}_{X_{2}}}\mathcal{M}_{1}\boxtimes_{\mathbf{C}}\mathcal{M}_{2}.$
  • Given $\mathcal{M}_{i}\in\mathrm{Mod}(\widehat{\mathcal{E}}_{X_{i}})$, $i=1,2$, we define
    $\mathcal{M}_{1}\stackrel{{\scriptstyle\mathcal{E}}}{{\boxtimes}}\mathcal{M}_{2}=\widehat{\mathcal{E}}_{X_{1}\times X_{2}}\bigotimes_{\widehat{\mathcal{E}}_{X_{1}}\boxtimes_{\mathbf{C}}\widehat{\mathcal{E}}_{X_{2}}}\mathcal{M}_{1}\boxtimes_{\mathbf{C}}\mathcal{M}_{2}.$
We denote by $\mathcal{D}_{\ast}$ the transfer bimodule associated to the addition map
$a\colon X_{1}\times\mathbf{C}_{t_{1}}\times X_{2}\times\mathbf{C}_{t_{2}}\longrightarrow X_{1}\times X_{2}\times\mathbf{C}_{t}$
given by $(x_{1},t_{2},x_{2},t_{2})\mapsto(x_{1},x_{2},t_{1}+t_{2})$. Explicitly,
$\mathcal{D}_{\ast}:=\frac{\mathcal{D}_{X_{1}\times\mathbf{C}_{t_{1}}\times X_{2}\times\mathbf{C}_{t_{2}}}}{(\partial_{t_{1}}-\partial_{t_{2}})\mathcal{D}_{X_{1}\times\mathbf{C}_{t_{1}}\times X_{2}\times\mathbf{C}_{t_{2}}}}$
This has the structure of a $(a^{-1}\mathcal{D}_{X_{1}\times X_{2}\times\mathbf{C}_{t}},\mathcal{D}_{X_{1}\times\mathbf{C}_{t_{1}}\times X_{2}\times\mathbf{C}_{t_{2}}})$-bimodule, where $t$ acts on the left via $t_{1}+t_{2}$. We have the associated functor
$\stackrel{{\scriptstyle\ast}}{{\boxtimes}}\colon\mathrm{Mod}(\mathcal{D}_{X_{1}\times\mathbf{C}_{t_{1}}})\times\mathrm{Mod}(\mathcal{D}_{X_{2}\times\mathbf{C}_{t_{2}}})\longrightarrow\mathrm{Mod}(\mathcal{D}_{X_{1}\times X_{2}\times\mathbf{C}_{t}}),$
defined by
$\mathcal{M}_{1}\stackrel{{\scriptstyle\ast}}{{\boxtimes}}\mathcal{M}_{2}:=a_{\ast}\left(\mathcal{D}_{\ast}\bigotimes_{\mathcal{D}_{X_{1}\times\mathbf{C}_{t_{1}}\times X_{2}\times\mathbf{C}_{t_{2}}}}\mathcal{M}_{1}\stackrel{{\scriptstyle\mathcal{D}}}{{\boxtimes}}\mathcal{M}_{2}\right).$
In this section, for any complex manifold $X$, we will consider ^$\widehat{\mathcal{E}}_{X\times\mathbf{C}}$ as a sheaf on $\mathrm{J}^{1}X\subseteq\mathrm{P}^{\ast}(X\times\mathbf{C})$. We denote by $\alpha\colon\mathrm{J}^{1}X_{1}\times\mathrm{J}^{1}X_{2}\to\mathrm{J}^{1}(X_{1}\times X_{2})$ the addition map. Consider the corresponding transfer $(\alpha^{-1}\widehat{\mathcal{E}}_{X_{1}\times X_{2}\times\mathbf{C}_{t}},\widehat{\mathcal{E}}_{X_{1}\times\mathbf{C}_{t_{1}}\times X_{2}\times\mathbf{C}_{t_{2}}})$-bimodule (see [SKK73, Section 1.3] for the general notion of transfer bimodules)
^:=^$\widehat{\mathcal{E}}_{\ast}:=\frac{\widehat{\mathcal{E}}_{X_{1}\times\mathbf{C}_{t_{1}}\times X_{2}\times\mathbf{C}_{t_{2}}}}{(\partial_{t_{1}}-\partial_{t_{2}})\widehat{\mathcal{E}}_{X_{1}\times\mathbf{C}_{t_{1}}\times X_{2}\times\mathbf{C}_{t_{2}}}}.$
Similarly, we have the associated functor
$\stackrel{{\scriptstyle\ast}}{{\boxtimes}}\colon\mathrm{Mod}(\widehat{\mathcal{E}}_{X_{1}\times\mathbf{C}_{t_{1}}})\times\mathrm{Mod}(\widehat{\mathcal{E}}_{X_{2}\times\mathbf{C}_{t_{2}}})\longrightarrow\mathrm{Mod}(\widehat{\mathcal{E}}_{X_{1}\times X_{2}\times\mathbf{C}_{t}}),$
defined by
$\mathcal{M}_{1}\stackrel{{\scriptstyle\ast}}{{\boxtimes}}\mathcal{M}_{2}:=\alpha_{\ast}\left(\widehat{\mathcal{E}}_{\ast}\bigotimes_{\widehat{\mathcal{E}}_{X_{1}\times\mathbf{C}_{t_{1}}\times X_{2}\times\mathbf{C}_{t_{2}}}}\mathcal{M}_{1}\stackrel{{\scriptstyle\mathcal{E}}}{{\boxtimes}}\mathcal{M}_{2}\right)$
Again, we have a commutative diagram
${\mathrm{Mod}(\mathcal{D}_{X_{1}\times\mathbf{C}})\times\mathrm{Mod}(\mathcal{D}_{X_{2}\times\mathbf{C}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\stackrel{{\scriptstyle\ast}}{{\boxtimes}}}$${\mathrm{Mod}(\mathcal{D}_{X_{1}\times X_{2}\times\mathbf{C}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{Mod}(\widehat{\mathcal{E}}_{X_{1}\times\mathbf{C}})\times\mathrm{Mod}(\widehat{\mathcal{E}}_{X_{2}\times\mathbf{C}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\stackrel{{\scriptstyle\ast}}{{\boxtimes}}}$${\mathrm{Mod}(\widehat{\mathcal{E}}_{X_{1}\times X_{2}\times\mathbf{C}}),}$
where the vertical functors $\mathrm{Mod}(\mathcal{D}_{X})\to\mathrm{Mod}(\widehat{\mathcal{E}}_{X})$ are given by ^$\mathcal{M}\mapsto\widehat{\mathcal{E}}_{X}\otimes_{\pi^{-1}\mathcal{D}_{X}}\pi^{-1}\mathcal{M}$.
The functors $\stackrel{{\scriptstyle\ast}}{{\boxtimes}}$ for $D$- and $E$-modules preserve regular holonomic modules. We will now use these functors to construct a Thom–Sebastiani isomorphism for the sheaf of vanishing cycles. We begin with a version of the Thom–Sebastiani isomorphism for $E$-modules.
Proposition 5.13.
For any $\lambda,\mu\in\mathbf{C}$ the morphism of ^$\widehat{\mathcal{E}}_{X_{1}\times X_{2}\times\mathbf{C}}$-modules
$\mathrm{TS}^{\mathcal{E}}_{f_{1},f_{2}}\colon\mathcal{C}^{\lambda+\mu-1/2}_{f_{1}\boxplus f_{2}}\longrightarrow\mathcal{C}^{\lambda}_{f_{1}}\stackrel{{\scriptstyle\ast}}{{\boxtimes}}\mathcal{C}^{\mu}_{f_{2}}$
given by
$\partial_{t}^{\lambda+\mu-1}\delta(t-f_{1}-f_{2})\mapsto\partial_{t_{1}}^{\lambda-1/2}\delta(t_{1}-f_{1})\otimes\partial_{t_{2}}^{\mu-1/2}\delta(t_{2}-f_{2})$
on the generators is an isomorphism.
Proof.
We have
$\mathcal{C}^{\lambda}_{f_{1}}\stackrel{{\scriptstyle\ast}}{{\boxtimes}}\mathcal{C}^{\mu}_{f_{2}}=\alpha_{*}\left(\widehat{\mathcal{E}}_{\ast}\otimes_{\widehat{\mathcal{E}}_{X_{1}\times\mathbf{C}\times X_{2}\times\mathbf{C}}}\mathcal{C}^{\lambda}_{f_{1}}\stackrel{{\scriptstyle\mathcal{E}}}{{\boxtimes}}\mathcal{C}^{\mu}_{f_{2}}\right).$
The restriction of $\alpha\colon\mathrm{J}^{1}X_{1}\times\mathrm{J}^{1}X_{2}\rightarrow\mathrm{J}^{1}(X_{1}\times X_{2})$ to the support of the sheaf ^^$\widehat{\mathcal{E}}_{\ast}\otimes_{\widehat{\mathcal{E}}_{X_{1}\times\mathbf{C}\times X_{2}\times\mathbf{C}}}\mathcal{C}^{\lambda}_{f_{1}}\stackrel{{\scriptstyle\mathcal{E}}}{{\boxtimes}}\mathcal{C}^{\mu}_{f_{2}}$ is a closed immersion, so $\mathcal{C}^{\lambda}_{f_{1}}\stackrel{{\scriptstyle\ast}}{{\boxtimes}}\mathcal{C}^{\mu}_{f_{2}}$ is a cyclic ^$\widehat{\mathcal{E}}_{X_{1}\times X_{2}\times\mathbf{C}}$-module with generator
$\partial_{t_{1}}^{\lambda-1/2}\delta(t_{1}-f_{1})\otimes\partial_{t_{2}}^{\mu-1/2}\delta(t_{2}-f_{2}).$
It is supported on a smooth Legendrian subvariety $j^{1}f_{1}\stackrel{{\scriptstyle\ast}}{{\times}}j^{1}f_{2}\cong j^{1}(f_{1}\boxplus f_{2})\subset\mathrm{J}^{1}(X_{1}\times X_{2})$, so it is simple. But $\mathcal{C}^{\lambda+\mu-1/2}_{f_{1}\boxplus f_{2}}$ is also a simple ^$\widehat{\mathcal{E}}_{X_{1}\times X_{2}\times\mathbf{C}}$-module along the same Legendrian $j^{1}(f_{1}\boxplus f_{2})$. Since the map between these two simple modules is nonzero (as it sends a generator to a generator), it is an isomorphism. ∎
Consider the Legendrians $\Lambda_{i}=j^{1}0\subset\mathrm{J}^{1}X_{i}$ and $\Lambda=j^{1}0\subset\mathrm{J}^{1}X$. The canonical $V$-filtrations on $\mathcal{C}_{f_{i}}$ along $\Lambda_{i}$ induce by convolution a filtration (which we also denote by $V$) on $\mathcal{C}_{f_{1}}\stackrel{{\scriptstyle\ast}}{{\boxtimes}}\mathcal{C}_{f_{2}}$. Consider also the canonical $V$-filtration on $\mathcal{C}_{f_{1}\boxplus f_{2}}$ along $\Lambda$.
Proposition 5.14.
The morphism
$\mathrm{TS}^{\mathcal{E}}_{f_{1},f_{2}}\colon(\mathcal{C}_{f_{1}\boxplus f_{2}},V)\longrightarrow(\mathcal{C}_{f_{1}}\stackrel{{\scriptstyle\ast}}{{\boxtimes}}\mathcal{C}_{f_{2}},V)$
is strictly compatible with the filtrations. In particular, after passing to the associated graded we obtain an isomorphism
$\mathrm{TS}^{\mathcal{D}}_{f_{1},f_{2}}\colon\phi^{\mathcal{D}}_{f_{1}\boxplus f_{2}}|_{f_{1}^{-1}(0)\times f_{2}^{-1}(0)}\longrightarrow\phi^{\mathcal{D}}_{f_{1}}|_{f_{1}^{-1}(0)}\stackrel{{\scriptstyle\mathcal{D}}}{{\boxtimes}}\phi^{\mathcal{D}}_{f_{2}}|_{f_{2}^{-1}(0)}$
in $\mathrm{Mod}_{\mathrm{rh},\mathrm{aut}}(\mathcal{D}_{f_{1}^{-1}(0)\times f_{2}^{-1}(0)})$.
Proof.
Let $\Sigma_{i}=\operatorname{Sing}f_{i}^{-1}(0)$. Possibly passing to open subsets of $X_{1},X_{2}$, we may assume that $\operatorname{Sing}(f_{1}\boxplus f_{2})^{-1}(0)=\Sigma_{1}\times\Sigma_{2}$.
Consider the algebraic microlocalization
~$\widetilde{\mathcal{B}}_{f_{i}}=\mathbf{C}[\partial_{t_{i}},\partial_{t_{i}}^{-1}]\otimes_{\mathbf{C}[\partial_{t_{i}}]}\mathcal{B}_{f_{i}}.$
The $V$-filtration on $\mathcal{B}_{f_{i}}$ induces one on ~$\widetilde{\mathcal{B}}_{f_{i}}$. Moreover, the natural maps $\pi^{-1}\mathcal{B}_{f_{i}}\rightarrow\pi^{-1}\widetilde{\mathcal{B}}_{f_{i}}\rightarrow\mathcal{C}_{f_{i}}$ induce isomorphisms
$\phi^{\mathcal{D}}_{f_{i}}=\operatorname{gr}^{0}_{V}\mathcal{B}_{f_{i}}\longrightarrow\operatorname{gr}^{0}_{V}\widetilde{\mathcal{B}}_{f_{i}}\longrightarrow\operatorname{gr}^{0}_{V}\mathcal{C}_{f_{i}}.$
Indeed, for the first map this is shown in [MSS20, Section 1.1] and for the composite this is (2).
The $V$-filtrations on ~$\widetilde{\mathcal{B}}_{f_{i}}$ induce by convolution a filtration on ~$\widetilde{\mathcal{B}}_{f_{1}}\stackrel{{\scriptstyle\mathcal{D}}}{{\boxtimes}}\widetilde{\mathcal{B}}_{f_{2}}$. It is shown in [MSS20, Theorem 1.2] and [SS, Lemma 11.8.2] that the natural map
~$\widetilde{\mathcal{B}}_{f_{1}}\stackrel{{\scriptstyle\mathcal{D}}}{{\boxtimes}}\widetilde{\mathcal{B}}_{f_{2}}\longrightarrow\widetilde{\mathcal{B}}_{f_{1}\boxplus f_{2}}$
defined by $\delta(t-f_{1})\otimes\delta(t-f_{2})\mapsto\delta(t-(f_{1}\boxplus f_{2}))$ is strictly compatible with the $V$-filtrations. Moreover, it induces an isomorphism after taking the cokernel of $\partial_{t_{1}}-\partial_{t_{2}}$ on the left-hand side. Applying $\operatorname{gr}^{0}_{V}$ we obtain the isomorphism $\mathrm{TS}^{\mathcal{D}}_{f_{1},f_{2}}$. ∎
Under the Riemann–Hilbert correspondence the Thom–Sebastiani isomorphism for $D$-modules goes to the Thom–Sebastiani isomorphism
(6) $\mathrm{TS}_{f_{1},f_{2}}\colon\phi_{f_{1}}|_{f_{1}^{-1}(0)}\boxtimes\phi_{f_{2}}|_{f_{2}^{-1}(0)}\cong\phi_{f_{1}\boxplus f_{2}}|_{f_{1}^{-1}(0)\times f_{2}^{-1}(0)}$
in $\mathrm{Perv}_{\mathrm{aut}}(f_{1}^{-1}(0)\times f_{2}^{-1}(0))$.
There is a natural morphism
(7) $(\mathcal{C}_{0}^{\mathrm{R}}\otimes^{\mathbb{L}}_{\widehat{\mathcal{W}}_{X_{1}}}\mathcal{C}_{f_{1}})\boxtimes(\mathcal{C}_{0}^{\mathrm{R}}\otimes^{\mathbb{L}}_{\widehat{\mathcal{W}}_{X_{2}}}\mathcal{C}_{f_{2}})\longrightarrow(\mathcal{C}_{0}\stackrel{{\scriptstyle\ast}}{{\boxtimes}}\mathcal{C}_{0})^{\mathrm{R}}\otimes^{\mathbb{L}}_{\widehat{\mathcal{W}}_{X_{1}\times X_{2}}}(\mathcal{C}_{f_{1}}\stackrel{{\scriptstyle\ast}}{{\boxtimes}}\mathcal{C}_{f_{2}})$
of differential perverse sheaves. The compatibility between the Thom–Sebastiani isomorphisms $\mathrm{TS}^{\mathcal{E}}_{f_{1},f_{2}}$ and $\mathrm{TS}_{f_{1},f_{2}}$ is then expressed by the following commutative diagram.
Proposition 5.15.
The diagram
[…]
commutes, where the vertical isomorphisms are given by the isomorphisms from 5.7. In particular, (7) is an isomorphism.
Proof.
Unwinding the construction of $M_{f}$ in 5.7 via microlocalization, the top line of the diagram under consideration may be computed as the image under $\widehat{\mathrm{RH}}^{-1}$ of a sequence of morphisms of perverse sheaves with automorphism:
$(K_{X_{1}}\otimes_{\mathcal{D}_{X_{1}}}^{\mathbb{L}}\phi_{f_{1}}^{\mathcal{D}})\boxtimes(K_{X_{2}}\otimes_{\mathcal{D}_{X_{2}}}^{\mathbb{L}}\phi_{f_{2}}^{\mathcal{D}})\to(K_{X_{1}}\stackrel{{\scriptstyle\mathcal{D}}}{{\boxtimes}}K_{X_{2}})\otimes_{\mathcal{D}_{X_{1}\times X_{2}}}^{\mathbb{L}}(\phi_{f_{1}}\stackrel{{\scriptstyle\mathcal{D}}}{{\boxtimes}}\phi_{f_{2}})\to K_{X_{1}\times X_{2}}\otimes_{\mathcal{D}_{X_{1}\times X_{2}}}^{\mathbb{L}}\phi_{f_{1}\boxplus f_{2}}.$
5.15 then implies that the morphism on the right is induced by $\mathrm{TS}_{f_{1},f_{2}}^{\mathcal{D}}$ (together with the natural identification of right $\mathcal{D}$-modules $K_{X_{1}}\stackrel{{\scriptstyle\mathcal{D}}}{{\boxtimes}}K_{X_{2}}\cong K_{X_{1}\times X_{2}}$, which is manifestly induced by $\mathrm{TS}_{0,0}^{\mathcal{E}}$ under microlocalization). ∎
6. DT sheaf on oriented d-critical loci
6.1. Exact two-forms on complex-analytic spaces
Recall that given a chain complex $V$ in degrees $\geq-1$ one can canonically associate to it a groupoid $|V|$ by a basic version of the Dold–Kan correspondence: its objects are closed elements $x\in V^{0}$ of degree 00 and morphisms from $x$ to $y$ are given by elements $h\in V^{-1}$ of degree $-1$ such that $dh=x-y$.
[…]
Definition 6.1.
The groupoid of exact two-forms of degree 00 is
[…]
The set of exact two-forms of degree $-1$ is
[…]
Example 6.2.
If $X$ is a complex manifold with an open cover $\{U_{i}\}$ by open balls, then $\mathcal{A}^{2,\mathrm{ex}}(X,0)$ has objects given by pairs $(f_{ij},\alpha_{i})$, where $f_{ij}$ is a function on $U_{ij}$ and $\alpha_{i}$ is a one-form on $U_{i}$ which together satisfy
$\alpha_{i}-\alpha_{j}=df_{ij},\qquad f_{ij}+f_{jk}+f_{ki}=0.$
Morphisms from $(f_{ij},\alpha_{i})$ to $(g_{ij},\beta_{i})$ are given by functions $h_{i}$ such that
$f_{ij}-g_{ij}=h_{i}-h_{j},\qquad\alpha_{i}-\beta_{i}=dh_{i}.$
Similarly, $\mathcal{A}^{2,\mathrm{ex}}(X,-1)$ is the set of locally constant functions on $X$.
These constructions have the following features:
  • If $Y\subset X$ is an analytic subset, there is a natural restriction map
    $\mathcal{A}^{2,\mathrm{ex}}(X,0)\longrightarrow\mathcal{A}^{2,\mathrm{ex}}(Y,0).$
  • One may identify
    $\mathrm{End}_{\mathcal{A}^{2,\mathrm{ex}}(X,0)}(0)\cong\mathcal{A}^{2,\mathrm{ex}}(X,-1).$
  • There is a forgetful map
    $\mathcal{A}^{2,\mathrm{ex}}(X,-1)\longrightarrow\mathrm{H}^{0}(X,\mathcal{O}_{X}),$
    so an exact two-form of degree $-1$ has an underlying function. Moreover, the restriction of this function to the reduced part $X^{\mathrm{red}}$ is locally constant.
6.2. d-critical loci
We will use the notions related to d-critical loci from [Bra+15, Joy15] which we now briefly recall. Let us first explain how to construct exact two-forms of degree $-1$ on critical loci.
Proposition 6.3.
Let $U$ be a complex manifold and $f\colon U\rightarrow\mathbf{C}$ is a holomorphic function. There is a natural exact two-form of degree $-1$ on the critical locus $\mathrm{Crit}(f)$. Its underlying function coincides with the restriction of $f$ to $\mathrm{Crit}(f)$.
Proof.
For a one-form $\alpha$ on $U$ we denote by $\Gamma_{\alpha}\subset\mathrm{T}^{*}U$ its graph. The critical locus fits into a commutative diagram
[…]
The Liouville one-form $\lambda$ on $\mathrm{T}^{*}U$ gives an element $\lambda\in\mathcal{A}^{ex,}(\mathrm{T}^{*}U,0)$. Its restriction to $\Gamma_{df}$ has a nullhomotopy provided by $f$ and similarly for $\Gamma_{0}$. Comparing the two nullhomotopies on $\mathrm{Crit}(f)$ we obtain an endomorphism $s\in\mathrm{End}_{\mathcal{A}^{2,\mathrm{ex}}(\mathrm{Crit}(f),0)}(0)\cong\mathcal{A}^{2,\mathrm{ex}}(\mathrm{Crit}(f),-1)$. ∎
The above structure on a critical locus is globalized using the notion of a d-critical locus.
Definition 6.4.
Let $X$ be a complex analytic space. A d-critical structure on $X$ is the data of an exact two-form $s\in\mathcal{A}^{2,\mathrm{ex}}(X,-1)$ of degree $-1$ with the following property:
  • For any point $x\in X$ there is a neighborhood $R\subset X$ together with a complex manifold $U$, a holomorphic function $f\colon U\rightarrow\mathbf{C}$ and a closed embedding $i\colon R\hookrightarrow U$ such that $i(R)=\mathrm{Crit}(f)$, $s|_{R}$ coincides with the canonical exact two-form of degree $-1$ on $\mathrm{Crit}(f)$ constructed in 6.3 and $f|_{R^{\mathrm{red}}}=0$. We call such a quadruple $(R,U,f,i)$ a critical chart of $X$.
if $X$ has a d-critical structure, we call $X$ a d-critical locus.
We will now recall important results on “transition maps” between different critical charts.
Definition 6.5.
Given a d-critical locus $(X,s)$, an embedding of critical charts $(R,U,f,i)\hookrightarrow(S,V,g,j)$, where $R\subset S$, is a locally closed embedding $\Xi\colon U\hookrightarrow V$ compatible with the rest of the data.
We have the following important theorems proved in [Joy15]:
  • Given an embedding $\Xi\colon(R,U,f,i)\hookrightarrow(S,V,g,j)$ of critical charts there is a canonical quadratic form $q_{\Xi}$ on the normal bundle $N_{\Xi}|_{R}$ of $U\subset V$. The quadratic vector bundle $N_{\Xi}|_{R}$ defines a $\mathbf{Z}/2$-torsor $P_{\Xi}$ on $R$ of orientations of $q_{\Xi}$. See [Joy15, Definition 2.26].
  • An embedding of critical charts $(R,U,f,i)\hookrightarrow(S,V,g,j)$ is, locally, of the form $(R,U,f,i)\hookrightarrow(R,U\times\mathbf{C}^{n},f\boxplus(z_{1}^{2}+\dots+z_{n}^{2}),j)$, where $U\hookrightarrow U\times\mathbf{C}^{n}$ is given by $u\mapsto(u,0)$. See [Joy15, Proposition 2.22].
  • Given critical charts $(R,U,f,i)$ and $(S,V,g,j)$ and an intersection point $x\in R\cap S$ we can find near $x$ another critical chart $(T,W,h,k)$ and local embeddings $(R,U,f,i)\hookrightarrow(T,W,h,k)$ and $(S,V,g,j)\hookrightarrow(T,W,h,k)$. See [Joy15, Theorem 2.20].
Another important structure on a d-critical locus $(X,s)$ is the virtual canonical bundle $K^{\mathrm{vir}}_{X}$, which is a holomorphic line bundle on $X^{\mathrm{red}}$ with the following properties:
  • Given a critical chart $(R,U,f,i)$ there is an isomorphism $K^{\mathrm{vir}}_{X}|_{R^{\mathrm{red}}}\cong i^{*}K_{U}^{\otimes 2}|_{R^{\mathrm{red}}}$.
  • Given an embedding $\Xi\colon(R,U,f,i)\hookrightarrow(S,V,g,j)$ of critical charts there is an isomorphism
    $i^{*}K_{U}|_{R^{\mathrm{red}}}\otimes\det(N_{\Xi}|_{R})\cong j^{*}K_{V}|_{R^{\mathrm{red}}}$
    whose square, together with a trivialization of $\det(N_{\Xi}|_{R})^{\otimes 2}$ given by the quadratic form $q_{\Xi}$ on $N_{\Xi}|_{R}$, gives rise to an isomorphism of the presentations of the virtual canonical bundle in the two charts.
Definition 6.6.
An orientation of a d-critical locus $(X,s)$ is the choice of a square root $(K^{\mathrm{vir}}_{X})^{1/2}$ of the virtual canonical bundle.
Suppose $(X,s)$ is an oriented d-critical locus. Then we have the following data in local charts:
  • Given a critical chart $(R,U,f,i)$, there is a natural $\mathbf{Z}/2$-torsor $Q_{R,U,f,i}$ on $R$ measuring the difference between $(K^{\mathrm{vir}}_{X})^{1/2}$ and $i^{*}K_{U}|_{R^{\mathrm{red}}}$.
  • Given an embedding $\Xi\colon(R,U,f,i)\hookrightarrow(S,V,g,j)$ of critical charts there is a canonical isomorphism
    (8) $\Lambda_{\Xi}\colon Q_{S,V,g,j}|_{R}\xrightarrow{\sim}P_{\Xi}\otimes_{\mathbf{Z}/2}Q_{R,U,f,i}$
    of $\mathbf{Z}/2$-torsors on $R$.
6.3. Lagrangian intersections
In this section we explain the main example of a d-critical locus in this paper given by a Lagrangian intersection. Let $S$ be a holomorphic symplectic manifold and $L,M\subset S$ two Lagrangian submanifolds. Consider the intersection $L\cap M$ as a complex analytic space. The following is a generalization of 6.3 proved in [Bus14].
Proposition 6.7.
There is a natural d-critical structure on $L\cap M$. Moreover, there is a natural isomorphism of holomorphic line bundles
$K^{\mathrm{vir}}_{L\cap M}\cong K_{L}|_{(L\cap M)^{\mathrm{red}}}\otimes K_{M}|_{(L\cap M)^{\mathrm{red}}}$
on $(L\cap M)^{\mathrm{red}}$. In particular, choices of square roots $K_{L}^{1/2}$ and $K_{M}^{1/2}$ determine an orientation of $L\cap M$.
Proof.
Consider the Lagrangian subvariety $L\cup M\subset S$. According to 1.12, by shrinking $S$ to an open neighborhood of $L\cup M$, there is a canonical contactification $\rho\colon Y\rightarrow S$ together with a lift of $L,M\subset S$ to Legendrian submanifolds $\Lambda_{L},\Lambda_{M}\subset S$ with the property that $\rho\colon\Lambda_{L}\cap\Lambda_{M}\rightarrow L\cap M$ is a homeomorphism.
The contactification $\rho\colon Y\rightarrow S$ defines an exact two-form of degree 00 on $S$: $s\in\mathcal{A}^{2,\mathrm{ex}}(S,0)$. The lift $\Lambda_{L}\subset Y$ of $L\subset S$ provides a nullhomotopy $h_{L}\colon s|_{L}\sim 0$ of the restriction of $s$ to $L$ and similarly for $M\subset S$. Consider a commutative diagram
${L\cap M\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${L\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${M\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${S}$
Restricting the two nullhomotopies $h_{L}$ and $h_{M}$ to $L\cap M$ we obtain an endomorphism
$s\in\mathrm{End}_{\mathcal{A}^{2,\mathrm{ex}}(L\cap M,0)}(0)\cong\mathcal{A}^{2,\mathrm{ex}}(L\cap M,-1).$
Let us now show that $s$ defines a d-critical structure on $L\cap M$.
Locally near a point $x\in L\cap M$ we may choose a polarization on $S$ transverse to $L$ and $M$. By 1.27 we have a local identification of $Y\rightarrow S$ with $\mathrm{J}^{1}L\rightarrow\mathrm{T}^{*}L$, $\Lambda_{L}\rightarrow L$ is identified with $j^{1}0\rightarrow\mathrm{T}^{*}_{L}L$ and $\Lambda_{M}\rightarrow M$ is identified with $j^{1}f\rightarrow\Gamma_{df}$. Thus, near $x$ we may identify $L\cap M\cong\mathrm{Crit}(f)$ and the d-critical structure $s$ reduces to the one on the critical locus $\mathrm{Crit}(f)$ constructed in 6.3. This proves that $s$ defines a d-critical structure on $L\cap M$.
Moreover, it is clear that this d-critical structure on $L\cap M$ coincides with the one constructed in [Bus14, Theorem 3.1], where the virtual canonical bundle is identified in the same way. ∎
Remark 6.8.
As complex analytic spaces $L\cap M\cong M\cap L$. The corresponding two $d$-critical structures defined in 6.7 differ by a sign.
6.4. DT sheaf
Let $(X,s)$ be an oriented d-critical locus. The paper [Bra+15] constructs a canonical perverse sheaf $\phi_{X}\in\mathrm{Perv}_{\mathrm{aut}}(X)$ equipped with an automorphism on $X$, which we call the DT sheaf, locally modeled on the sheaf of vanishing cycles. Let us briefly recall its construction.
For a complex manifold $U$ with a holomorphic function $f\colon U\rightarrow\mathbf{C}$ let
$\mathcal{PV}_{U,f}=(\phi_{f},(-1)^{\dim U}T)\in\mathrm{Perv}_{\mathrm{aut}}(\mathrm{Crit}(f))$
be the sheaf of vanishing cycles with the monodromy operator twisted by a sign.
For a vector space $W$ with a quadratic form $q\colon W\rightarrow\mathbf{C}$ the sheaf of vanishing cycles $\phi_{q}$ is supported at the origin, where it is given by a one-dimensional vector space, and has monodromy automorphism $(-1)^{\dim W}$. We consider its explicit trivialization as follows: there is a unique isomorphism
(9) $T_{q}\colon(\mathcal{PV}_{U,q})|_{0}\cong P_{q}\otimes_{\mathbf{Z}/2}\mathbf{C}$
of vector spaces equipped with automorphisms such that the composite
(10) $P_{q}\otimes_{\mathbf{Z}/2}\mathbf{C}(\!(\hbar)\!)\xrightarrow{\mathrm{RH}^{-1}(T_{q})}\mathrm{RH}^{-1}(\mathcal{PV}_{W,q})|_{0}\xleftarrow{M_{q}}(\mathcal{C}^{\mathrm{R}}_{0}\otimes_{\widehat{\mathcal{W}}_{W}}\mathcal{C}^{(\dim W+1)/2}_{q})|_{0}$
is given by $\omega\otimes 1\mapsto\omega\delta(t)\otimes\partial_{t}^{\dim W/2}\delta(t-q)$ (see Example 5.10 for the composite isomorphism).
Remark 6.9.
Any two choices of the isomorphisms $T_{q}$ compatible with the Thom–Sebastiani isomorphism (6) differ by a multiplication $\alpha^{\dim W}$ for $\alpha\in\mathbf{C}$.
Let $\Xi\colon(R,U,f,i)\hookrightarrow(S,V,g,j)$ be an embedding of critical charts. Then [Bra+15, Theorem 5.4] shows that there is the stabilization isomorphism
(11) $\Theta_{\Xi}\colon\mathcal{PV}_{U,f}\longrightarrow(\mathcal{PV}_{V,g})|_{\mathrm{Crit}(f)}\otimes_{\mathbf{Z}/2}P_{\Xi}$
of perverse sheaves equipped with automorphisms which has the following description:
  • Locally near a point of $R$ identify $V\cong R\times W$ and $g\mapsto f\boxplus q$ for a quadratic form $q\colon W\rightarrow\mathbf{C}$ using [Joy15, Proposition 2.22], so that under this isomorphism the embedding becomes $U\hookrightarrow U\times W$ given by $u\mapsto(u,0)$. In this local model the isomorphism $\Theta_{\Xi}$ is equal to the composite
    $(\phi_{f\boxplus q})|_{\mathrm{Crit}(f)}\cong\phi_{f}\otimes_{\mathbb{Q}}(\phi_{q}|_{0})\cong\phi_{f}\otimes_{\mathbf{Z}/2}P_{q},$
    where the first isomorphism is the Thom–Sebastiani isomorphism (6) and the second isomorphism is $T_{q}$.
Combining the isomorphisms $\Lambda_{\Xi}$ (8) and $\Theta_{\Xi}$ (11) we obtain an isomorphism of perverse sheaves equipped with automorphisms
(12) $\Theta_{\Xi}\otimes{\Lambda_{\Xi}}\colon\mathcal{PV}_{U,f}\otimes_{\mathbf{Z}/2}Q_{R,U,f,i}\longrightarrow(\mathcal{PV}_{S,V,g,j}\otimes_{\mathbf{Z}/2}Q_{S,v,g,j})|_{\mathrm{Crit}(f)}$
for any critical embedding $\Phi\colon(R,U,f,i)\hookrightarrow(S,V,g,j)$.
Let $(X,s)$ be an oriented $d$-critical locus. The DT sheaf $\phi_{X}$ is a perverse sheaf equipped with an automorphism on $X$ with the following property:
  • In any critical chart $(R,U,f,i)$ there is an isomorphism $\phi_{X}|_{R}\cong\mathcal{PV}_{U,f}\otimes_{\mathbf{Z}/2}Q_{R,U,f,i}$.
  • By [Joy15, Theorem 2.20] any two critical charts $(R,U,f,i)$, $(S,V,g,j)$ are locally related by a zig-zag $(R,U,f,i)\hookrightarrow(T,W,h,k)\hookleftarrow(S,V,g,j)$ of critical embeddings. The isomorphism of the local models for $\phi_{X}$ in charts $(R,U,f,i)$ and $(S,V,g,j)$ is given by applying the stabilization isomorphism (12) to the two critical embeddings.
Remark 6.10.
We have chosen the isomorphism $T_{q}$ to define the sheaf $\phi_{X}$. Given two such isomorphisms differing by a multiplication by $\alpha^{\dim W}$ the resulting sheaves $\phi_{X}$ are canonically isomorphic: the isomorphism is given by a multiplication by $\alpha^{\dim U}$ in each critical chart $(R,U,f,i)$.
7. Comparison of the two perverse sheaves
Throughout this section $S$ denotes a complex symplectic manifold, $\rho\colon Y\rightarrow S$ its contactification, $L,M\subset S$ Lagrangian submanifolds equipped with orientation data and $\Lambda_{L},\Lambda_{M}\subset Y$ their Legendrian lifts. The Lagrangian intersection $L\cap M$ has a structure of an oriented d-critical locus by 6.7. The goal of this section is to prove the following result.
Theorem 7.1.
Let $\lambda,\mu\in\mathbf{C}$. There is an isomorphism of differential perverse sheaves
$\mathsf{M}^{\lambda}_{L}\otimes^{\mathbb{L}}_{\mathsf{W}_{S}}\mathsf{M}^{\mu}_{M}\cong\mathrm{RH}^{-1}(\phi_{L\cap M}\otimes\mathbf{C}_{\mu+\lambda}),$
where $\mathbf{C}_{\mu+\lambda}$ is the trivial vector space with the monodromy automorphism $T=\exp(2\pi i(\mu+\lambda))$.
Let $\dim L=\dim M=n$.
7.1. d-critical charts
Both sides of 7.1 involve gluing local models. So, we have to construct an isomorphism in local charts and show that the isomorphism is compatible with gluing data.
The manifold $S$ admits local polarizations. We use them to get a local description of $(Y\rightarrow S,\Lambda_{L}\rightarrow L,\Lambda_{M}\rightarrow M)$ following the terminology from [Bus14, Section 2]:
  • An $L$-chart is an open subset $U\subset S$ equipped with a polarization, where we use 1.27 to identify $\rho^{-1}(U)\rightarrow U$ with an open subset of $\mathrm{J}^{1}L\rightarrow\mathrm{T}^{*}L$, $\Lambda_{L}\rightarrow L$ with $j^{1}0\rightarrow\Gamma_{0}$ and $\Lambda_{M}\rightarrow M$ with $j^{1}f\rightarrow\Gamma_{df}$ for a locally defined function $f\colon L\rightarrow\mathbf{C}$ such that $f|_{(L\cap M)^{\mathrm{red}}}=0$. This defines a critical chart $(\mathrm{Crit}(f),L,f,i)$.
  • An $M$-chart is an open subset $U\subset S$ equipped with a polarization, where we use 1.27 to identify $\rho^{-1}(U)\rightarrow U$ with an open subset of $\mathrm{J}^{1}M\rightarrow\mathrm{T}^{*}M$, $\Lambda_{M}\rightarrow M$ with $j^{1}0\rightarrow\Gamma_{0}$ and $\Lambda_{L}\rightarrow L$ with $j^{1}g\rightarrow\Gamma_{dg}$ for a locally defined function $g\colon M\rightarrow\mathbf{C}$ such that $g|_{(L\cap M)^{\mathrm{red}}}=0$. This defines a critical chart $(\mathrm{Crit}(g),M,-g,i)$.
To simplify the notation, throughout this section we will distinguish two such charts with polarizations $\pi_{1}$ and $\pi_{2}$, respectively; we make the first one into an $L$-chart and the second one into an $M$-chart. We assume the polarizations on the $L$-chart and the $M$-chart are transverse on the intersection. Then we may describe the intersection using an $LM$-chart as follows:
  • By 1.29 there is a locally defined function $h\colon L\times M\rightarrow\mathbf{C}$ such that $\Gamma_{dh}$ gives (locally) the symplectomorphism between $\mathrm{T}^{*}L$ and $\mathrm{T}^{*}M$. The $L$-chart is embedded into the $LM$-chart via $\Xi\colon L\hookrightarrow L\times M$ under which $h$ restricts to $f$. The $M$-chart is embedded into the $LM$-chart via $\Upsilon\colon M\hookrightarrow L\times M$ under which $h$ restricts to $-g$. This defines a critical chart $(\mathrm{Crit}(h),L\times M,h,j)$.
We have the following description of $\phi_{L\cap M}$ in these charts:
  • Consider the $\mathbf{Z}/2$-torsor
    $Q_{L}=Q_{\mathrm{Crit}(f),L,f,i}.$
    The fiber of this $\mathbf{Z}/2$-torsor at $x\in L\cap M$ can be explicitly identified with the $\mathbf{Z}/2$-torsor of isomorphisms $K_{L,x}^{1/2}\rightarrow K_{M,x}^{1/2}$ squaring to the isomorphism
    $C(\mathrm{T}_{L},T_{\pi_{1}},T_{M})\colon K_{L,x}\rightarrow K_{M,x}.$
    In other words, we have an isomorphism
    $Q_{L}\cong P_{T_{L},T_{\pi_{1}},T_{M}},$
    where the $\mathbf{Z}/2$-torsor on the right-hand side is introduced in 1.18. The perverse sheaf $\phi_{L\cap M}$ in the $L$-chart is
    $\phi_{f}\otimes_{\mathbf{Z}/2}Q_{L}.$
  • Consider the $\mathbf{Z}/2$-torsor
    $Q_{M}=Q_{\mathrm{Crit}(g),M,-g,i}\cong P_{T_{M},T_{\pi_{2}},T_{L}}.$
    The perverse sheaf $\phi_{L\cap M}$ in the $M$-chart is
    $\phi_{-g}\otimes_{\mathbf{Z}/2}Q_{M}.$
  • Consider the $\mathbf{Z}/2$-torsor
    $Q_{LM}=Q_{\mathrm{Crit}(h),L\times M,h,}\cong P_{T_{L},T_{\pi_{1}},T_{\pi_{2}},T_{M}}.$
    The perverse sheaf $\phi_{L\cap M}$ in the $LM$-chart is
    $\phi_{h}\otimes_{\mathbf{Z}/2}Q_{LM}.$
  • Consider the embedding $\Xi\colon L\hookrightarrow L\times M$ of the $L$-chart into the $LM$-chart. Using 1.30 we have
    $P_{\Xi}\cong P_{T_{M},T_{\pi_{1}},T_{\pi_{2}},T_{M}}.$
    The isomorphism
    $\Lambda_{\Xi}\colon Q_{L}\longrightarrow P_{\Xi}\otimes_{\mathbf{Z}/2}Q_{LM}$
    is given by composition as in 1.16. We have the stabilization isomorphism
    $\Theta_{\Xi}\colon\phi_{f}\longrightarrow\phi_{h}\otimes_{\mathbf{Z}/2}P_{\Xi},$
    so that the isomorphism between the $L$-chart model and the $LM$-chart model is
    $\Theta_{\Xi}\otimes\Lambda_{\Xi}\colon\phi_{f}\otimes_{\mathbf{Z}/2}Q_{L}\longrightarrow\phi_{h}\otimes_{\mathbf{Z}/2}Q_{LM}.$
  • Consider the embedding $\Upsilon\colon M\hookrightarrow L\times M$ of the $M$-chart into the $LM$-chart. We have
    $P_{\Upsilon}\cong P_{T_{L},T_{\pi_{1}},T_{\pi_{2}},T_{L}}$
    and the isomorphism
    $\Lambda_{\Upsilon}\colon Q_{M}\longrightarrow P_{\Upsilon}\otimes_{\mathbf{Z}/2}Q_{LM}$
    is given by composition as in 1.16. We have the stabilization isomorphism
    $\Theta_{\Upsilon}\colon\phi_{-g}\longrightarrow\phi_{h}\otimes_{\mathbf{Z}/2}P_{\Upsilon},$
    so that the isomorphism between the $M$-chart model and the $LM$-chart model is
    $\Theta_{\Upsilon}\otimes\Lambda_{\Upsilon}\colon\phi_{-g}\otimes_{\mathbf{Z}/2}Q_{M}\longrightarrow\phi_{h}\otimes_{\mathbf{Z}/2}Q_{LM}.$
Let us also describe the orientation data on $L$ and $M$ in these charts. We use the following convention. Consider a complex manifold $U$ with an orientation data, i.e. it has a line bundle $K_{U}^{1/2}$ together with an isomorphism $(K_{U}^{1/2})^{\otimes 2}\cong K_{U}$. Locally $K_{U}^{1/2}$ admits a trivialization; once we choose a nonzero section of $K_{U}^{1/2}$, the isomorphism $(K_{U}^{1/2})^{\otimes 2}\cong K_{U}$ is determined by a volume form $\omega_{U}$. In this case we write $\sqrt{\omega_{U}}$ as the section of $K_{U}^{1/2}$. Changing the trivialization of $K_{U}^{1/2}$ by an invertible function $s$ the volume form changes by $\omega_{U}\mapsto s^{2}\omega_{U}$.
So, we will choose the following additional data:
  • Choose a volume form $\omega_{L}$ on $L$, so that $\sqrt{\omega_{L}}$ defines a nonvanishing section of $K_{L}^{1/2}$.
  • Choose a volume form $\omega_{M}$ on $M$, so that $\sqrt{\omega_{M}}$ defines a nonvanishing section of $K_{M}^{1/2}$.
Using these choices the $\mathbf{Z}/2$-torsors $Q_{L}$, $Q_{M}$, $Q_{LM}$ have the following explicit description:
  • $Q_{L}$ parametrizes functions $s$ on $L$ such that
    $s^{2}=\frac{(\pi_{1}^{S\rightarrow M})^{*}\omega_{M}}{\omega_{L}},$
    where we identify $M$ and $L$ via the composite $\pi_{1}^{S\rightarrow M}\colon L\hookrightarrow S\xrightarrow{\pi_{1}}M$.
  • $Q_{M}$ parametrizes functions $s$ on $M$ such that
    $s^{2}=\frac{(\pi_{2}^{S\rightarrow L})^{*}\omega_{L}}{\omega_{M}},$
    where we identify $L$ and $M$ via the composite $\pi_{1}^{S\rightarrow L}\colon M\hookrightarrow S\xrightarrow{\pi_{2}}L$.
  • $Q_{LM}$ parametrizes functions $u$ on $L\times M$ such that
    $u^{2}=\frac{(\pi_{12}^{S\times S\rightarrow S})^{*}\mathrm{vol}_{S}}{\omega_{L}\wedge\omega_{M}},$
    where we identify $S$ with $L\times M$ via the composite $L\times M\hookrightarrow S\times S\xrightarrow{\pi_{1}\times\pi_{2}}S$.
  • $P_{\Xi}$ parametrizes volume forms $\omega$ on $M$ such that $q_{\Xi}(\omega)=1$. The isomorphism
    $\Lambda_{\Xi}\colon Q_{L}\longrightarrow Q_{LM}\otimes_{\mathbf{Z}/2}P_{\Xi}$
    is given by
    $s\mapsto u\otimes\frac{u\omega_{M}}{s}.$
  • $P_{\Upsilon}$ parametrizes volume forms $\omega$ on $L$ such that $q_{\Upsilon}(\omega)=1$. The isomorphism
    $\Lambda_{\Upsilon}\colon Q_{M}\longrightarrow Q_{LM}\otimes_{\mathbf{Z}/2}P_{\Upsilon}$
    is given by
    $s\mapsto u\otimes\frac{u\omega_{L}}{s}.$
7.2. Local description of DQ modules
Let us now give a local description of deformation quantization algebras and modules in the $L$- and $M$-charts from before.
  • In the $L$-chart we have $\mathsf{E}_{Y}=\widehat{\mathcal{E}}^{\sqrt{v}}_{L\times\mathbf{C}}$ and $\mathsf{W}_{S}=\widehat{\mathcal{W}}^{\sqrt{v}}_{L}$. The modules are
    $\mathsf{M}^{\lambda}_{L}=\mathcal{C}^{\lambda,\sqrt{L}}_{0},\qquad\mathsf{M}^{\mu}_{M}=\mathcal{C}^{\mu,\sqrt{M}}_{f}.$
    We consider simple generators
    $u_{L}^{L}=\sqrt{\omega_{L}}\partial_{t}^{\lambda-1/2}\delta(t)\in\mathcal{C}^{\lambda,\sqrt{L}}_{0},\qquad u_{M}^{L}=\sqrt{\omega_{M}}\partial_{t}^{\mu-1/2}\delta(t-f)\in\mathcal{C}^{\mu,\sqrt{M}}_{f},$
    so that
    $\sigma_{\Lambda_{L}}(u_{L}^{L})=\tau^{\lambda-1/2}\sqrt{d\tau\wedge\omega_{L}},\qquad\sigma_{\Lambda_{M}}(u_{M}^{L})=\tau^{\mu-1/2}\sqrt{d\tau\wedge\omega_{M}}.$
  • In the $M$-chart we have $\mathsf{E}_{Y}=\widehat{\mathcal{E}}^{\sqrt{v}}_{M\times\mathbf{C}}$ and $\mathsf{W}_{S}=\widehat{\mathcal{W}}^{\sqrt{v}}_{M}$. The modules are
    $\mathsf{M}^{\lambda}_{L}=\mathcal{C}^{\lambda,\sqrt{L}}_{g},\qquad\mathsf{M}^{\mu}_{M}=\mathcal{C}^{\mu,\sqrt{M}}_{0}.$
    We consider simple generators
    $u_{L}^{M}=\sqrt{\omega_{L}}\partial_{t}^{\lambda-1/2}\delta(t-g)\in\mathcal{C}^{\lambda,\sqrt{L}}_{g},\qquad u_{M}^{M}=\sqrt{\omega_{M}}\partial_{t}^{\mu-1/2}\delta(t)\in\mathcal{C}^{\mu,\sqrt{M}}_{0},$
    so that
    $\sigma_{\Lambda_{L}}(u_{L}^{M})=\tau^{\lambda-1/2}\sqrt{d\tau\wedge\omega_{L}},\qquad\sigma_{\Lambda_{M}}(u_{M}^{M})=\tau^{\mu-1/2}\sqrt{d\tau\wedge\omega_{M}}.$
  • On the intersection we have a local contactomorphism $\psi\colon\mathrm{J}^{1}L\xrightarrow{\sim}\mathrm{J}^{1}M$ covering the local symplectomorphism $\phi\colon\mathrm{T}^{*}L\xrightarrow{\sim}\mathrm{T}^{*}M$. We choose a $\ast$-preserving QCT $\Psi\colon\widehat{\mathcal{E}}^{\sqrt{v}}_{M\times\mathbf{C}}\xrightarrow{\sim}\psi_{*}\widehat{\mathcal{E}}^{\sqrt{v}}_{L\times\mathbf{C}}$ restricting to a $\ast$-preserving QST $\Phi\colon\widehat{\mathcal{W}}^{\sqrt{v}}_{M}\xrightarrow{\sim}\phi_{*}\widehat{\mathcal{W}}^{\sqrt{v}}_{L}$. There is an isomorphism
    (13) ${}_{\Psi}\mathcal{C}^{\lambda,\sqrt{L}}_{0}\cong\mathcal{C}^{\lambda,\sqrt{L}}_{g}$
    of ^$\widehat{\mathcal{E}}^{\sqrt{v}}_{M\times\mathbf{C}}$-modules uniquely determined by the condition that there is an element $b^{LM}_{L}\in\widehat{\mathcal{E}}^{\sqrt{v}}_{M\times\mathbf{C}}(0)$ such that under this isomorphism $u_{L}^{L}\mapsto b^{LM}_{L}u_{L}^{M}$ and $\sigma_{0}(b^{LM}_{L})|_{\Lambda_{L}}=1$. There is also an isomorphism
    (14) $\mathcal{C}^{\mu,\sqrt{M}}_{f}\cong{}_{\Psi^{-1}}\mathcal{C}^{\mu,\sqrt{M}}_{0}$
    of ^$\widehat{\mathcal{E}}^{\sqrt{v}}_{L\times\mathbf{C}}$-modules uniquely determined by the condition that there is an element $b^{LM}_{M}\in\widehat{\mathcal{E}}^{\sqrt{v}}_{L\times\mathbf{C}}(0)$ such that under this isomorphism $u_{M}^{L}\mapsto b^{LM}_{M}u_{M}^{M}$ and $\sigma_{0}(b^{LM}_{M})|_{\Lambda_{M}}=1$.
Using 3.15 we encode the QCT $\Psi$ using a ^$\widehat{\mathcal{E}}^{\sqrt{v}}_{L\times\mathbf{C}\times M\times\mathbf{C}}$-module $\mathcal{M}$ simple along the Legendrian $\mathrm{P}^{*}_{Z}(L\times\mathbf{C}\times M\times\mathbf{C})$, where $Z=\{t_{L}-t_{M}=h\}\subset L\times\mathbf{C}\times M\times\mathbf{C}$, together with a simple generator $K\in\mathcal{M}$. Using 4.6 we encode the QST $\Phi$ using a ^$\widehat{\mathcal{E}}^{\sqrt{v}}_{L\times M\times\mathbf{C}}$-module $\mathcal{M}^{\partial}\subset\mathcal{M}$ simple along the Legendrian $j^{1}h\subset\mathrm{J}^{1}(L\times M)$, again with a simple generator $K\in\mathcal{M}^{\partial}$. We have $\sigma_{j^{1}h}(K)=\sqrt{\tau^{n}\mathrm{vol}_{S}\wedge d\tau}$, so we may locally identify
$\mathcal{M}^{\partial}\cong\mathcal{C}^{(n+1)/2,\sqrt{S}}_{h}$
as ^$\widehat{\mathcal{E}}^{\sqrt{v}}_{L\times M\times\mathbf{C}}$-modules. Let
$u_{LM}=\sqrt{\mathrm{vol}_{S}}\partial_{t}^{n/2}\delta(t-h)\in\mathcal{C}^{(n+1)/2,\sqrt{S}}_{h}$
be the standard simple generator. As $K$ and $u_{LM}$ are two simple generators with the same symbol, we have $K=b_{LM}u_{LM}$, where $b_{LM}\in\widehat{\mathcal{E}}^{\sqrt{v}}_{L\times M\times\mathbf{C}}(0)$ with $\sigma_{0}(b_{LM})|_{j^{1}h}$ locally constant. We rescale $K$ so that $\sigma_{0}(b_{LM})|_{j^{1}h}=1$.
Using $\mathcal{M}$ we may rewrite isomorphisms (13) and (14) as
$\Delta_{\Upsilon}\colon\mathcal{C}_{0}^{\lambda,\sqrt{L}}\otimes_{\widehat{\mathcal{W}}^{\sqrt{v}}_{L}}\mathcal{M}^{\partial}\cong\mathcal{C}_{0}^{\lambda,\sqrt{L}}\otimes_{\widehat{\mathcal{E}}^{\sqrt{v}}_{L\times\mathbf{C}}}\mathcal{M}\cong\mathcal{C}^{\lambda,\sqrt{L}}_{-g},$
an isomorphism of left ^$\widehat{\mathcal{E}}^{\sqrt{v}}_{M\times\mathbf{C}}$-modules, uniquely determined by the fact that it sends $u_{L}^{M}\otimes b_{L}^{LM}K\mapsto u_{L}^{L}$ and
$\Delta_{\Xi}\colon\mathcal{C}_{0}^{\mu,\sqrt{M}}\otimes_{\widehat{\mathcal{W}}^{\sqrt{v}}_{M}}\mathcal{M}^{\partial}\cong\mathcal{C}_{0}^{\mu,\sqrt{M}}\otimes_{\widehat{\mathcal{E}}^{\sqrt{v}}_{M\times\mathbf{C}}}\mathcal{M}\cong\mathcal{C}_{f}^{\mu,\sqrt{M}},$
an isomorphism of left ^$\widehat{\mathcal{E}}^{\sqrt{v}}_{L\times\mathbf{C}}$-modules, uniquely determined by the fact that it sends $u_{M}^{M}\otimes b_{M}^{LM}K\mapsto u_{M}^{L}$.
We will use the following isomorphisms to change the orientation data for DQ modules:
  • A section of $Q_{L}$, i.e. a function $s$ on $L|_{L\cap M}$ such that $s^{2}=(\pi_{1}^{S\rightarrow M})^{*}\omega_{M}/\omega_{L}$, determines an isomorphism $K_{L}^{1/2}\rightarrow K_{M}^{1/2}$ given by sending $s\sqrt{\omega_{L}}\mapsto\sqrt{\omega_{M}}$. Therefore, we obtain an isomorphism $\mathcal{C}^{\mu,\sqrt{L}}_{f}\otimes Q_{L}\cong\mathcal{C}^{\mu,\sqrt{M}}_{f}$ under which
    $s\sqrt{\omega_{L}}\partial_{t}^{\mu-1/2}\delta(t-f)\otimes s\mapsto\sqrt{\omega_{M}}\partial_{t}^{\mu-1/2}\delta(t-f)=u^{L}_{M}.$
  • A section of $Q_{M}$, i.e. a function $s$ on $M|_{L\cap M}$ such that $s^{2}=(\pi_{2}^{S\rightarrow L})^{*}\omega_{L}/\omega_{M}$, determines an isomorphism $K_{M}^{1/2}\rightarrow K_{L}^{1/2}$ given by sending $s\sqrt{\omega_{M}}\mapsto\sqrt{\omega_{L}}$. Therefore, we obtain an isomorphism $\mathcal{C}^{\lambda,\sqrt{M}}_{g}\otimes Q_{M}\cong\mathcal{C}^{\lambda,\sqrt{L}}_{g}$ under which
    $su_{L}^{M}\otimes s=s\sqrt{\omega_{L}}\partial_{t}^{\lambda-1/2}\delta(t-g)\otimes s\mapsto\sqrt{\omega_{M}}\partial_{t}^{\lambda-1/2}\delta(t-g).$
7.3. Proof of 7.1
We use the gluing data for the perverse sheaf $\phi_{L\cap M}$ as described in [Bus14, Section 2.3]. Cover $S$ by open sets which admit polarizations and consider them as $L$-charts. Given an intersection of two such $L$-charts $U_{1},U_{2}$, we may cover $U_{1}\cap U_{2}$ by open sets $\{V_{a}\}$ with polarizations such that for every $a$ the polarization on $V_{a}$ is transverse both to the polarization on $U_{1}$ and on $U_{2}$. We regard $V_{a}$ as $M$-charts.
Thus, we have to provide an isomorphism of differential perverse sheaves as in 7.1 in each $L$- and $M$-chart and show that the isomorphisms (merely of perverse sheaves) are equal on the intersections.
Step 1. Consider an $L$-chart. By the explicit description of the modules given in the previous sections we have to construct an isomorphism of differential perverse sheaves
$\mathcal{C}^{\lambda,\sqrt{L}}_{0}\otimes^{\mathbb{L}}_{\widehat{\mathcal{W}}^{\sqrt{v}}_{L}}\mathcal{C}^{\mu,\sqrt{M}}_{f}\cong\mathrm{RH}^{-1}(\phi_{f}\otimes\mathbf{C}_{\mu+\lambda})\otimes_{\mathbf{Z}/2}Q_{L}.$
This isomorphism is given by the composite of the isomorphism $\mathcal{C}^{\sqrt{M}}_{f}\cong\mathcal{C}^{\sqrt{L}}_{f}\otimes_{\mathbf{Z}/2}Q_{L}$, the isomorphism
$\mathcal{C}^{\lambda,\sqrt{L}}_{0}\otimes^{\mathbb{L}}_{\widehat{\mathcal{W}}^{\sqrt{v}}_{L}}\mathcal{C}^{\mu,\sqrt{L}}_{f}\cong\mathcal{C}^{\lambda,\mathrm{R}}_{0}\otimes^{\mathbb{L}}_{\widehat{\mathcal{W}}_{L}}\mathcal{C}^{\mu}_{f}$
and the isomorphism $M_{f}$ constructed in 5.7.
Step 2. Consider an $M$-chart. In this case we have to construct an isomorphism of differential perverse sheaves
$\mathcal{C}^{\lambda,\sqrt{L}}_{g}\otimes^{\mathbb{L}}_{\widehat{\mathcal{W}}^{\sqrt{v}}_{M}}\mathcal{C}^{\mu,\sqrt{M}}_{0}\cong\mathrm{RH}^{-1}(\phi_{-g})\otimes\mathbf{C}_{\mu+\lambda}\otimes_{\mathbf{Z}/2}Q_{M}.$
This isomorphism is given by the composite of the isomorphism
$\mathcal{C}^{\lambda,\sqrt{L}}_{g}\otimes^{\mathbb{L}}_{\widehat{\mathcal{W}}^{\sqrt{v}}_{M}}\mathcal{C}^{\mu,\sqrt{M}}_{0}\cong\mathcal{C}^{\mu,\sqrt{M}}_{0}\otimes^{\mathbb{L}}_{\widehat{\mathcal{W}}^{\sqrt{v}}_{M}}\mathcal{C}^{\lambda,\sqrt{L}}_{-g}$
from (4) as well as the isomorphism constructed as in Step 1.
Step 3. We have to show that thus defined isomorphisms are equal on intersections of an $L$-chart and an $M$-chart. We describe the intersection using an $LM$-chart. In this case the isomorphism of DT sheaves in the local models is given by the composite
$\phi_{f}\otimes_{\mathbf{Z}/2}Q_{L}\xrightarrow{\Theta_{\Xi}\otimes\Lambda_{\Xi}}\phi_{h}\otimes_{\mathbf{Z}/2}Q_{LM}\xleftarrow{\Theta_{\Upsilon}\otimes\Lambda_{\Upsilon}}\phi_{-g}\otimes_{\mathbf{Z}/2}Q_{M}.$
Similarly, the isomorphism of the relative tensor product of DQ modules is given by the composite (we do not write the orders as we will be checking an equality of isomorphisms)
$\mathcal{C}^{\sqrt{L}}_{0}\otimes^{\mathbb{L}}_{\widehat{\mathcal{W}}^{\sqrt{v}}_{L}}\mathcal{C}^{\sqrt{M}}_{f}\xrightarrow{\mathrm{id}\otimes\Delta_{\Xi}}(\mathcal{C}^{\sqrt{L}}_{0}\stackrel{{\scriptstyle\ast}}{{\boxtimes}}\mathcal{C}^{\sqrt{M}}_{0})\otimes^{\mathbb{L}}_{\widehat{\mathcal{W}}^{\sqrt{v}}_{L}\boxtimes\widehat{\mathcal{W}}^{\sqrt{v}}_{M}}\mathcal{C}^{\sqrt{S}}_{h}\xleftarrow{\mathrm{id}\otimes\Delta_{\Upsilon}}\mathcal{C}^{\sqrt{M}}_{0}\otimes^{\mathbb{L}}_{\widehat{\mathcal{W}}^{\sqrt{v}}_{M}}\mathcal{C}^{\sqrt{L}}_{-g}\cong\mathcal{C}^{\sqrt{L}}_{g}\otimes^{\mathbb{L}}_{\widehat{\mathcal{W}}^{\sqrt{v}}_{M}}\mathcal{C}^{\sqrt{M}}_{0}.$
We will prove a stronger claim that the diagram
[…]
commutes and similarly for $\Upsilon$. The two proofs are identical, so we will only give it for $\Xi$.
Untwisting the modules, we have to show that the diagram
[…]
is commutative. Here
$\Delta^{\prime}_{\Xi}\colon\mathcal{C}_{f}\otimes_{\mathbf{Z}/2}Q_{L}\longrightarrow\mathcal{C}^{\mathrm{R}}_{0}\otimes_{\widehat{\mathcal{W}}_{M}}\mathcal{C}_{h}\otimes_{\mathbf{Z}/2}Q_{LM}$
is the isomorphism of left ^$\widehat{\mathcal{E}}_{L\times\mathbf{C}}$-modules simple along $j^{1}f$ under which
$\partial_{t}^{\mu-1/2}\delta(t-f)\otimes s\mapsto\frac{u\omega_{M}}{s}\partial_{t}^{\mu-1/2}\delta(t)\otimes P\partial_{t}^{n/2}\delta(t-h)\otimes u$
for $P=b^{LM}_{M}b_{LM}\in\widehat{\mathcal{E}}_{L\times M\times\mathbf{C}}(0)$ with $\sigma_{0}(P)|_{j^{1}h}=1$. Applying the isomorphism $\Lambda_{\Xi}\colon Q_{L}\cong Q_{LM}\otimes_{\mathbf{Z}/2}P_{\Xi}$ we are reduced to showing that the diagram
[…]
commutes, where
$\Delta^{\prime\prime}_{\Xi}\colon\mathcal{C}_{f}\longrightarrow\mathcal{C}^{\mathrm{R}}_{0}\otimes_{\widehat{\mathcal{W}}_{M}}\mathcal{C}_{h}\otimes_{\mathbf{Z}/2}P_{\Xi}$
is the isomorphism of left ^$\widehat{\mathcal{E}}_{L\times\mathbf{C}}$-modules simple along $j^{1}f$ under which
$\partial_{t}^{\mu-1/2}\delta(t-f)\mapsto\omega\partial_{t}^{\mu-1/2}\delta(t)\otimes P\partial_{t}^{n/2}\delta(t-h)\otimes\omega.$
It is enough to check commutativity of the diagram locally. Consider a point $x\in L\cap M$. We have that $\Xi\colon(L\cap M,L,f,i)\rightarrow(L\cap M,L\times M,h,i)$ is a critical embedding. Therefore, by [Joy15, Proposition 2.22] in a neighborhood of $x$ there is a local isomorphism $L\times M\cong L\times W$, where $W=\mathrm{T}_{x}M$ is equipped with the Maslov quadratic form $q=q(\mathrm{T}_{M},\mathrm{T}_{\pi_{1}},\mathrm{T}_{\pi_{2}})\colon W\rightarrow\mathbf{C}$, such that under this isomorphism $h\mapsto f\boxplus q$ and the embedding $\Xi\colon L\rightarrow L\times M$ becomes $(\mathrm{id}\times 0)\colon L\rightarrow L\times W$. The $\mathbf{Z}/2$-torsor $P_{\Xi}$ in a neighborhood of $x$ canonically identifies with $P_{q}$.
Using these identification we obtain Thom–Sebastiani isomorphisms
$\mathrm{TS}^{\mathcal{E}}_{q,f}\colon\mathcal{C}_{q}\stackrel{{\scriptstyle\ast}}{{\boxtimes}}\mathcal{C}_{f}\cong\mathcal{C}_{h}$
from 5.13 and
$\mathrm{TS}_{q,f}\colon\phi_{q}\boxtimes\phi_{f}\cong\phi_{h}$
from (6).
Consider the composite isomorphism
(18) $\mathcal{C}_{f}\xrightarrow{\Delta^{\prime\prime}_{\Xi}}\mathcal{C}^{\mathrm{R}}_{0}\otimes_{\widehat{\mathcal{W}}_{M}}\mathcal{C}_{h}\otimes_{\mathbf{Z}/2}P_{q}\xleftarrow{\mathrm{TS}^{\mathcal{E}}_{q,f}}(\mathcal{C}^{\mathrm{R}}_{0}\otimes_{\widehat{\mathcal{W}}_{W}}\mathcal{C}_{q})\boxtimes\mathcal{C}_{f}\otimes_{\mathbf{Z}/2}P_{q}$
We claim that under this composite we have
$\partial_{t}^{\mu-1/2}\delta(t-f)\mapsto\omega\delta(t)\otimes\partial_{t}^{n/2}\delta(t-q)\otimes\partial_{t}^{\mu-1/2}\delta(t-f)\otimes\omega.$
In other words, $\Delta^{\prime\prime}_{\Xi}=\Delta^{\prime\prime\prime}_{q}\otimes\mathrm{id}$, where
$\Delta^{\prime\prime\prime}_{q}\colon\mathbf{C}(\!(\hbar)\!)\longrightarrow(\mathcal{C}^{\mathrm{R}}_{0}\otimes_{\widehat{\mathcal{W}}_{W}}\mathcal{C}_{q})\otimes_{\mathbf{Z}/2}P_{q}$
is given by
$1\mapsto\omega\delta(t)\otimes\partial_{t}^{n/2}\delta(t-q)\otimes\omega.$
Indeed, both sides of (18) are ^$\widehat{\mathcal{E}}_{L\times\mathbf{C}}$-modules simple along $j^{1}f$. Any two isomorphisms of simple modules differ by a locally-constant $\mathbf{C}$-valued function and the claim follows from the equality $\sigma_{0}(P)_{j^{1}h}=1$.
Therefore, we are reduced to showing that the diagram
[…]
is commutative. The commutativity of the square on the left follows from the definition of $T_{q}$ (see (10)). The commutativity of the square on the right follows from a compatibility of the Thom–Sebastiani isomorphisms $\mathrm{TS}^{\mathcal{E}}$ and $\mathrm{TS}$ (see 5.15). This finishes the proof of 7.1.
7.4. Example: a clean Lagrangian intersection
In this section we consider a simple application of 7.1. Namely, we compute the perverse sheaf $\phi_{L\cap M}$ on the right-hand side when the Lagrangians $L,M$ intersect cleanly, i.e. when $L\cap M$ is smooth.
Since the d-critical locus $L\cap M$ is smooth, it defines a critical chart $(L\cap M,L\cap M,0,\mathrm{id})$ with respect to the zero potential. The orientation data on the Lagrangians induces an orientation data on the d-critical locus $L\cap M$. Since it is also a critical chart, we obtain an isomorphism
$\phi_{L\cap M}\cong Q\otimes_{\mathbf{Z}/2}\mathbf{C},$
where $Q=Q_{L\cap M,L\cap M,0,\mathrm{id}}$ is the $\mathbf{Z}/2$-torsor on the critical chart measuring the difference between the orientation of the d-critical locus and the natural orientation of the critical chart.
By definition, $Q$ parametrizes square roots of the natural isomorphism
$\iota_{L\cap M,L\cap M,0,\mathrm{id}}\colon K^{\mathrm{vir}}_{L\cap M}\longrightarrow(K_{L\cap M})^{\otimes 2}$
described in [Joy15, Theorem 2.28 (i)] for a general d-critical locus. By [Bus14, Theorem 3.1] we may additionally identify
$K^{\mathrm{vir}}_{L\cap M}\cong K_{L}|_{L\cap M}\otimes K_{M}|_{L\cap M}.$
The resulting isomorphism
$\iota^{\prime}\colon K_{L}|_{L\cap M}\otimes K_{M}|_{L\cap M}\cong(K_{L\cap M})^{\otimes 2}$
has the following description. We have a long exact sequence
$0\longrightarrow\mathrm{T}_{L\cap M}\longrightarrow\mathrm{T}_{L}|_{L\cap M}\oplus\mathrm{T}_{M}|_{L\cap M}\longrightarrow T_{S}|_{L\cap M}\longrightarrow\mathrm{T}^{*}_{L\cap M}\longrightarrow 0$
Taking its determinant and trivializing $\det(\mathrm{T}_{S})$ using the symplectic volume form $\mathrm{vol}_{S}$ we get the isomorphism $\iota^{\prime}$.

### v2

@@
 Kashiwara, Polesello, Schapira and D’Agnolo defined canonical deformation quantizations of a holomorphic symplectic manifold and a holomorphic Lagrangian submanifold equipped with an orientation data.
-The goal of this paper is to describe the RHom complex between two such deformation quantization modules associated to a pair of Lagrangian submanifolds in terms of the derived geometry of the Lagrangian intersection.
+The goal of this paper is to use deformation quantization modules to construct a Fukaya-like category of holomorphic Lagrangians, resolving a conjecture of Joyce.
+Our main result describes the RHom complex between two such deformation quantization modules associated to a pair of Lagrangian submanifolds in terms of the derived geometry of the Lagrangian intersection.
 Namely, we identify the RHom complex with the DT sheaf associated to the d-critical structure on the Lagrangian intersection.
+Via a Riemann–Hilbert correspondence this describes the RHom complex in the category of microsheaves of sheaf quantizations of conic holomorphic Lagrangians.
 Introduction
-Deformation quantization modules
-Let $Y$ be a holomorphic contact manifold.
-Kashiwara [Kas96] has defined a filtered $\mathbf{C}$-algebroid $\mathsf{E}_{Y}$ on $Y$.
-It provides a deformation quantization of $Y$ in the sense that $\operatorname{gr}\mathsf{E}_{Y}$ is isomorphic to $\oplus_{m\in\mathbf{Z}}\mathcal{O}_{Y}(m)$ as a sheaf of Poisson algebras.
-Locally $Y$ is isomorphic to an open subset of the projectivized cotangent bundle $\mathrm{P}^{*}X$ of a complex manifold $X$ in which case $\mathsf{E}_{Y}$ reduces to the algebra of microdifferential operators introduced by Sato, Kawai and Kashiwara [SKK73].
-Given a holomorphic Legendrian submanifold $\Lambda\subset Y$ there is, moreover, a classification of regular holonomic $\mathsf{E}_{Y}$-modules supported on $\Lambda$ (see [Kas96, Proposition 4] and its corrected version [DS07, Corollary 6.4]) in terms of twisted local systems on the $\mathbf{C}^{\times}$-torsor $\tilde{\Lambda}\rightarrow\Lambda$ constructed from the symplectization $\tilde{Y}\rightarrow Y$.
-These results were generalized from the contact to the symplectic setting in [PS04, DS07].
-Namely, given a holomorphic symplectic manifold $S$ Polesello and Schapira [PS04] have defined a $\mathbf{C}(\!(\hbar)\!)$-algebroid $\mathsf{W}_{S}$ together with a $\mathbf{C}[\![\hbar]\!]$-lattice $\mathsf{W}_{S}(0)\subset\mathsf{W}_{S}$.
-It again provides a deformation quantization of $S$ in the sense that $\mathsf{W}_{S}(0)/\hbar$ is isomorphic to $\mathcal{O}_{S}$ as a sheaf of Poisson algebras.
-The construction of $\mathsf{W}_{S}$ is based on the following property: given a contactification $\rho\colon Y\rightarrow S$ of the symplectic manifold $S$, there is an embedding $\mathsf{W}_{S}\subset\rho_{*}\mathsf{E}_{Y}$ which identifies $\mathsf{W}_{S}$ with microdifferential operators $\mathsf{E}_{Y}$ on $Y$ commuting with $\partial_{t}$ (the translation vector field along the fibers of $\rho$).
-Under this correspondence $\hbar=\partial_{t}^{-1}$.
-Given a holomorphic Lagrangian submanifold $L\subset S$ equipped with a twisted local system and a number $\lambda\in\mathbf{C}/\mathbf{Z}$ there is a regular holonomic $\mathsf{W}_{S}$-module supported on $L$.
-In a contactification $\rho\colon Y\rightarrow S$ with a lift of $L\subset S$ to a holomorphic Legendrian $\Lambda\subset Y$ it corresponds to a regular holonomic $\mathsf{E}_{Y}$-module supported on $\Lambda$ associated to a twisted local system on $\tilde{\Lambda}\cong L\times\mathbf{C}^{\times}$ with the monodromy $\exp(2\pi i\lambda)$ along the $\mathbf{C}^{\times}$ direction.
-In particular, if $L$ is equipped with a square root $K_{L}^{1/2}$ of the canonical bundle, there is a simple $\mathsf{W}_{S}$-module $\mathsf{M}^{\lambda}_{L}$ supported on $L$.
-Kashiwara and Schapira [KS12] have developed a general theory of deformation quantization modules.
-In particular, they have shown (see [KS12, Theorem 7.2.3]) that for any regular holonomic $\mathsf{W}_{S}$-modules $\mathcal{M}_{1},\mathcal{M}_{2}$ the complexes of sheaves $\mathbb{R}\mathcal{H}om_{\mathsf{W}_{S}}(\mathcal{M}_{1},\mathcal{M}_{2})[\dim S/2]$ and $\mathcal{M}_{1}\otimes^{\mathbb{L}}_{\mathsf{W}_{S}}\mathcal{M}_{2}$ are perverse sheaves of $\mathbf{C}(\!(\hbar)\!)$-vector spaces.
-The main goal of this paper is to describe these perverse sheaves in terms of the derived geometry of the Lagrangian intersection $L\cap M$ of the supports of $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ when both $L$ and $M$ are smooth holomorphic Lagrangian submanifolds of $S$.
-Microlocal sheaves and Fukaya categories
-In this section we outline connections between the theory of deformation quantization modules and Fukaya categories, on the one hand, and microlocal sheaves, on the other hand, to place our results in a broader context.
-As the first goal let us describe the relevant category of deformation quantization modules associated to a holomorphic symplectic manifold $S$ (we review this construction in Section 4).
-Fix a (possibly singular) complex Lagrangian subvariety $L\subset S$.
-There is a canonical contactification $\rho\colon Y\rightarrow S$ of a neighborhood of $L$ with a lift of $L\subset S$ to a Legendrian subvariety $\Lambda\subset Y$ (see 1.12).
-As in [DK11] we define the abelian stack $\mathrm{RH}_{S,L}$ (i.e.
-a sheaf of $\mathbf{C}$-linear abelian categories) of regular holonomic DQ modules on $S$ supported on $L$ to be the stack of regular holonomic $\mathsf{E}_{Y}$-modules supported on $\Lambda$, i.e.
-$\mathrm{RH}_{S,L}=\rho_{*}\mathrm{Mod}_{\Lambda,\mathrm{rh}}(\mathsf{E}_{Y}).$
-Lagrangian subvarieties of $S$ form a filtered system and we define $\mathrm{RH}_{S}$ as the filtered colimit
-$\mathrm{RH}_{S}=\operatorname*{colim}_{L}\mathrm{RH}_{S,L}.$
-In particular, if $L\subset S$ is a smooth Lagrangian submanifold equipped with a square root line bundle $K_{L}^{1/2}$ (orientation data) and $\lambda\in\mathbf{C}/\mathbf{Z}$ we have the objects $\mathsf{M}^{\lambda}_{L}\in\mathrm{RH}_{S,L}$ mentioned before.
-Let us now describe a topological avatar of the stack $\mathrm{RH}_{S}$.
-Let $W$ be an exact holomorphic symplectic manifold.
-The authors of [NS20, Côt+22] have introduced a sheaf of stable $\infty$-categories $\mu\mathrm{sh}_{W,\mathbf{C}\text{-}c}$ on $W$ by globalizing the microlocal theory of sheaves of Kashiwara and Schapira [KS94].
-For a holomorphic contact manifold $Y$ one then defines the stack $\mathbb{P}\mu\mathrm{sh}_{Y,\mathbf{C}\text{-}c}=\gamma_{*}\mu\mathrm{sh}_{\tilde{Y},\mathbf{C}\text{-}c}$, where $\gamma\colon\tilde{Y}\rightarrow Y$ is the symplectization.
-One has
-$\mathbb{P}\mu\mathrm{sh}_{Y,\mathbf{C}\text{-}c}=\operatorname*{colim}_{\Lambda}\mathbb{P}\mu\mathrm{sh}_{\Lambda},$
-where $\mathbb{P}\mu\mathrm{sh}_{\Lambda}\subset\mathbb{P}\mu\mathrm{sh}_{Y,\mathbf{C}\text{-}c}$ is the subcategory of objects supported on a Legendrian subvariety $\Lambda\subset Y$.
-Moreover, the authors of [Côt+22] construct a $t$-structure on $\mathbb{P}\mu\mathrm{sh}_{Y,\mathbf{C}\text{-}c}$ (respectively, $\mathbb{P}\mu\mathrm{sh}_{\Lambda}$) whose heart is the stack $\mu\mathrm{Perv}_{Y}$ (respectively, $\mu\mathrm{Perv}_{Y,\Lambda}$) of microlocal perverse sheaves as in [And96, Was04].
-For $Y=\mathrm{P}^{*}X$ a projectivized cotangent bundle and $\Lambda\subset Y$ a fixed Legendrian subvariety Waschkies [Was05] has defined a microlocal Riemann–Hilbert correspondence which identifies
-$\mu\mathrm{Perv}_{Y}\cong\mathrm{Mod}_{\mathrm{rh}}(\mathsf{E}_{Y}),\qquad\mu\mathrm{Perv}_{Y,\Lambda}\cong\mathrm{Mod}_{\Lambda,\mathrm{rh}}(\mathsf{E}_{Y}),$
-which explains a connection between the stack $\mathrm{RH}_{S}$ and the stacks of microlocal sheaves.
-If the exact holomorphic symplectic manifold $\tilde{Y}$ is Weinstein, then Ganatra, Pardon and Shende [GPS18] identify $\mu\mathrm{sh}_{\tilde{Y}}$ with the ind-completion of the wrapped Fukaya category of $\tilde{Y}$.
-Thus, we have three independent ways to approach the stack $\mathrm{RH}_{S}$: in terms of deformation quantization modules, in terms of microlocal sheaf theory and in terms of Fukaya categories.
-D-critical loci and DQ modules
-The relevant derived structure on a Lagrangian intersection that we will need is the structure of a d-critical locus introduced in [Joy15].
-We recall the definition in Section 6.2, but let us simply mention that a d-critical locus is locally isomorphic to the critical locus $\mathrm{Crit}(f)$ of a holomorphic function $f\colon U\rightarrow\mathbf{C}$ on a complex manifold $U$ as a complex-analytic space.
-Any d-critical locus has an orientation $\mathbf{Z}/2$-gerbe whose trivialization is the choice of an orientation data.
-The key result of [Bra+15] is a construction of a canonical perverse sheaf $\phi_{X}$ (of $\mathbf{C}$-vector spaces), which we call the DT sheaf, on an oriented d-critical locus globalizing the sheaf of vanishing cycles $\phi_{f}$.
-(More canonically, we can think of $\phi_{X}$ as a sheaf twisted by the orientation gerbe similar to the fundamental class of a closed manifold being a homology class twisted by the orientation local system.)
-As explained in [Bus14], if $S$ is a holomorphic symplectic manifold and $L,M\subset S$ are holomorphic Lagrangian submanifolds, the Lagrangian intersection $L\cap M$ has a canonical structure of a d-critical locus.
-Moreover, if $L$ and $M$ are equipped with square roots $K_{L}^{1/2},K_{M}^{1/2}$ of the canonical bundles, then the d-critical locus $L\cap M$ is oriented; in particular, there is the DT sheaf $\phi_{L\cap M}$ on the Lagrangian intersection.
-We may then state the main result (7.1) as follows.
-Theorem A.
-For any $\lambda,\mu\in\mathbf{C}$ there is an isomorphism of $\mathbf{C}(\!(\hbar)\!)$-linear perverse sheaves
-$\mathsf{M}^{\lambda}_{L}\otimes^{\mathbb{L}}_{\mathsf{W}_{S}}\mathsf{M}^{\mu}_{M}\cong\phi_{L\cap M}\otimes_{\mathbf{C}}\mathbf{C}(\!(\hbar)\!).$
-When the Lagrangians $L,M$ intersect cleanly, the right-hand side reduces to an explicit $\mathbf{Z}/2$-torsor described in Section 7.4.
-This fact should be compared with a similar computation of Lagrangian Floer homology of clean intersections [Poź99, Sch16].
-In fact, we prove a strengthening of the above statement which explains the origin of the $\hbar$ parameter.
-For this, recall that the DT sheaf $\phi_{L\cap M}$ carries a canonical monodromy automorphism.
-The Riemann–Hilbert correspondence identifies perverse sheaves $P$ of $\mathbf{C}$-vector spaces with an automorphism and perverse sheaves $\mathrm{RH}^{-1}(P)$ of $\mathbf{C}(\!(\hbar)\!)$-vector spaces equipped with a connection in the $\hbar$-direction with regular singularities (see 2.11).
-We call perverse sheaves of $\mathbf{C}(\!(\hbar)\!)$-vector spaces equipped with connections in the $\hbar$-direction differential perverse sheaves.
-We also equip the derived tensor product $\mathsf{M}^{\lambda}_{L}\otimes^{\mathbb{L}}_{\mathsf{W}_{S}}\mathsf{M}^{\mu}_{M}$ with a natural connection in the $\hbar$-direction coming from the fact that the $\mathsf{W}_{S}$-modules $\mathsf{M}^{\lambda}_{L},\mathsf{M}^{\mu}_{M}$ locally come from $\mathsf{E}_{Y}$-modules on a contactification.
-The stronger statement of the main result (7.1) is then as follows.
-Theorem B.
-For any $\lambda,\mu\in\mathbf{C}$ there is an isomorphism of differential perverse sheaves
-$\mathsf{M}^{\lambda}_{L}\otimes^{\mathbb{L}}_{\mathsf{W}_{S}}\mathsf{M}^{\mu}_{M}\cong\mathrm{RH}^{-1}(\phi_{L\cap M}\otimes\mathbf{C}_{\mu+\lambda}),$
-where $\mathbf{C}_{\mu+\lambda}$ is the trivial one-dimensional vector space equipped with the automorphism $\exp(2\pi i(\mu+\lambda))$.
-One motivation for this statement is a proposal [Bra+15, Remark 6.15] for a Fukaya-like category associated to a holomorphic symplectic manifold $S$:
+Suppose we are given a holomorphic symplectic manifold $S$ and a pair of Lagrangian submanifolds $L,M$ each equipped with a choice of square root of the canonical bundle.
+Associated to this data we have the following two perverse sheaves on the intersection $L\cap M$:
+  • By [Bussi, BBDJS], the intersection $L\cap M$ admits the structure of a $d$-critical locus, and thus carries a canonical $\mathbf{C}$-linear perverse sheaf $\phi_{L\cap M}$ (called the DT sheaf), equipped with a canonical automorphism $T_{L\cap M}$ (called the monodromy), obtained by gluing together vanishing cycle sheaves on local charts.
+  • By [PoleselloSchapira, DAgnoloSchapira, KashiwaraSchapira], there is a canonical $\mathbf{C}(\!(\hbar)\!)$-algebroid $\widehat{\mathsf{W}}_{S}$ together with modules $\mathsf{M}_{L}$, $\mathsf{M}_{M}$, and
+    $\mathsf{M}_{L}\otimes^{L}_{\widehat{\mathsf{W}}_{S}}\mathsf{M}_{M}\cong\mathbb{R}\mathcal{H}om_{\widehat{\mathsf{W}}_{S}}(\mathsf{M}_{L},\mathsf{M}_{M})[\dim S]$
+    is a $\mathbf{C}(\!(\hbar)\!)$-linear perverse sheaf supported on $L\cap M$.
+Our primary goal of this paper is to compare these two perverse sheaves.
+To put the two sides on the same footing, we use a version of the Riemann-Hilbert correspondence to reinterpret the pair $(\phi_{L\cap M},T_{L\cap M})$ as a differential perverse sheaf $\mathrm{RH}^{-1}(\phi_{L\cap M},T_{L\cap M})$ (informally, a $\mathbf{C}(\!(\hbar)\!)$-linear perverse sheaf equipped with a regular $\hbar$-connection).
+Theorem A (7.1).
+There is a natural differential structure on the perverse sheaf $\mathsf{M}_{L}\otimes^{L}_{\widehat{\mathsf{W}}_{S}}\mathsf{M}_{M}$, and an isomorphism of differential perverse sheaves
+$\mathsf{M}_{L}\otimes^{\mathbb{L}}_{\widehat{\mathsf{W}}_{S}}\mathsf{M}_{M}\cong\mathrm{RH}^{-1}(\phi_{L\cap M},T_{L\cap M}).$
+It follows from A that there is an isomorphism of $\mathbf{C}(\!(\hbar)\!)$-linear perverse sheaves
+$\mathbb{R}\mathcal{H}om_{\widehat{\mathsf{W}}_{S}}(\mathsf{M}_{L},\mathsf{M}_{M})[\dim S]\cong\phi_{L\cap M}\otimes_{\mathbf{C}}\mathbf{C}(\!(\hbar)\!)$
+While such a comparison has been predicted for some time in the community (see [Bussi, Section 4]), this is the first proof of this result known to the authors and also the first statement that properly accounts for the monodromy and discrepancy of coefficient fields.
+As we explain below, the extra structure corresponding to the monodromy appears naturally in our approach and is a necessary ingredient in our proof of any such comparison.
+A has a number of consequences and potential applications, which we explore in the rest of this introduction: to skein modules, microlocal sheaves and Fukaya categories.
+A holomorphic Fukaya category
+One source of motivation for our work is a proposal [BBDJS, Remark 6.15] for a Fukaya-like category associated to a holomorphic symplectic manifold $S$:
   • Holomorphic Lagrangian submanifolds $L\subset S$ equipped with square roots $K_{L}^{1/2}$ should define objects.
   • The graded Hom space between objects associated to oriented Lagrangians $L,M$ is given by the hypercohomology $\mathrm{H}^{\bullet-\dim S/2}(L\cap M,\phi_{L\cap M})$.
-  • Composition of Homs is given by a conjectural functoriality of the DT sheaf, see [JS19, Conjecture 1.1].
-We refer to [SV19] and [DR22, Section B.2] for a connection between this category and Lagrangian Floer homology.
-7.1 shows that the stack $\mathrm{RH}_{S}$ provides such a category.
-Indeed, there is an isomorphism $\mathcal{M}_{1}\otimes^{\mathbb{L}}_{\mathsf{W}_{S}}\mathcal{M}_{2}\cong\mathbb{R}\mathcal{H}om_{\mathsf{W}_{S}}(\mathbb{D}(\mathcal{M}_{1}),\mathcal{M}_{2})[\dim S/2]$ valid for any holonomic $\mathsf{W}_{S}$-modules $\mathcal{M}_{1},\mathcal{M}_{2}$, where the duality functor is $\mathbb{D}(\mathcal{M}_{1})=\mathbb{R}\mathcal{H}om(\mathcal{M}_{1},\mathsf{W}_{S})[\dim S/2]$.
-Moreover, the modules $\mathsf{M}^{\lambda}_{L}$ are self-dual.
-The format of the statement of 7.1 is inspired by a related result [SS14] of Sabbah and Saito (see also [Sab10]) relating the twisted de Rham complex
-$\mathrm{DR}_{f}=(\Omega^{\bullet}_{U}(\!(\hbar)\!),d+\hbar^{-1}df)|_{f^{-1}(0)}$
-and the sheaf of vanishing cycles $\phi_{f}$.
-(Note that we restrict to the zero set of $f$ to simplify the presentation.) Namely, there is a natural connection on $\mathrm{DR}_{f}$ in the $\hbar$ direction given by $\partial_{\hbar}-\hbar^{-2}f$, which equips $\mathrm{DR}_{f}$ with the structure of a differential perverse sheaf.
-Sabbah and Saito [SS14] identify the cohomology sheaves $\mathcal{H}^{k}\mathrm{DR}_{f}$ and $\mathcal{H}^{k}\mathrm{RH}^{-1}(\phi_{f})$.
-We show that their isomorphism lifts to an isomorphism of differential perverse sheaves (see 5.11).
-Theorem C.
-Let $f\colon U\rightarrow\mathbf{C}$ be a holomorphic function on a complex manifold $U$. Then there is an isomorphism of differential perverse sheaves
-$(\mathrm{DR}_{f},\hbar\partial_{\hbar}-\hbar^{-1}f)\cong\mathrm{RH}^{-1}(\phi_{f}).$
-A closely related result has recently been shown by Schefers in [Sch23, Proposition 9.20] using different methods.
+  • Composition of Homs is given by a conjectural functoriality of the DT sheaf, see [JoyceSafronov, Conjecture 1.1].
+Our results provide the first construction of such a category (at least over $\mathbf{C}$).
+Let $\mathrm{Rep}(\mathbf{Z})$ be the $\infty$-category of complexes of $\mathbf{C}$-vector spaces equipped with an automorphism.
+Theorem B (8.5).
+Let $S$ be a holomorphic symplectic manifold.
+There is a sheaf $\mathbf{D}\mathsf{RH}_{S}$ of $\mathrm{Rep}(\mathbf{Z})$-linear $\infty$-categories on $S$, equipped with a $t$-structure, which satisfies the following properties:
+  • For each Lagrangian submanifold $L$ of $S$ equipped with orientation data, there is a canonical object $\mathsf{M}_{L}$.
+  • Given a pair of oriented Lagrangian submanifolds $L,M$, there is an isomorphism
+    $\underline{\mathcal{H}om}_{\mathbf{D}\mathsf{RH}_{S}}(\mathsf{M}_{M},\mathsf{M}_{L})[\dim S/2]\cong(\phi_{L\cap M},T_{L\cap M})$
+    of perverse sheaves equipped with an automorphism.
 Skein modules and perverse sheaves
-Another motivation for this work is a conjectural relationship between skein modules and the DT sheaf formulated in [GJS23].
+Another motivation for this work is a conjectural relationship between skein modules and the DT sheaf formulated in [GJS].
 Let us recall the statement.
-Let $(M,x)$ be a connected closed oriented 3-manifold and $G$ a reductive algebraic group equipped with some extra data described in [GJS23, Section 1.3].
+Let $M$ be a connected closed oriented 3-manifold with a basepoint $x\in M$ and $G$ a reductive algebraic group equipped with some extra data described in [GJS, Section 1.3].
 Then one can define the skein module $\mathrm{Sk}^{\mathrm{gen}}_{G}(M)$ of $M$ for generic quantum parameters, which is a $\mathbf{C}(q^{1/d})$-vector space for some integer $d$.
@@
 carries a natural structure of a d-critical locus.
-In the case $G=\mathrm{SL}_{2}$ Abouzaid and Manolescu [AM20] denote the hypercohomology of the DT sheaf $\phi_{R_{G}(M)}$ on $R_{G}(M)$ by $\mathrm{HP}^{\bullet}_{\sharp}(M)$, the framed complexified instanton Floer homology of $M$.
+In the case $G=\mathrm{SL}_{2}$ Abouzaid and Manolescu [AbouzaidManolescu] denote the hypercohomology of the DT sheaf $\phi_{R_{G}(M)}$ on $R_{G}(M)$ by $\mathrm{HP}^{\bullet}_{\sharp}(M)$, the framed complexified instanton Floer homology of $M$.
 The conjectural relationship between the two objects is as follows.
-Conjecture D.
+Conjecture C.
 There is an isomorphism of $\mathbf{C}(q^{1/d})$-vector spaces
 $\mathrm{Sk}^{\mathrm{gen}}_{G}(M)\cong\mathrm{H}^{0}(R_{G}(M),\phi_{R_{G}(M)})\otimes_{\mathbf{C}}\mathbf{C}(q^{1/d}).$
-We refer to [DKS23] for some checks of this conjecture for the Kauffman bracket skein module.
+We refer to [DKS] for some checks of this conjecture for the Kauffman bracket skein module.
 The above conjecture is related to A as follows.
@@
   • The representation variety $R_{G}(M)$ becomes an intersection $L_{1}\cap L_{2}$ of two complex algebraic Lagrangian submanifolds $L_{1},L_{2}$ (representation varieties of the handlebodies $N_{1},N_{2}$) in a complex symplectic variety $S$ (an open subset of the representation variety of the open surface $\Sigma^{*}=\Sigma\setminus D$).
-We refer to [GJS23, Section 5.3] for details.
+We refer to [GJS, Section 5.3] for details.
   • The generic skein module $\mathrm{Sk}^{\mathrm{gen}}_{G}(M)\otimes_{\mathbf{C}(q^{1/d})}\mathbf{C}(\!(\hbar)\!)$ defined with respect to the formal variable $\hbar$ with $q=\exp(\hbar)$ is a relative tensor product
@@
     of algebraic deformation quantization modules $\mathrm{Sk}^{\mathrm{int}}_{G}(N_{i})$ quantizing $L_{i}\subset S$ over the deformation quantization algebra $\mathrm{SkAlg}^{\mathrm{int}}_{G}(\Sigma^{*})$ providing an algebraic deformation quantization of the symplectic variety $S$.
-Thus, in order to prove D using A, one must in particular relate the tensor product of the algebraic deformation quantization provided by skein theory to the corresponding tensor product of the canonical analytic deformation quantization modules.
+Thus, in order to prove C using A, one must in particular relate the tensor product of the algebraic deformation quantization modules provided by skein theory to the corresponding tensor product of the canonical analytic deformation quantization modules.
 We plan to address this problem in future work.
+Vanishing cycles and the twisted de Rham complex
+Consider the case when $S=\mathrm{T}^{\ast}X$ is the cotangent bundle of a complex manifold $X$, $L=\Gamma_{0}$ is the zero-section and $M=\Gamma_{df}$ is the graph of $df$, where $f\colon X\to\mathbf{C}$ is a holomorphic function.
+This can be considered as a local model for the setup of A, in the sense that every holomorphic Lagrangian intersection is locally isomorphic to such an intersection.
+In this case, the DT sheaf $\phi_{\Gamma_{0}\cap\Gamma_{df}}$ is identified with the sheaf of vanishing cycles $\phi_{f}$.
+The format of the statement of A is inspired by a related result [SabbahSaito] of Sabbah and Saito (see also [Sabbah]) relating the twisted de Rham complex
+$\mathrm{DR}_{f}=(\Omega^{\bullet}_{U}(\!(\hbar)\!),d+\hbar^{-1}df)|_{f^{-1}(0)}$
+and the sheaf of vanishing cycles $\phi_{f}$.
+(Note that we restrict to the zero set of $f$ to simplify the presentation.) Namely, there is a natural connection on $\mathrm{DR}_{f}$ in the $\hbar$ direction given by $\partial_{\hbar}-\hbar^{-2}f$, which equips $\mathrm{DR}_{f}$ with the structure of a differential perverse sheaf.
+Sabbah and Saito [SabbahSaito] identify the cohomology sheaves $\mathcal{H}^{k}\mathrm{DR}_{f}$ and $\mathcal{H}^{k}\mathrm{RH}^{-1}(\phi_{f})$.
+We show that their isomorphism lifts to an isomorphism of differential perverse sheaves.
+Theorem D (5.11).
+Let $f\colon U\rightarrow\mathbf{C}$ be a holomorphic function on a complex manifold $U$. Then there is an isomorphism of differential perverse sheaves
+$(\mathrm{DR}_{f},\hbar\partial_{\hbar}-\hbar^{-1}f)\cong\mathrm{RH}^{-1}(\phi_{f}).$
+A closely related result has recently been shown by Schefers in [Schefers, Proposition 9.20] using different methods.
+Contactification and symplectization
+The results and techniques in this paper are rooted in the tight interplay between symplectic and contact geometry in the holomorphic setting.
+Understanding this interplay also leads to a conceptual explanation for the structure appearing in the holomorphic Fukaya category of B.
+Given a pair $(S,L)$ of a symplectic manifold together with a Lagrangian subvariety (for example, a finite union of Lagrangian submanifolds), a contactification of $(S,L)$ is a pair $(Y,\rho,\Lambda)$ where $\rho:Y\to S$ is a principal $\mathbf{C}$-bundle with connection whose curvature is the symplectic form, $\Lambda$ is a Legendrian subvariety of $Y$ (with respect to the canonical contact structure), and $\rho|_{\Lambda}$ is a homeomorphism onto $L$, analytic on the smooth part of $\Lambda$.
+Given an arbitrary contact manifold $Y$, there is a canonical $\mathbf{C}^{\times}$-principal bundle $\gamma:\widetilde{Y}\to Y$ called the symplectization.
+In the case when $\rho:Y\to S$ is the contactification of a symplectic manifold $S$, the symplectization is canonically trivial: $\widetilde{Y}=Y\times\mathbf{C}^{\times}$.
+Moreover, if $L$ is a Lagrangian subvariety of $S$, and $(Y,\Lambda)$ a contactification of $(S,L)$, then $\widetilde{\Lambda}=\gamma^{-1}(\Lambda)\cong\Lambda\times\mathbf{C}^{\times}$ is an exact Lagrangian subvariety in $\widetilde{Y}$.
+Note that in our main results we do not assume that our symplectic manifold $S$ is exact.
+In particular, there may not exist a global contactification $\rho:Y\to S$.
+Nevertheless, we have the following fundamental result (see [DAgnoloKashiwara, Proposition 2.5.3]).
+Proposition (Unique Contactification Lemma, 1.12).
+Any pair $(S,L)$ of a symplectic manifold and Lagrangian subvariety admits a contactification $(Y,\rho,\Lambda)$ over a neighborhood of $L$ in $S$.
+Moreover, this data is unique up to a suitable notion of isomorphism.
+An important consequence of this result is that, given an arbitrary Lagrangian intersection $L\cap M$ in a symplectic manifold $S$, there is a canonical Legendrian intersection $\Lambda_{L}\cap\Lambda_{M}$ in a contactification $Y$ of $S$ (indeed, consider the unique contactification of $(S,L\cup M)$).
+In particular, $(L\cap M)\times\mathbf{C}^{\times}$ is homeomorphic to the intersection $\widetilde{\Lambda}_{L}\cap\widetilde{\Lambda}_{M}$ of $\mathbf{C}^{\times}$-invariant (exact) Lagrangians in the symplectization $\widetilde{Y}$.
+As explained in Section 8, the various perverse sheaves associated to the Lagrangian intersection $L\cap M$ in $S$ can be realized naturally as perverse sheaves on $\widetilde{\Lambda}_{L}\cap\widetilde{\Lambda}_{M}\cong(L\cap M)\times\mathbf{C}^{\times}$ which are lisse along the $\mathbf{C}^{\times}$-direction.
+We are now in a position to recall the construction, due to D’Agnolo–Kashiwara [DAgnoloKashiwara], of the category $\mathsf{RH}_{S}$ on an arbitrary symplectic manifold $S$.
+Suppose $L\subseteq S$ is a Lagrangian subvariety and let $(Y,\Lambda)$ be a contactification in a neighborhood of $L$ (which exists by the unique contactification lemma 1.12).
+Kashiwara [KashiwaraContact] has defined a certain filtered $\mathbf{C}$-algebroid $\mathsf{E}_{Y}$ on a general contact manifold $Y$, locally modelled on the algebra of microdifferential operators $\mathcal{E}_{X}$ on the projectivized cotangent bundle $\mathrm{P}^{\ast}X$ of a complex manifold $X$.
+We denote by $\mathsf{RH}_{S,L}$ the category of regular holonomic $\mathsf{E}_{Y}$-modules along $\Lambda$ (understood as a sheaf of categories on $L$ by pushing forward along the homeomorphism $\rho|_{\Lambda}$).
+We then define $\mathsf{RH}_{S}$ by taking the filtered colimit of $\mathsf{RH}_{S,L}$ along the diagram of all Lagrangian subvarieties $L$ in $S$.
+The uniqueness part of the unique contactification lemma tells us that this construction is well-defined.
+Microlocal sheaves
+For an exact holomorphic symplectic manifold $W$ the authors of [NadlerShende, CKNS1] have introduced a sheaf of stable $\infty$-categories $\mu\mathrm{sh}_{W}$ on $W$ of microsheaves by globalizing the microlocal theory of sheaves of Kashiwara and Schapira [KashiwaraSchapiraSheaves].
+Moreover, for a holomorphic contact manifold $Y$ the authors of [CKNS1] have defined a sheaf of abelian categories $\mathbb{P}\mathrm{erv}_{Y}\subset\gamma_{*}\mu\mathrm{sh}_{\tilde{Y}}$ of perverse microsheaves following [Andronikof2, Waschkies1].
+Perverse microsheaves relate to deformation quantization modules via a microlocal Riemann–Hilbert correspondence [Andronikof3, Waschkies2, CKNS2]: namely, there is an equivalence of sheaves of abelian categories
+$\mu\mathrm{RH}\colon\mathbb{P}\mathrm{erv}_{Y}\xrightarrow{\sim}\mathrm{Mod}_{\mathrm{rh}}(\mathsf{E}_{Y}).$
+Given a holomorphic Legendrian submanifold $\Lambda\subset Y$ equipped with a square root line bundle $K_{\Lambda}^{1/2}$ there is a perverse microsheaf $\mathcal{M}_{\Lambda}\in\mathbb{P}\mathrm{erv}_{Y}$ supported on $\Lambda$ which provides a sheaf quantization of $\Lambda$ in the sense that its microlocalization to $\Lambda$ is a trivial (twisted) local system of rank $1$.
+Equivalently, $\mu\mathrm{RH}(\mathcal{M}_{\Lambda})=\mathsf{M}_{\Lambda}$ is a simple $\mathsf{E}_{Y}$-module along $\Lambda$.
+B allows us to compute Hom complexes between two such sheaf quantizations.
+Theorem E (8.8).
+Let $Y$ be a holomorphic contact manifold and $\Lambda_{1},\Lambda_{2}\subset Y$ two holomorphic Legendrian submanifolds equipped with square root line bundles $K_{\Lambda_{1}}^{1/2},K_{\Lambda_{2}}^{1/2}$.
+Let $\tilde{\Lambda}_{1},\tilde{\Lambda}_{2}$ be the corresponding homogeneous Lagrangian submanifolds of the symplectization $\tilde{Y}\rightarrow Y$.
+There is an isomorphism of complexes of sheaves
+$\mathcal{H}om_{\mu\mathrm{sh}_{\tilde{Y}}}(\mathcal{M}_{\Lambda_{1}},\mathcal{M}_{\Lambda_{2}})[\dim\tilde{Y}/2]\cong\phi_{\tilde{\Lambda}_{1}\cap\tilde{\Lambda}_{2}}$
+on $\tilde{Y}$.
+Fukaya categories
+Consider, as before, a holomorphic contact manifold $Y$ with a symplectization $\tilde{Y}\rightarrow Y$.
+In particular, $\tilde{Y}$ is a real exact symplectic manifold.
+If we assume that it is Weinstein, then Ganatra, Pardon and Shende [GPS] identify $\mu\mathrm{sh}_{\tilde{Y}}$ with the ind-completion of the wrapped Fukaya category of $\tilde{Y}$.
+In particular, in this case the left-hand side of E may be rewritten in terms of FLoer homology.
+Via the microlocal Riemann–Hilbert correspondence, this relates the sheaf of $\infty$-categories $\mathbf{D}\mathsf{RH}_{S}$ on a (non-exact) holomorphic symplectic manifold $S$ constructed in B to Floer theory in the symplectization $\tilde{Y}$ of a local contactification $Y\rightarrow S$.
+The relationship between the sheaf of $\infty$-categories $\mathbf{D}\mathsf{RH}_{S}$ and Lagrangian Floer theory in $S$ is more subtle, see the discussion in [DoanRezchikov, Section B.2] and [KontsevichSoibelman, Section 6.3].
+In this direction, let us mention the locality results for holomorphic Floer theory established by Solomon and Verbitsky in [SolomonVerbitsky] as well as a relationship between Lagrangian Floer theory in $S$ and that on the symplectization of a real contactification established by Kuwagaki, Petr and Shende in [KuwagakiPetrShende].
 Conventions
-  • For a complex manifold $X$ and a ring $R$ we denote by $\mathbf{D}(X;R)$ the derived category of sheaves of complexes of $R$-modules.
-  • $\mathbf{D}_{c}(X;R)\subset\mathbf{D}(X;R)$ denotes the derived category of constructible sheaves of $R$-modules on $X$.
+  • For a complex manifold $X$ and a ring $R$ we denote by $\mathbf{D}(X;R)$ the derived $\infty$-category of sheaves of complexes of $R$-modules.
+  • $\mathbf{D}_{c}(X;R)\subset\mathbf{D}(X;R)$ denotes the subcategory of constructible complexes of $R$-modules on $X$.
   • $\mathrm{Perv}(X;R)\subset\mathbf{D}_{c}(X;R)$ denotes the subcategory of perverse sheaves.
@@
   • We consider the subset $G=\{-1<\operatorname{Re}\lambda\leq 0\}\subset\mathbf{C}$ which defines a splitting of the projection $\mathbf{C}\rightarrow\mathbf{C}/\mathbf{Z}$.
+  • A stack of categories is the same as a sheaf of categories. When there is no confusion, we will denote its global sections by the same letter.
 Acknowledgements
-We would like to thank Claude Sabbah, Pierre Schapira and Nick Sheridan for useful conversations.
+We would like to thank Claude Sabbah, Pierre Schapira and Nick Sheridan for useful conversations. S.G. was partially supported by NSF grant DMS-2202363.
 1. Background on symplectic geometry
@@
 Remark 1.5.
-It is not in general possible to find a contactification with $\rho\colon\Lambda\rightarrow L$ an analytic isomorphism everywhere, see [DK11, Example 2.5.4].
+It is not in general possible to find a contactification with $\rho\colon\Lambda\rightarrow L$ an analytic isomorphism everywhere, see [DAgnoloKashiwara, Example 2.5.4].
 Given a symplectomorphism $\phi\colon S_{1}\rightarrow S_{2}$ and contactifications $Y_{1}\rightarrow S_{1}$ and $Y_{2}\rightarrow S_{2}$, an isomorphism of contactifications is a diffeomorphism $\psi\colon Y_{1}\rightarrow Y_{2}$ which is $\mathbf{C}$-equivariant and which intertwines the connection one-forms $\alpha_{i}$.
@@
 Proof.
-The existence of such a contactification is shown in [DK11, Proposition 2.5.3].
+The existence of such a contactification is shown in [DAgnoloKashiwara, Proposition 2.5.3].
 Let us now show that any two such contactifications are uniquely isomorphic.
-By [Gie64] we may choose a triangulation of $S$ with $L\subset S$ as a simplicial subcomplex.
+By [Giesecke] we may choose a triangulation of $S$ with $L\subset S$ as a simplicial subcomplex.
 Therefore, there is a neighborhood $U$ of $L$ which deformation retracts to $L$.
@@
 Note that $d(g_{i}-f_{i})|_{L_{\mathrm{reg}}}=0$, i.e.
-$g_{i}-f_{i}$ is locally constant on $L_{\mathrm{reg}}\cap U_{i}$ and hence, by density, on $L\cap U_{i}$ (see [DK11, Sublemma 2.5.2]).
+$g_{i}-f_{i}$ is locally constant on $L_{\mathrm{reg}}\cap U_{i}$ and hence, by density, on $L\cap U_{i}$ (see [DAgnoloKashiwara, Sublemma 2.5.2]).
 Thus, under the restriction map $\mathrm{H}^{1}(U;\mathbf{C})\rightarrow\mathrm{H}^{1}(L;\mathbf{C})$ the class $[g_{i}-g_{j}-\varphi_{ij}]$ restricts to zero.
@@
 $q(L_{1},L_{2},L_{3})\colon L_{1}\rightarrow k$
-which we call the Maslov quadratic form. For $k=\mathbf{R}$ its signature is the Maslov index $\tau(L_{1},L_{2},L_{3})$ of the triple of Lagrangians as defined in [KS94, Chapter A.3].
+which we call the Maslov quadratic form.
+For $k=\mathbf{R}$ its signature is the Maslov index $\tau(L_{1},L_{2},L_{3})$ of the triple of Lagrangians as defined in [KashiwaraSchapiraSheaves, Chapter A.3].
 Example 1.15.
@@
 Proof.
-For the statement on the level of symplectic manifolds, see [Wei71, Theorem 7.1].
+For the statement on the level of symplectic manifolds, see [WeinsteinLagrangian, Theorem 7.1].
 By Example 1.7 $j^{1}0\subset\mathrm{J}^{1}L$ is a contactification of the zero section $L\subset\mathrm{T}^{*}L$ and by 1.12 contactifications of Lagrangians are unique.
@@
 Remark 1.28.
-It is not in general possible to arrange $f$ to be zero on $L\cap M$ when this intersection is non-reduced, see [Joy15, Example 2.13].
+It is not in general possible to arrange $f$ to be zero on $L\cap M$ when this intersection is non-reduced, see [JoycedCrit, Example 2.13].
 Finally, we will need to describe the change of local models under a change of polarizations.
@@
 By construction the critical loci of $f$ on $L$ and $h$ on $L\times M$ are both $L\cap M$.
-So, by [Joy15, Proposition 2.25] there is a canonical quadratic form $q_{\Xi}$ on the normal bundle $N_{\Xi}|_{L\cap M}$ of the embedding $\Xi\colon L\rightarrow L\times M$.
+So, by [JoycedCrit, Proposition 2.25] there is a canonical quadratic form $q_{\Xi}$ on the normal bundle $N_{\Xi}|_{L\cap M}$ of the embedding $\Xi\colon L\rightarrow L\times M$.
 It has the following alternative description.
@@
 For brevity we call such a module a regular holonomic $\mathcal{D}_{X}[\theta]$-module.
-The following statement is standard.
+The following statement is standard (see e.g.
+[Brylinski, Proposition 7.12]).
 Proposition 2.3.
 A $D$-module $\mathcal{M}\in\mathrm{Mod}_{\mathrm{rh}}(\mathcal{D}_{Y})$ is monodromic if, and only if, either of the following conditions holds:
-  • $\mathcal{M}$ is lisse along the orbits of the $\mathbf{C}^{\times}$-action.
+  • $\mathcal{M}$ is locally constant along the orbits of the $\mathbf{C}^{\times}$-action.
   • The natural map
@@
   • The stack of perverse sheaves $\mathcal{F}$ on $X$ equipped with an endomorphism $\theta\colon\mathcal{F}\rightarrow\mathcal{F}$ whose eigenvalues lie in $G$.
-  • The stack of perverse sheaves on $X\times\mathbf{C}^{\times}$ lisse along $\mathbf{C}^{\times}$.
+  • The stack of perverse sheaves on $X\times\mathbf{C}^{\times}$, which are locally constant along the $\mathbf{C}^{\times}$ orbits.
 Given a regular holonomic left $\mathcal{D}_{X}$-module $\mathcal{L}$ and a regular holonomic right $\mathcal{D}_{X}$-module $\mathcal{M}$, the complex of sheaves $\mathcal{M}\otimes^{\mathbb{L}}_{\mathcal{D}_{X}}\mathcal{L}$ is a perverse sheaf on $X$.
@@
     $\mathcal{M}^{\prime}\otimes^{\mathbb{L}}_{\mathcal{D}_{Y}}\mathcal{L}^{\prime}$
-    on $Y$ is lisse along $\mathbf{C}^{\times}$.
+    on $Y$ is locally constant along the $\mathbf{C}^{\times}$ orbits.
   • There is an isomorphism of perverse sheaves
@@
 Then it is equal to the monodromy of the sheaf on the right along the $\mathbf{C}^{\times}$ direction.
-We will also consider perverse sheaves of $\mathbf{C}[\![\hbar]\!]$-modules as in [DGS11, Section 5].
+We will also consider perverse sheaves of $\mathbf{C}[\![\hbar]\!]$-modules as in [DAgnoloGuillermouSchapira, Section 5].
 We say a perverse sheaf $\mathcal{L}\in\mathrm{Perv}(X;\mathbf{C}[\![\hbar]\!])$ is torsion-free if the map $\hbar\colon\mathcal{L}\rightarrow\mathcal{L}$ is a monomorphism in the abelian category $\mathrm{Perv}(X;\mathbf{C}[\![\hbar]\!])$.
@@
 If $\mathcal{M}\in\mathrm{Perv}_{\mathrm{diff}}(X)$ has a lattice, we say that it has regular singularities.
-2.3. The Riemann-Hilbert functor
+We will need the following Yoneda lemma for monodromic perverse sheaves. Consider a $\mathbf{C}^{\times}$-torsor $\gamma\colon\tilde{Y}\rightarrow Y$ and let
+$\mathrm{Perv}_{\mathbf{C}^{\times}}(\tilde{Y})\subset\mathrm{Perv}(\tilde{Y})$
+be the full substack of perverse sheaves, which are locally constant along the fibers of $\gamma$, and
+$\mathrm{LocSys}(\tilde{Y})\subset\mathrm{Perv}(\tilde{Y})$
+the full substack of local systems.
+There is a natural action of the symmetric monoidal stack $\mathrm{LocSys}(Y)$ on $\mathrm{LocSys}(\tilde{Y})$ (via $\gamma^{*}$) and $\mathrm{Perv}(Y)$.
+For $\mathcal{F}\in\mathrm{Perv}_{\mathbf{C}^{\times}}(\tilde{Y})$ the derived pushforward $\gamma_{*}\mathcal{F}$ has cohomology only perverse degrees $-1$ and 00.
+Lemma 2.7.
+Consider the functor
+$\gamma_{*}\mathrm{Perv}_{\mathbf{C}^{\times}}(\tilde{Y})\longrightarrow\mathrm{Fun}_{\mathrm{LocSys}(Y)}(\gamma_{*}\mathrm{LocSys}(\tilde{Y}),\mathrm{Perv}(Y))$
+of stacks on $Y$ given by sending $\mathcal{F}\in\mathrm{Perv}_{\mathbf{C}^{\times}}(\tilde{Y})$ to the $\mathrm{LocSys}(Y)$-linear functor $V\in\mathrm{LocSys}(\tilde{Y})\mapsto{}^{p}\mathcal{H}^{-1}(\gamma_{*}(V\otimes\mathcal{F}))$.
+It induces an isomorphism after applying $\mathbb{R}\mathcal{H}om(-,-)$: i.e.
+for every $\mathcal{F}_{1},\mathcal{F}_{2}\in\gamma_{*}\mathrm{Perv}_{\mathbf{C}^{\times}}(\tilde{Y})$ the natural morphism
+$\mathbb{R}\mathcal{H}om(\mathcal{F}_{1},\mathcal{F}_{2})\longrightarrow\mathbb{R}\mathcal{H}om({}^{p}\mathcal{H}^{-1}(\gamma_{*}((-)\otimes\mathcal{F}_{1})),{}^{p}\mathcal{H}^{-1}(\gamma_{*}((-)\otimes\mathcal{F}_{2})))$
+is an isomorphism of complexes of sheaves on $Y$.
+Proof.
+The question is local on $Y$, so we may assume that $\tilde{Y}=Y\times\mathbf{C}^{\times}\rightarrow Y$ is a trivial $\mathbf{C}^{\times}$-torsor and $Y$ is contractible.
+By 2.1 both $\gamma_{*}\mathrm{Perv}_{\mathbf{C}^{\times}}(\tilde{Y})\cong\mathrm{Perv}_{\mathrm{aut}}(Y)$ and $\gamma_{*}\mathrm{LocSys}(\tilde{Y})$ split into blocks according to the generalized eigenvalue of the monodromy operator.
+This allows us to reduce the statement to a similar statement for the substack of perverse sheaves $\mathrm{Perv}_{\mathrm{aut}}^{0}(Y)\subset\mathrm{Perv}_{\mathrm{aut}}(Y)$ consisting of pairs $(\mathcal{F},T)\in\mathrm{Perv}_{\mathrm{aut}}(Y)$, where the automorphism $T\colon\mathcal{F}\rightarrow\mathcal{F}$ is unipotent, i.e.
+$(T-\mathrm{id})^{N}=0\colon\mathcal{F}\rightarrow\mathcal{F}$ for some $N$.
+Moreover, since $Y$ is assumed to be contractible, $\gamma_{*}\mathrm{LocSys}(\tilde{Y})$ is given by the stack of $\mathbf{C}[T^{\pm 1}]$-modules.
+For a local system $V$ of $\mathbf{C}[T^{\pm 1}]$-modules and $\mathcal{F}\in\mathrm{Perv}_{\mathrm{aut}}^{0}(Y)$ we have
+${}^{p}\mathcal{H}^{-1}(\gamma_{*}(V\otimes\mathcal{F}))\cong(V\otimes\mathcal{F})^{T},$
+where on the right we consider $T$-invariants. This sheaf vanishes if the $T$-action on $V$ is not unipotent, so we will further restrict to such modules.
+So, for $\mathcal{F}_{1},\mathcal{F}_{2}\in\mathrm{Perv}_{\mathrm{aut}}^{0}(Y)$ we have to show that the natural morphism
+$\mathbb{R}\mathcal{H}om_{\mathbf{C}[T^{\pm 1}]}(\mathcal{F}_{1},\mathcal{F}_{2})\longrightarrow\mathbb{R}\mathcal{H}om_{\mathbf{C}}(((-)\otimes\mathcal{F}_{1})^{T},((-)\otimes\mathcal{F}_{2})^{T})$
+is an isomorphism. Fix a natural number $N$, so that $(T-1)^{N}$ annihilates both $\mathcal{F}_{1}$ and $\mathcal{F}_{2}$. In this case the left-hand side becomes isomorphic to
+$\mathbb{R}\mathcal{H}om_{\mathbf{C}[T^{\pm 1}]/(T-1)^{N}}(\mathcal{F}_{1},\mathcal{F}_{2})$
+and the right-hand side becomes isomorphic to
+$\mathbb{R}\mathcal{H}om_{\mathbf{C}[T^{\pm 1}]/(T-1)^{N}}((\mathbf{C}[T^{\pm 1}]/(T-1)^{N}\otimes\mathcal{F}_{1})^{T},(\mathbf{C}[T^{\pm 1}]/(T-1)^{N}\otimes\mathcal{F}_{2})^{T}),$
+which is again isomorphic to $\mathbb{R}\mathcal{H}om_{\mathbf{C}[T^{\pm 1}]/(T-1)^{N}}(\mathcal{F}_{1},\mathcal{F}_{2})$ by an explicit computation of the invariants. ∎
+2.3. The Riemann–Hilbert functor
 Our next goal is to define an inverse Riemann–Hilbert functor
-$\widehat{\mathrm{RH}}^{-1}\colon\mathrm{Perv}_{\mathrm{aut}}(X)\to\mathrm{Perv}_{\mathrm{diff}}(X)$
+$\mathrm{RH}^{-1}\colon\mathrm{Perv}_{\mathrm{aut}}(X)\to\mathrm{Perv}_{\mathrm{diff}}(X)$
 relating differential perverse sheaves and perverse sheaves equipped with an automorphism.
@@
 Proof.
-By [KS12, Proposition 7.1.7] a constructible complex $\mathcal{F}$ is cohomologically complete, i.e.
+By [KashiwaraSchapira, Proposition 7.1.7] a constructible complex $\mathcal{F}$ is cohomologically complete, i.e.
 $\mathbb{R}\mathcal{H}om_{\mathbf{C}_{X}[\![\hbar]\!]}(\mathbf{C}_{X}(\!(\hbar)\!),\mathcal{F})=0.$
@@
 $\mathcal{F}\longrightarrow\lim_{n}\mathcal{F}/\hbar^{n}$
-is an isomorphism is the content of [PSY14, Theorem 7.8] in the case $X=\mathrm{pt}$ and the same proof works for sheaves. ∎
-We are now ready to define the functor $\widehat{\mathrm{RH}}^{-1}$.
+is an isomorphism is the content of [PSY, Theorem 7.8] in the case $X=\mathrm{pt}$ and the same proof works for sheaves. ∎
+We are now ready to define the functor $\mathrm{RH}^{-1}$.
 Given a perverse sheaf $\mathcal{F}$ equipped with an automorphism $T$ we can write $T=\exp(-2\pi iM)$ for a unique endomorphism $M\colon\mathcal{F}\rightarrow\mathcal{F}$ whose eigenvalues lie in $G\subset\mathbf{C}$.
-Then we define $\widehat{\mathrm{RH}}^{-1}(\mathcal{F})$ to be the pair $(\mathcal{F}(\!(\hbar)\!),D)$, where
+Then we define $\mathrm{RH}^{-1}(\mathcal{F})$ to be the pair $(\mathcal{F}(\!(\hbar)\!),D)$, where
 $D(f\otimes\eta)=f\otimes(M\eta)+\hbar\frac{df}{d\hbar}\otimes\eta$
@@
 Proposition 2.12.
-The functor $\widehat{\mathrm{RH}}^{-1}\colon\mathrm{Perv}_{\mathrm{aut}}(X)\rightarrow\mathrm{Perv}_{\mathrm{diff}}(X)$ is fully faithful.
-Moreover, if $\mathcal{M}\in\mathrm{Perv}_{\mathrm{diff}}(U)$ is a differential perverse sheaf with a lattice $\mathcal{L}\subset\mathcal{M}$, then RH^1(/)superscriptnormal-^normal-$\mathcal{M}\cong\widehat{\mathrm{RH}}^{-1}(\mathcal{L}/\hbar)$.
-In particular, the essential image of $\widehat{\mathrm{RH}}^{-1}$ consists of differential perverse sheaves with regular singularities.
+The functor $\mathrm{RH}^{-1}\colon\mathrm{Perv}_{\mathrm{aut}}(X)\rightarrow\mathrm{Perv}_{\mathrm{diff}}(X)$ is fully faithful.
+Moreover, if $\mathcal{M}\in\mathrm{Perv}_{\mathrm{diff}}(U)$ is a differential perverse sheaf with a lattice $\mathcal{L}\subset\mathcal{M}$, then $\mathcal{M}\cong\mathrm{RH}^{-1}(\mathcal{L}/\hbar)$.
+In particular, the essential image of $\mathrm{RH}^{-1}$ consists of differential perverse sheaves with regular singularities.
 Proof.
-Given a morphism $\mathcal{F}_{1}\rightarrow\mathcal{F}_{2}$ in $\mathrm{Perv}_{\mathrm{aut}}(X)$, a morphism $\widehat{\mathrm{RH}}^{-1}(\mathcal{F}_{1})\rightarrow\widehat{\mathrm{RH}}^{-1}(\mathcal{F}_{2})$ of $\mathbf{C}(\!(\hbar)\!)$-linear perverse sheaves is uniquely specified by its restriction $\mathcal{F}_{1}\rightarrow\mathcal{F}_{2}(\!(\hbar)\!)$ to $\mathcal{F}_{1}\hbar^{0}\subset\mathcal{F}_{1}(\!(\hbar)\!)$.
+Given a morphism $\mathcal{F}_{1}\rightarrow\mathcal{F}_{2}$ in $\mathrm{Perv}_{\mathrm{aut}}(X)$, a morphism $\mathrm{RH}^{-1}(\mathcal{F}_{1})\rightarrow\mathrm{RH}^{-1}(\mathcal{F}_{2})$ of $\mathbf{C}(\!(\hbar)\!)$-linear perverse sheaves is uniquely specified by its restriction $\mathcal{F}_{1}\rightarrow\mathcal{F}_{2}(\!(\hbar)\!)$ to $\mathcal{F}_{1}\hbar^{0}\subset\mathcal{F}_{1}(\!(\hbar)\!)$.
 Let $\mathcal{F}_{1}^{\lambda}\subset\mathcal{F}_{1}$ be a generalized eigenspace of $D$ with eigenvalue $m\in G$, where $\exp(-2\pi im)=\lambda$.
@@
 is an isomorphism, so this identifies $\mathcal{L}\cong\mathcal{F}[\![\hbar]\!]$.
-Therefore, $\mathcal{E}\cong\mathcal{F}(\!(\hbar)\!)=\widehat{\mathrm{RH}}^{-1}(\mathcal{F})$.
+Therefore, $\mathcal{E}\cong\mathcal{F}(\!(\hbar)\!)=\mathrm{RH}^{-1}(\mathcal{F})$.
 Under this isomorphism the action of $D$ on $\mathcal{E}$ coincides with the action of $D$ coming from the monodromy operator $T$ on $\mathcal{F}$.
@@
 In this section we recall some algebras of microdifferential operators that we will use.
-We refer to [Kas03, Kas86] for more details.
+We refer to [KashiwaraDmodules, KashiwaraIntroduction] for more details.
 Recall the punctured cotangent bundle $\mathring{\mathrm{T}}^{*}X=\mathrm{T}^{*}X\setminus\mathrm{T}^{*}_{X}X$ and the projectivized cotangent bundle $\mathrm{P}^{*}X=\mathring{\mathrm{T}}^{*}X/\mathbf{C}^{\times}$ with $\gamma\colon\mathring{\mathrm{T}}^{*}X\rightarrow\mathrm{P}^{*}X$ the natural projection.
 We denote the projections $\mathrm{T}^{*}X\rightarrow X$ and $\mathrm{P}^{*}X\rightarrow X$ by the same letter $\pi$.
-Consider the sheaf ^$\widehat{\mathcal{E}}_{X}$ on $\mathrm{P}^{*}X$ of formal microdifferential operators introduced in [SKK73]. It has a filtration
+Consider the sheaf ^$\widehat{\mathcal{E}}_{X}$ on $\mathrm{P}^{*}X$ of formal microdifferential operators introduced in [SatoKawaiKashiwara]. It has a filtration
 ^$\widehat{\mathcal{E}}_{X}=\bigcup_{m\in\mathbf{Z}}\widehat{\mathcal{E}}_{X}(m)$
@@
 Remark 3.1.
-Sometimes one considers the sheaf $\mathcal{E}_{X}$ of microdifferential operators; the sheaf ^$\widehat{\mathcal{E}}_{X}$ is obtained from $\mathcal{E}_{X}$ by completion with respect to the order.
-The theory of regular holonomic $\mathcal{E}_{X}$- and ^$\widehat{\mathcal{E}}_{X}$-modules coincide (see e.g.
-[KK83, Theorem 6.3]), so the difference is inessential for our purposes.
+In Section 8 we will also encounter the sheaf $\mathcal{E}_{X}$ of non-formal microdifferential operators, so that ^$\widehat{\mathcal{E}}_{X}$ is obtained by completion from $\mathcal{E}_{X}$.
 By construction there is a natural embedding
@@
 $V^{k}_{Z}\mathcal{D}_{X}=\{P\in\mathcal{D}_{X}|_{Z}\mid PJ_{Z}^{j}\subset J_{Z}^{j+k},\ \forall j,j+k\geq 0\}.$
-As shown in [Kas83], there is a natural embedding
+As shown in [KashiwaraVanishing], there is a natural embedding
 $\operatorname{gr}_{Z}\mathcal{D}_{X}\subset\tau_{*}\mathcal{D}_{\mathrm{T}_{Z}X},$
@@
 Denote by $\operatorname{gr}_{Z}\mathcal{M}$ the associated graded module with respect to the canonical $V$-filtration $V^{k}_{Z}\mathcal{M}$.
-We refer to [Kas83, Mon96] for details on the following functors.
+We refer to [KashiwaraVanishing, MonteiroFernandes] for details on the following functors.
 Definition 3.5.
@@
 $K_{\tilde{\Lambda}|X}=K_{\tilde{\Lambda}}\otimes\pi^{*}K_{X}^{-1}$
-on $\tilde{\Lambda}$. By [KK81, Lemma 1.5.1] there is a natural embedding
+on $\tilde{\Lambda}$. By [KashiwaraKawaiHolonomic, Lemma 1.5.1] there is a natural embedding
 $L\colon\operatorname{gr}_{\Lambda}\widehat{\mathcal{E}}_{X}\hookrightarrow\gamma_{*}\mathcal{D}_{\tilde{\Lambda}}^{K^{1/2}_{\tilde{\Lambda}|X}},$
@@
 Since $\tilde{\Lambda}\subset\mathrm{T}^{*}X$ is homogeneous, the Lie derivative along the Euler vector field defines an element $\theta^{\prime}\in\mathcal{D}_{\tilde{\Lambda}}^{K^{1/2}_{\tilde{\Lambda}|X}}$.
-We denote by $\theta^{\prime}\in\widehat{\mathcal{E}}_{X}(1)|_{\Lambda}$ any preimage (such a lift is characterized by [KK81, Equations (1.5.5) and (1.5.6)]).
-If $\mathcal{M}$ is a regular holonomic ^$\widehat{\mathcal{E}}_{X}$-module and $u\in\mathcal{M}$ is a local nonvanishing section, then by [KK80, Theorem 4.1.1] there is a polynomial $b_{u}(s)\in\mathbf{C}[s]$ such that $b_{u}(\theta^{\prime})u\in\widehat{\mathcal{E}}_{\Lambda/X}(-1)u$.
+We denote by $\theta^{\prime}\in\widehat{\mathcal{E}}_{X}(1)|_{\Lambda}$ any preimage (such a lift is characterized by [KashiwaraKawaiHolonomic, Equations (1.5.5) and (1.5.6)]).
+If $\mathcal{M}$ is a regular holonomic ^$\widehat{\mathcal{E}}_{X}$-module and $u\in\mathcal{M}$ is a local nonvanishing section, then by [KashiwaraKawaiMicrolocalization, Theorem 4.1.1] there is a polynomial $b_{u}(s)\in\mathbf{C}[s]$ such that $b_{u}(\theta^{\prime})u\in\widehat{\mathcal{E}}_{\Lambda/X}(-1)u$.
 The order $\mathrm{ord}_{\Lambda}(u)$ of $u\in\mathcal{M}$ is the set of zeros of its $b$-function.
@@
 Then we may consider the $V$-filtration on $\mathcal{M}$ along $Z$ or its microlocal version on $\tilde{\mathcal{M}}$ along $\Lambda$.
-They are related as follows (see [Mon96, Remark 1.2.4]): the natural map $\pi^{-1}\mathcal{M}\rightarrow\tilde{\mathcal{M}}$ induces an isomorphism
+They are related as follows (see [MonteiroFernandes, Remark 1.2.4]): the natural map $\pi^{-1}\mathcal{M}\rightarrow\tilde{\mathcal{M}}$ induces an isomorphism
 (2) $\pi^{-1}\operatorname{gr}^{k}_{Z}\mathcal{M}\cong K_{Z|X}^{-1}\otimes_{\pi^{-1}\mathcal{O}_{X}}\operatorname{gr}^{k}_{\Lambda}\tilde{\mathcal{M}}$
@@
   • The order of simple ^$\widehat{\mathcal{E}}^{\sqrt{v}}_{X}$-module is $\mathrm{ord}(\mathcal{M})\in\mathbf{C}/\mathbf{Z}$ defined to be the equivalence class of $\mathrm{ord}(u)$ for a simple generator $u$.
-We have the following important classification result of microdifferential modules regular along a Legendrian (see [Kas86, Theorem 10.3.1]). Let
+We have the following important classification result of microdifferential modules regular along a Legendrian. Let
 $\mathrm{Mod}_{\mathrm{lisse}}\left(\mathcal{D}^{\sqrt{v}}_{\tilde{\Lambda}}\right)\subset\mathrm{Mod}_{\mathrm{rh}}\left(\mathcal{D}^{\sqrt{v}}_{\tilde{\Lambda}}\right)$
-be the full substack of twisted $D$-modules on $\tilde{\Lambda}$ which are lisse, i.e. whose underlying $\mathcal{O}$-module is a finite rank vector bundle.
+be the full substack of twisted $D$-modules on $\tilde{\Lambda}$ which are local systems, i.e.
+whose underlying $\mathcal{O}$-module is a finite rank vector bundle.
+The following statement combines [KashiwaraIntroduction, Theorem 10.3.1] and [KashiwaraKawaiMicrolocal, Theorem 7.9].
 Theorem 3.10.
@@
 $\mu_{\Lambda}^{dR}\colon\mathrm{Mod}_{\Lambda,\mathrm{rh}}(\widehat{\mathcal{E}}^{\sqrt{v}}_{X})\xrightarrow{\sim}\gamma_{*}\mathrm{Mod}_{\mathrm{lisse}}\left(\mathcal{D}^{\sqrt{v}}_{\tilde{\Lambda}}\right)$
-between the stack of ^$\widehat{\mathcal{E}}^{\sqrt{v}}_{X}$-modules regular along $\Lambda$ and lisse twisted $D$-modules on $\tilde{\Lambda}$.
+between the stack of ^$\widehat{\mathcal{E}}^{\sqrt{v}}_{X}$-modules regular along $\Lambda$ and twisted local systems on $\tilde{\Lambda}$.
+Moreover, if $\mathcal{M}$ is a regular holonomic ^$\widehat{\mathcal{E}}^{\sqrt{v}}_{X}$-module and $\mathcal{L}$ a ^$\widehat{\mathcal{E}}^{\sqrt{v}}_{X}$-module regular along $\Lambda$, then microlocalization defines an isomorphism
+$\mathbb{R}\mathcal{H}om_{\widehat{\mathcal{E}}^{\sqrt{v}}_{X}}(\mathcal{M},\mathcal{L})\xrightarrow{\sim}\gamma_{*}\mathbb{R}\mathcal{H}om_{\mathcal{D}^{\sqrt{v}}_{\tilde{\Lambda}}}(\mu_{\Lambda}^{dR}(\mathcal{M}),\mu_{\Lambda}^{dR}(\mathcal{L})).$
 Remark 3.11.
@@
 3.5. Quantized contact transformations
-To define the stack of microdifferential modules for a general contact manifold $Y$, Kashiwara in [Kas96] uses the theory of quantized contact transformations developed in [SKK73] that we now recall.
+To define the stack of microdifferential modules for a general contact manifold $Y$, Kashiwara in [KashiwaraContact] uses the theory of quantized contact transformations developed in [SatoKawaiKashiwara] that we now recall.
 Definition 3.12.
@@
 of sheaves of filtered $\mathbf{C}$-algebras which coincides with $\psi$ after passing to the associated graded algebras.
-By [Kas03, Theorem 8.11] quantized contact transformations exist locally. We will also need a half-twisted version as follows.
+By [KashiwaraDmodules, Theorem 8.11] quantized contact transformations exist locally. We will also need a half-twisted version as follows.
 Definition 3.13.
 Suppose $V_{X}\subset\mathrm{P}^{*}X$ and $V_{Y}\subset\mathrm{P}^{*}Y$ are open subsets together with a contactomorphism $\psi\colon V_{X}\xrightarrow{\sim}V_{Y}$.
-A normal-$\ast$-preserving quantized contact transformation over $\psi$ is an isomorphism
+A $\ast$-preserving quantized contact transformation over $\psi$ is an isomorphism
 $\Psi\colon\widehat{\mathcal{E}}^{\sqrt{v}}_{Y}|_{V_{Y}}\xrightarrow{\sim}\psi_{*}\widehat{\mathcal{E}}^{\sqrt{v}}_{X}|_{V_{X}}$
@@
 Proposition 3.14.
-Consider a normal-$\ast$-preserving quantized contact transformation $\Psi$ over $\psi\colon V_{X}\xrightarrow{\sim}V_{Y}$ as in the above definition.
+Consider a $\ast$-preserving quantized contact transformation $\Psi$ over $\psi\colon V_{X}\xrightarrow{\sim}V_{Y}$ as in the above definition.
 Suppose $\Lambda_{X}\subset V_{X}$ and $\Lambda_{Y}\subset V_{Y}$ are two Legendrians such that $\Lambda_{Y}=\psi(\Lambda_{X})$.
-Then a normal-$\ast$-preserving quantized contact transformation $\Psi\colon\widehat{\mathcal{E}}^{\sqrt{v}}_{Y}|_{V_{Y}}\xrightarrow{\sim}\psi_{*}\widehat{\mathcal{E}}^{\sqrt{v}}_{X}|_{V_{X}}$ restricts to an isomorphism
+Then a $\ast$-preserving quantized contact transformation $\Psi\colon\widehat{\mathcal{E}}^{\sqrt{v}}_{Y}|_{V_{Y}}\xrightarrow{\sim}\psi_{*}\widehat{\mathcal{E}}^{\sqrt{v}}_{X}|_{V_{X}}$ restricts to an isomorphism
 $\Psi\colon\widehat{\mathcal{E}}^{\sqrt{v}}_{\Lambda_{Y}/Y}|_{V_{Y}}\xrightarrow{\sim}\psi_{*}\widehat{\mathcal{E}}^{\sqrt{v}}_{\Lambda_{X}/X}|_{V_{X}}.$
@@
 $\Psi(Q)K=Q^{*}K.$
-If $\Psi$ is a normal-$\ast$-preserving quantized contact transformation, then
+If $\Psi$ is a $\ast$-preserving quantized contact transformation, then
 $\sigma_{\Lambda}(K)=\sqrt{\mathrm{vol}_{\mathrm{T}^{*}Y}},\qquad\mathrm{ord}(K)=(\dim Y)/2.$
 Proof.
-The correspondence between quantized contact transformations and pairs $(\mathcal{M},K)$ is established in [Kas03, Chapter 8.2].
+The correspondence between quantized contact transformations and pairs $(\mathcal{M},K)$ is established in [KashiwaraDmodules, Chapter 8.2].
 Now assume $\Psi$ is a $\ast$-preserving quantized contact transformation. Consider $Q\in\widehat{\mathcal{E}}^{\sqrt{v}}_{Y}(m+1)$. By definition we have
 $(\Psi(Q)-Q^{*})K=0.$
-To show that $\sigma_{\Lambda}(K)=\sqrt{\mathrm{vol}_{\mathrm{T}^{*}Y}}$, using the notation of [Kas03, Chapter 8.6] we have to prove that
+To show that $\sigma_{\Lambda}(K)=\sqrt{\mathrm{vol}_{\mathrm{T}^{*}Y}}$, using the notation of [KashiwaraDmodules, Chapter 8.6] we have to prove that
 $(H_{\sigma_{m+1}(\Psi(Q))}+\sigma_{m}(\Psi(Q))-H_{\sigma_{m+1}(Q^{*})}-\sigma_{m}(Q^{*}))\sqrt{\mathrm{vol}_{\mathrm{T}^{*}Y}}=0.$
@@
 Let $Y$ be a contact manifold with a symplectization $\gamma\colon\tilde{Y}\rightarrow Y$.
-The papers [Kas96, PS04] have defined the algebroid $\mathsf{E}_{Y}$ over $Y$ uniquely determined by the following properties:
+The papers [KashiwaraContact, PoleselloSchapira] have defined the algebroid $\widehat{\mathsf{E}}_{Y}$ over $Y$ uniquely determined by the following properties:
   • If $U\subset Y$ is an open subset together with a contact embedding $U\hookrightarrow\mathrm{P}^{*}X$ into a projectivized cotangent bundle, then
-    $\mathsf{E}_{Y}|_{U}\cong\widehat{\mathcal{E}}^{\sqrt{v}}_{X}|_{U}.$
+    $\widehat{\mathsf{E}}_{Y}|_{U}\cong\widehat{\mathcal{E}}^{\sqrt{v}}_{X}|_{U}.$
   • Given an intersection of two such charts, the isomorphism of the algebras of half-twisted microdifferential operators is given by a $\ast$-preserving quantized contact transformation.
 Consider a Legendrian submanifold $\Lambda\subset Y$.
-As the notion of regularity of ^$\widehat{\mathcal{E}}^{\sqrt{v}}_{X}$-modules is local and invariant under quantized contact transformations, it extends to $\mathsf{E}_{Y}$-modules.
+As the notion of regularity of ^$\widehat{\mathcal{E}}^{\sqrt{v}}_{X}$-modules is local and invariant under quantized contact transformations, it extends to $\widehat{\mathsf{E}}_{Y}$-modules.
 Moreover, using 3.14 we obtain a global version of the microlocalization functor
-$\mu^{dR}_{\Lambda}\colon\mathrm{Mod}_{\mathrm{rh}}(\mathsf{E}_{Y})\longrightarrow\gamma_{*}\mathrm{Mod}_{\mathrm{rh}}\left(\mathcal{D}^{\sqrt{v}}_{\tilde{\Lambda}}\right).$
-Using this functor one obtains the following classification result (see [DS07, Corollary 6.4]).
+$\mu^{dR}_{\Lambda}\colon\mathrm{Mod}_{\mathrm{rh}}(\widehat{\mathsf{E}}_{Y})\longrightarrow\gamma_{*}\mathrm{Mod}_{\mathrm{rh}}\left(\mathcal{D}^{\sqrt{v}}_{\tilde{\Lambda}}\right).$
+Using this functor one obtains the following classification result generalizing 3.10 (see [DAgnoloSchapira, Corollary 6.4]).
 Theorem 3.16.
 Let $\Lambda\subset Y$ be a Legendrian submanifold. Microlocalization defines an equivalence of stacks
-$\mu_{\Lambda}^{dR}\colon\mathrm{Mod}_{\Lambda,\mathrm{rh}}(\mathsf{E}_{Y})\xrightarrow{\sim}\gamma_{*}\mathrm{Mod}_{\mathrm{lisse}}\left(\mathcal{D}^{\sqrt{v}}_{\tilde{\Lambda}}\right)$
-between the stack of $\mathsf{E}_{Y}$-modules regular along $\Lambda$ and the stack of lisse twisted $D$-modules on $\tilde{\Lambda}$.
-The antiinvolution $\ast$ on ^$\widehat{\mathcal{E}}^{\sqrt{v}}_{X}$ globalizes to an isomorphism of algebroids $\mathsf{E}_{Y}\cong\mathsf{E}_{Y}^{\mathrm{op}}$ for any contact manifold $Y$.
-Thus, any left $\mathsf{E}_{Y}$-module $\mathsf{M}$ can be considered as a right $\mathsf{E}_{Y}$-module and vice versa.
+$\mu_{\Lambda}^{dR}\colon\mathrm{Mod}_{\Lambda,\mathrm{rh}}(\widehat{\mathsf{E}}_{Y})\xrightarrow{\sim}\gamma_{*}\mathrm{Mod}_{\mathrm{lisse}}\left(\mathcal{D}^{\sqrt{v}}_{\tilde{\Lambda}}\right)$
+between the stack of $\widehat{\mathsf{E}}_{Y}$-modules regular along $\Lambda$ and the stack of twisted local systems on $\tilde{\Lambda}$.
+Moreover, if $\mathcal{M}$ is a regular holonomic $\widehat{\mathsf{E}}_{Y}$-module and $\mathcal{L}$ is an $\widehat{\mathsf{E}}_{Y}$-module regular along $\Lambda$, then microlocalization defines an isomorphism
+$\mathbb{R}\mathcal{H}om_{\widehat{\mathsf{E}}_{Y}}(\mathcal{M},\mathcal{L})\xrightarrow{\sim}\gamma_{*}\mathbb{R}\mathcal{H}om_{\mathcal{D}^{\sqrt{v}}_{\tilde{\Lambda}}}(\mu_{\Lambda}^{dR}(\mathcal{M}),\mu_{\Lambda}^{dR}(\mathcal{L})).$
+The antiinvolution $\ast$ on ^$\widehat{\mathcal{E}}^{\sqrt{v}}_{X}$ globalizes to an isomorphism of algebroids $\widehat{\mathsf{E}}_{Y}\cong\widehat{\mathsf{E}}_{Y}^{\mathrm{op}}$ for any contact manifold $Y$.
+Thus, any left $\widehat{\mathsf{E}}_{Y}$-module $\mathsf{M}$ can be considered as a right $\widehat{\mathsf{E}}_{Y}$-module and vice versa.
 4. DQ modules
@@
 $\widehat{\mathcal{W}}_{X}\subset\rho_{*}(\widehat{\mathcal{E}}_{X\times\mathbf{C}})$
-of $\mathbf{C}(\!(\hbar)\!)$-algebras of formal WKB-differential operators introduced in [PS04], where $\rho\colon\mathrm{J}^{1}X\rightarrow\mathrm{T}^{*}X$ and $\hbar=\partial_{t}^{-1}$.
+of $\mathbf{C}(\!(\hbar)\!)$-algebras of formal WKB-differential operators introduced in [PoleselloSchapira], where $\rho\colon\mathrm{J}^{1}X\rightarrow\mathrm{T}^{*}X$ and $\hbar=\partial_{t}^{-1}$.
 Remark 4.1.
@@
 We have that $\rho_{*}(\widehat{\mathcal{E}}_{X\times\mathbf{C}}(0))$ is $\hbar$-complete and has no $\hbar$-torsion.
-Therefore, by [KS12, Corollary 1.5.7] it is cohomologically $\hbar$-complete.
-By [KS12, Theorem 1.6.6] in this case it is enough to prove flatness modulo $\hbar$, i.e.
+Therefore, by [KashiwaraSchapira, Corollary 1.5.7] it is cohomologically $\hbar$-complete.
+By [KashiwaraSchapira, Theorem 1.6.6] in this case it is enough to prove flatness modulo $\hbar$, i.e.
 it is enough to prove that $\rho_{*}\mathcal{O}_{\mathrm{J}^{1}X}$ is flat as a $\mathcal{O}_{\mathrm{T}^{*}X}$-module.
-Since the map $\mathrm{J}^{1}X\rightarrow\mathrm{T}^{*}X$ is open, it is flat by [BS76, Theorem V.2.13].
+Since the map $\mathrm{J}^{1}X\rightarrow\mathrm{T}^{*}X$ is open, it is flat by [BanicaStanasila, Theorem V.2.13].
 ∎
@@
 Recall that in this case $\psi$ is compatible with the Hamiltonian $\mathbf{C}$-actions.
-The following is [PS04, Lemma 8.5].
+The following is [PoleselloSchapira, Lemma 8.5].
 Proposition 4.5.
-Consider an isomorphism of contactifications $\psi\colon V_{X}=\rho^{-1}(U_{X})\xrightarrow{\sim}V_{Y}=\rho^{-1}(U_{Y})$ as above.
-Then there is a normal-$\ast$-preserving quantized contact transformation
+Consider an isomorphism of contactifications $\psi\colon V_{X}=\rho^{-1}(U_{X})\xrightarrow{\sim}V_{Y}=\rho^{-1}(U_{Y})$ as above. Then there is a $\ast$-preserving quantized contact transformation
 $\Psi\colon\widehat{\mathcal{E}}^{\sqrt{v}}_{Y\times\mathbf{C}}|_{V_{Y}}\xrightarrow{\sim}\psi_{*}\widehat{\mathcal{E}}^{\sqrt{v}}_{X\times\mathbf{C}}|_{V_{X}}$
 compatible with the Hamiltonian $\mathbf{C}$-actions, i.e. $\Psi(\partial_{t})=\partial_{t}$ and $\Psi$ is $\mathbf{C}$-equivariant.
-Consider the setting of 4.5. Then we may restrict the $\ast$-preserving quantized contact transformation $\Psi$ to a normal-$\ast$-preserving quantized symplectic transformation
+Consider the setting of 4.5. Then we may restrict the $\ast$-preserving quantized contact transformation $\Psi$ to a $\ast$-preserving quantized symplectic transformation
 $\Phi\colon\widehat{\mathcal{W}}^{\sqrt{v}}_{Y}|_{U_{Y}}\xrightarrow{\sim}\phi_{*}\widehat{\mathcal{W}}^{\sqrt{v}}_{X}|_{U_{X}}.$
@@
 4.2. Globalization
-Let $S$ be a symplectic manifold. The authors of [PS04] have constructed a $\mathbf{C}(\!(\hbar)\!)$-algebroid $\mathsf{W}_{S}$ on $S$ with the following properties:
+Let $S$ be a symplectic manifold. The authors of [PoleselloSchapira] have constructed a $\mathbf{C}(\!(\hbar)\!)$-algebroid $\widehat{\mathsf{W}}_{S}$ on $S$ with the following properties:
   • If $U\subset S$ is an open subset together with a symplectic embedding $U\hookrightarrow\mathrm{T}^{*}X$ into a cotangent bundle, then
-    $\mathsf{W}_{S}|_{U}\cong\widehat{\mathcal{W}}^{\sqrt{v}}_{X}|_{U}.$
+    $\widehat{\mathsf{W}}_{S}|_{U}\cong\widehat{\mathcal{W}}^{\sqrt{v}}_{X}|_{U}.$
   • Given an intersection of two such charts, the isomorphism of the algebras of half-twisted WKB-differential operators is given by a $\ast$-preserving quantized symplectic transformation.
   • If $Y\rightarrow S$ is a contactification, then there is a natural embedding
-    $\mathsf{W}_{S}\subset\rho_{*}\mathsf{E}_{Y}.$
-The antiinvolution $\ast$ on $\widehat{\mathcal{W}}^{\sqrt{v}}_{X}$ globalizes to a $\mathbf{C}$-linear isomorphism of algebroids $\mathsf{W}_{S}\cong\mathsf{W}^{\mathrm{op}}_{S}$ which sends $\hbar\mapsto-\hbar$.
-As for $\mathsf{E}_{Y}$-modules, we freely pass between left and right $\mathsf{W}_{S}$-modules using this antiinvolution.
+    $\widehat{\mathsf{W}}_{S}\subset\rho_{*}\widehat{\mathsf{E}}_{Y}.$
+The antiinvolution $\ast$ on $\widehat{\mathcal{W}}^{\sqrt{v}}_{X}$ globalizes to a $\mathbf{C}$-linear isomorphism of algebroids $\widehat{\mathsf{W}}_{S}\cong\widehat{\mathsf{W}}^{\mathrm{op}}_{S}$ which sends $\hbar\mapsto-\hbar$.
+As for $\widehat{\mathsf{E}}_{Y}$-modules, we freely pass between left and right $\widehat{\mathsf{W}}_{S}$-modules using this antiinvolution.
 Definition 4.7.
-A $\mathsf{W}_{S}$-module is holonomic if its support is a Lagrangian subvariety of $S$.
-For a Lagrangian subvariety $L\subset S$ we denote by $\mathsf{W}_{L/S}\subset\mathsf{W}_{S}|_{L}$ generated by
-$I_{L}=\{P\in\mathsf{W}_{S}(1)|_{L}\mid\sigma_{1}(P)|_{L}=0\}.$
+A $\widehat{\mathsf{W}}_{S}$-module is holonomic if its support is a Lagrangian subvariety of $S$.
+For a Lagrangian subvariety $L\subset S$ we denote by $\widehat{\mathsf{W}}_{L/S}\subset\widehat{\mathsf{W}}_{S}|_{L}$ generated by
+$I_{L}=\{P\in\widehat{\mathsf{W}}_{S}(1)|_{L}\mid\sigma_{1}(P)|_{L}=0\}.$
 Definition 4.8.
-A holonomic $\mathsf{W}_{S}$-module $\mathcal{M}$ is regular holonomic if locally there is a coherent $\mathsf{W}_{S}(0)$-module $\mathcal{M}_{0}\subset\mathcal{M}$ and a Lagrangian subvariety $L\subset S$ containing $\operatorname{supp}\mathcal{M}$ such that $I_{L}\mathcal{M}_{0}\subset\mathcal{M}_{0}$ and $\mathsf{W}_{S}\mathcal{M}_{0}=\mathcal{M}$.
+A holonomic $\widehat{\mathsf{W}}_{S}$-module $\mathcal{M}$ is regular holonomic if locally there is a coherent $\widehat{\mathsf{W}}_{S}(0)$-module $\mathcal{M}_{0}\subset\mathcal{M}$ and a Lagrangian subvariety $L\subset S$ containing $\operatorname{supp}\mathcal{M}$ such that $I_{L}\mathcal{M}_{0}\subset\mathcal{M}_{0}$ and $\widehat{\mathsf{W}}_{S}\mathcal{M}_{0}=\mathcal{M}$.
 Example 4.9.
 Let $L\subset S$ be a Lagrangian subvariety and $\Lambda\subset Y$ be its contactification with $\rho\colon Y\rightarrow S$ the projection map.
-If $\mathcal{M}$ is a $\mathsf{E}_{Y}$-module supported on $\Lambda$, then $\mathcal{M}$, regarded as a $\mathsf{W}_{S}$-module via the embedding $\mathsf{W}_{S}\subset\rho_{*}\mathsf{E}_{Y}$, is supported on $L$ and is, therefore, holonomic.
+If $\mathcal{M}$ is a $\widehat{\mathsf{E}}_{Y}$-module supported on $\Lambda$, then $\mathcal{M}$, regarded as a $\widehat{\mathsf{W}}_{S}$-module via the embedding $\widehat{\mathsf{W}}_{S}\subset\rho_{*}\widehat{\mathsf{E}}_{Y}$, is supported on $L$ and is, therefore, holonomic.
 This gives a forgetful functor
-$\rho_{*}\mathrm{Mod}_{hol}(\mathsf{E}_{Y})\longrightarrow\mathrm{Mod}_{hol}(\mathsf{W}_{S})$
-from holonomic $\mathsf{E}_{Y}$-modules to holonomic $\mathsf{W}_{S}$-modules. Since $\mathsf{W}_{L/S}\subset\rho_{*}\mathsf{E}_{\Lambda/Y}$, this forgetful functor restricts to a functor
-$\rho_{*}\mathrm{Mod}_{\mathrm{rh}}(\mathsf{E}_{Y})\longrightarrow\mathrm{Mod}_{\mathrm{rh}}(\mathsf{W}_{S}).$
-The authors of [DK11] have introduced the stack $\mathrm{RH}_{S}$ of regular holonomic modules on $S$ as follows:
-  • Let $L\subset S$ be a Lagrangian subvariety and $\Lambda\subset Y$ its contactification with projection $\rho\colon Y\rightarrow S$. We define
-    $\mathrm{RH}_{S,L}=\rho_{*}\mathrm{Mod}_{\Lambda,\mathrm{rh}}(\mathsf{E}_{Y})=\mathrm{Mod}_{\Lambda,\mathrm{rh}}(\rho_{*}\mathsf{E}_{Y}),$
-    the stack of regular holonomic $\mathsf{E}_{Y}$-modules supported on $\Lambda$.
-  • As the contactification of $L\subset S$ exists locally and is unique by 1.12, the stack $\mathrm{RH}_{S,L}$ only depends on $L\subset S$.
-  • Lagrangian subvarieties of $S$ form a filtered system and we define
-    $\mathrm{RH}_{S}=\operatorname*{colim}_{L}\mathrm{RH}_{S,L}.$
-Example 4.10.
-Let $L\subset S$ be a Lagrangian submanifold.
-The stack $\mathrm{RH}_{S,L}$ consists of $\mathsf{E}_{Y}$-modules regular along a contactification $\Lambda\subset Y$ of $L\subset S$.
-By 3.16 microlocalization establishes an equivalence
-$\mathrm{RH}_{S,L}\cong\gamma_{*}\mathrm{Mod}_{\mathrm{lisse}}(\mathcal{D}^{\sqrt{v}}_{L\times\mathbf{C}^{\times}}),$
-where $\gamma\colon L\times\mathbf{C}^{\times}\rightarrow L$ is the projection on the first factor.
-Half-twisted $D$-modules along $\mathbf{C}^{\times}$ whose underlying $\mathcal{O}$-module is a line bundle are classified by their order $\lambda\in\mathbf{C}/\mathbf{Z}$.
-In particular, if $L$ is equipped with an orientation data and given $\lambda\in\mathbf{C}/\mathbf{Z}$, there is a canonical module
-$\mathsf{M}^{\lambda}_{L}\in\mathrm{RH}_{S,L}\subset\mathrm{RH}_{S}$
-simple along $L$ with order $\lambda$.
+$\rho_{*}\mathrm{Mod}_{hol}(\widehat{\mathsf{E}}_{Y})\longrightarrow\mathrm{Mod}_{hol}(\widehat{\mathsf{W}}_{S})$
+from holonomic $\widehat{\mathsf{E}}_{Y}$-modules to holonomic $\widehat{\mathsf{W}}_{S}$-modules.
+Since $\widehat{\mathsf{W}}_{L/S}\subset\rho_{*}\widehat{\mathsf{E}}_{\Lambda/Y}$, this forgetful functor restricts to a functor
+$\rho_{*}\mathrm{Mod}_{\mathrm{rh}}(\widehat{\mathsf{E}}_{Y})\longrightarrow\mathrm{Mod}_{\mathrm{rh}}(\widehat{\mathsf{W}}_{S}).$
 4.3. Tensor product of DQ modules
-Consider two objects $\mathcal{L},\mathcal{M}\in\mathrm{RH}_{S}$ which are regular holonomic $\rho_{*}\mathsf{E}_{Y}$-modules.
-In this section we will study the complex of $\mathbf{C}(\!(\hbar)\!)$-linear sheaves $\mathcal{L}\otimes^{\mathbb{L}}_{\mathsf{W}_{S}}\mathcal{M}$ on $X$.
+Consider two regular holonomic $\widehat{\mathsf{E}}_{Y}$-modules $\mathcal{L},\mathcal{M}$.
+In this section we will study the complex of $\mathbf{C}(\!(\hbar)\!)$-linear sheaves $\mathcal{L}\otimes^{\mathbb{L}}_{\widehat{\mathsf{W}}_{S}}\mathcal{M}$ on $X$.
 Our first goal is to endow it with extra structure.
-Choose a contactification $\rho\colon Y\rightarrow S$ and assume that $\mathsf{E}_{Y}$ and $\mathsf{W}_{S}$ are represented by sheaves of algebras, there is an element $\hbar\in\mathsf{E}_{Y}(-1)$ with principal symbol $\tau^{-1}$ satisfying $\hbar^{*}=-\hbar$ and $\mathsf{W}_{S}$ is the subalgebra of $\rho_{*}\mathsf{E}_{Y}$ of elements commuting with $\hbar$.
+Choose a contactification $\rho\colon Y\rightarrow S$ and assume that $\widehat{\mathsf{E}}_{Y}$ and $\widehat{\mathsf{W}}_{S}$ are represented by sheaves of algebras, there is an element $\hbar\in\widehat{\mathsf{E}}_{Y}(-1)$ with principal symbol $\tau^{-1}$ satisfying $\hbar^{*}=-\hbar$ and $\widehat{\mathsf{W}}_{S}$ is the subalgebra of $\rho_{*}\widehat{\mathsf{E}}_{Y}$ of elements commuting with $\hbar$.
 Consider
-$\mathsf{B}_{Y}=(\rho_{*}\mathsf{E}_{Y})\otimes_{\mathsf{W}_{S}}(\rho_{*}\mathsf{E}_{Y}),$
-where the right action of $\mathsf{W}_{S}$ on the first copy of $\rho_{*}\mathsf{E}_{Y}$ is via $P\in\mathsf{W}_{S},Q\in\rho_{*}\mathsf{E}_{Y}\mapsto P^{*}Q$.
-It carries a natural bimodule structure as follows: $\rho_{*}\mathsf{E}_{Y}\otimes\rho_{*}\mathsf{E}_{Y}$ acts in an obvious way on the right and $\hbar$ acts on $Q_{1}\otimes Q_{2}\in\mathsf{B}_{Y}$ as $Q_{1}\otimes\hbar Q_{2}=-\hbar Q_{1}\otimes Q_{2}$.
+$\mathsf{B}_{Y}=(\rho_{*}\widehat{\mathsf{E}}_{Y})\otimes_{\widehat{\mathsf{W}}_{S}}(\rho_{*}\widehat{\mathsf{E}}_{Y}),$
+where the right action of $\widehat{\mathsf{W}}_{S}$ on the first copy of $\rho_{*}\widehat{\mathsf{E}}_{Y}$ is via $P\in\widehat{\mathsf{W}}_{S},Q\in\rho_{*}\widehat{\mathsf{E}}_{Y}\mapsto P^{*}Q$.
+It carries a natural bimodule structure as follows: $\rho_{*}\widehat{\mathsf{E}}_{Y}\otimes\rho_{*}\widehat{\mathsf{E}}_{Y}$ acts in an obvious way on the right and $\hbar$ acts on $Q_{1}\otimes Q_{2}\in\mathsf{B}_{Y}$ as $Q_{1}\otimes\hbar Q_{2}=-\hbar Q_{1}\otimes Q_{2}$.
 Proposition 4.10.
-$\mathsf{B}_{Y}$ has a canonical structure of a $(\mathbf{C}(\!(\hbar)\!)\langle D\rangle,\rho_{*}\mathsf{E}_{Y}\otimes\rho_{*}\mathsf{E}_{Y})$-bimodule, where $D\hbar=\hbar(D+1)$.
+$\mathsf{B}_{Y}$ has a canonical structure of a $(\mathbf{C}(\!(\hbar)\!)\langle D\rangle,\rho_{*}\widehat{\mathsf{E}}_{Y}\otimes\rho_{*}\widehat{\mathsf{E}}_{Y})$-bimodule, where $D\hbar=\hbar(D+1)$.
 Proof.
-Let $\partial_{t}=\hbar^{-1}\in\mathsf{E}_{Y}(1)$. Locally we may choose an element $t\in\mathsf{E}_{Y}(0)$ satisfying
+Let $\partial_{t}=\hbar^{-1}\in\widehat{\mathsf{E}}_{Y}(1)$. Locally we may choose an element $t\in\widehat{\mathsf{E}}_{Y}(0)$ satisfying
 $[\partial_{t},t]=1.$
@@
 $[D,\hbar]=\hbar,\qquad[D^{*},\hbar]=-\hbar.$
-This allows us to enhance $\mathsf{B}_{Y}$ to a $(\mathbf{C}(\!(\hbar)\!)\langle D\rangle,\rho_{*}\mathsf{E}_{Y}\otimes\rho_{*}\mathsf{E}_{Y})$-bimodule, where $D=\partial_{t}t$ acts on $Q_{1}\otimes Q_{2}\in\mathsf{B}_{Y}$ as $-D^{*}Q_{1}\otimes Q_{2}+Q_{1}\otimes DQ_{2}$.
+This allows us to enhance $\mathsf{B}_{Y}$ to a $(\mathbf{C}(\!(\hbar)\!)\langle D\rangle,\rho_{*}\widehat{\mathsf{E}}_{Y}\otimes\rho_{*}\widehat{\mathsf{E}}_{Y})$-bimodule, where $D=\partial_{t}t$ acts on $Q_{1}\otimes Q_{2}\in\mathsf{B}_{Y}$ as $-D^{*}Q_{1}\otimes Q_{2}+Q_{1}\otimes DQ_{2}$.
 Let us show that this enhancement is independent of the choice of $t$.
-If $t^{\prime}$ is another element of $\mathsf{E}_{Y}(0)$ satisfying $[\partial_{t},t^{\prime}]=1$, then $[\partial_{t},t-t^{\prime}]=0$, i.e.
+If $t^{\prime}$ is another element of $\widehat{\mathsf{E}}_{Y}(0)$ satisfying $[\partial_{t},t^{\prime}]=1$, then $[\partial_{t},t-t^{\prime}]=0$, i.e.
 $[\hbar,t-t^{\prime}]=0$.
-Thus, there is an element $R\in\mathsf{W}_{S}(0)$ such that
+Thus, there is an element $R\in\widehat{\mathsf{W}}_{S}(0)$ such that
 $t^{\prime}=t+R.$
@@
 $-(\partial_{t}t^{\prime})^{*}Q_{1}\otimes Q_{2}+Q_{1}\otimes\partial_{t}t^{\prime}Q_{2}=-(\partial_{t}t)^{*}Q_{1}\otimes Q_{2}+Q_{1}\otimes\partial_{t}tQ_{2}-(\partial_{t}R)^{*}Q_{1}\otimes Q_{2}+Q_{1}\otimes\partial_{t}RQ_{2}.$
-But $\partial_{t}R\in\mathsf{W}_{S}$ and hence the last two terms cancel. ∎
+But $\partial_{t}R\in\widehat{\mathsf{W}}_{S}$ and hence the last two terms cancel. ∎
 Using the previous proposition we obtain the following result.
 Theorem 4.11.
-For any $\mathcal{L},\mathcal{M}\in\mathrm{RH}_{S}$ the sheaf $\mathcal{L}\otimes^{\mathbb{L}}_{\mathsf{W}_{S}}\mathcal{M}$ has a canonical structure of a differential perverse sheaf.
+For any regular holonomic $\rho_{*}\widehat{\mathsf{E}}_{Y}$-modules $\mathcal{L},\mathcal{M}$ the sheaf $\mathcal{L}\otimes^{\mathbb{L}}_{\widehat{\mathsf{W}}_{S}}\mathcal{M}$ has a canonical structure of a differential perverse sheaf.
 Proof.
-Consider the duality functor $\mathbb{D}(\mathcal{L})=\mathbb{R}\mathcal{H}om_{\mathsf{W}_{S}}(\mathcal{L},\mathsf{W}_{S})[\dim S/2]$.
-By [KS12, Lemma 7.2.2] $\mathbb{D}(\mathcal{L})$ is a holonomic $\mathsf{W}_{S}$-module concentrated in degree 0 and we have
-$\mathbb{R}\mathcal{H}om_{\mathsf{W}_{S}}(\mathbb{D}(\mathcal{L}),\mathcal{M})[\dim S/2]\cong\mathcal{L}\otimes^{\mathbb{L}}_{\mathsf{W}_{S}}\mathcal{M}.$
-It is then shown in [KS12, Theorem 7.2.3] that $\mathbb{R}\mathcal{H}om_{\mathsf{W}_{S}}(\mathbb{D}(\mathcal{L}),\mathcal{M})[\dim S/2]$ is a perverse sheaf of $\mathbf{C}(\!(\hbar)\!)$-vector spaces.
-$\rho_{*}\mathsf{E}_{Y}$ is flat as a $\mathsf{W}_{S}$-module by Lemma 4.4. Therefore,
-$\mathcal{L}\otimes^{\mathbb{L}}_{\mathsf{W}_{S}}\mathcal{M}\cong\mathsf{B}_{Y}\otimes^{\mathbb{L}}_{\rho_{*}\mathsf{E}_{Y}\otimes\rho_{*}\mathsf{E}_{Y}}(\mathcal{L}\otimes\mathcal{M})$
-which has a differential structure by 4.11. ∎
+Consider the duality functor $\mathbb{D}(\mathcal{L})=\mathbb{R}\mathcal{H}om_{\widehat{\mathsf{W}}_{S}}(\mathcal{L},\widehat{\mathsf{W}}_{S})[\dim S/2]$.
+By [KashiwaraSchapira, Lemma 7.2.2] $\mathbb{D}(\mathcal{L})$ is a holonomic $\widehat{\mathsf{W}}_{S}$-module concentrated in degree 0 and we have
+$\mathbb{R}\mathcal{H}om_{\widehat{\mathsf{W}}_{S}}(\mathbb{D}(\mathcal{L}),\mathcal{M})[\dim S/2]\cong\mathcal{L}\otimes^{\mathbb{L}}_{\widehat{\mathsf{W}}_{S}}\mathcal{M}.$
+It is then shown in [KashiwaraSchapira, Theorem 7.2.3] that $\mathbb{R}\mathcal{H}om_{\widehat{\mathsf{W}}_{S}}(\mathbb{D}(\mathcal{L}),\mathcal{M})[\dim S/2]$ is a perverse sheaf of $\mathbf{C}(\!(\hbar)\!)$-vector spaces.
+$\rho_{*}\widehat{\mathsf{E}}_{Y}$ is flat as a $\widehat{\mathsf{W}}_{S}$-module by Lemma 4.4. Therefore,
+$\mathcal{L}\otimes^{\mathbb{L}}_{\widehat{\mathsf{W}}_{S}}\mathcal{M}\cong\mathsf{B}_{Y}\otimes^{\mathbb{L}}_{\rho_{*}\widehat{\mathsf{E}}_{Y}\otimes\rho_{*}\widehat{\mathsf{E}}_{Y}}(\mathcal{L}\otimes\mathcal{M})$
+which has a differential structure by 4.10. ∎
 Our next goal is to show that the differential structure has regular singularities.
 Theorem 4.12.
-Consider two regular holonomic modules $\mathcal{L},\mathcal{M}\in\mathrm{RH}_{S}$ and assume $\mathcal{L}$ is simple along a Lagrangian submanifold $L\subset S$. Then
-$\mathcal{L}\otimes^{\mathbb{L}}_{\mathsf{W}_{S}}\mathcal{M}$
+Consider two regular holonomic $\rho_{*}\widehat{\mathsf{E}}_{Y}$-modules and assume $\mathcal{L}$ is simple along a Lagrangian submanifold $L\subset S$. Then
+$\mathcal{L}\otimes^{\mathbb{L}}_{\widehat{\mathsf{W}}_{S}}\mathcal{M}$
 is a differential perverse sheaf with regular singularities.
@@
 Let $\rho\colon Y\rightarrow S$ be a contactification defined in a neighborhood of the union of supports of $\mathcal{L},\mathcal{M}$ and $\Lambda\subset Y$ a Legendrian lift of $L\subset S$.
-To define the differential structure we may choose $t\in\mathsf{E}_{Y}(0)$ whose principal symbol $\sigma_{0}(t)\in\mathcal{O}_{Y}$ vanishes along $\Lambda$.
-With this choice $D=\partial_{t}t=\hbar\partial_{\hbar}\in\mathsf{E}_{\Lambda/Y}$.
-By assumption $\mathcal{L}$ is an $\mathsf{E}_{Y}$-module simple along $\Lambda$.
-Thus, there is an $\mathsf{E}_{\Lambda/Y}$-lattice $\mathcal{L}_{0}\subset\mathcal{L}$ generating $\mathcal{L}$ over $\mathsf{E}_{Y}$ and $D$ acts on
-¯$\overline{\mathcal{L}}=\mathcal{L}_{0}/\mathsf{E}_{\Lambda/Y}(-1)\mathcal{L}_{0}$
+To define the differential structure we may choose $t\in\widehat{\mathsf{E}}_{Y}(0)$ whose principal symbol $\sigma_{0}(t)\in\mathcal{O}_{Y}$ vanishes along $\Lambda$.
+With this choice $D=\partial_{t}t=\hbar\partial_{\hbar}\in\widehat{\mathsf{E}}_{\Lambda/Y}$.
+By assumption $\mathcal{L}$ is an $\widehat{\mathsf{E}}_{Y}$-module simple along $\Lambda$.
+Thus, there is an $\widehat{\mathsf{E}}_{\Lambda/Y}$-lattice $\mathcal{L}_{0}\subset\mathcal{L}$ generating $\mathcal{L}$ over $\widehat{\mathsf{E}}_{Y}$ and $D$ acts on
+¯$\overline{\mathcal{L}}=\mathcal{L}_{0}/\widehat{\mathsf{E}}_{\Lambda/Y}(-1)\mathcal{L}_{0}$
 via multiplication by $\lambda\in\mathbf{C}$.
-Let $\mathcal{M}_{0}\subset\mathcal{M}|_{\Lambda}$ be a $\mathsf{E}_{\Lambda/Y}$-lattice such that no two eigenvalues of $D$ on
-¯$\overline{\mathcal{M}}=\mathcal{M}_{0}/\mathsf{E}_{\Lambda/Y}(-1)\mathcal{M}_{0}$
+Let $\mathcal{M}_{0}\subset\mathcal{M}|_{\Lambda}$ be a $\widehat{\mathsf{E}}_{\Lambda/Y}$-lattice such that no two eigenvalues of $D$ on
+¯$\overline{\mathcal{M}}=\mathcal{M}_{0}/\widehat{\mathsf{E}}_{\Lambda/Y}(-1)\mathcal{M}_{0}$
 differ by an integer.
-By [KS12, Lemma 7.1.3]
-$\mathcal{L}_{0}\otimes^{\mathbb{L}}_{\mathsf{W}_{L/S}}\mathcal{M}_{0}$
-is a $\mathbf{C}[\![\hbar]\!]$-lattice in $\mathcal{L}\otimes^{\mathbb{L}}_{\mathsf{W}_{S}}\mathcal{M}$.
-Since $D\in\mathsf{E}_{\Lambda/Y}$, this lattice is preserved by $D$.
+By [KashiwaraSchapira, Lemma 7.1.3]
+$\mathcal{L}_{0}\otimes^{\mathbb{L}}_{\widehat{\mathsf{W}}_{L/S}}\mathcal{M}_{0}$
+is a $\mathbf{C}[\![\hbar]\!]$-lattice in $\mathcal{L}\otimes^{\mathbb{L}}_{\widehat{\mathsf{W}}_{S}}\mathcal{M}$.
+Since $D\in\widehat{\mathsf{E}}_{\Lambda/Y}$, this lattice is preserved by $D$.
 Moreover, no two eigenvalues of $D$ on the $\mathbf{C}$-linear perverse sheaf
-$(\mathcal{L}_{0}\otimes^{\mathbb{L}}_{\mathsf{W}_{L/S}}\mathcal{M}_{0})/\hbar\cong\overline{\mathcal{L}}_{0}\otimes_{\mathsf{W}_{L/S}/\hbar}\overline{\mathcal{M}}_{0}$
+$(\mathcal{L}_{0}\otimes^{\mathbb{L}}_{\widehat{\mathsf{W}}_{L/S}}\mathcal{M}_{0})/\hbar\cong\overline{\mathcal{L}}_{0}\otimes_{\widehat{\mathsf{W}}_{L/S}/\hbar}\overline{\mathcal{M}}_{0}$
 differ by an integer. ∎
 We will now give a description of the above differential perverse sheaf using microlocalization.
-Consider the setting of 4.13 and suppose $L\subset S$ is lifted to a Legendrian $\Lambda\subset Y$ in a contactification $\rho\colon Y\rightarrow S$.
+Consider the setting of 4.12 and suppose $L\subset S$ is lifted to a Legendrian $\Lambda\subset Y$ in a contactification $\rho\colon Y\rightarrow S$.
 In this case we may identify $\tilde{\Lambda}\cong L\times\mathbf{C}^{\times}$.
@@
 $\mu^{dR}_{\Lambda}\mathcal{L}\otimes^{\mathbb{L}}_{\mathcal{D}^{\sqrt{v}}_{\tilde{\Lambda}}}\mu^{dR}_{\Lambda}\mathcal{M}$
-is a perverse sheaf on $\tilde{\Lambda}=L\times\mathbf{C}^{\times}$, lisse in the $\mathbf{C}^{\times}$-direction. Its restriction to $L\times\{1\}$ defines a perverse sheaf
+is a perverse sheaf on $\tilde{\Lambda}=L\times\mathbf{C}^{\times}$, which is locally constant along the $\mathbf{C}^{\times}$ orbits. Its restriction to $L\times\{1\}$ defines a perverse sheaf
 $(\mu^{dR}_{\Lambda}\mathcal{L}\otimes^{\mathbb{L}}_{\mathcal{D}^{\sqrt{v}}_{\tilde{\Lambda}}}\mu^{dR}_{\Lambda}\mathcal{M})|_{L}[-1]$
@@
 Theorem 4.13.
-Consider the setting of 4.13. There is an isomorphism of differential perverse sheaves
-$\mathcal{L}\otimes^{\mathbb{L}}_{\mathsf{W}_{S}}\mathcal{M}\cong\mathrm{RH}^{-1}\left((\mu^{dR}_{\Lambda}\mathcal{L}\otimes^{\mathbb{L}}_{\mathcal{D}^{\sqrt{v}}_{\tilde{\Lambda}}}\mu^{dR}_{\Lambda}\mathcal{M})|_{L}[-1]\right)$
+Consider the setting of 4.12. There is an isomorphism of differential perverse sheaves
+$\mathcal{L}\otimes^{\mathbb{L}}_{\widehat{\mathsf{W}}_{S}}\mathcal{M}\cong\mathrm{RH}^{-1}\left((\mu^{dR}_{\Lambda}\mathcal{L}\otimes^{\mathbb{L}}_{\mathcal{D}^{\sqrt{v}}_{\tilde{\Lambda}}}\mu^{dR}_{\Lambda}\mathcal{M})|_{L}[-1]\right)$
 on $L$.
 Proof.
-As in the proof of 4.13 we choose $\mathsf{E}_{\Lambda/Y}$-submodules $\mathcal{L}_{0}\subset\mathcal{L}$ and $\mathcal{M}_{0}\subset\mathcal{M}|_{\Lambda}$ generating these modules over $\mathsf{E}_{Y}|_{\Lambda}$ as well as an element $D=\partial_{t}t\in\mathsf{E}_{\Lambda/Y}$.
+As in the proof of 4.12 we choose $\widehat{\mathsf{E}}_{\Lambda/Y}$-submodules $\mathcal{L}_{0}\subset\mathcal{L}$ and $\mathcal{M}_{0}\subset\mathcal{M}|_{\Lambda}$ generating these modules over $\widehat{\mathsf{E}}_{Y}|_{\Lambda}$ as well as an element $D=\partial_{t}t\in\widehat{\mathsf{E}}_{\Lambda/Y}$.
 We may identify
-$\operatorname{gr}^{0}_{\Lambda}\mathsf{E}_{\Lambda/Y}\cong\mathcal{D}^{\sqrt{v}}_{L}[\theta],$
+$\operatorname{gr}^{0}_{\Lambda}\widehat{\mathsf{E}}_{\Lambda/Y}\cong\mathcal{D}^{\sqrt{v}}_{L}[\theta],$
 where $\theta$ is the Euler operator along $\Lambda$ which is the image of $D$ under the projection
-$\mathsf{E}_{\Lambda/Y}\longrightarrow\operatorname{gr}^{0}_{\Lambda}\mathsf{E}_{\Lambda/Y}.$
+$\widehat{\mathsf{E}}_{\Lambda/Y}\longrightarrow\operatorname{gr}^{0}_{\Lambda}\widehat{\mathsf{E}}_{\Lambda/Y}.$
 Consider
 ¯$\overline{\mathcal{L}}_{0}=\mathcal{L}_{0}/\hbar,\qquad\overline{\mathcal{M}}_{0}=\mathcal{M}_{0}/\hbar.$
-As $\hbar\mathsf{E}_{\Lambda/Y}(0)=\mathsf{E}_{\Lambda/Y}(-1)$, these are exactly the associated graded sheaves with respect to the microlocal $V$-filtration
+As $\hbar\widehat{\mathsf{E}}_{\Lambda/Y}(0)=\widehat{\mathsf{E}}_{\Lambda/Y}(-1)$, these are exactly the associated graded sheaves with respect to the microlocal $V$-filtration
 ¯$\overline{\mathcal{L}}_{0}\cong\operatorname{gr}^{0}_{\Lambda}\mathcal{L},\qquad\overline{\mathcal{M}}_{0}\cong\operatorname{gr}^{0}_{\Lambda}\mathcal{M}.$
 The complex of sheaves
-$\mathcal{L}_{0}\otimes^{\mathbb{L}}_{\mathsf{W}_{L/S}}\mathcal{M}_{0}$
-defines a $\mathbf{C}[\![\hbar]\!]$-lattice in $\mathcal{L}\otimes_{\mathsf{W}_{S}}\mathcal{M}$ preserved by $D$. Therefore, by 2.11 the claim will follow once we establish an isomorphism
+$\mathcal{L}_{0}\otimes^{\mathbb{L}}_{\widehat{\mathsf{W}}_{L/S}}\mathcal{M}_{0}$
+defines a $\mathbf{C}[\![\hbar]\!]$-lattice in $\mathcal{L}\otimes_{\widehat{\mathsf{W}}_{S}}\mathcal{M}$ preserved by $D$. Therefore, by 2.12 the claim will follow once we establish an isomorphism
 $\operatorname{gr}^{0}_{\Lambda}\mathcal{L}\otimes^{\mathbb{L}}_{\mathcal{D}^{\sqrt{v}}_{L}}\operatorname{gr}^{0}_{\Lambda}\mathcal{M}\cong(\mu^{dR}_{\Lambda}\mathcal{L}\otimes^{\mathbb{L}}_{\mathcal{D}^{\sqrt{v}}_{L\times\mathbf{C}^{\times}}}\mu^{dR}_{\Lambda}\mathcal{M})|_{L}[-1]$
@@
 $(-)^{\prime}\colon\mathsf{B}_{Y}\longrightarrow\mathsf{B}_{Y}$
-given by $Q_{1}\otimes Q_{2}\mapsto Q_{2}^{\prime}\otimes Q_{1}^{\prime}$ of $(\mathbf{C}(\!(\hbar)\!)\langle D\rangle,\rho_{*}\mathsf{E}_{Y}\otimes\rho_{*}\mathsf{E}_{Y})$-bimodules.
-For an object $\mathcal{M}\in\mathrm{RH}_{\mathrm{T}^{*}X}$ denote by $\mathcal{M}^{\prime}$ the same module with the ^$\widehat{\mathcal{E}}^{\sqrt{v}}_{X\times\mathbf{C}}$-action twisted by the involution $(-)^{\prime}$.
+given by $Q_{1}\otimes Q_{2}\mapsto Q_{2}^{\prime}\otimes Q_{1}^{\prime}$ of $(\mathbf{C}(\!(\hbar)\!)\langle D\rangle,\rho_{*}\widehat{\mathsf{E}}_{Y}\otimes\rho_{*}\widehat{\mathsf{E}}_{Y})$-bimodules.
+For a regular holonomic ^$\widehat{\mathcal{E}}^{\sqrt{v}}_{X\times\mathbf{C}}$-module $\mathcal{M}$ denote by $\mathcal{M}^{\prime}$ the same module with the ^$\widehat{\mathcal{E}}^{\sqrt{v}}_{X\times\mathbf{C}}$-action twisted by the involution $(-)^{\prime}$.
 Then the above involution of $\mathsf{B}_{Y}$ induces an isomorphism
@@
 Let $X$ be a complex manifold equipped with a holomorphic function $f\colon X\rightarrow\mathbf{C}$.
-Recall the perverse sheaf $\phi_{f}\in\mathrm{Perv}(f^{-1}(0);\mathbf{C})$ of vanishing cycles of the constant perverse sheaf $\mathbf{C}_{X}[\dim X]\in\mathrm{Perv}(X)$ [Dim04], which carries a monodromy automorphism $T$.
+Recall the perverse sheaf $\phi_{f}\in\mathrm{Perv}(f^{-1}(0);\mathbf{C})$ of vanishing cycles of the constant perverse sheaf $\mathbf{C}_{X}[\dim X]\in\mathrm{Perv}(X)$ [Dimca], which carries a monodromy automorphism $T$.
 In this section we give a microlocal description of this sheaf.
@@
 $\sigma_{\Lambda}(\sqrt{\omega}\partial_{t}^{\lambda-1/2}\delta(t-f))=\tau^{\lambda-1/2}\sqrt{d\tau\wedge\omega},$
-see [Kas03, Lemma 8.23]. Its order is
+see [KashiwaraDmodules, Lemma 8.23]. Its order is
 $\mathrm{ord}(\sqrt{\omega}\partial_{t}^{\lambda-1/2}\delta(t-f))=\lambda.$
-The standard projection $\rho\colon\mathrm{J}^{1}X\rightarrow\mathrm{T}^{*}X$ provides a contactification, so we may regard $\mathcal{C}^{\lambda,\sqrt{X}}_{f}$ as an object of $\mathrm{RH}_{\mathrm{T}^{*}X}$.
-This is a simple module along the Lagrangian $\Gamma_{df}\subset\mathrm{T}^{*}X$ of order $\lambda$.
 Consider the $V$-filtration along the hypersurface $\Gamma_{0}:t=0$ in $X\times\mathbf{C}$ and let
@@
 $\phi_{0}^{\mathcal{D}}$ is isomorphic to $\mathcal{O}_{X}$ equipped with the endomorphism $\theta=0$.
-The following statement is due to Kashiwara and Malgrange, see [Mal83, Théorème 6.2.6] and [Kas83, Theorem 2].
+The following statement is due to Kashiwara and Malgrange, see [Malgrange, Théorème 6.2.6] and [KashiwaraVanishing, Theorem 2].
 Proposition 5.5.
@@
 The second claim is merely the definition of the microlocalization functor. ∎
-Using this microlocal perspective we can prove a relationship between vanishing cycles and DQ modules. Recall from 4.13 that
+Using this microlocal perspective we can prove a relationship between vanishing cycles and DQ modules. Recall from 4.12 that
 $\mathcal{C}^{\mathrm{R}}_{0}\otimes^{\mathbb{L}}_{\widehat{\mathcal{W}}_{X}}\mathcal{C}_{f}$
@@
 For any $\lambda,\mu\in\mathbf{C}$ there is an isomorphism of differential perverse sheaves
-$M_{f}\colon\mathcal{C}^{\lambda,\mathrm{R}}_{0}\otimes^{\mathbb{L}}_{\widehat{\mathcal{W}}_{X}}\mathcal{C}^{\mu}_{f}\cong\widehat{\mathrm{RH}}^{-1}(\phi_{f}\otimes\mathbf{C}_{\mu+\lambda}),$
+$M_{f}\colon\mathcal{C}^{\lambda,\mathrm{R}}_{0}\otimes^{\mathbb{L}}_{\widehat{\mathcal{W}}_{X}}\mathcal{C}^{\mu}_{f}\cong\mathrm{RH}^{-1}(\phi_{f}\otimes\mathbf{C}_{\mu+\lambda}),$
 where $\mathbf{C}_{\mu+\lambda}$ is the trivial one-dimensional vector space equipped with the automorphism $\exp(2\pi i(\mu+\lambda))$.
@@
 $\mathcal{C}_{0}$ as a $\widehat{\mathcal{W}}_{X}$-module is simple along $L$.
-Using 4.14 we get an isomorphism
-$\mathcal{C}^{\lambda,\mathrm{R}}_{0}\otimes^{\mathbb{L}}_{\widehat{\mathcal{W}}_{X}}\mathcal{C}^{\mu}_{f}\cong\widehat{\mathrm{RH}}^{-1}\left((\mu^{dR}_{\Lambda}\mathcal{C}^{\lambda,\mathrm{R}}_{0}\otimes^{\mathbb{L}}_{\mathcal{D}_{\tilde{\Lambda}}}\mu^{dR}_{\Lambda}\mathcal{C}^{\mu}_{f})|_{L}[-1]\right)$
+Using 4.13 we get an isomorphism
+$\mathcal{C}^{\lambda,\mathrm{R}}_{0}\otimes^{\mathbb{L}}_{\widehat{\mathcal{W}}_{X}}\mathcal{C}^{\mu}_{f}\cong\mathrm{RH}^{-1}\left((\mu^{dR}_{\Lambda}\mathcal{C}^{\lambda,\mathrm{R}}_{0}\otimes^{\mathbb{L}}_{\mathcal{D}_{\tilde{\Lambda}}}\mu^{dR}_{\Lambda}\mathcal{C}^{\mu}_{f})|_{L}[-1]\right)$
 of differential perverse sheaves.
@@
 Under this morphism the action of $D=\partial_{t}t=\hbar^{-1}t$ on the left goes to the action of $D=\hbar^{-1}(\hbar^{2}\partial_{\hbar}-f)=\hbar\partial_{\hbar}-\hbar^{-1}f$ on the right.
-By [SS14, Proposition 2] the induced morphism
+By [SabbahSaito, Proposition 2] the induced morphism
 $K_{X}(\!(\hbar)\!)\otimes^{\mathbb{L}}_{\mathcal{D}_{X}(\!(\hbar)\!)}\mathcal{C}_{f}|_{X}\longrightarrow\mathrm{DR}_{f}$
@@
 There is an isomorphism of differential perverse sheaves
-$(\mathrm{DR}_{f},\hbar\partial_{\hbar}-\hbar^{-1}f)\cong\widehat{\mathrm{RH}}^{-1}(\phi_{f}).$
+$(\mathrm{DR}_{f},\hbar\partial_{\hbar}-\hbar^{-1}f)\cong\mathrm{RH}^{-1}(\phi_{f}).$
 Proof.
@@
 Remark 5.12.
-Such an isomorphism was established in [Sab10, SS14] on the level of cohomology sheaves and was recently upgraded in [Sch23] to an isomorphism of differential perverse sheaves, by different methods than above.
+Such an isomorphism was established in [Sabbah, SabbahSaito] on the level of cohomology sheaves and was recently upgraded in [Schefers] to an isomorphism of differential perverse sheaves, by different methods than above.
 5.3. Thom–Sebastiani theorem
@@
 We denote by $\alpha\colon\mathrm{J}^{1}X_{1}\times\mathrm{J}^{1}X_{2}\to\mathrm{J}^{1}(X_{1}\times X_{2})$ the addition map.
-Consider the corresponding transfer $(\alpha^{-1}\widehat{\mathcal{E}}_{X_{1}\times X_{2}\times\mathbf{C}_{t}},\widehat{\mathcal{E}}_{X_{1}\times\mathbf{C}_{t_{1}}\times X_{2}\times\mathbf{C}_{t_{2}}})$-bimodule (see [SKK73, Section 1.3] for the general notion of transfer bimodules)
+Consider the corresponding transfer $(\alpha^{-1}\widehat{\mathcal{E}}_{X_{1}\times X_{2}\times\mathbf{C}_{t}},\widehat{\mathcal{E}}_{X_{1}\times\mathbf{C}_{t_{1}}\times X_{2}\times\mathbf{C}_{t_{2}}})$-bimodule (see [SatoKawaiKashiwara, Section 1.3] for the general notion of transfer bimodules)
 ^:=^$\widehat{\mathcal{E}}_{\ast}:=\frac{\widehat{\mathcal{E}}_{X_{1}\times\mathbf{C}_{t_{1}}\times X_{2}\times\mathbf{C}_{t_{2}}}}{(\partial_{t_{1}}-\partial_{t_{2}})\widehat{\mathcal{E}}_{X_{1}\times\mathbf{C}_{t_{1}}\times X_{2}\times\mathbf{C}_{t_{2}}}}.$
@@
 $\phi^{\mathcal{D}}_{f_{i}}=\operatorname{gr}^{0}_{V}\mathcal{B}_{f_{i}}\longrightarrow\operatorname{gr}^{0}_{V}\widetilde{\mathcal{B}}_{f_{i}}\longrightarrow\operatorname{gr}^{0}_{V}\mathcal{C}_{f_{i}}.$
-Indeed, for the first map this is shown in [MSS20, Section 1.1] and for the composite this is (2).
+Indeed, for the first map this is shown in [MaximSaitoSchurmann, Section 1.1] and for the composite this is (2).
 The $V$-filtrations on ~$\widetilde{\mathcal{B}}_{f_{i}}$ induce by convolution a filtration on ~$\widetilde{\mathcal{B}}_{f_{1}}\stackrel{{\scriptstyle\mathcal{D}}}{{\boxtimes}}\widetilde{\mathcal{B}}_{f_{2}}$.
-It is shown in [MSS20, Theorem 1.2] and [SS, Lemma 11.8.2] that the natural map
+It is shown in [MaximSaitoSchurmann, Theorem 1.2] and [SabbahSchnell, Lemma 11.8.2] that the natural map
 ~$\widetilde{\mathcal{B}}_{f_{1}}\stackrel{{\scriptstyle\mathcal{D}}}{{\boxtimes}}\widetilde{\mathcal{B}}_{f_{2}}\longrightarrow\widetilde{\mathcal{B}}_{f_{1}\boxplus f_{2}}$
@@
 Proof.
-Unwinding the construction of $M_{f}$ in 5.7 via microlocalization, the top line of the diagram under consideration may be computed as the image under $\widehat{\mathrm{RH}}^{-1}$ of a sequence of morphisms of perverse sheaves with automorphism:
+Unwinding the construction of $M_{f}$ in 5.7 via microlocalization, the top line of the diagram under consideration may be computed as the image under $\mathrm{RH}^{-1}$ of a sequence of morphisms of perverse sheaves with automorphism:
 $(K_{X_{1}}\otimes_{\mathcal{D}_{X_{1}}}^{\mathbb{L}}\phi_{f_{1}}^{\mathcal{D}})\boxtimes(K_{X_{2}}\otimes_{\mathcal{D}_{X_{2}}}^{\mathbb{L}}\phi_{f_{2}}^{\mathcal{D}})\to(K_{X_{1}}\stackrel{{\scriptstyle\mathcal{D}}}{{\boxtimes}}K_{X_{2}})\otimes_{\mathcal{D}_{X_{1}\times X_{2}}}^{\mathbb{L}}(\phi_{f_{1}}\stackrel{{\scriptstyle\mathcal{D}}}{{\boxtimes}}\phi_{f_{2}})\to K_{X_{1}\times X_{2}}\otimes_{\mathcal{D}_{X_{1}\times X_{2}}}^{\mathbb{L}}\phi_{f_{1}\boxplus f_{2}}.$
@@
 6.2. d-critical loci
-We will use the notions related to d-critical loci from [Bra+15, Joy15] which we now briefly recall. Let us first explain how to construct exact two-forms of degree $-1$ on critical loci.
+We will use the notions related to d-critical loci from [BBDJS, JoycedCrit] which we now briefly recall. Let us first explain how to construct exact two-forms of degree $-1$ on critical loci.
 Proposition 6.3.
@@
 Given a d-critical locus $(X,s)$, an embedding of critical charts $(R,U,f,i)\hookrightarrow(S,V,g,j)$, where $R\subset S$, is a locally closed embedding $\Xi\colon U\hookrightarrow V$ compatible with the rest of the data.
-We have the following important theorems proved in [Joy15]:
+We have the following important theorems proved in [JoycedCrit]:
   • Given an embedding $\Xi\colon(R,U,f,i)\hookrightarrow(S,V,g,j)$ of critical charts there is a canonical quadratic form $q_{\Xi}$ on the normal bundle $N_{\Xi}|_{R}$ of $U\subset V$.
 The quadratic vector bundle $N_{\Xi}|_{R}$ defines a $\mathbf{Z}/2$-torsor $P_{\Xi}$ on $R$ of orientations of $q_{\Xi}$.
-See [Joy15, Definition 2.26].
+See [JoycedCrit, Definition 2.26].
   • An embedding of critical charts $(R,U,f,i)\hookrightarrow(S,V,g,j)$ is, locally, of the form $(R,U,f,i)\hookrightarrow(R,U\times\mathbf{C}^{n},f\boxplus(z_{1}^{2}+\dots+z_{n}^{2}),j)$, where $U\hookrightarrow U\times\mathbf{C}^{n}$ is given by $u\mapsto(u,0)$.
-See [Joy15, Proposition 2.22].
+See [JoycedCrit, Proposition 2.22].
   • Given critical charts $(R,U,f,i)$ and $(S,V,g,j)$ and an intersection point $x\in R\cap S$ we can find near $x$ another critical chart $(T,W,h,k)$ and local embeddings $(R,U,f,i)\hookrightarrow(T,W,h,k)$ and $(S,V,g,j)\hookrightarrow(T,W,h,k)$.
-See [Joy15, Theorem 2.20].
+See [JoycedCrit, Theorem 2.20].
 Another important structure on a d-critical locus $(X,s)$ is the virtual canonical bundle $K^{\mathrm{vir}}_{X}$, which is a holomorphic line bundle on $X^{\mathrm{red}}$ with the following properties:
@@
 Consider the intersection $L\cap M$ as a complex analytic space.
-The following is a generalization of 6.3 proved in [Bus14].
+The following is a generalization of 6.3 proved in [Bussi].
 Proposition 6.7.
@@
 This proves that $s$ defines a d-critical structure on $L\cap M$.
-Moreover, it is clear that this d-critical structure on $L\cap M$ coincides with the one constructed in [Bus14, Theorem 3.1], where the virtual canonical bundle is identified in the same way. ∎
+Moreover, it is clear that this d-critical structure on $L\cap M$ coincides with the one constructed in [Bussi, Theorem 3.1], where the virtual canonical bundle is identified in the same way. ∎
 Remark 6.8.
@@
 Let $(X,s)$ be an oriented d-critical locus.
-The paper [Bra+15] constructs a canonical perverse sheaf $\phi_{X}\in\mathrm{Perv}_{\mathrm{aut}}(X)$ equipped with an automorphism on $X$, which we call the DT sheaf, locally modeled on the sheaf of vanishing cycles.
+The paper [BBDJS] constructs a canonical perverse sheaf $\phi_{X}\in\mathrm{Perv}_{\mathrm{aut}}(X)$ equipped with an automorphism on $X$, which we call the DT sheaf, locally modeled on the sheaf of vanishing cycles.
 Let us briefly recall its construction.
@@
 Any two choices of the isomorphisms $T_{q}$ compatible with the Thom–Sebastiani isomorphism (6) differ by a multiplication $\alpha^{\dim W}$ for $\alpha\in\mathbf{C}$.
-Let $\Xi\colon(R,U,f,i)\hookrightarrow(S,V,g,j)$ be an embedding of critical charts. Then [Bra+15, Theorem 5.4] shows that there is the stabilization isomorphism
+Let $\Xi\colon(R,U,f,i)\hookrightarrow(S,V,g,j)$ be an embedding of critical charts. Then [BBDJS, Theorem 5.4] shows that there is the stabilization isomorphism
 (11) $\Theta_{\Xi}\colon\mathcal{PV}_{U,f}\longrightarrow(\mathcal{PV}_{V,g})|_{\mathrm{Crit}(f)}\otimes_{\mathbf{Z}/2}P_{\Xi}$
 of perverse sheaves equipped with automorphisms which has the following description:
-  • Locally near a point of $R$ identify $V\cong R\times W$ and $g\mapsto f\boxplus q$ for a quadratic form $q\colon W\rightarrow\mathbf{C}$ using [Joy15, Proposition 2.22], so that under this isomorphism the embedding becomes $U\hookrightarrow U\times W$ given by $u\mapsto(u,0)$.
+  • Locally near a point of $R$ identify $V\cong R\times W$ and $g\mapsto f\boxplus q$ for a quadratic form $q\colon W\rightarrow\mathbf{C}$ using [JoycedCrit, Proposition 2.22], so that under this isomorphism the embedding becomes $U\hookrightarrow U\times W$ given by $u\mapsto(u,0)$.
 In this local model the isomorphism $\Theta_{\Xi}$ is equal to the composite
@@
   • In any critical chart $(R,U,f,i)$ there is an isomorphism $\phi_{X}|_{R}\cong\mathcal{PV}_{U,f}\otimes_{\mathbf{Z}/2}Q_{R,U,f,i}$.
-  • By [Joy15, Theorem 2.20] any two critical charts $(R,U,f,i)$, $(S,V,g,j)$ are locally related by a zig-zag $(R,U,f,i)\hookrightarrow(T,W,h,k)\hookleftarrow(S,V,g,j)$ of critical embeddings.
+  • By [JoycedCrit, Theorem 2.20] any two critical charts $(R,U,f,i)$, $(S,V,g,j)$ are locally related by a zig-zag $(R,U,f,i)\hookrightarrow(T,W,h,k)\hookleftarrow(S,V,g,j)$ of critical embeddings.
 The isomorphism of the local models for $\phi_{X}$ in charts $(R,U,f,i)$ and $(S,V,g,j)$ is given by applying the stabilization isomorphism (12) to the two critical embeddings.
@@
 Given two such isomorphisms differing by a multiplication by $\alpha^{\dim W}$ the resulting sheaves $\phi_{X}$ are canonically isomorphic: the isomorphism is given by a multiplication by $\alpha^{\dim U}$ in each critical chart $(R,U,f,i)$.
+We denote by $\phi^{\mathcal{D}}_{X}\in\mathrm{Mod}_{\mathrm{rh},\mathrm{aut}}(\mathcal{D}_{X})$ the $D$-module corresponding to $\phi_{X}$ under the Riemann–Hilbert correspondence.
 7. Comparison of the two perverse sheaves
-Throughout this section $S$ denotes a complex symplectic manifold, $\rho\colon Y\rightarrow S$ its contactification, $L,M\subset S$ Lagrangian submanifolds equipped with orientation data and $\Lambda_{L},\Lambda_{M}\subset Y$ their Legendrian lifts.
+7.1. Statement
+Throughout this section $S$ denotes a complex symplectic manifold, $\rho\colon Y\rightarrow S$ its contactification, $L,M\subset S$ Lagrangian submanifolds equipped with orientation data $K_{L}^{1/2},K_{M}^{1/2}$ and $\Lambda_{L},\Lambda_{M}\subset Y$ their Legendrian lifts.
 The Lagrangian intersection $L\cap M$ has a structure of an oriented d-critical locus by 6.7.
+By 3.16 microlocalization establishes an equivalence
+$\mu^{dR}_{\Lambda_{L}}\colon\mathrm{Mod}_{\Lambda,\mathrm{rh}}(\widehat{\mathsf{E}}_{Y})\xrightarrow{\sim}\gamma_{*}\mathrm{Mod}_{\mathrm{lisse}}(\mathcal{D}^{\sqrt{v}}_{\tilde{\Lambda}_{L}}).$
+We may canonically identify $\tilde{\Lambda}_{L}\cong\Lambda_{L}\times\mathbf{C}^{\times}\cong L\times\mathbf{C}^{\times}$.
+Let $i_{L}\colon L\rightarrow\tilde{\Lambda}_{L}$ be the inclusion of $1\times\mathbf{C}^{\times}$ and similarly for $i_{M}\colon M\rightarrow\tilde{\Lambda}_{M}$.
+As explained in Section 2.1, for a monodromic (twisted) $D$-module $M$ on $\tilde{\Lambda}_{L}$ the pullback $i_{L}^{*}M$ is a (twisted) $D$-module on $L$ equipped with an automorphism.
+Fix $\lambda\in\mathbf{C}$. Then, up to isomorphism, there is a unique $\widehat{\mathsf{E}}_{Y}$-module $\mathsf{M}^{\lambda}_{L}$ simple along $\Lambda_{L}$ such that
+$i_{L}^{*}\mu_{\Lambda_{L}}^{dR}(\mathsf{M}^{\lambda}_{L})\cong K_{L}^{1/2}\otimes\mathbf{C}_{\lambda}.$
+Similarly, for $\mu\in\mathbf{C}$ there is a unique $\widehat{\mathsf{E}}_{Y}$-module $\mathsf{M}^{\mu}_{M}$ simple along $\Lambda_{M}$ such that
+$i_{M}^{*}\mu_{\Lambda_{M}}^{dR}(\mathsf{M}^{\mu}_{M})\cong K_{M}^{1/2}\otimes\mathbf{C}_{\mu}.$
 The goal of this section is to prove the following result.
@@
 Let $\lambda,\mu\in\mathbf{C}$. There is an isomorphism of differential perverse sheaves
-$\mathsf{M}^{\lambda}_{L}\otimes^{\mathbb{L}}_{\mathsf{W}_{S}}\mathsf{M}^{\mu}_{M}\cong\mathrm{RH}^{-1}(\phi_{L\cap M}\otimes\mathbf{C}_{\mu+\lambda}),$
+$\mathsf{M}^{\lambda}_{L}\otimes^{\mathbb{L}}_{\widehat{\mathsf{W}}_{S}}\mathsf{M}^{\mu}_{M}\cong\mathrm{RH}^{-1}(\phi_{L\cap M}\otimes\mathbf{C}_{\mu+\lambda}),$
 where $\mathbf{C}_{\mu+\lambda}$ is the trivial vector space with the monodromy automorphism $T=\exp(2\pi i(\mu+\lambda))$.
-Let $\dim L=\dim M=n$.
+Corollary 7.2.
+Let $\mu\in\mathbf{C}$. There is an isomorphism
+$i_{L}^{*}\mu^{dR}_{\Lambda_{L}}\mathsf{M}^{\mu}_{M}\cong\phi^{\mathcal{D}}_{L\cap M}\otimes\mathbf{C}_{\mu}$
+of $\mathcal{D}_{L}$-modules equipped with automorphisms, where we identify $\mathcal{D}^{\sqrt{v}}_{L}$-modules with $\mathcal{D}_{L}$-modules using the orientation data $K_{L}^{1/2}$.
+Proof.
+Pick any $\lambda\in\mathbf{C}$. By 7.1 we have
+$\mathsf{M}^{\lambda}_{L}\otimes^{\mathbb{L}}_{\widehat{\mathsf{W}}_{S}}\mathsf{M}^{\mu}_{M}\cong\mathrm{RH}^{-1}(\phi_{L\cap M}\otimes\mathbf{C}_{\mu+\lambda}).$
+By 4.13 we have
+$\mathsf{M}^{\lambda}_{L}\otimes^{\mathbb{L}}_{\widehat{\mathsf{W}}_{S}}\mathsf{M}^{\mu}_{M}\cong\mathrm{RH}^{-1}\left((\mu^{dR}_{\Lambda_{L}}\mathsf{M}^{\lambda}_{L}\otimes^{\mathbb{L}}_{\mathcal{D}^{\sqrt{v}}_{\tilde{\Lambda}_{L}}}\mu^{dR}_{\Lambda_{L}}\mathsf{M}^{\mu}_{M})|_{L}[-1]\right).$
+By definition $i^{*}\mu^{dR}_{\Lambda_{L}}\mathsf{M}^{\lambda}_{L}\cong K_{L}^{1/2}$ equipped with the automorphism $\exp(2\pi i\lambda)$. Therefore, we obtain an isomorphism
+$\mathbf{C}_{\mu+\lambda}\otimes\phi_{L\cap M}\cong\mathbf{C}_{\lambda}\otimes(K_{L}\otimes^{\mathbb{L}}_{\mathcal{D}_{\tilde{\Lambda}_{L}}}\mu^{dR}_{\Lambda_{L}}\mathsf{M}^{\mu}_{M})|_{L}[-1]$
+of perverse sheaves on $L$ equipped with an automorphism. Identifying $D$-modules with perverse sheaves via the Riemann–Hilbert correspondence we get the result. ∎
+Throughout this section we let $\dim L=\dim M=n$.
 7.2. d-critical charts
@@
 The manifold $S$ admits local polarizations.
-We use them to get a local description of $(Y\rightarrow S,\Lambda_{L}\rightarrow L,\Lambda_{M}\rightarrow M)$ following the terminology from [Bus14, Section 2]:
+We use them to get a local description of $(Y\rightarrow S,\Lambda_{L}\rightarrow L,\Lambda_{M}\rightarrow M)$ following the terminology from [Bussi, Section 2]:
   • An $L$-chart is an open subset $U\subset S$ equipped with a polarization, where we use 1.27 to identify $\rho^{-1}(U)\rightarrow U$ with an open subset of $\mathrm{J}^{1}L\rightarrow\mathrm{T}^{*}L$, $\Lambda_{L}\rightarrow L$ with $j^{1}0\rightarrow\Gamma_{0}$ and $\Lambda_{M}\rightarrow M$ with $j^{1}f\rightarrow\Gamma_{df}$ for a locally defined function $f\colon L\rightarrow\mathbf{C}$ such that $f|_{(L\cap M)^{\mathrm{red}}}=0$.
@@
 Let us now give a local description of deformation quantization algebras and modules in the $L$- and $M$-charts from before.
-  • In the $L$-chart we have $\mathsf{E}_{Y}=\widehat{\mathcal{E}}^{\sqrt{v}}_{L\times\mathbf{C}}$ and $\mathsf{W}_{S}=\widehat{\mathcal{W}}^{\sqrt{v}}_{L}$. The modules are
+  • In the $L$-chart we have $\widehat{\mathsf{E}}_{Y}=\widehat{\mathcal{E}}^{\sqrt{v}}_{L\times\mathbf{C}}$ and $\widehat{\mathsf{W}}_{S}=\widehat{\mathcal{W}}^{\sqrt{v}}_{L}$. The modules are
     $\mathsf{M}^{\lambda}_{L}=\mathcal{C}^{\lambda,\sqrt{L}}_{0},\qquad\mathsf{M}^{\mu}_{M}=\mathcal{C}^{\mu,\sqrt{M}}_{f}.$
@@
     $\sigma_{\Lambda_{L}}(u_{L}^{L})=\tau^{\lambda-1/2}\sqrt{d\tau\wedge\omega_{L}},\qquad\sigma_{\Lambda_{M}}(u_{M}^{L})=\tau^{\mu-1/2}\sqrt{d\tau\wedge\omega_{M}}.$
-  • In the $M$-chart we have $\mathsf{E}_{Y}=\widehat{\mathcal{E}}^{\sqrt{v}}_{M\times\mathbf{C}}$ and $\mathsf{W}_{S}=\widehat{\mathcal{W}}^{\sqrt{v}}_{M}$. The modules are
+  • In the $M$-chart we have $\widehat{\mathsf{E}}_{Y}=\widehat{\mathcal{E}}^{\sqrt{v}}_{M\times\mathbf{C}}$ and $\widehat{\mathsf{W}}_{S}=\widehat{\mathcal{W}}^{\sqrt{v}}_{M}$. The modules are
     $\mathsf{M}^{\lambda}_{L}=\mathcal{C}^{\lambda,\sqrt{L}}_{g},\qquad\mathsf{M}^{\mu}_{M}=\mathcal{C}^{\mu,\sqrt{M}}_{0}.$
@@
 7.4. Proof of 7.1
-We use the gluing data for the perverse sheaf $\phi_{L\cap M}$ as described in [Bus14, Section 2.3].
+We use the gluing data for the perverse sheaf $\phi_{L\cap M}$ as described in [Bussi, Section 2.3].
 Cover $S$ by open sets which admit polarizations and consider them as $L$-charts.
@@
 We have that $\Xi\colon(L\cap M,L,f,i)\rightarrow(L\cap M,L\times M,h,i)$ is a critical embedding.
-Therefore, by [Joy15, Proposition 2.22] in a neighborhood of $x$ there is a local isomorphism $L\times M\cong L\times W$, where $W=\mathrm{T}_{x}M$ is equipped with the Maslov quadratic form $q=q(\mathrm{T}_{M},\mathrm{T}_{\pi_{1}},\mathrm{T}_{\pi_{2}})\colon W\rightarrow\mathbf{C}$, such that under this isomorphism $h\mapsto f\boxplus q$ and the embedding $\Xi\colon L\rightarrow L\times M$ becomes $(\mathrm{id}\times 0)\colon L\rightarrow L\times W$.
+Therefore, by [JoycedCrit, Proposition 2.22] in a neighborhood of $x$ there is a local isomorphism $L\times M\cong L\times W$, where $W=\mathrm{T}_{x}M$ is equipped with the Maslov quadratic form $q=q(\mathrm{T}_{M},\mathrm{T}_{\pi_{1}},\mathrm{T}_{\pi_{2}})\colon W\rightarrow\mathbf{C}$, such that under this isomorphism $h\mapsto f\boxplus q$ and the embedding $\Xi\colon L\rightarrow L\times M$ becomes $(\mathrm{id}\times 0)\colon L\rightarrow L\times W$.
 The $\mathbf{Z}/2$-torsor $P_{\Xi}$ in a neighborhood of $x$ canonically identifies with $P_{q}$.
@@
 $\iota_{L\cap M,L\cap M,0,\mathrm{id}}\colon K^{\mathrm{vir}}_{L\cap M}\longrightarrow(K_{L\cap M})^{\otimes 2}$
-described in [Joy15, Theorem 2.28 (i)] for a general d-critical locus. By [Bus14, Theorem 3.1] we may additionally identify
+described in [JoycedCrit, Theorem 2.28 (i)] for a general d-critical locus. By [Bussi, Theorem 3.1] we may additionally identify
 $K^{\mathrm{vir}}_{L\cap M}\cong K_{L}|_{L\cap M}\otimes K_{M}|_{L\cap M}.$
@@
 $0\longrightarrow\mathrm{T}_{L\cap M}\longrightarrow\mathrm{T}_{L}|_{L\cap M}\oplus\mathrm{T}_{M}|_{L\cap M}\longrightarrow T_{S}|_{L\cap M}\longrightarrow\mathrm{T}^{*}_{L\cap M}\longrightarrow 0$
-Taking its determinant and trivializing $\det(\mathrm{T}_{S})$ using the symplectic volume form $\mathrm{vol}_{S}$ we get the isomorphism $\iota^{\prime}$.
+Taking its determinant and trivializing $\det(\mathrm{T}_{S})$, using the symplectic volume form $\mathrm{vol}_{S}$ we get the isomorphism $\iota^{\prime}$.
+8. A category associated to a holomoprhic symplectic manifold
+In this section we will define a sheaf of dg categories associated to a holomorphic symplectic manifold which behaves as a “Fukaya category” of holomorphic Lagrangians as conjectured by Joyce.
+8.1. Holomorphic microlocal operators and their modules
+Let $X$ be a complex manifold.
+In Section 3.1 we have discussed the sheaf ^$\widehat{\mathcal{E}}_{X}$ on $\mathrm{P}^{*}X$ of formal microdifferential operators.
+In this section we discuss several of its versions:
+  • There is a sheaf $\mathcal{E}_{X}$ of microdifferential operators on $\mathrm{P}^{*}X$ introduced in [SatoKawaiKashiwara]. It is endowed with a filtration
+    $\mathcal{E}_{X}=\cup_{m\in\mathbf{Z}}\mathcal{E}_{X}(m)$
+    by the order of the microdifferential operator, so that
+    ^$\widehat{\mathcal{E}}_{X}(m)=\lim_{k}\mathcal{E}_{X}(m)/\mathcal{E}_{X}(k)$
+    is the completion.
+In particular, there is a natural embedding $\mathcal{E}_{X}\rightarrow\widehat{\mathcal{E}}_{X}$ which is faithfully flat by [SatoKawaiKashiwara, Chapter II, Theorem 3.4.1].
+There is a half-twisted version $\mathcal{E}^{\sqrt{v}}_{X}$ of the sheaf $\mathcal{E}_{X}$ defined analogously to ^$\widehat{\mathcal{E}}^{\sqrt{v}}_{X}$.
+  • There is a sheaf $\mathcal{E}^{\mathbf{R},f}_{X}$ of finite order holomorphic microlocal operators on $\mathring{\mathrm{T}}^{*}X$ introduced in [Andronikof1] with a natural embedding $\gamma^{-1}\mathcal{E}_{X}\rightarrow\mathcal{E}^{\mathbf{R},f}_{X}$ which is faithfully flat by [Andronikof1, Théorème 5.3.4] and which identifies $\mathcal{E}_{X}\cong\gamma_{*}\mathcal{E}^{\mathbf{R},f}_{X}$.
+There is also a half-twisted version $\mathcal{E}^{\sqrt{v},\mathbf{R},f}_{X}$ of $\mathcal{E}^{\mathbf{R},f}_{X}$.
+There is a notion of a (regular) holonomic $\mathcal{E}_{X}$-module defined analogously to the notion of (regular) holonomic ^$\widehat{\mathcal{E}}_{X}$-modules from 3.3.
+The induction functor ^$\mathcal{M}\mapsto\widehat{\mathcal{E}}_{X}\otimes_{\mathcal{E}_{X}}\mathcal{M}$ sends (regular) holonomic $\mathcal{E}_{X}$-modules to (regular) holonomic ^$\widehat{\mathcal{E}}_{X}$-modules.
+The following is [KashiwaraKawaiMicrolocal, Theorem 6.3].
+Proposition 8.1.
+Let $X$ be a complex manifold and $\mathcal{L},\mathcal{M}$ two regular holonomic $\mathcal{E}_{X}$-modules. Then the natural morphism
+$\mathbb{R}\mathcal{H}om_{\mathcal{E}_{X}}(\mathcal{L},\mathcal{M})\longrightarrow\mathbb{R}\mathcal{H}om_{\mathcal{E}_{X}}(\mathcal{L},\widehat{\mathcal{E}}_{X}\otimes_{\mathcal{E}_{X}}\mathcal{M})$
+is an isomorphism.
+Quantized contact transformations from Section 3.5 work equally well for $\mathcal{E}_{X}$ and $\mathcal{E}^{\mathbf{R},f}_{X}$ (see [SatoKawaiKashiwara] for the former and [Andronikof1, Section 5.1] for the latter).
+Namely, suppose $V_{X_{1}}\subset\mathrm{P}^{*}X_{1}$ and $V_{X_{2}}\subset\mathrm{P}^{*}X_{2}$ are open subsets together with a contactomorphism $\psi\colon V_{X_{1}}\xrightarrow{\sim}V_{X_{2}}$ and denote by $\tilde{\psi}\colon\tilde{V}_{X_{1}}\xrightarrow{\sim}\tilde{V}_{X_{2}}$ the corresponding homogeneous symplectomorphism of open subsets $\tilde{V}_{X_{1}}\subset\mathring{\mathrm{T}}^{*}X_{1}$ and $\tilde{V}_{X_{2}}\subset\mathring{\mathrm{T}}^{*}X_{2}$.
+Then there is a $\ast$-preserving quantized contact transformation over $\psi$
+$\Psi\colon\widehat{\mathcal{E}}^{\sqrt{v}}_{X_{2}}|_{V_{X_{2}}}\xrightarrow{\sim}\psi_{*}\widehat{\mathcal{E}}^{\sqrt{v}}_{X_{1}}|_{V_{X_{1}}}$
+which restricts to an isomorphism
+$\Psi\colon\mathcal{E}^{\sqrt{v}}_{X_{2}}|_{V_{X_{2}}}\xrightarrow{\sim}\psi_{*}\mathcal{E}^{\sqrt{v}}_{X_{1}}|_{V_{X_{1}}}$
+and extends to an isomorphism
+$\tilde{\Psi}\colon\mathcal{E}^{\sqrt{v},\mathbf{R},f}_{X_{2}}|_{\tilde{V}_{X_{2}}}\xrightarrow{\sim}\tilde{\psi}_{*}\mathcal{E}^{\sqrt{v},\mathbf{R},f}_{X_{1}}|_{\tilde{V}_{X_{1}}}.$
+In particular, if $Y$ is a holomorphic contact manifold with $\gamma\colon\tilde{Y}\rightarrow Y$ its symplectization, then there is an algebroid $\mathsf{E}_{Y}$ on $Y$ and an algebroid $\mathsf{E}^{\mathbf{R},f}_{\tilde{Y}}$ on $\tilde{Y}$, locally given by $\mathcal{E}^{\sqrt{v}}_{X}$ and $\mathcal{E}^{\sqrt{v},\mathbf{R},f}_{X}$, together with embeddings
+$\mathsf{E}_{Y}\subset\widehat{\mathsf{E}}_{Y},\qquad\gamma^{-1}\mathsf{E}_{Y}\subset\mathsf{E}^{\mathbf{R},f}_{\tilde{Y}}.$
+Analogously to $\widehat{\mathsf{E}}_{Y}$-modules, for a Legendrian subvariety $\Lambda\subset Y$ one may define a microlocalization functor
+$\mu_{\Lambda}^{dR}\colon\mathrm{Mod}_{\mathrm{rh}}(\mathsf{E}_{Y})\longrightarrow\gamma_{*}\mathrm{Mod}_{\mathrm{rh}}(\mathcal{D}^{\sqrt{v}}_{\tilde{\Lambda}})$
+either by precomposing the microlocalization functor $\mu_{\Lambda}^{dR}$ of $\widehat{\mathsf{E}}_{Y}$-modules with the induction functor along the inclusion $\mathsf{E}_{Y}\subset\widehat{\mathsf{E}}_{Y}$ or by constructing the subalgebroid $\mathsf{E}_{\Lambda/Y}\subset\mathsf{E}_{Y}|_{\Lambda}$ analogously to $\widehat{\mathsf{E}}_{\Lambda/Y}\subset\widehat{\mathsf{E}}_{Y}|_{\Lambda}$.
+Moreover, if $\Lambda\subset Y$ is a Legendrian submanifold, 3.16 still holds for $\mathsf{E}_{Y}$-modules: microlocalization defines an equivalence of stacks
+$\mu_{\Lambda}^{dR}\colon\mathrm{Mod}_{\Lambda,\mathrm{rh}}(\mathsf{E}_{Y})\xrightarrow{\sim}\gamma_{*}\mathrm{Mod}_{\mathrm{lisse}}(\mathcal{D}^{\sqrt{v}}_{\tilde{\Lambda}}).$
+For an $\mathsf{E}_{Y}$-module $\mathcal{M}$ we denote
+$\mathcal{M}^{\mathbf{R},f}=\mathsf{E}^{\mathbf{R},f}_{\tilde{Y}}\otimes_{\gamma^{-1}\mathsf{E}_{Y}}\gamma^{-1}\mathcal{M}$
+the corresponding $\mathsf{E}^{\mathbf{R},f}_{\tilde{Y}}$-module. We have the following result.
+Proposition 8.2.
+Let $Y$ be a contact manifold and $\mathcal{L}$ and $\mathcal{M}$ two regular holonomic $\mathsf{E}_{Y}$-modules. Then
+$\mathbb{R}\mathcal{H}om_{\mathsf{E}^{\mathbf{R},f}_{\tilde{Y}}}(\mathcal{M}^{\mathbf{R},f},\mathcal{L}^{\mathbf{R},f})[(\dim Y+1)/2]$
+is a perverse sheaf on $\tilde{Y}$, which is locally constant along the fibers of $\gamma\colon\tilde{Y}\rightarrow Y$.
+Proof.
+The statement is local on $Y$, so we may assume that $Y$ is an open subset of $\mathrm{P}^{*}X$ for some complex manifold $X$.
+By [KashiwaraKawaiMicrolocal, Theorem 7.11] and [Andronikof1, Corollaire 5.6.4] there is an isomorphism
+$\mathbb{R}\mathcal{H}om_{\mathsf{E}^{\mathbf{R},f}_{\tilde{Y}}}(\mathcal{M}^{\mathbf{R},f},\mathcal{L}^{\mathbf{R},f})\cong\mathbb{R}\mathcal{H}om_{\mathcal{D}^{\sqrt{v}}_{\mathring{\mathrm{T}}^{*}_{X}(X\times X)}}(\Phi(\mathcal{L},\mathcal{M}^{*}),\mathcal{O}_{\mathring{\mathrm{T}}^{*}_{X}(X\times X)})[\dim X],$
+where
+$\Phi(\mathcal{L},\mathcal{M}^{*})=\mu^{dR}_{\overline{\Delta}}(\mathcal{L}\boxtimes^{\mathcal{E}}\mathcal{M}^{*})$
+is the microlocalization of the $\mathcal{E}^{\sqrt{v}}_{X\times X}$-module $\mathcal{L}\boxtimes^{\mathcal{E}}\mathcal{M}^{*}$ to the Legendrian $\overline{\Delta}=\mathrm{P}^{*}_{X}(X\times X)\subset\mathrm{P}^{*}(X\times X)$.
+By construction the microlocalization functor produces a regular holonomic $D$-module on $\mathring{\mathrm{T}}^{*}_{X}(X\times X)$ which is monodromic along the fibers of $\gamma\colon\mathring{\mathrm{T}}^{*}_{X}(X\times X)\rightarrow\mathrm{P}^{*}_{X}(X\times X)$.
+Applying the solution functor $\mathbb{R}\mathcal{H}om_{\mathcal{D}^{\sqrt{v}}_{\mathring{\mathrm{T}}^{*}_{X}(X\times X)}}(-,\mathcal{O}_{\mathring{\mathrm{T}}^{*}_{X}(X\times X)})[2\dim X]$ we therefore obtain a perverse sheaf, which is locally constant along the fibers of $\gamma$.
+∎
+Corollary 8.3.
+Let $Y$ be a contact manifold and $\mathcal{L}$ and $\mathcal{M}$ two bounded complexes of $\mathsf{E}_{Y}$-modules with regular holonomic cohomology sheaves.
+Then $\mathbb{R}\mathcal{H}om_{\mathsf{E}^{\mathbf{R},f}_{\tilde{Y}}}(\mathcal{M}^{\mathbf{R},f},\mathcal{L}^{\mathbf{R},f})$ is a constructible complex on $\tilde{Y}$, which is locally constant along the fibers of $\gamma\colon\tilde{Y}\rightarrow Y$.
+When one of the modules is regular along a Legendrian submanifold, the corresponding $\mathbb{R}\mathcal{H}om$ can be computed by microlocalization.
+Proposition 8.4.
+Let $Y$ be a contact manifold, $\mathcal{M}$ a regular holonomic $\mathsf{E}_{Y}$-module and $\mathcal{L}$ an $\mathsf{E}_{Y}$-module regular along a Legendrian submanifold $\Lambda\subset Y$.
+Then there is an isomorphism
+$\mathbb{R}\mathcal{H}om_{\mathsf{E}^{\mathbf{R},f}_{\tilde{Y}}}(\mathcal{M}^{\mathbf{R},f},\mathcal{L}^{\mathbf{R},f})|_{\tilde{\Lambda}}\cong\mathbb{R}\mathcal{H}om_{\mathcal{D}^{\sqrt{v}}_{\tilde{\Lambda}}}(\mu_{\Lambda}^{dR}(\mathcal{M}),\mu_{\Lambda}^{dR}(\mathcal{L})).$
+Proof.
+By 8.2
+$\mathcal{F}_{1}=\mathbb{R}\mathcal{H}om_{\mathsf{E}^{\mathbf{R},f}_{\tilde{Y}}}(\mathcal{M}^{\mathbf{R},f},\mathcal{L}^{\mathbf{R},f})|_{\tilde{\Lambda}}[\dim\tilde{\Lambda}]$
+is a perverse sheaf on $\tilde{\Lambda}$, which is locally constant along the fibers of $\gamma\colon\tilde{\Lambda}\rightarrow\Lambda$.
+Similarly, since $\mu_{\Lambda}^{dR}(\mathcal{M})$ and $\mu_{\Lambda}^{dR}(\mathcal{L})$ are $\mathcal{D}^{\sqrt{v}}_{\tilde{\Lambda}}$-modules, which are locally constant along the fibers of $\gamma$,
+$\mathcal{F}_{2}=\mathbb{R}\mathcal{H}om_{\mathcal{D}^{\sqrt{v}}_{\tilde{\Lambda}}}(\mu_{\Lambda}^{dR}(\mathcal{M}),\mu_{\Lambda}^{dR}(\mathcal{L}))[\dim\tilde{\Lambda}]$
+is a perverse sheaf on $\tilde{\Lambda}$, which is locally constant along the fibers of $\gamma$.
+Let $V\in\mathrm{LocSys}(\tilde{\Lambda})$ be a local system on $\tilde{\Lambda}$.
+By the classification of $\mathsf{E}_{Y}$-modules regular along $\Lambda$ we may find a $\mathsf{E}_{Y}$-module $\mathcal{L}_{V}$ regular along $\Lambda$ such that
+$\mu_{\Lambda}^{dR}(\mathcal{L}_{V})\cong\mu_{\Lambda}^{dR}(\mathcal{L})\otimes V,\qquad(\mathcal{L}_{V})^{\mathbf{R},f}\cong\mathcal{L}^{\mathbf{R},f}\otimes V.$
+Therefore, by 3.16 we obtain an isomorphism
+$\gamma_{*}(\mathcal{F}_{1}\otimes V)\cong\gamma_{*}(\mathcal{F}_{2}\otimes V)$
+of constructible complexes on $\Lambda$, natural in $V$.
+In particular, we obtain an isomorphism ${}^{p}\mathcal{H}^{-1}(\gamma_{*}(\mathcal{F}_{1}\otimes V))\cong{}^{p}\mathcal{H}^{-1}(\gamma_{*}(\mathcal{F}_{2}\otimes V))$ of perverse sheaves on $\Lambda$, natural in $V$.
+By Lemma 2.7 it lifts to an isomorphism $\mathcal{F}_{1}\cong\mathcal{F}_{2}$ of perverse sheaves on $\tilde{\Lambda}$.
+∎
+8.2. Construction of the category
+Recall that for a monoidal $\infty$-category $\mathcal{V}$, one has the notion of a $\mathcal{V}$-enriched $\infty$-category $\mathcal{C}$, see [GepnerHaugseng].
+Given two objects $x,y\in\mathcal{C}$ in a $\mathcal{V}$-enriched $\infty$-category we have the Hom object $\mathrm{Hom}_{\mathcal{C}}(x,y)\in\mathcal{V}$.
+When $R$ is a commutative ring and $\mathcal{V}=\mathbf{D}(R)$ is the derived $\infty$-category of chain complexes of $R$-modules, an $R$-linear $\infty$-category is, by definition, a $\mathcal{V}$-enriched $\infty$-category.
+If $\mathcal{C}$ is a sheaf of $R$-enriched $\infty$-categories on a complex manifold $X$, for two global objects $x,y\in\mathcal{C}$ we may define both the usual Hom object $\mathrm{Hom}_{\mathcal{C}}(x,y)\in\mathbf{D}(R)$ as well as the sheaf Hom object $\mathcal{H}om_{\mathcal{C}}(x,y)\in\mathbf{D}(X;R)$.
+In this section for a holomorphic symplectic manifold $S$ we define a sheaf of $\mathbf{C}$-linear $\infty$-categories $\mathbf{D}\mathsf{RH}_{S}$ on $S$ by slightly generalizing the construction of [DAgnoloKashiwara].
+We begin by observing that the set $\mathrm{Lag}_{S}$ of Lagrangian subvarieties in $S$ is directed: given two Lagrangian subvarieties $L_{1},L_{2}\subset S$, their union $L_{1}\cup L_{2}\subset S$ is again a Lagrangian subvariety and it contains both $L_{1}$ and $L_{2}$.
+Let $\mathrm{Lag}^{\mathrm{contact}}_{S}$ be the category of triples $(L,Y,\Lambda)$ of a Lagrangian subvariety $L\subset S$ a contactification $\rho\colon Y\rightarrow U$ of a neighborhood $U$ of $L$ and a lift of $L$ to a Legendrian subvariety $\Lambda\subset Y$.
+By 1.12 the forgetful functor $\mathrm{Lag}^{\mathrm{contact}}_{S}\rightarrow\mathrm{Lag}_{S}$ is an equivalence.
+First, let us fix an object $(L,Y,\Lambda)\in\mathrm{Lag}_{S}^{\mathrm{contact}}$.
+Consider the $\mathbf{C}$-linear algebroid $\mathsf{E}^{\mathbf{R},f}_{\tilde{Y}}$ on $\tilde{Y}$.
+Then the sheaf of presentable $\infty$-categories $\mathbf{D}(\mathsf{E}^{\mathbf{R},f}_{\tilde{Y}})$ on $\tilde{Y}$ is tensored over $\mathbf{D}(\mathbf{C})$, so by [GepnerHaugseng, Corollary 7.4.13] it is $\mathbf{C}$-linear.
+Since $\gamma_{*}\mathsf{E}^{\mathbf{R},f}_{\tilde{Y}}\cong\mathsf{E}_{Y}$, we have a fully faithful functor of $\mathbf{C}$-linear $\infty$-categories
+$\mathbf{D}(\mathsf{E}_{Y})\longrightarrow\gamma_{*}\mathbf{D}(\mathsf{E}^{\mathbf{R},f}_{\tilde{Y}})$
+given by $\mathcal{M}\mapsto\mathcal{M}^{\mathbf{R},f}$.
+Thus, $\mathbf{D}(\mathsf{E}_{Y})$ canonically enhances to a $\mathbf{D}(\mathbf{C}^{\times};\mathbf{C})$-enriched $\infty$-category, so we have two versions of sheaf Hom objects:
+$\mathcal{H}om(\mathcal{M},\mathcal{L})=\mathbb{R}\mathcal{H}om_{\mathsf{E}_{Y}}(\mathcal{M},\mathcal{L}),\qquad\underline{\mathcal{H}om}(\mathcal{M},\mathcal{L})=\mathbb{R}\mathcal{H}om_{\mathsf{E}^{\mathbf{R},f}_{Y}}(\mathcal{M}^{\mathbf{R},f},\mathcal{L}^{\mathbf{R},f}).$
+Let $\mathbf{D}_{\mathrm{rh}}(\mathsf{E}_{Y})\subset\mathbf{D}(\mathsf{E}_{Y})$ be the full subcategory of objects of bounded cohomological amplitude with cohomology sheaves being regular holonomic $\mathsf{E}_{Y}$-modules.
+Let $\mathbf{D}_{\Lambda,\mathrm{rh}}(\mathsf{E}_{Y})\subset\mathbf{D}_{\mathrm{rh}}(\mathsf{E}_{Y})$ be the subcategory of objects supported on a Lagrangian subvariety $\Lambda\subset Y$.
+Set
+$\mathbf{D}\mathsf{RH}_{S,L}=\rho_{*}\mathbf{D}_{\Lambda,\mathrm{rh}}(\mathsf{E}_{Y}),$
+which, as before, is a sheaf of $\mathbf{D}(\mathbf{C}^{\times};\mathbf{C})$-enriched $\infty$-categories.
+Using 8.3 we see that, in fact, $\mathbf{D}\mathsf{RH}_{S,L}$ is enriched over the full subcategory $\mathbf{D}_{\mathrm{lisse}}(\mathbf{C}^{\times};\mathbf{C})$ of local systems on $\mathbf{C}^{\times}$.
+Taking monodromy provides an equivalence of symmetric monoidal $\infty$-categories
+$\mathbf{D}_{\mathrm{lisse}}(\mathbf{C}^{\times};\mathbf{C})\cong\mathbf{D}(\mathbf{C}[T^{\pm 1}])^{\mathrm{conv}}$
+with respect to the convolution monoidal structure on the right-hand side. Since the forgetful functor
+$\mathbf{D}(\mathbf{C}[T^{\pm 1}])^{\mathrm{conv}}\longrightarrow\mathbf{D}(\mathbf{C})$
+is monoidal, we may also regard $\mathbf{D}\mathsf{RH}_{S,L}$ as a sheaf of $\mathbf{C}$-linear $\infty$-categories.
+So far, for a fixed object $(L,Y,\Lambda)\in\mathrm{Lag}^{\mathrm{contact}}_{S}$, we have defined a sheaf of $\mathbf{D}(\mathbf{C}[T^{\pm 1}])^{\mathrm{conv}}$-enriched $\infty$-categories $\mathbf{D}\mathsf{RH}_{S,L}$ on $L$.
+The construction is obviously functorial with respect to inclusions of Lagrangian/Legendrian subvarieties and we set
+$\mathbf{D}\mathsf{RH}_{S}=\operatorname*{colim}_{(L,Y,\Lambda)\in\mathrm{Lag}^{\mathrm{contact}}_{S}}i_{*}\mathbf{D}\mathsf{RH}_{S,L},$
+where $i\colon L\rightarrow S$ is the natural inclusion. We may summarize the constructions of this section as follows.
+Theorem 8.5.
+Let $S$ be a symplectic manifold. There is a sheaf $\mathbf{D}\mathsf{RH}_{S}$ of $\mathbf{D}(\mathbf{C}[T^{\pm 1}])^{\mathrm{conv}}$-enriched $\infty$-categories on $S$ with the following properties:
+  1. (1)
+    It is equipped with a $t$-structure; we denote by $\mathsf{RH}_{S}\subset\mathbf{D}\mathsf{RH}_{S}$ the heart.
+  2. (2)
+    For any two objects $\mathcal{M},\mathcal{L}\in\mathbf{D}\mathsf{RH}_{S}$ the internal Hom sheaf $\underline{\mathcal{H}om}_{\mathbf{D}\mathsf{RH}_{S}}(\mathcal{M},\mathcal{L})\in\mathbf{D}(S;\mathbf{C}[T^{\pm 1}])$ is constructible as a sheaf of complexes of $\mathbf{C}$-vector spaces.
+  3. (3)
+    For any two objects $\mathcal{M},\mathcal{L}\in\mathsf{RH}_{S}$ the internal Hom sheaf $\underline{\mathcal{H}om}_{\mathbf{D}\mathsf{RH}_{S}}(\mathcal{M},\mathcal{L})[\dim S/2]$ is perverse.
+  4. (4)
+    If $L\subset S$ is a Lagrangian submanifold, there is a full subcategory $\gamma_{*}\mathrm{Mod}_{\mathrm{lisse}}(\mathcal{D}^{\sqrt{v}}_{L\times\mathbf{C}^{\times}})\subset\mathsf{RH}_{S}|_{L}$ of half-twisted $D$-modules on $L\times\mathbf{C}^{\times}$, where $\gamma\colon L\times\mathbf{C}^{\times}\rightarrow L$ is the projection.
+In particular, if $L$ is equipped with an orientation data $K_{L}^{1/2}$ and $\lambda\in\mathbf{C}/\mathbf{Z}$, there is a canonical object $\mathsf{M}^{\lambda}_{L}\in\mathsf{RH}_{S}$ supported on $L$ corresponding to the trivial local system on $L$ with monodromy $\exp(2\pi i\lambda)$ around $\mathbf{C}^{\times}$.
+  5. (5)
+    Let $L,M\subset S$ be Lagrangian submanifolds equipped with orientation data $K_{L}^{1/2},K_{M}^{1/2}$ and $\lambda,\mu\in\mathbf{C}/\mathbf{Z}$. Then there is an isomorphism
+    $\underline{\mathcal{H}om}_{\mathbf{D}\mathsf{RH}_{S}}(\mathsf{M}^{\mu}_{M},\mathsf{M}^{\lambda}_{L})[\dim S/2]\cong\phi_{L\cap M}\otimes\mathbf{C}_{\lambda-\mu}$
+    of perverse sheaves equipped with a monodromy operator, where $\mathbf{C}_{\lambda-\mu}$ is the trivial one-dimensional local system equipped with the monodromy operator $\exp(2\pi i(\lambda-\mu))$.
+Proof.
+Let us explain why the sheaf $\mathbf{D}\mathsf{RH}_{S}$ constructed in this section satisfies the asserted properties:
+  1. (1)
+    The natural $t$-structure on the derived $\infty$-category $\mathbf{D}(\mathsf{E}_{Y})$ restricts to a $t$-structure on $\mathbf{D}\mathsf{RH}_{S,L}=\rho_{*}\mathbf{D}_{\Lambda,\mathrm{rh}}(\mathsf{E}_{Y})$.
+Inclusion $i\colon L_{1}\rightarrow L_{2}$ of Lagrangian subvarieties of $S$ induces a $t$-exact functor $i_{*}\mathbf{D}\mathsf{RH}_{S,L_{1}}\rightarrow\mathbf{D}\mathsf{RH}_{S,L_{2}}$, so the $t$-structure descends to the colimit $\mathbf{D}\mathsf{RH}_{S}$.
+The heart $\mathrm{RH}_{S,L}$ of the $t$-structure on $\mathbf{D}\mathsf{RH}_{S,L}$ is given by the sheaf $\rho_{*}\mathrm{Mod}_{\Lambda,\mathrm{rh}}(\mathsf{E}_{Y})$ of regular holonomic $\mathsf{E}_{Y}$-modules supported on $\Lambda$.
+  2. (2)
+    By definition, we may find an object $(L,Y,\Lambda)\in\mathrm{Lag}^{\mathrm{contact}}_{S}$, such that $\mathcal{M},\mathcal{L}\in\mathbf{D}_{\Lambda,\mathrm{rh}}(\mathsf{E}_{Y})$.
+Moreover, we have
+    $\underline{\mathcal{H}om}_{\mathbf{D}\mathsf{RH}_{S}}(\mathcal{M},\mathcal{L})=\mathbb{R}\mathcal{H}om_{\mathsf{E}^{\mathbf{R},f}_{\tilde{Y}}}(\mathcal{M}^{\mathbf{R},f},\mathcal{L}^{\mathbf{R},f})|_{\Lambda\times\{1\}}[-1].$
+    By 8.3 the sheaf of complexes of $\mathbf{C}$-vector spaces $\mathbb{R}\mathcal{H}om_{\mathsf{E}^{\mathbf{R},f}_{\tilde{Y}}}(\mathcal{M}^{\mathbf{R},f},\mathcal{L}^{\mathbf{R},f})$ is constructible, which implies the result.
+  3. (3)
+    Unpacking the notation as in part (2) the relevant claim follows from 8.2.
+  4. (4)
+    Consider the full subcategory $\mathrm{RH}_{S,L}=\rho_{*}\mathrm{Mod}_{\Lambda,\mathrm{rh}}(\mathsf{E}_{Y})\subset\mathrm{RH}_{S}|_{L}$. Then by 3.16 we have an equivalence
+    $\mu_{\Lambda}^{dR}\colon\mathsf{RH}_{S,L}\xrightarrow{\sim}\gamma_{*}\mathrm{Mod}_{\mathrm{lisse}}(\mathcal{D}^{\sqrt{v}}_{L\times\mathbf{C}^{\times}}).$
+  5. (5)
+    Consider a contactification $\rho\colon Y\rightarrow S$ of a neighborhood of $L\cup M$ and the corresponding Legendrian submanifolds $\Lambda_{L},\Lambda_{M}\subset S$ lifting $L,M\subset S$.
+We will identify $\mathcal{D}_{L}^{\sqrt{v}}$-modules with $\mathcal{D}_{L}$-modules using the orientation data $K_{L}^{1/2}$.
+Then by construction we have
+    (20) $i_{L}^{*}\mu_{\Lambda_{L}}^{dR}(\mathsf{M}^{\lambda}_{L})\cong\mathcal{O}_{L}\otimes\mathbf{C}_{\lambda}.$
+    Moreover, by 7.2 we have
+    (21) $i_{L}^{*}\mu_{\Lambda_{L}}^{dR}(\mathsf{M}^{\mu}_{M})\cong\phi^{\mathcal{D}}_{L\cap M}\otimes\mathbf{C}_{\mu}.$
+    We have
+    $\underline{\mathcal{H}om}_{\mathbf{D}\mathsf{RH}_{S}}(\mathsf{M}^{\mu}_{M},\mathsf{M}^{\lambda}_{L})=\mathbb{R}\mathcal{H}om_{\mathsf{E}^{\mathbf{R},f}_{\tilde{Y}}}(\mathsf{M}^{\mu,\mathbf{R},f}_{M},\mathsf{M}^{\lambda,\mathbf{R},f}_{L})|_{L\times\{1\}}[-1].$
+    Plugging in (20) and (21) into 8.4 we have
+    $\mathbb{R}\mathcal{H}om_{\mathsf{E}^{\mathbf{R},f}_{\tilde{Y}}}(\mathsf{M}^{\mu,\mathbf{R},f}_{M},\mathsf{M}^{\lambda,\mathbf{R},f}_{L})|_{L\times\{1\}}[-1]\cong\mathbb{R}\mathcal{H}om_{\mathcal{D}_{L}}(\phi^{\mathcal{D}}_{L\cap M}\otimes\mathbf{C}_{\mu},\mathcal{O}_{L}\otimes\mathbf{C}_{\lambda}).$
+    Using the Verdier self-duality of $\phi^{\mathcal{D}}_{L\cap M}$ (see [BBDJS, Equation (6.5)]) we get the result.
+Remark 8.6.
+For $\mathcal{M},\mathcal{L}\in\mathbf{D}\mathsf{RH}_{S}$ we have
+$\mathcal{H}om_{\mathbf{D}\mathsf{RH}_{S}}(\mathcal{M},\mathcal{L})\cong\mathrm{Hom}_{\mathbf{C}[T^{\pm 1}]}(\mathbf{C}_{0},\underline{\mathcal{H}om}_{\mathbf{D}\mathsf{RH}_{S}}(\mathcal{M},\mathcal{L}))[1],$
+where $\mathbf{C}_{0}$ is the one-dimensional $\mathbf{C}[T^{\pm 1}]$-module, where $T$ acts by the identity, and the $\mathrm{Hom}$ on the right is implicitly derived.
+The authors of [DAgnoloKashiwara] consider the underlying homotopy category of $\mathbf{D}\mathsf{RH}_{S}$ (as well as $\mathrm{RH}_{S}$) considered as a $\mathbf{C}$-linear category via $\mathcal{H}om$.
+8.3. Microlocal Riemann–Hilbert correspondence
+For a contact manifold $Y$ we denote by $\mu\mathrm{sh}_{\tilde{Y}}$ the sheaf of $\mathbf{C}$-linear $\infty$-categories of microsheaves on $\tilde{Y}$ constructed in [NadlerShende, CKNS1].
+We denote by $\mathbb{P}\mathrm{erv}_{Y}\subset\gamma_{*}\mu\mathrm{sh}_{\tilde{Y}}$ the subcategory of perverse microsheaves constructed in [Waschkies1, CKNS1].
+Note that we view $\mathbb{P}\mathrm{erv}_{Y}$ as a sheaf of abelian categories on $Y$.
+The microlocal Riemann–Hilbert equivalence established in [Andronikof3, Waschkies2, CKNS2] produces an equivalence
+$\mu\mathrm{RH}\colon\mathbb{P}\mathrm{erv}^{\mathrm{op}}_{Y}\longrightarrow\mathrm{Mod}_{\mathrm{rh}}(\mathsf{E}_{Y}).$
+Proposition 8.7.
+Let $\mathcal{F},\mathcal{G}$ be two perverse microsheaves on $Y$. Then there is an isomorphism
+$\mathcal{H}om_{\mu\mathrm{sh}_{\tilde{Y}}}(\mathcal{F},\mathcal{G})\cong\mathbb{R}\mathcal{H}om_{\mathsf{E}^{\mathbf{R},f}_{\tilde{Y}}}(\mu\mathrm{RH}(\mathcal{G})^{\mathbf{R},f},\mu\mathrm{RH}(\mathcal{F})^{\mathbf{R},f})$
+of complexes of sheaves on $\tilde{Y}$.
+Proof.
+Consider a cover of $Y$ by open subsets $V_{i}\subset\mathrm{P}^{*}X_{i}$, where $X_{i}$ is a complex manifold.
+The right-hand side of the required isomorphism is a shift of a perverse sheaf by 8.2.
+Therefore, it is enough to construct the isomorphism on each $V_{i}$ and check that these isomorphisms are compatible on each double intersection $V_{ij}=V_{i}\cap V_{j}$.
+As before, denote $\tilde{V}_{i}=\gamma^{-1}(V_{i})\subset\tilde{Y}$ and $\tilde{V}_{ij}=\gamma^{-1}(V_{ij})\subset\tilde{Y}$.
+To simplify the notation, we assume that each $X_{i}$ is equipped with a choice of the square root line bundle $K^{1/2}_{X_{i}}$.
+Consider the algebra $\mathcal{E}^{\sqrt{v},\mathbf{R}}_{X_{i}}\supset\mathcal{E}^{\sqrt{v},\mathbf{R},f}_{X_{i}}$ of holomorphic microlocal operators constructed in [SatoKawaiKashiwara].
+For a $\mathcal{E}^{\sqrt{v}}_{X_{i}}$-module $\mathcal{M}$ let
+$\mathcal{M}^{\mathbf{R}}=\mathcal{E}^{\sqrt{v},\mathbf{R}}_{X_{i}}\otimes_{\gamma^{-1}\mathcal{E}^{\sqrt{v}}_{X_{i}}}\gamma^{-1}\mathcal{M}$
+be the corresponding $\mathcal{E}^{\sqrt{v},\mathbf{R}}_{X_{i}}$-module.
+Consider also the microsheaf $\mu K^{1/2}_{X_{i}}\in\mu\mathrm{sh}_{\mathring{\mathrm{T}}^{*}X_{i}}$ obtained by microlocalizing the sheaf of holomorphic sections of $K_{X_{i}}^{1/2}$.
+It is naturally a $\mathcal{E}^{\sqrt{v},\mathbf{R}}_{X_{i}}$-module.
+To simplify the notation, we will drop certain obvious restrictions to open subsets.
+By [CKNS1, Example 5.3] on $V_{ij}$ we may obtain a quantized contact transformation given by a microsheaf kernel
+$\mathcal{K}_{ij}\in\mu\mathrm{sh}_{\overline{\tilde{V_{i}}}\times\tilde{V_{j}}}$
+and by [Waschkies2, Theorem 3.4.2], possibly refining the cover $\{V_{i}\}$, we may choose a nondegenerate section of the $(\mathcal{E}^{\sqrt{v},\mathbf{R}}_{X_{j}},\mathcal{E}^{\sqrt{v},\mathbf{R}}_{X_{i}})$-bimodule
+$\mathcal{H}_{ij}=\mathcal{H}om_{\mu\mathrm{sh}_{\overline{\tilde{V_{i}}}\times\tilde{V_{j}}}}(\mathcal{K}_{ij}[-\dim X_{i}],K^{1/2}_{X_{i}}\boxtimes K^{1/2}_{X_{j}})$
+which provides isomorphisms
+$\mathcal{E}^{\sqrt{v},\mathbf{R}}_{X_{j}}\xrightarrow{\sim}\mathcal{E}^{\sqrt{v},\mathbf{R}}_{X_{i}}$
+of algebras on $\tilde{V}_{ij}$ and
+$s_{ij}\colon\mathcal{K}_{ij}\circ\mu K^{1/2}_{X_{i}}\xrightarrow{\sim}\mu K^{1/2}_{X_{j}}$
+of $\mathcal{E}^{\sqrt{v},\mathbf{R}}_{X_{j}}$-modules on $\tilde{V}_{ij}$.
+Using quantized contact transformations we may glue the sheaves of algebras $\mathcal{E}^{\sqrt{v},\mathbf{R}}_{X_{i}}$ into an algebroid $\mathsf{E}^{\mathbf{R}}_{\tilde{Y}}$ on $\tilde{Y}$.
+Moreover, for any two regular holonomic $\mathsf{E}_{Y}$-modules $\mathcal{M}_{1},\mathcal{M}_{2}$, the natural morphism
+$\mathbb{R}\mathcal{H}om_{\mathsf{E}^{\mathbf{R},f}_{\tilde{Y}}}(\mathcal{M}_{1}^{\mathbf{R},f},\mathcal{M}_{2}^{\mathbf{R},f})\longrightarrow\mathbb{R}\mathcal{H}om_{\mathsf{E}^{\mathbf{R}}_{\tilde{Y}}}(\mathcal{M}_{1}^{\mathbf{R}},\mathcal{M}_{2}^{\mathbf{R}})$
+is an isomorphism: the claim is local, so it reduces to the same claim on $V_{i}$, where it is shown in [Andronikof1, Corollaire 5.6.4]. Thus, we have to construct an isomorphism
+$\mathcal{H}om_{\mu\mathrm{sh}_{\tilde{Y}}}(\mathcal{F},\mathcal{G})\cong\mathbb{R}\mathcal{H}om_{\mathsf{E}^{\mathbf{R}}_{\tilde{Y}}}(\mu\mathrm{RH}(\mathcal{G})^{\mathbf{R}},\mu\mathrm{RH}(\mathcal{F})^{\mathbf{R}}).$
+The perverse microsheaves $\mathcal{F},\mathcal{G}\in\mathbb{P}\mathrm{erv}_{Y}$ are given in terms of the cover $\{V_{i}\}$ as follows:
+  • On $V_{i}$ we have objects $\mathcal{F}_{i},\mathcal{G}_{i}\in\mu\mathrm{sh}_{\mathring{\mathrm{T}}^{*}X_{i}}$.
+  • On $V_{ij}$ we are given isomorphisms $f_{ij}\colon\mathcal{K}_{ij}\circ\mathcal{F}_{i}\xrightarrow{\sim}\mathcal{F}_{j}$ and $g_{ij}\colon\mathcal{K}_{ij}\circ\mathcal{G}_{i}\xrightarrow{\sim}\mathcal{G}_{j}$.
+By definition (see [CKNS2, Theorem 7.7]) we may describe the $\mathsf{E}^{\mathbf{R}}_{\tilde{Y}}$-module $\mu\mathrm{RH}(\mathcal{F})^{\mathbf{R}}$ (and similarly for $\mu\mathrm{RH}(\mathcal{G})^{\mathbf{R}}$) in terms of the cover $\{V_{i}\}$ as follows:
+  • On $V_{i}$ we have the $\mathcal{E}^{\sqrt{v},\mathbf{R}}_{X_{i}}$-module $\mathcal{H}om_{\mu\mathrm{sh}_{\mathring{\mathrm{T}}^{*}X_{i}}}(\mathcal{F}_{i},\mu K^{1/2}_{X_{i}})$.
+  • On $V_{ij}$ we are given the isomorphism
+    $\mathcal{H}om_{\mu\mathrm{sh}_{\mathring{\mathrm{T}}^{*}X_{i}}}(\mathcal{F}_{i},\mu K^{1/2}_{X_{i}})\xrightarrow{\sim}\mathcal{H}om_{\mu\mathrm{sh}_{\mathring{\mathrm{T}}^{*}X_{j}}}(\mathcal{K}_{ij}\circ\mathcal{F}_{i},\mathcal{K}_{ij}\circ\mu K^{1/2}_{X_{i}})\xrightarrow{(f_{ij},s_{ij})}\mathcal{H}om_{\mu\mathrm{sh}_{\mathring{\mathrm{T}}^{*}X_{j}}}(\mathcal{F}_{j},\mu K^{1/2}_{X_{j}}).$
+We are ready to construct the required isomorphism of (shifted) perverse sheaves. On $V_{i}$ we have a natural morphism of complexes of sheaves
+$\mathcal{H}om_{\mu\mathrm{sh}_{\mathring{\mathrm{T}}^{*}X_{i}}}(\mathcal{F}_{i},\mathcal{G}_{i})\longrightarrow\mathbb{R}\mathcal{H}om_{\mathcal{E}^{\mathbf{R}}_{X_{i}}}(\mathcal{H}om_{\mu\mathrm{sh}_{\mathring{\mathrm{T}}^{*}X_{i}}}(\mathcal{G}_{i},\mu K^{1/2}_{X_{i}}),\mathcal{H}om_{\mu\mathrm{sh}_{\mathring{\mathrm{T}}^{*}X_{i}}}(\mathcal{F}_{i},\mu K^{1/2}_{X_{i}}))$
+given by the functoriality of $\mathcal{H}om_{\mu\mathrm{sh}_{\mathring{\mathrm{T}}^{*}X_{i}}}(-,\mu K^{1/2}_{X_{i}})$.
+The claim that this morphism is an isomorphism is shown in the proof of [Waschkies2, Theorem 3.6.5] by reducing to the case of ordinary perverse sheaves on $X_{i}$ using a generic position argument.
+The compatibility of these isomorphisms on $V_{ij}$ follows from the commutative diagram
+${\mathcal{H}om(\mathcal{F}_{i},\mathcal{G}_{i})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathbb{R}\mathcal{H}om_{\mathcal{E}^{\mathbf{R}}_{X_{i}}}(\mathcal{H}om(\mathcal{G}_{i},\mu K^{1/2}_{X_{i}}),\mathcal{H}om(\mathcal{F}_{i},\mu K^{1/2}_{X_{i}}))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{H}om(\mathcal{K}_{ij}\circ\mathcal{F}_{i},\mathcal{K}_{ij}\circ\mathcal{G}_{i})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{(f_{ij},g_{ij})}$${\mathbb{R}\mathcal{H}om_{\mathcal{E}^{\mathbf{R}}_{X_{i}}}(\mathcal{H}om(\mathcal{K}_{ij}\circ\mathcal{G}_{i},\mathcal{K}_{ij}\circ\mu K^{1/2}_{X_{i}}),\mathcal{H}om(\mathcal{K}_{ij}\circ\mathcal{F}_{i},\mathcal{K}_{ij}\circ\mu K^{1/2}_{X_{i}}))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{(f_{ij},g_{ij})}$${\mathcal{H}om(\mathcal{F}_{j},\mathcal{G}_{j})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathbb{R}\mathcal{H}om_{\mathcal{E}^{\mathbf{R}}_{X_{j}}}(\mathcal{H}om(\mathcal{G}_{j},\mathcal{K}_{ij}\circ\mu K^{1/2}_{X_{i}}),\mathcal{H}om(\mathcal{F}_{j},\mathcal{K}_{ij}\circ\mu K^{1/2}_{X_{j}}))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{s_{ij}}$${\mathbb{R}\mathcal{H}om_{\mathcal{E}^{\mathbf{R}}_{X_{j}}}(\mathcal{H}om(\mathcal{G}_{j},\mu K^{1/2}_{X_{j}}),\mathcal{H}om(\mathcal{F}_{j},\mu K^{1/2}_{X_{j}})).}$
+Now let $\Lambda\subset Y$ be a Legendrian submanifold equipped with a square root line bundle $K_{\Lambda}^{1/2}$.
+In particular, the corresponding homogeneous Lagrangian $\tilde{\Lambda}\subset\tilde{Y}$ in the symplectization is equipped with a square root line bundle $K_{\tilde{\Lambda}}^{1/2}$ and so, by the classification of simple $\mathsf{E}_{Y}$-modules given in 3.16, there is a simple $\mathsf{E}_{Y}$-module supported $\mathsf{M}_{\Lambda}$ on $\Lambda$ with
+$\mu_{\Lambda}^{dR}(\mathsf{M}_{\Lambda})\cong K_{\tilde{\Lambda}}^{1/2}$
+the trivial local system. Let $\mathcal{M}_{\Lambda}\in\mathbb{P}\mathrm{erv}_{Y}$ be the perverse microsheaf such that
+$\mu\mathrm{RH}(\mathcal{M}_{\Lambda})\cong\mathsf{M}_{\Lambda}.$
+Theorem 8.8.
+Let $Y$ be a contact manifold and $\Lambda_{1},\Lambda_{2}\subset Y$ two Legendrian submanifolds equipped with square root line bundles $K_{\Lambda_{i}}^{1/2}$. There is an isomorphism
+$\mathcal{H}om_{\mu\mathrm{sh}_{\tilde{Y}}}(\mathcal{M}_{\Lambda_{1}},\mathcal{M}_{\Lambda_{2}})[\dim\tilde{Y}/2]\cong\phi_{\tilde{\Lambda}_{2}\cap\tilde{\Lambda}_{1}}.$
+Proof.
+We use a dummy variable trick as in [KashiwaraKawaiHolonomic, Appendix A] to reduce the case of an arbitrary contact manifold to the case of a contactification of a symplectic manifold.
+The symplectic manifold $\tilde{Y}$ is exact and so it admits a contactification $\tilde{\rho}\colon\tilde{Y}\times\mathbf{C}\rightarrow\tilde{Y}$ whose symplectization is $\tilde{Y}\times\mathrm{T}^{*}\mathbf{C}^{\times}$.
+The Lagrangian submanifolds $\tilde{\Lambda}_{1},\tilde{\Lambda}_{2}\subset\tilde{Y}$ lift to Legendrians $\tilde{\Lambda}_{1}\times\{0\},\tilde{\Lambda}_{2}\times\{0\}\subset\tilde{Y}\times\mathbf{C}$ and hence lift to homogeneous Lagrangian submanifolds $\tilde{\Lambda}_{1}\times\mathbf{C}^{\times},\tilde{\Lambda}_{2}\times\mathbf{C}^{\times}\subset\tilde{Y}\times\mathrm{T}^{*}\mathbf{C}^{\times}$.
+The shifted constant local system $\mathbf{C}_{\mathbf{C}^{\times}}[1]$ on $\mathbf{C}^{\times}$ defines a perverse microsheaf on $\mathrm{T}^{*}\mathbf{C}^{\times}$ and
+$\mathcal{M}_{\tilde{\Lambda}_{i}\times\mathbf{C}}\cong\mathcal{M}_{\Lambda_{i}}\boxtimes\mathbf{C}_{\mathbf{C}^{\times}}[1].$
+Moreover, we have
+$\mathcal{H}om_{\mu\mathrm{sh}_{\tilde{Y}}}(\mathcal{M}_{\Lambda_{1}},\mathcal{M}_{\Lambda_{2}})\cong\mathcal{H}om_{\mu\mathrm{sh}_{\tilde{Y}\times\mathrm{T}^{*}\mathbf{C}^{\times}}}(\mathcal{M}_{\tilde{\Lambda}_{1}\times\mathbf{C}},\mathcal{M}_{\tilde{\Lambda}_{2}\times\mathbf{C}})|_{\tilde{Y}}[-1].$
+By 8.7
+$\mathcal{H}om_{\mu\mathrm{sh}_{\tilde{Y}\times\mathrm{T}^{*}\mathbf{C}^{\times}}}(\mathcal{M}_{\tilde{\Lambda}_{1}\times\mathbf{C}},\mathcal{M}_{\tilde{\Lambda}_{2}\times\mathbf{C}})\cong\mathbb{R}\mathcal{H}om_{\mathsf{E}^{\mathbf{R},f}_{\tilde{Y}\times\mathrm{T}^{*}\mathbf{C}^{\times}}}(\mathsf{M}_{\tilde{\Lambda}_{2}\times\mathbf{C}^{\times}},\mathsf{M}_{\tilde{\Lambda}_{1}\times\mathbf{C}^{\times}}).$
+By 8.5(5) we obtain an isomorphism
+$\mathbb{R}\mathcal{H}om_{\mathsf{E}^{\mathbf{R},f}_{\tilde{Y}\times\mathrm{T}^{*}\mathbf{C}^{\times}}}(\mathsf{M}_{\tilde{\Lambda}_{2}\times\mathbf{C}^{\times}},\mathsf{M}_{\tilde{\Lambda}_{1}\times\mathbf{C}^{\times}})[\dim\tilde{Y}/2+1]\cong\phi_{\tilde{\Lambda}_{2}\cap\tilde{\Lambda}_{1}}$
+which concludes the proof. ∎
+[…]